A Complete Bibliography of Publications in Transactions of the American Fisheries

Society: 1970–1979

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

09 June 2023 Version 1.02

Title word cross-reference

131 [Tar71]. ^{14} [GPT70, GTP71, GTP72, MRSK70]. _{2} [RWC73]. _{50} [RS75]. α [FMS79]. F_{1} [Gun73]. × [KBR71, Wil75].

-ethanol [Mar74]. -Methyltestosterone [FMS79].

1-dichloro-2 [DZ72]. 1.6-ha [Ell74]. 1016 [HSF75]. 1108 [Hog70]. 1254 [SHF74]. 1967 [Hig70]. 1969 [Ano70c, Ano70f]. 1970 [Ano71a]. 1971 [Ano72c]. 1972 [Ano73a]. 1973 [Ano74c, Ano74d, Ano74f]. 1974 [Ano74b, Ano75c, Ano75d, MP75]. 1975 [Ano75b, Ano75a, Ric76b].

2 [CWW70, DM79]. 2- [Mar74]. 2-bis [DZ72]. 21.3-Meter [EP73]. 222 [CB72].

4 [BBAN75, CWW70, DM79]. 4-D [CWW70, DM79]. 4-naphthoquinone [Mar70]. 4’-Tetrachlorobiphenyl [BBAN75].

5-hydroxy-1 [Mar70].

6.1-Meter [KJ78].

8-Tetrachlorodibenzo-p-dioxin [HN77, MNL79]. 80-Hectare [CSG^{+}78].

96-Hour [RS75].

Abater(R) [Tsa78]. abdominalis [Med79]. Ability [Ell74]. Abnormalities [Dah70]. Absence [HJB^{+}78]. Absorption [LS73b]. Abudefduf [Med79]. Abundance [AS71b, Bar71, BH75, Bro77, Cla77, CCN75, Elr74, For74, GS72a, HW77, HS75c, HCP76, IWR78, KTJ75, LM71b, Mou74, Ogd70, PB73, SKK79, Sta79b, TM74, TC72, WH75]. Accelerated [BBAN75]. Acclimation [AS71a, HBM79, HH79c, MER77, Ott74, RS79]. Accumulation [Har78, JS77, KS71, PGP77, Tar71, WWLC77, WW71b]. Accurately [HKB76]. Acetylcholinesterase [GHC^{+}79, Hog71]. AChE [Cop72]. Acid [Bow78, HGJ75, KB75, LC70, MA75, Sim72, SDW78, Wea74, YS79]. Acidosis [Cai71]. Acids [DH70]. Acipenser [HGM78]. Acoustic [Abb72, Cha76]. Acquisition [HC73b]. Acrylamide [BR75, Par70a]. Acrylic [LM74]. Action [Han74]. Activated [DMB76]. Active [DZ72]. Activities [Zau70]. Activity [Byr71, Cam71, DWB75, DK75, EBM^{+}75, Ell76, LS73a, MSS73, MD72, MS70b, RC76c, RW79, SPSF72, SN76, WH71]. Actual [Rad73]. Acute [JAKG78, Kaw73, LL70, LHL77, RK76, TRS78, Voy75, MBN72]. Acutely [CS78, MJ72]. Adaptability [MG70]. Adaption [Zau70]. Added [Par73a]. Addendum [Ano70a, Fis70]. Adhesiveness [Wal76]. Adipose [OC75]. Adirondack [Hut71]. Adult [Abb74, Bea74, BCSS76, Bul70, CC78, CS78, Cha78a, DMT77, Gar73, GH77, Rei75, RW79, Sla71, SHH76, Tar71, WL70]. Adult-to-Smolt [Cha78a]. Adults [HSF75, HBD78]. Aegla [Bur72a]. Aeration [BT79, Fas73]. Aerially [CP75]. Aerially-Applied [CP75]. Aerococcus [RGBL79]. Aeromonas [LA71]. aestivalis [LL77a, Bur74, Chi72b, MP78]. Affecting [AS71b, NBS76, Nev75a]. affinis [CRC^{+}76, Smi72b, SMDN76]. Africa [TG79]. After [OC74, Bai78, BLCD75, HH79c, Hög73, MLL71, MBT70, PKP74, Spe74]. Agassiz [Mat73]. Age [Alt76, AB70, Arm71, ALTA77, Bai72, Ban73, Bar78, BBR79, BTL75, Gri78b, HH79b, HP74, IWF75, Joh70, Kai73a, KM74, Kos78, LL75b, MH77, MH78a, MR74a, MD70, Nic71, Ogd70, OK78, OM75, PW75b, Pri77, Pri71, Pri75a, Ric73, Ric76b, Sav78, SS75, SH78, Sta78, Sta72, Sta79b, TL76, TD76, Wol77, WM78, WW76b]. Age-0 [BBR79, Sta79b]. Age-Class [Sav78]. Age-Length [LL75b]. Agent [Wal76]. Ages [Hav73c, Jen70].

Aggregation [Abb72, Smi72a]. Aggregations [Dav71, JH77]. Agnes [Hoo75]. Agricultural [AB75, GA77]. Air [NBS76, NB76]. Air-Supersaturated [NBS76, NB76]. Alabama [SDKT79, CL75, Elr71, Mat73, MR74b, PSG^{+}74, Tat73]. Alaska [Alt71, Alt79, Arm71, CH76, Gar73, KB77, LL77b, Qua72, RM71, Val74]. Alaskan [Alt76, MRK79]. Albama [KDS79]. Alberta [DME77, SML75, TD76]. Albinism [HR72]. Aldrin [KS71]. Aleutian [Val74]. Alevins [DB70, Ell70, HBD78]. Alewife [Coo78, DM73, Eds70a, Gan76, Hut71, Kis74, OKR76, SPSF72, SB78, Smi70a, SC71]. Alewives [Dom73, Hav73b, Lin76, Ric75a, Ric75b, RW79, Spe74, Tyu74, Wag72]. Algae [LS73b]. Algorithm [TL78]. Alkaline [LC70]. Alkalinity [RA72b]. Alkylate [HS71b]. All-Male [Gue75]. Allele [Eck73]. Allelic [UF78]. Allen [Mag73, Str73]. Alloperla [Ell70]. Along [PSG^{+}74, Mes76]. Alosa [Bur74, Chi72a, Chi72b, Coo78, Dom73, DM73, Gan76, Hav73b, Hut71, Kis74, Leg72, Lin76, LL77a, MP78, OKR76, Ric75a, RW79, SC71, Tyu74]. alosae [KG72]. Alpine [WV73]. alpinus [MLK77]. Alteration [Odu70]. Alterations [KO74]. Altered [SK71]. Alternate [For74]. Altitude [RWC73]. Alum [Boy79a]. Aluminum [Boy79a, FE71, Fre73]. alutus [CH76]. Amargosa [Nai75]. Ambicoloration [SW75]. Ambient [WCH71]. Ambloplites [Bou72, GH79, LWO76]. ambloplitis [Ant74]. Amchitka [Val74]. America [MG70]. American [Ano70c, Ano70f, Ano71a, Ano72c, Ano73a, Ano74e, Ano74f, MP75, Ano70b, Ano72a, Bri74, BBR78, Bur72a, Chi72a, Coo70, CCN75, Gun71, Hur73, Leg72, MJN72, Mat78, Mor71, Mos70, Ogd70, RGBL79, Wat70, WAMR79]. americana [MR74a, MUR^{+}76, MP78, Rei72]. americanus [BPL75, BCP76, CE76, Coo70, CCN75, Fra73a, Fra73b, Fra74a, Lev74, LPH70, Lux73, Mor71, RGBL79, UM72, VD79, WAMR79]. Amino [DH70]. Aminonaphthylamidase [Bou79]. Ammodytes [WAMR79]. Ammonia [Boy79b, BK77, CLB^{+}73, RS79, TRS78, WHW78]. Among [Moy73, Spa70, BCSS76, ES76, Man72b, ROH79]. Amount [WFS78]. Amounts [AFS79]. Amphipods [All72]. Amplifier [DD74, GCD^{+}77]. Anabolic [BS73]. Anadromous [Kis74, MRK79, MC78, Ric75a, Tyu74]. Analogy [EBM^{+}75]. analostanus [GB78]. Analyses [Che75]. Analysis [BR75, CK71, DPRC78, ES76, Eck73, HL78, HL76, Hog70, Jen71, LKT79, MS78, Nym70, PHF^{+}77, PLZ75, Set70, TB79, You76]. Anchoa [MMS75]. Anchor [BM72, WL70]. Anchovies [Sch70]. Anchovy [MMS75, Tal73]. Anderson [Mag73, Str73]. Androgen [SH74a]. Androgenetic [Sta76]. Androgens [Gue75]. Anesthetic [Mor70a]. Angler [Cla75, KC72]. Anglers [DH78, You74, Reg71a]. Angling [Elr71, PB73, Rad73]. Anguilla [Hur73, Ogd70]. Anhydrous [CLB^{+}73]. Animals [JAKG78, SND71]. Anne [DME77]. Annotated [KB71]. Annual [Ano70c, Ano70f, Ano71a, Ano72c, Ano73a, Ano74f, Etn72, KH72, LS73a, MP75, Mom78, PW75a, ST77]. annularis [AR71, LMG76]. Annulus [Cob70, Gri75, KGJ74]. Anomalies [Gun73, HBH74, OC75]. Antigenic [Arm74]. Antigens [SP76, TSW72]. antillarum [Dav71]. Antimycin [Kaw73, LDM71, MD72, MB75a, RF74]. Antipredator [Far77]. apache [Alc76]. Apparatus [Jah72, WCH71]. Apparent [SL73]. Application [All73, And73, Beh70, Cob73, JKS78, Mag73, OMS78, Par70b, Smi73b, Smi73a, Str73, UF78, WH75]. Applications [Lac75, Lay71, SH75, TB78]. Applied [CP75, HM74b]. Approach [MV70]. April [Ano75a]. Aqua [DM79]. Aqua-Kleen [DM79]. Aquarium [MLT76, MGW76, PS76, WP76]. Aquarium-Held [MLT76]. Aquatic [BS75, CR73, DZ72, GC72, HGJ75, JAKG78, KDR74, Kem70, LJH76, MS70a, MWHH72, Mit78, Neb72, RA72b, Rod73]. Archives [Gun71]. Archoplites [IWF75, MG70, MMB74]. Arctic [KS77, RM71]. arcticus [KS77, RM71]. Area [Bar71, EP73, WJ78]. Areas [RC76a, Smi76]. Arenicola [SN76]. Arguello [SM76a]. Argulus [KG72]. argus [Dav71, Dav75, Dav78]. Arizona [Joh70, Joh71, MJRW70]. Arkansas [Clo75, NK73, OC74, OK78]. Arlington [Ano73a]. Aroclor(R) [HSF75, SHF74]. array [GS74]. artedii [AS71b, EC70b, Eng76, SM76b]. Artificial [Ben74, DCA71, Fas73, Fas71, GGG75, Hea72, Jac71, KW71, Lac73a, LJH76, SRM76, Wal76, WWO73]. Asellus [Ose77]. Ash [CRC^{+}76, DB78]. Asian [BBR78]. Aspects [ER72, MS74b, WAMR79]. Assessing [Hor77]. Assessment [CH78a, GM79, Jen78, MW79]. Assimilation [Bow78, FB79]. Assiniboine [Gle78]. Associated [BKS76, Bur72b, LM71b, Wal72, Wor75]. Associations [ES76]. Assumptions [PHF^{+}77, Wal75a]. Astoria [Ano71a]. Asymmetry [AFS79]. Asymptomatic [AP72]. Athabasca [SML75]. Atherinidae [CM70a]. Atkins [MR79]. Atlantic [MW79, RH74, AS79, ASH75, Bar78, Dra70, FMD77, Fui79, Gib78, Ket75, KH78, KG72, KG73, KGJ74, LM71b, MMS75, Mes76, MM77, Nic71, RSL70, Rob70, SH72, WWLC77, War76, War78, WJ78, Wea74, Wis72, YHP77]. Atmospheric [Sch74a]. Atomic [MHM77]. ATPase [AZM75, Zau70]. atratulus [FT76]. atromaculatus [BM71]. Attachment [Par73a, PH72, VS79]. Attempt [Fra74b]. Attempts [RI71]. Attracting [Wic73, WWO73]. Attraction [KW71]. Attractors [Wil78]. audens [CM70a, EB76, HSS71, LMG76, MH78b]. Auditor’s [Ano72a]. auratus [BS73, GPT70, GTP71, JBR71, SH74a, SNDM77]. aurea [DS78, Gue75, RS79]. auritus [BM78b]. aurorubens [Bor71, Gri78b]. Automated [Rod73]. Availability [Fra74a]. Available [Tan70]. Avoid [Fra74b]. Avoidance [Dea78, Kyn74, LCL78]. aztecus [CP75, DLM72, GZ73, RA72a, RB70, Wat76].

B [Mag73, Str73, LL75b, AS79]. Back [RM73]. Bacteria [Sch74b]. Bacterial [SP76]. Bacteriophage [TSW72]. Bags [Kem70]. Bahama [Lev73]. baileyi [MR74b, NE74]. bairdi [FW77, LL75b]. Bait [BS75, ED73, HA76]. Bait-Size [HA76]. Baiting [Jes77]. Baja [SM76a]. Balance [SC71, Sta74, Tan70]. bangwelensis [Gri77]. Bangweulu [Gri75, Gri77, TG79]. Bank [Lux73]. Barrows [HD70, Hav74a]. Based [LKT79]. Basic [FE71]. Basin [Sim72]. Basins [HS75c]. Basis [Pat77].

Basket [DCA71, Sel71]. Basket-Type [DCA71]. Bass [Ant74, BH76a, BBR79, Bec72, BG72, BG75, Bou72, Bul70, CH72, Car73, CCM79, Cla75, Cla77, Clo75, Cob71, Cob72, DC79, DE75, DMT77, ES72, Ell76, Eng77, ED73, FB77, FB78, FL78a, FF74, GH79, GB72, HL78, Hai73, HC73a, KF74, KBR71, Kil78, KDS79, KW72b, LWO76, LHK^{+}74, LCG77, MR79, MB75b, MKK73, MUR^{+}76, MP78, MH78b, Mun70, MM77, Nev75a, NK73, OHC79, OK78, Ort79, OM75, Pri70, Pri71, Pri75a, Rue77, Sav78, SDKT79, SCR78, SA77, Smi76, Ste77, TC72, VW74, Wal73a, Wal75a, WL75, Win77, WH71, Wis74, ZD76, HA75]. Basses [VR75]. batrachus [Mus77]. Bay [AB70, Ban73, Dev78, KF74, LS73a, MMS75, MF75, Nob72, Pec74, PT70, RSD74, ST77, WH75, EP73, LM76, MMS75, Mes75, MRK79, MB75b, MKK73, MP77, SC79]. Bayer [Kaw73, KDR74]. Bays [DP73, RA72a]. Beach [GW72]. Bean [SSH72]. Bear [GC72]. Bearing [SM76a]. Beaver [NK73]. Beds [CA73, RH75]. Beef [AC72]. before [Bai78]. Beginnings [Reg74]. Behavior [BM77b, CH76, Coo70, Deg70, DR75, Ell74, Far77, GM72, Gib78, GGB77, Gri78a, HL78, Har78, Hea72, ING70, Koc73, Kyn74, MG72a, MS77, RC76b, Ree71a, Ric75a, Sal73]. Behavioral [DM79, FT76, KO74, Man79, NM74, Smi76]. Below [GG75, TG74, Tsa73]. Bemidji [HGO^{+}77]. Benthic [Hil71a, Hud70, Jac71, LM71a, ST75]. Benthological [Sel71]. Benthos [BH75, GT75, GGG75]. Benzene [EEW77, MRK79]. Benzie [Hil71a]. Between [Alt71, Bul70, For74, GH79, Nym70, Rad73, SPSF72, Ste71, Zau70, GG75, Tan70, TB74, TB77, ZD76]. Beverage [KM72]. Biased [Car72]. Bibliography [KB71]. Big [AC72, BM77b, FF74, Kus74a]. Bight [MMD73, MM77]. Bigmouth [SA70b]. Billfish [TW75]. Binding [RGBL79]. Binomial [HL78]. Bioassay [Fre71, LSN70, MD72]. Bioassays [BN75, FM78, Jen72, MMM72, RA72b]. Biochemical [Beh70, Fuj70, UH72]. Bioconcentration [BBAN75]. Bioeconomic [GG79]. Biofacies [HM74b]. Biofilter [LB76]. Biological [Boy73, GCD^{+}77, HE76, MS70a, MD72, Mit78, Pat77, Pat73, Sch76, TL78, Wal75a]. Biologically [DZ72, HWB71]. Biologists [Bri74]. Biology [Eip73, Ell78, Fra73a, GD79, HH79a, Hoy71, Kay77, Kir76, MR74b, Ree71b, SD73, WAMR79, CvW78b]. Biomass [Cla75, HM74a, KKM^{+}74, Mom78]. Biostatistical [HC73b]. Biota [HGJ75]. biotelemetry [HGO^{+}77]. Biphenyl [NPD74]. Biphenyls [NP74]. Birds [Sch70]. bis [DZ72]. Bismarck [BM77a, Wat70]. Bivalve [CJS71, MD71]. Black [Bow70, CH78b, Coo71, DB72b, LMG76, MM77, NE74, Ric73, SH77, VR75, Wil70]. Blackbanded [Mat73]. Blacknose [FT76, TT76]. Blacktail [BM77a]. Blackwell [Ort79, SA77]. Bladder [FW77]. Bloater [EB78]. Bloch [Smi72a]. Blood [BCS76, CC78, CB72, FB78, JP76, Kir74, LL71]. Blue [Abb74, CV72, Kel73, LCL78, LS73b, PT70, RCL^{+}78, RP75]. Blue-green [LS73b]. Blueback [Bur74, Chi72b, LL77a, MP78]. Bluefin [Dra70, SC73]. Bluefish [LM70, NRMJ74, Ric76b, Sil75]. Bluegill [AMB^{+}77, BK74, BM79a, Ben75, Bey71, BCC78, Clo75, Cob70, Cob72, CWW70, DB78, Eat74, FL78a, GM72, GB79, HD73, HS71b, JP76, KSB78, Laa71, NS76, Par73a, PW74, Pri77, PGP77, Sar77, Sav71, Sav73, SS75, Sie72, SOKEK76, WW73, Wil75]. Bluegills [BBR79, BMC72, CMS74, ECCC76, MJ72, OS72, Shi74]. board [GW72]. Boats [GW72]. Body [Ade78, AS72, EIAJ^{+}76, Hil70, ROH79, Sav71, SNDM77]. Body-Scale [Hil70]. Book [Ano79a, Ano79b, Ano79c, Ano79d, Ano79e, Ano79f]. Books [Ano74c, Ano74d, Ano74b, Ano74a, Ano75c, Ano75d, Ano75b, Ano75a]. Boothbay [CCN75]. Bottom [MHM77, TG74]. Bovine [Ade78]. Box [ECCC76]. Brachydanio [SP76]. Brackish [Tat73]. Brain [Cop72, GHC^{+}79, Hog71, Par70a]. Branch [Nev75a, SA70a]. Branchiuran [KG72]. Branding [RMG73]. Brands [RMG73]. Breakdown [MBN72]. breve [CJS71]. Brevoortia [FB72, HO78, KG72, LM71b, MMS75]. Bride [Kis74]. Bridge [Lab71]. Brine [Wea74]. British [ET73, All73]. Broad [Alt76, RH73]. Brook [Ada75, AS78, AJ75, Bél74, Ben74, BLCF76, Deg70, DOB74, Eck71, Eck73, Fra74b, Fra78, Gib78, Gri74, HH70, HW74, HC76, Her70, HBL79, Hoo75, HJB^{+}78, Jen71, KM79, MH76, Mil74, OC75, Phi75, SA70a, SS76, Spe74, SDW78, WE76, WA70, WBG75, Zim70]. Brown [BM77a, CP75, DW78, DLM72, GZ73, Han75, Kay77, MSW79, Men75, RA72a, RI78, RB70, SA70a, Tau76, Wat70]. Brush [VR75]. Brushed [KS73]. Bubble [BCSS76]. Buccal [BH74]. buccata [Hoy71, Wal73b]. Buena [Ano74f]. Buffalo [SA70b]. Buffalofishes [MJRW70]. Bull [VR75]. Bulletins [Ano74c, Ano74d, Ano74b, Ano74a, Ano75c, Ano75d, Ano75b, Ano75a]. Bullfrog [KF77]. Bullfrogs [WL70]. Bullhead [DB72b, OC79, RI78]. Bullheads [Wil70]. Buoyancy [GM72]. Burbot [Bai72]. Burkenroad [Bea70, VIS70]. Burrowing [Deg70, SN76]. Burundi [Ell78]. butoxyethanol [DM79]. Butte [Jes76]. Butterfish [MW79]. Buttonwood [Bea70]. By-Products [GA77]. Bypassing [Lis71].

C [Ano70f, GPT70, GTP71, GTP72, LL78b, MRSK70, MMC78, WWLC77]. C-DDT [GPT70, MRSK70, WWLC77]. C-Dieldrin [GTP71, GTP72, MRSK70]. C. [WW76b]. Cadmium [BLCF76, CS78, Cha78b, Eat74, KF74, SLD78, Voy75]. caeruleum [Coo79]. Cage [Tat73]. Calcium [Gre71]. Calculation [Hil70, Sav71]. California [SM76a, AB70, Ban73, Bur72b, CM70a, DL78, DJD79, EB76, Erm73, GS72a, IWR78, LMG76, MS74b, MP71, MH78b, SG72, SM76a, Ste77, Tal73, TW75, WBG75]. californiensis [GMS72]. Callinectes [Abb74, CV72, LCL78]. Caloric [Sav71]. campechanus [MBT70, WAWR79]. camtschatica [MY71]. Canal [Bea70, KW72b, RI78]. Candidate [Mar72a, Mar74]. Cannibalism [Che73, Tar74]. Cans [KM72]. Canyon [IWR78]. Capacity [KH71, RGBL79]. Cape [Mor71, VS79]. Captive [SMDN76]. Capture [Vin77]. Captured [Cai71, RH74]. Carangid [Wat73]. Caranx [Wat73]. Carapace [PT70]. Carassius [BS73, GPT70, GTP71, JBR71, SH74a, SNDM77]. Carbaryl [Kor73]. Carbon [DMB76, NBS76]. Carbonate [RA72b]. Card [Bro77]. Caribbean [MBT70]. Carl [Ort79, SA77]. Carolina [Bat72, BG72, DL71, Dev78, DDA72, Gri78b, HCD78, LM71b, MS71, MGW76, OC79, RH74, SH78, TL76, Tyu74, VS79]. carolinus [BA70, BA71a]. Carophyllaeidae [Wil74]. Carp [Ade78, Bai78, BBR79, BTR70, CSR78, CSG^{+}78, EC77, FA78, FL78a, GG78, Hai73, HG74, JH77, Kil78, LRIF78, Lew78, Men74, Mit78, NM78, Pfl78, Ril78, SCR78, Sta74, Sta76, SMS78, SL78, SK71, UF78, WC79]. carpio [Ade78, HG74, JH77]. Carpiodes [Mad71, WW76b]. Carpsucker [SK71]. Carpsuckers [WW76b]. Carrier [AP72]. Case [Ach75, BH75, HR72, RM76, Sim76]. Cases [Spa70]. Castle [WBG75]. cataractae [BLA78, Vin77]. Catch [BM79b, FB72, Jes77, KC72, Kim76, KJ78, MH78a, Reg71a, Wat76]. Catchability [Jes73, WC79, ZD76]. Catchable [HL76]. Catches [CV72, HCD78, MF75]. Catching [EP73]. Catchout [BBR76]. catenatus [MTM72]. catesbeiana [KF77, WL70]. Catfish [AS71a, AS72, AMG73, AP75, AM75, BBR76, BMTD72, BRJ78, BIJ70, Bul72, BS73, BKS76, CSH74, DTD76, DH70, EIAJ^{+}76, Elr74, GAR76a, GAR76b, HS75a, Hei75, HGS^{+}74, JAR76, JAR78, Kil78, Kir74, Kor73, LL75a, LL71, Lew78, LS73b, Lov78, LL78a, LL78b, MMC78, MA75, Mus77, NDN76, PGSP75, PR77, RC76a, RC76b, RB79, SL73, SS71, Tar71, TK72, TB78, TBM79, Wor75, YS79]. Cations [NB72]. Catostomid [Fui79, FW79, Koe70, Wil74]. Catostomus [Bea72, Cam71, FW79, SML75]. Caught [DDA72, MH70, Pet77]. Caught-and-Released [Pet77]. caurinus [Pri75b]. Cause [Cai71]. Caused [Plu72]. Caution [Bow78]. Cautionary [Wal71]. cavalla [Dev78]. Cavity [BH74]. Cayuga [Bor74]. CDC [BA72]. Cement [KS73]. Central [FEL79, TG79, Bec72, CLB^{+}73, FEL79, GD79, Hoo75, Joh70, Joh71, MJRW70, Ric76b, SDC78]. Centrarchid [AFS79, Mer71a, RMG73]. Centrarchidae [Etn71]. Centrarchids [Mer71b]. Centrifugation [BMS71]. Centropomus [MS71]. Centropristis [MM77]. cepedianum [Coo78, OFTM77]. cephalus [SN75, TB71]. cerebralis [Hof74, O’G79]. Certain [DB78, Kir74, Ogd70]. Cestoda [RA72a, Wil74]. Chain [CV72, Mci70, MG72b, War73]. Chamber [LHBS79, Ric75b, Sal73, SND71, ST75]. Chance [Mag76]. Change [CMS74]. Changes [BCP76, CCM79, Erm73, FB78, FN77, FL78b, KME70, Kel73, Lev74, Mos70, Nic71, RMG73, Ric76a, UM72]. Channel [AS71a, AS72, AMG73, AM75, BBR76, BMTD72, BRJ78, BIJ70, BS73, BKS76, CSH74, DTD76, DH70, EIAJ^{+}76, Elr74, GAR76a, GAR76b, Hei75, HGS^{+}74, JAR76, JAR78, Kil78, Kir74, Kor73, LL75a, LL71, Lew78, LS73b, Lov78, LL78a, LL78b, MMC78, NDN76, PGSP75, PR77, RC76a, RC76b, RB79, SL73, Tar71, Tho75, TK72, TBM79, Wor75, YS79]. Channelized [GS78]. Channels [DW78]. Chanos [Tsa78]. Chaoborus [PCP73]. Chapman [Roy73a]. Characteristics [JAR78, Koc73, Man79, MM76, Spe73, Tar74, ZD76]. Chasmistes [Koc73]. Chatanika [Alt71]. Chelonia [BNMC76]. Chemical [Boy73, Boy79b, EBM^{+}75, Gre70, LDM71]. Chemicals [DMB76, MMM72]. Chemistry [AB75, Boy76, BCS76, FB78, LRIF78, UM72]. Cherryfin [HB75]. Chesapeake [MKK73, KW72b, MB75b, MP77]. Chile [Kai73a, Kai73b]. Chinook [Arm74, BR76, Ber78, BD72, BH76b, Cha78b, CBP76, CA77, DB70, Ebe70, GGB77, GH77, GT72, HL75, HBD78, KM76, Ray79, Sch74a, SW74, SH74b, SSE78, Tho75, Van70]. Chironomidae [DZ72]. Chironomids [Tsa78]. Chironomus [DZ72, KDR74]. Chloramines [LHL77, TT76]. Chlorinated [GPT70, GTP71, HWB71]. Chlorination [BS77, SB78]. Chlorine [GE75b, HBM79, JWA77, MP77, MP78, SH74b, TT76]. Chlorine-Induced [GE75b, JWA77]. chlorophenyl [DZ72]. Choice [Far78, WE76]. Chorion [MVM77]. Chromic [Bow78, FB79]. Chromium [RS75]. Chromosome [Set70]. Chromosomes [Rob70]. Chronic [Ben75, Bru71, CWW70, DB70, Eat74, Fre71, GHC^{+}79, HN77, HBD78, KSB78, LL70, OS74, SLD78]. Chronically [BD72]. Chronology [And70, AR71, DE78, Mat73]. Chrosomus [Sta78]. chrysops [NK73, Pri70, VW74, WH71, Wis74]. Cichlid [Gri77, Que75]. Circular [ST75]. Cisco [EC70b, Eng74, Eng76]. City [Ano71a, Ano72c]. Clam [Mac77]. Clarias [Mus77]. Clark [Cro76, Rue77, SW70, Wal72, DB72a]. clarki [Arm71, DL78, DK75, Gri74, Han71, Lab71, MH70, ML72]. clarkii [GA77]. Class [Has70, HCP76, KDS79, Lux73, Par71, Par72, Sav78, SDKT79, Ste77, WH75, Wal72]. Classes [For71]. Classification [Cop70]. Clay [Boy79a]. Cleaning [Spa70]. Clear [BSI76, DB72b, EB76, Lab71, LMG76, MH78b]. Clearwater [PW75b, Pet77]. Cleithra [HH79b]. Climatic [MDS79]. Clipped [Cob72]. Clipping [Cob71]. Clips [MC79, MH76]. Clupea [EEW77]. clupeaformis [Bro75]. Cluster [CK71]. CO [RWC73]. Coast [Kai73b, Mes76, Mor70a, PSG^{+}74, TB71]. Coastal [CH76, KW71, KH78, KM76, Nic71, NRMJ74, Sek74, SH78, SP71, SHH76, Wic73]. Cod [Jen70, Mor71]. cognatus [OR77]. Cohabiting [BSI76]. Coho [BS79, BJ79, BH76b, CS78, DWB75, Eng76, FMS79, Gun73, Hil71a, LHL77, MS74a, MNL79, Pec70, Pec74, PLCM75, SB78, SS76, Syl73, Van70, Zau70, ZT74]. Cold [BGS77, RMG73, Tsa72]. Collected [Cla74, Hig70]. Collection [PF72]. Color [Jes73, Jes77, MG72b]. Colorado [HS75c, Bab71, EC77, IWF75, Lac73a, TB74]. Colors [LM74]. Columbia [Bec72, CBP76, GD79, All73, BGS77, BE70, BCSS76, ET73, HGM78, JS73, KM76, RR73]. Column [Hoa77]. Combined [WW71a]. Comment [Rei73]. Comments [MM77, Roy73a, Zau70]. Commercial [Hur73, Jes76, PS70, Tur77]. Commercially [FN71]. commersoni [Bea72, Cam71, SML75]. Common [ABW72, BCS76, DJD79, Fra74b, How72, MM70]. Communicating [Hil71b, Kea71]. Communication [Ste71]. Communications [Cha71]. Communities [Ada76, EH76a, RH73]. Community [Bro77, Clo75, LC70]. Comparative [Ame74, BGS77, FSM75, HS71a, Pet77]. Compared [Wil78]. Comparison [Bai78, ET73, GM78, HH79b, How71, Rad73, SGC^{+}78, Sim72, VIS70, Wal75a, ZD76]. Comparisons [BCC78, Coo78, Fui79, FW79, Wil75]. Compass [GB79, LCG77]. Competition [Fra78, HM74a, Pow74, Smi76]. Complete [Mea71]. Completion [BR76]. Complex [Spi73]. Complexes [FE71]. Component [EEW77]. Components [LJH76]. Composition [Ade78, AS72, GB72, MH78a, ROH79, Ric76a, Wea74, YS79]. Compounds [DZ72, DOB74, LS73b, NPD74]. Comprehensive [OE75]. Computer [GCD^{+}77, HM74a, HL76, Hog70, Lac75, Ort79, SE73, TL74]. Computer-Implemented [TL74]. Concentration [BJ76, Neb72, RS75, Tho75, Voy75]. Concentrations [BE70, BRJ78, BKS76, CSH74, DB78, DOB74, DE75, GW77, Jes77, KS71, Mus77, PPSF79, ROLH78, SH77, SSE78]. Concept [Lar77]. Conclusions [DC79]. Concrete [CNHT76]. Condition [BA75, BA71b, Eng74, MP74, RM71, Sek74, SK71]. Conditioning [Abb72, FB78]. Conditions [AM75, CH72, CE76, ED73, Fra73a, Fra73b, Gul78, LDM71, LSN70]. Conentrations [EEW77]. Conference [Ano70g]. Confined [SSE78]. Connecticut [MR74a, Kis74, KG72]. Conowingo [AR71, MMN79]. Consequences [TB78]. Considerations [Sch76]. Constant [Bro75, Bru71, Fre71, HH79c, Med79]. Constituents [FT76]. Constitution [Ano70b]. Construction [Bur72b, HGJ75]. Consumption [AM75, FN77, HP76, KDR74, WAWR79]. Containers [MJ76]. Containing [Lew78]. Contaminated [GPT70, GTP71, Hof74]. Contamination [AMB^{+}77]. Contemporary [Ano74f]. Content [Boy70, EIAJ^{+}76, FN77, Laa71]. Contents [BMS71, BF76, Mci70]. Context [Ach75]. Continuous [Cha78a, Eis75, PR77]. Continuously [KS71]. Contrasting [KK78, VD79]. Contribution [Byr71, KC72, LL77a, SD73]. Contributions [Alt79, PB78]. Control [Bow75, HWB71, Mit78, Ose77, RC71, RA72b, SH75, SDC78, Tsa78, TB78]. Controllable [TGR74]. Controlled [PB78]. Controller [BN75]. Controlling [WCH71, Wic73]. Conversion [AS72, Fra73b, LKT79]. Cooling [BG75]. Cope [Hoy71, Wal73b, Koc73]. Copepod [BWH79, GMS72, VS79]. Copepoda [BH76b]. Copper [Ben75, CS78, Cha78b, KF74, LL71, OS72, Pat73, Pri75b, TB78]. Coregonine [Eds70b]. Coregonus [Alt71, Alt76, AS71b, Bro75, EC70b, EB78, Eng76, SM76b]. cornutus [BCS76, ABW72, MSS73]. corporalis [Ree71b]. Corrections [Ano74e]. Correlations [EIAJ^{+}76, SPSF72]. Corticoid [SSG77]. Cortisol [SSE78]. Corydalus [MSS73]. Coryphaena [BH74, RH74]. Cost [PW74]. Cotton [Hög73]. Cottonwood [DL78]. Cottus [FW77, LL75b, NE74, OR77, PL72, Tho73a]. Cough [DOB74]. Council [Ano70f]. Counting [ST75]. County [Cro76, DB72a, DH78, EC77, Hav73b, Hil71a, MP71, OC74, Wil70]. Courting [ING70]. Cove [MF75, Nev75b, Pat75, TM74]. Cover [Bjo71, DW78, GZ73, VR75]. Coverage [Bri74]. Crab [Abb74, CV72, PPSF79, PH72]. Crabs [Bur72a, LCL78, MY71, Pri75b, PT70, RCL^{+}78]. Cranberry [Ree71a]. Crappie [AR71, CH78b, Coo71, LMG76, MMN79, SH77]. Crawfish [FA78, HA76]. Crayfish [BGM75, GA77, Mom78, Rei72]. Creek [BM77b, DL78, Erm73, GS72a, GC72, KM74, Mat73, MR74b, Mil74, OC74, SG72, EL75]. Creeks [Lab71]. Creel [HS71a, MDS78, MDS79]. Crevice [GG77, GB78]. cristata [SN76]. Criteria [Smi73a]. Critical [FB77, FB78, Wal75a]. Critique [Kem70]. Croaker [MMS75]. Croakers [KH78, YHP77]. cromis [Ric73]. Crop [HW74]. Crops [HD70, Hav74a]. crosbyi [MSS73]. Crowded [BH76a]. Crude [EK75, EEW77, MRK79, SC79]. Cryptic [Bro77]. crysoleucas [LL71, MSS73]. Crystal [CA73]. Ctenopharyngodon [CSG^{+}78, NM78, Pfl78, Sta74]. Cui [Koc73]. Cui-ui [Koc73]. cujus [Koc73]. Culaea [Deg70]. Culture [AP75, AM75, Boy79b, HC73a, Lay71, LYSB78, RC76a, RC76b, RBC78, RHW75, Smi72b, Tan70, Tat73]. Cultured [SW75]. Cultures [Boy70, LS73b]. Current [MLL71, MG70, Sch70]. Currents [FM70, Ric75b]. Curriculum [Roy73b]. Curves [BMC72]. Cutlips [FL78b]. Cutthroat [Arm71, Bow75, DL78, DK75, Gri74, Han71, Hog71, Jah72, Lab71, MH70, ML72, RC71, TRS78, TB74]. cyanellus [Ell74, HP76, KH72, Tau77, Wil75]. Cyanide [KSB78]. Cycle [EB76, FEL79, KH72, LS73a, Rue77]. Cycles [BG75, RW79, Sta79b, WCH71]. Cyclic [Med79, Ott74]. Cycling [HH79c]. Cyclopoida [BH76b]. cylindraceum [ALTA77]. Cynoglossidae [Sti76]. Cynoscion [JWA77]. Cypress [Kus74a]. Cypretta [Kaw73]. cyprinacaea [BWH79]. cyprinacea [WL70]. cyprinellus [SA70b]. Cyprinid [And72, BM71, DL71, HBH74]. Cyprinidae [Moy73, Ree71b, Vin77]. Cyprinodon [GHC^{+}79, HSF75, LM75b, Mar72b, MP71, RD73, SHF74]. Cyprinodontidae [Mar72b]. Cyprinodontids [dVKP78]. Cyprinus [Ade78, HG74, JH77, Mad71, WW76b]. Cytochalasin [AS79]. Cytogenetic [Beh70]. Cytogenetics [Hub70].

D [Ano70f, Roy73a, CWW70, DM79, LL78a]. Dace [BLA78, FT76, SKFM72, Sta78, TT76, Vin77]. Daily [DE78, LKT79]. Dakota [Elr74, HGAF70, Has70, HP74, MS77, SA70b, Wal72]. Dam [TL74, TG74, BE70, Dud74, PF72]. Damage [MJ72]. damaged [HW74]. Dams [BR76, Lis71, Ray79]. Dana [BGM75]. Daphnia [NP74]. Dark [Cam71]. Dart [Sto72a]. Darter [Mat73, Tsa72]. darters [Coo79]. Data [Cla74, HC73b, LKT79, MMD77, Wal75a, WH75]. Day [BE70, Sar77]. Daytime [DW78]. DDE [DZ72, HW76]. DDT [AJ75, GPT70, Han72, LSN70, LH72, MS70a, MRSK70, NB72, Sta79a, WWLC77, WW71b]. Dead [Par70b]. Death [Cop72]. December [Ano74d, Ano75d]. Decimated [BG70]. Decision [PLZ75, TB79]. Decision-making [PLZ75]. Decline [RBC78]. Decrease [Bro70]. Deep [FM78]. Defined [NS76]. Definition [All71]. Degradation [AJ75]. Degree [HWB71]. Dehydrogenase [HB78, MSW79, Mor70b]. Delaware [DP73, KW72b, LM76, MWK71]. Delayed [BR76]. Delivery [BM70, Fre71]. Delta [SKK79]. Demand [Boy73, Boy79b]. Demersal [Qua72]. Demonstrated [Hes70]. densa [Sta74]. Density [BMS71, Bjo71, Dom73, Men75, PCP73, Ree71a, vWDB78]. Density-dependent [vWDB78]. Density-Gradient [BMS71]. dentatus [SW75]. dentium [GHH77]. dependent [vWDB78]. Deposition [LJH76, MY71]. Depth [Fas73, Fas71, GS74, GH77, Lin76, OFTM77, Smi73a]. Descaled [BS79]. Description [Coo79, KTJ75, SD78]. Descriptions [Fui79, FW79]. Desert [Pis74]. Design [GH73, Spe73]. Designed [GCD^{+}77]. Destratification [Fas71, Lac73a]. Detecting [DD74, Goo72, LA71, SD72b]. Detection [KS73]. Determination [Boy79b, Cas74, LHL77, MS70b, MKK73, SL73]. Determinations [Gan76]. Determine [MP76]. Determined [HGO^{+}77, Jen70, MD72, NM78, WL75]. Determining [GS74, VS76, vWDB78]. Deterrent [KF77]. Detoxification [LDM71]. Detoxifying [MB75a]. Detritus [BMS71]. Developing [Wal73a]. Development [ASH75, BD72, Bro75, BM78a, BM78b, BM79c, Coo71, Eds70a, For71, FL78b, Gal73, Has70, Hea71, HB77, Hun76, Jah72, Kin79, Mer71a, MP78, NRMJ74, OS78, Sch76, Sil75, Smi73a, SC79, Tau77]. Developmental [HSS71, SS76]. Device [Fre71, HKB76, MMM72, Rod73]. Devices [PS76]. Diablo [IWR78]. Diadema [Dav71]. Diameters [Han74]. Diamond [LKT79]. Diaphorase [MSW79]. Diazinon(R) [GHC^{+}79]. Dibrom [MSS73]. dichloro [DZ72]. Diel [ADSB77, BK74, CH78b, DE78, DD70, Ell76, EB76, FB72]. Dieldrin [BKS76, GTP71, GTP72, KS71, LL70, LJH76, MRSK70]. Diet [GPT70, GTP71, JBR71, Kel73, Pri75b, SH78, WFS78]. Dietary [DS78, Ket75, Lov78, MRSK70, WWLC77, YS79]. Diethylstilbestrol [Bul72]. Diets [Bro77, DTD76, Gle78, KK78, LL75a, LL78a, LL78b, MA75, ROLH78]. Difference [BNS70]. Differences [Arm74, RR73]. Different [Bro75, DM73, DOB74, GTP72, Han74, Hav73c, Mil74, PGSP75]. Differential [Gan76, KW72a]. digeranylamino [Mar74]. Digestibility [SL73, WFS78]. Digestible [KW72a]. Digestion [Che75, Gan76]. Dimorphism [Tho73b]. Dionda [BNS70]. Dioxide [NBS76]. dioxin [HN77, MNL79]. Diplostomatidae [Vin77]. Diptera [DZ72, PCP73]. Direct [MLL71]. Discharge [LS73a, RI78, Szl74]. Discriminant [DPRC78]. Discrimination [SMDN76]. Discussion [Hea70, Hil71b, Mag73, Str73]. Disease [BCSS76, HL75, Hof74, MMD73, O’G79]. Disinfection [Hof74]. Disney [Ano74f]. dispersant [EK75]. Displaced [MLK77]. Displacing [Rei75]. Disposable [KM72]. Disposal [SS74]. Disputes [MV70]. Dissolved [AMG73, BE70, BRJ78, CSH74, CH78b, CCM79, DE75, FM78, RH75, RBC78, RB79, Sch74a, SH77, Voy75]. Distinguish [RR73]. Distributed [OMS78]. Distribution [BK74, Bea70, Bur74, Bur72a, Cam71, DB72a, DZ72, DL78, DLM72, Edw78, Fas73, Fas71, FM70, GH77, GTP71, HL78, HW77, HGAF70, HS75c, Lin76, MRSK70, Mar72b, MP74, Mun70, MM77, NRMJ74, OFTM77, Pfl78, PCP73, RM73, Ric73, SKK79, SDC78, SM76b, TC72, WBG75]. Distributional [NM74]. Distributions [DC79, GW77, MS78, Nic71]. Disturbances [Cop70]. Diurnal [Bou72, CMS74]. Diversity [Edw78, Fra74a, HM74b, Ohn70]. Diverting [ML71]. DNA [BCC78, Mus77]. Do [GCB73]. Document [GH73]. dolomieui [RC76c, GH79, MR79, Nev75a, Wal73a]. Dolphin [BH74, RH74]. Domestic [FT76]. Domesticated [GMS72]. Dominant [For71]. Dorosoma [Coo78, Gri78a, HB74, Joh70, Joh71, OFTM77]. Dorsal [SM74]. Doses [Far77]. Dosing [Eis75]. Dover [MS74b, SM76a]. Downstream [EL75, Sta72]. Dragonfly [WW71b]. Drainage [Fui79, Ric75a, Sim72, NE74]. Drawdown [BH75]. Dredge [SS74]. Dredges [Hud70]. Dress [EIAJ^{+}76]. Dress-out [EIAJ^{+}76]. Drift [BM74, GM78, GG75, Gri74, Lar74, WW70]. Drop [Hoa77, KTJ75]. Drop-net [KTJ75]. Drum [BA75, BA71b, Ric73, SW70, TL76]. Dry [Dav75]. Drying [Her73]. Due [Ade78, KO74]. duorara [RI71]. duorarum [Bea70, CNHT76, Bea70, CNHT76, RI71, VIS70, Wat76]. Duplication [Ohn70]. duquesnei [Bow70]. Duration [CV72, HBM79]. during [Cha78a, CJS71, DLM72, FL78b, KH72, Leg72, Reh78, SA77, Wal76]. Dursban(R) [MWHH72]. Duskystripe [HD73]. dwelling [War78]. Dye [LM71a]. Dying [KG72]. Dynamic [LSN70, Sch70]. Dynamics [DM79, Elr71, Eng76, HM74a, HA73, Hav74b, Joh70, KKM^{+}74, Koe70, MC78, Ort79, RB79, Sav78].

Early [BRJ78, ER72, GS72b, Has70, Kir76, LHL77, OS78, PS79, SS76, SW70, Tau77]. Earthy [LS73b]. Earthy-musty [LS73b]. East [MW71, Ric76b, WC79]. East-Central [Ric76b]. Eastern [BLA78, Jos70, Tsa72, Wis72, WM78]. Eaten [MP76]. Echo [Cha76, SRT73]. Ecological [DB72a, Dic73, HGAF70, HB75, MT72, Moy73, RH73, TB74]. Ecology [Ada76, DL71, MM77, NM74, Sar77]. Economic [Cru73, Fis70, GCB73, TBM79]. Economics [Sil77]. Ecosystem [HW76, Pat75, Tag76]. Editorial [Rei73]. Educational [Eip73]. Eel [Hur73, Ogd70]. Eelgrass [Ada76]. Effect [AS70, Bab71, BE70, BCS76, Bro75, CH78b, CV72, CH75, CP75, CB72, Dea78, Dom73, Ebe70, Eds70a, Edw78, Etn72, Far78, Fra78, Gre71, Han72, Han74, Hil71a, Hof74, Hoo75, Hut71, HGS^{+}74, Jen71, KDR74, LL71, LMG76, Neb72, NPD74, NP74, RGBL79, SN76, SC79, SKFM72, Syl73, Tag76, Tho75, Voy75, Wal75b, WFS78, You76]. Effective [Mar72a]. Effectiveness [How71, LM74, MB75a]. Effects [Ada75, AP75, AB75, ASH75, BH73, Ben75, BLCF76, BH75, Bor74, Bou72, BS75, BS77, Bru71, BS73, BK77, Bur72b, CC78, CSH74, CLB^{+}73, Cha78a, Cla75, Cob71, Coo70, CWW70, DM73, DB70, DTD76, Dud74, EH76a, EB76, Eng74, FB77, Fas73, Fas71, FA78, FL78a, FF74, GM72, Gan76, GC72, Goo72, Gri78a, Hai73, HSF75, Han75, Har78, Hav73a, Hav74a, HBD78, HS75b, HBL79, HP76, How72, HCS76, Jes77, KH72, Kin76, KDS79, Kir74, KM79, Kus74a, Lac73a, LL70, LDM71, LRIF78, LH72, MSS73, MMD77, MBN72, MUR^{+}76, MP78, MH78b, NB76, NC73, NB72, OE75, OS78, Ott74, OFTM77, PLCM75, PB73, RMG73, Ray79, Rei75, RH75, RB79, RK76, SSH72, Sav71, SH72, SHF74, SB78, SBCG79, SG72, SS76, SLD78, SNDM77, SC71, Sta79a, Sto72a]. Effects [Sto72b, Szl74, TG74, TBM79, VS79, WW71a, Wal73a, WH71, WW70, YS79]. Efficacy [SA71, WWO73]. Efficiencies [EP73, KJ78]. Efficiency [Fra73b, Hög73, LKT79, Wat76]. Effluent [BG72, CRC^{+}76, GGB77, MJN72, NM74]. Effort [BM79b, MDS79]. Egeria [Sta74]. Egg [Bro75, Mer71b, PH72, Tho75]. Eggs [AS70, Ame74, AJ75, ASH75, BH73, Bon76, BK77, CH75, Coo78, Coo79, DB70, Eds70a, Ell70, HBD78, HG74, JWA77, MY71, MUR^{+}76, MP77, NRMJ74, Ose77, PKP74, Sta79a, SRM76, SN75, Wal76, WW73]. Ekman [How71]. elassodon [MW71]. Electivity [Str79]. Electric [ABW72, GM79, Nug70]. Electrical [DD74, EP73, Wat76]. Electrode [SND71, SD72b]. Electrofishing [SVL73]. Electrolyte [SC71]. Electrolytes [Kir74]. Electrophoresis [BR75]. Electrophoretic [Bou72, MKK73, Nym70]. Elemental [GB72]. Elements [DB78, vWDB78]. Elephant [Jes76]. Elevated [BGM75, Bru71, HBM79, Nug70, OS78]. Elevation [AZM75, Dom73]. Eliminate [Wal76]. Elimination [GPT70, GTP72, MRSK70, PGP77]. Elodea [Sta74]. Elops [Sek74]. Elutriator [Ste75]. Embiotocidae [AB70]. Embryology [ZT74]. Embryonic [Bro75, Gal73, OS78, SC79]. Embryos [CSH74, DE75, ES72, EEW77, HSF75, HMJ73, Mer71a, SHF74, WW71a]. Emerald [Sie72]. Emergence [HC76, Neb72, PLCM75]. Emergency [BT79, WP76]. Emergent [HT70]. Emigration [YHP77]. Emphasis [CR73, Hof74]. Empirical [Mat78, TG79]. Enclosure [BJ70]. Endrin [LSN70]. Ends [Reg74]. Endurance [OS72]. Energetics [CE76, EEW77]. Energy [ABW72, Bur74, FN77, HE76, Kel73, OHC79, PW74, WAWR79, Wis74]. England [LM70]. English [OP73]. Engraulid [LM75a]. Enormous [Ohn70]. Enrichment [Hai73]. Entrance [Dom73, FB72]. Entrance-Pool [Dom73]. Entrapment [SKK79]. Environment [CH76, DZ72, GE75a, NC73, Odu70, PB78, SK71, WNH79, War76]. Environmental [AS72, CCM79, DM71, Has70, Hor77, LHBS79, MS74b, OHC79, SPSF72, SDW78]. Environments [GH73, HWB71, KK78, Kem70, Kus74b, WCH71]. Epinephelus [Smi72a]. Epitaph [Lar77]. Epizootic [Plu72]. Epizootiology [MW71]. Equilibria [HWB71]. Eradicated [Phi75]. Ergasilus [BH76b]. Ericymba [Hoy71, Wal73b]. Erie [Coo79, DH78, KF74, Par71, Par72, ST77, Wol77]. Erosion [MS74b]. Errata [Ano76a]. Erratum [Ano75e, Ano75f]. erythrurum [FW79]. Escanaba [KC72]. Escaped [SMS78]. Escapement [MM76]. Esocidae [WA77]. Esox [Cas74, Cas75, DME77, Gal73, GE75a, GHH77, HH79a, HMJ73, MC79, MT72, Mci70, MG72b, MLMS74, War73]. Essential [DH70, JBR71]. Essentiality [LL78a]. Established [CM70a]. ester [DM79]. Esterases [Koe70, RSL70]. Estimate [CSG^{+}78, You76]. Estimates [Car72, Cha76, MDS79, Str79, TG79, Wis72]. Estimating [Bar71, BM79b, KH71, Kim76, KTJ75, Reg71a, WH75]. Estimation [BJ76, You72, You74]. Estimator [MLO77, MS78]. Estrogen [SH74a]. Estuaries [SHD72]. Estuarine [BH73, CA73, Cop70, Dah70, Far78, Fra73a, GGG75, HCP76, KJ78, Odu70, Sti76, Tag76, VS79]. Estuary [RM73, Ste77, DPRC78, Hea70, JS73, TC72]. ethanol [Mar74]. Etheostoma [Coo79, Tsa72]. ethylene [DZ72]. Euphausiids [All72]. European [KP75]. Euryhaline [MMD73]. Euthynnus [SMDN76]. Eutrophic [Lin76]. Eva [Arm71]. Evacuation [DE78, Nob73, PH74, WKN^{+}76]. Evaluate [Sim72]. Evaluation [BM74, Dav78, DCA71, FSM75, GCB73, Han71, HC73a, Hun76, LB76, LHA78, MDS78, MC79, MS78, Mit78, Pow74, RF74, SD72a, Tan70, WC79]. Events [GCD^{+}77]. Everglades [Bea70, Cla74, Hea70]. Evidence [GMS72]. Evolution [Eve71]. Evolutionary [Koe70]. Example [Hea70, Hor77, Reg71a]. Exchange [HWB71]. Excretion [Sav73, TKM70]. Exercises [SE73]. Exhaust [Ada75]. Exoglossum [FL78b]. Exotic [Ano70g, CR73]. Exotics [Mag76]. Expenditure [WAWR79]. Experience [BBR76, Que75]. Experiment [Gre71, OM75]. Experimental [CE76, ED73, Fra73b, HT70, MJ76, Ric75b, SW74, Tag76, WA77, WCH71]. Experimentally [BG70, BS79, RI71]. Experiments [CNHT76, DC79, KP75, Ohn70, PGSP75]. Exploitation [You76]. Exploited [Cla75, Fox70, HS75b]. Exponential [Bro70, Fox70]. Exposed [BMC72, DOB74, EEW77, KS71, MJ72, TT76]. Exposure [BK77, Cha78a, DB70, Far77, HBM79, HBD78, KO74, LA71, MLL71, OS72, TT76]. Exposures [HBL79]. Extended [DM71, HA76]. Extensive [Bai78, Ohn70, RHW75]. External [Cas74]. Extremes [KS75]. Eye [SC73].

Fabricius [DZ72]. Facilities [Spe73]. Factor [BA75, BA71b, Che73, ES76, Eng74, GE75a, GZ73, Mer71b, Que75, Sek74, WE76]. Factorial [OM75]. Factors [AS71b, Bow75, BKS76, Cob71, Gar73, HD70, HS71b, Mar72b, NBS76, Nev75a, RC76b, RM71, SDW78, Wal72]. Fall [BH75, Ebe70, GT72, KBvL77]. Fall-Spawning [KBvL77]. Fallfish [Ree71b]. falling [Hoa77]. False [Cob70]. Family [Beh70, MM70]. Fantail [Coo79]. Farm [LJH76, Shi74]. Fathead [And70, Bro70, Bru71, HP74, MVM77, NPD74, RS75, SM74]. Fatigue [MH70]. Fatty [Wea74, YS79]. Fauna [Cro76, TG74]. Faunal [Bro77]. Fear [VS79]. Fecundity [Bat72, BTL75, EB78, GB78, Hea71, Joh71, Kai73b, LL75b, Man72a, MMN79, MLL71, Mes76, NK73, PS77, Rue77, ST77, Tau76, WW76b]. Fed [CE76, GPT70, GTP71, Sta74, Wea74, WFS78]. Feed [Bul72, GA77, HCS76]. Feeding [Ada76, ADSB77, AS72, AP75, AR71, BK74, Bul70, Bur74, CNHT76, DR75, DPRC78, DE78, Ell76, EB76, Fra74a, Gan76, Har78, Jud73, LKT79, Lev74, LH71, Mat73, MS77, MMB74, MH78b, PPSF79, RC76b, Rod73, RR73, Sar77, TD76, TBM79, Wor75]. Feeds [SL73]. Female [Abb74, Mad71, Pet77, Tar71, TK72]. Ferric [BCS76]. Fertilization [BS78, DB76, LB77]. Fertilized [BK77]. Fiber [LL75a]. Field [Dav78, LKT79, LCL78, Pit73, Wal75a]. Filament [Han74]. Fin [Cob71, Cob72, LPH70, MMD73, MC79, MS74b, MH76, NC73, WW75]. Finescale [Sta78]. Fingerling [BIJ70, FF74, HC73a, Lew78, MH76, OHC79, Van70]. Fingerlings [CSR78, DTD76, KC72]. Fins [OC75]. Fiords [CH76]. Firehole [Erm72, KK78, Kay77]. First [Ano72c, Che73, DC79, EBM^{+}75, KGJ74, NS75, SA77, Sie72, Tar74, TC79]. First-Order [EBM^{+}75]. First-Year [NS75, DC79]. Fish [Ada76, AP72, AFS79, AC72, And73, AB75, Bai78, BF76, BR75, BM71, Boy70, Boy79a, Boy79b, BT79, Bro77, BM70, BSI76, BWH79, Byr71, CK71, Che75, CL75, Cro76, DL71, DB78, Dea78, Dom73, DD74, Edw78, ET73, EBM^{+}75, Erm73, Etn72, Fox70, Fui79, FW79, GM78, GS74, GB71, Gri77, GE75b, GM79, GCD^{+}77, Hai73, HM74b, HS75b, Hof74, Hor77, Hub70, Hut71, IWR78, KME70, KB71, KH71, Kim76, KKM^{+}74, KJ78, KM72, Kus74a, Kus74b, Lay71, LA71, LB76, LYSB78, LSN70, LM76, Lis71, LM74, MS70a, MM70, MWHH72, Mar70, MD72, Mar74, MB75a, MH70, MS70b, Mat78, MMD77, MLO77, Mel76, OE75, Ohn70, OC74, Ose77, Par73a, Par70a, Par70b, PH74, PCP73, Qua72, Que75, RMG73, RH73, Ric76a, Sal73, Sch74b, SRT73, SRM76, SDC78]. Fish [Szl74, Tan70, TG79, TB79, Tsa73, Vin77, Wal71, WV73, Wat73, Wic73, WWO73, Wil78, WFS78]. Fish-Food [Par73a]. Fish-Holding [BM70]. Fisheries [Ach75, All71, Ano70b, Ano70c, Ano70f, Ano71a, Ano72a, Ano72c, Ano73a, Ano74e, Ano74f, BLCD75, CLB^{+}73, Cob73, Dic73, DH78, Eve71, GCB73, Gun71, HL76, HC73b, KB71, Lac73b, Lac75, Mag73, MP75, MM77, PLZ75, RH73, Roy73b, SH75, SE73, Str73, Tal73]. Fisherman [Erm72]. Fishermen [Man79]. Fishery [Ach75, Bab71, Bri74, CH73, Car72, Cha73, Cru73, Dut76, Eip73, FF74, GG79, GS78, Gul78, HM74a, Hög73, HCD78, Hur73, Jen78, KB77, MV70, Mea71, OMS78, PS70, Reh78, Roy73a, SS74, Smi73b, SD72a, TW75, WH75]. Fishes [Ano70g, BH73, BCSS76, BBR78, CH78a, CA73, CL75, CR73, Dah70, DB72a, Dea76, DP73, DJD79, EC70a, ES76, Elr71, FM78, GW77, HW77, Hav74a, Her73, Hil70, HS75c, Jes73, Jes76, Jes77, JS77, KW71, Koe70, LHBS79, MMD73, MHM77, MF75, MLT76, MP74, MBN72, Mil72, MRK79, MP77, Mor70a, NM74, Par73b, Pis74, PSG^{+}74, RM73, Ril78, SBCG79, Spa70, Sti76, Tan70, TM74, TC79, VS79, Wag72, Wil74, WHW78]. Fishing [BA72, Jes77, KDS79, MDS79, SH72, Sil77, WK79, Yeh77, You72, vWDB78]. Fishkills [BLCD75]. Fishway [Dom73, SPSF72]. Fitted [GS74]. Fitting [GQ79]. Five [BSI76, Eck73, GW77, MP77]. flabellare [Coo79]. Flagfish [NPD74, SLD78]. Flat [OC79]. Flathead [HS75a, MW71]. flavescens [BTL75, CH75, LWO76, NS75, Nob73, Nob75]. Flavor [LS73b]. flesh [Mus77]. Flies [WJ78]. Floating [SS74]. Flood [HW74, SDC78]. Flood-damaged [HW74]. Flooding [Hoo75, HA76]. Floodplain [Dud74]. Floods [SG72, Kir76, MLK77]. Florida [Ano74f, PSG^{+}74, BH76a, Bea70, Bro77, BH74, CA73, Cla74, CSG^{+}78, CL75, CR73, Dav75, Kus74a, MH78b, Ril78, TB71, VIS70, Wil78, ZD76]. floridae [SLD78]. floridanus [MH78b, SHD72]. Flotation [LM71a]. Flounder [BCP76, CE76, Fra73a, Fra73b, Fra74a, HC75, HCP76, Lev74, LPH70, Lux73, PH77, PS79, SW75, UM72, VD79]. Flow [BN75, Bjo71, EL75, Fre71, HH70, Ste77, TG74, TC72, MMM72]. Flow-through [BN75]. Flows [Hav74a, Van70]. Fluctuation [CH78b, Wal72]. Fluctuations [FB72]. Fluorescent [And72, Eng77, HT70, RD73]. Fluvial [BE79]. fluviatilis [TKM70]. following [Erm73, OC74, Par70b]. fontinalis [Ada75, AS78, Ben74, BLCF76, DOB74, Eck71, Eck73, Gib78, Gri74, HC76, HBL79, HJB^{+}78, Mil74, OC75, WE76]. Food [AS72, Ant74, Arm71, ALTA77, AR71, Bai72, BA75, BA71a, BG72, Bjo71, BMTD72, BA71b, BSI76, CH76, CA73, CSR78, DPRC78, Dev78, Dra70, EC77, ET73, Eng76, Etn71, FN77, FB79, GE75a, GH79, HM74a, HP74, IWF75, Jud73, KW72a, KB75, LKT79, LHK^{+}74, LL77b, Mat73, MS74a, MTM72, Mil74, MS77, Nai75, NE74, Ogd70, Par73a, Par71, PL72, Pec74, PH74, PB78, PW74, PP73, Pow74, Pri70, Ric76b, RH74, RH77, SP71, Sie72, SHD72, Spe74, SA70b, Sti76, Str79, VW74, Wal71, War73, Wil70, WH71, WM78, WBG75, Zim70]. Foods [BM71, MJRW70, Sek74, Tan70]. Forage [Rei72, Str79]. Forbes [BNS70]. Foregut [BMS71]. Fork [DDA72, NE74, PW75b]. Formalin [Eng74, Sto72b]. Formation [Cob70, Gri75, KGJ74]. Forming [Dut76]. Formulated [HCS76]. Fort [Clo75, OK78]. FORTRAN [BA72, Hog70]. fossilis [ADSB77]. Four [Cha78b, Kyn74, LJH76, Man79, UH72, Win77]. Fourth [MP75]. Fraction [SC79, You74]. Fractional [GG77, GB78]. Fractions [Bou72, KW72a]. Fragile [LM75a]. Francis [BH75]. Fraser [SHH76]. Free [Hoa77, SND71, SMDN76, TT76]. Free-falling [Hoa77]. Free-Swimming [SMDN76, SND71]. Freedom [MBT70]. Freezing [Eng74]. Frequencies [Dah70]. Frequency [AP75, Eck73, GB78, GW77, MS78]. Fresh [BS79, NF70, SC71]. Fresh-water [NF70]. Freshwater [Boy70, Bur72a, Cha78a, Cob73, DB78, Dea76, FA78, GW77, Hea70, HBD78, JS77, KS71, Kus74b, Leg72, Mag73, Mil72, MRK79, OR77, RK76, Spa70, Str73, SW70, TC79]. frown [Lev73]. Fry [AS70, AJ75, BMTD72, Bow75, BK77, Ell76, EL75, GT72, HSF75, HT70, Mil74, MA75, Nev75a, Nob72, PLCM75, RC71, SHF74, Sta79a, Tho75, TRS78, WW73, YS79]. Function [MM76, vWDB78]. Functional [Koe70, SM74]. Functions [DPRC78]. Fundulus [MTM72, Voy75, WW75]. Fundy [Mes75]. Fungus [Ose77]. furcatus [Ban73]. Further [Byr71, ML72, PH77].

G [Mag73, Str73]. Gag [MH78a, MBT70]. Gain [SNDM77]. gairdneri [Abb72, AZM75, Bou79, BK77, CC78, Del74, DM71, Fas73, GMS72, KW72a, MER77, NC73, OFTM77, Pet77, PB78, Plu72, ROLH78, Tat73, Van74, WKN^{+}76, WFS78, WWN76]. Galveston [FB72, LS73a, MMS75, PT70, RSD74]. Gambusia [CRC^{+}76]. Game [Hai73, Mag76, TL74]. Gamma [BD72, DB70, HBD78]. Gamma-Irradiated [BD72]. Gammarus [NP74, OS74, Ose77, Smi73c]. Gar [KM74, MW72]. Gars [ER72]. Gas [BCSS76, FM78, GH77, Sch74a]. Gas-Supersaturated [GH77]. Gases [PS76]. Gastric [DE78, WKN^{+}76]. Gatewells [Lis71, PF72]. Gear [Yeh77]. Gel [Par70a]. Gene [Ohn70]. Genera [HBH74]. General [MMD77]. Generations [BLCF76, KBvL77]. Genetic [Bow75, KM76, RC71, SDW78, UH72]. Genetics [Fuj70, KM79]. Genome [Ohn70]. Genus [MJRW70]. Geographical [LPH70]. Geographically [NS76]. Geographically-Defined [NS76]. Georges [Lux73]. Georgia [Dah70, DMT77, KDS79, SDKT79, SHD72, Sti76]. georgianus [ES72]. Geothermally [KK78, Kay77]. gibbosus [Ree71a]. Gill [Han74, Hög73, Jes77, Zau70]. Gillnets [Jes73]. Gills [BH74]. Girard [IWF75, PW75b, PP73, Ste73]. Given [Mag73, Str73]. Gizzard [Coo78, CM70b, Jud73, MP76, OFTM77]. Glands [Laa71]. Glycogen [CB72]. Gobius [Hea71]. Goby [Hea71]. Gogebic [SEC77]. Golden [LL71, MSS73, Orc79]. Goldfish [BS73, GPT70, GTP71, GTP72, JBR71, SH74a, SNDM77]. Gonad [Hea71]. Gonadal [BD72, Kai73b, SNDM77]. Gonads [KH72]. Goode [MMS75, SSH72]. Goose [BR76]. Goosefish [BPL75]. Goosequill [EC77]. gorbuscha [HT70, SH74b]. Gorge [Dud74]. Gourami [RP75]. Grabs [How71]. gracilis [OR77]. Gradient [BMS71, SD72b]. Grand [KME70]. Grass [Bai78, BBR79, CSR78, CSG^{+}78, FA78, FL78a, GG78, HBM79, Kil78, LRIF78, Lew78, Mit78, NM78, Pfl78, Ril78, SCR78, Sta74, Sta76, SMS78, SL78, UF78, WC79]. Gravel [PLCM75, KS77]. Gravity [Bon76]. Grayling [KS77, RM71]. Great [Man72b, Smi73b, HS71a, HGJ75, Law70, Smi70a]. Green [BNMC76, Ell74, HP76, KH72, Tau77, Wil75, LS73b, WH75]. Gross [Gun73, MLO77]. Grounds [OP73]. Groundwater [Han75]. Group [PH77]. Grouper [Smi72a]. Growth [Ade78, Alt76, AB70, AS72, AMG73, ALTA77, Bai72, Ban73, Bar78, BA75, BM79a, BA70, Ben75, Ber78, BMTD72, BTL75, BS73, Bur74, CNHT76, CH76, CvW78a, Cob71, Coo70, Coo71, DS78, DC79, DTD76, Dud74, DH70, Elr74, FMS79, Fra78, GE75a, GQ79, Gle78, Gri75, Gri77, Gri78b, GS72b, HS71a, HH79b, Hav73a, Hei75, HP74, HA75, Hil70, HC75, HCS76, IWF75, Joh70, KK78, Kai73a, KBR71, Kil78, KBvL77, KM74, KH78, Kos78, KGJ74, LHBS79, LMG76, MH77, MH78a, MR74a, MLL71, MD70, Med79, Mes75, Mou74, NS75, NC73, Nob75, OK78, Ose77, OM75, PW75b, Pri77, Pri71, Pri75a, ROH79, Ric73, Ric76b, Ric75a, RH77, SSH72, Sav78, SS75, SDKT79, SH78, SA77, SNDM77, Sta78, Sta72, SK71, SM76b, TL76, TD76, WW71a]. Growth [WAWR79, Wat73, WA77, Wol77, WM78, WW76b, Zim70, ZD76].

Gulf [Bor71, Mes76, BPL75, FN71, FB72, GG79, Jen70, LM75b, MMS75, PSG^{+}74]. Günther [Joh70, Joh71]. Guppies [MNL79]. Gut [BF76]. Gymnodinium [CJS71]. Gynogenetic [Sta76]. gyrinus [HR72].

H [Mag73, Str73]. Habitat [BE79, DME77, GH79, GZ73, Hun76, PL72, Que75, Win77]. Habitats [Pis74]. Habits [Ant74, ALTA77, AR71, BA75, BA71a, BMTD72, BA71b, CH76, CA73, DPRC78, EC77, Eng76, Fra74a, GG77, HP74, IWF75, Jud73, KB75, LL77b, Mat73, MS74a, MMB74, MH78b, Nai75, NE74, Ogd70, PL72, Pec74, PP73, RR73, RH74, SHD72, Spe74, Sti76, VW74]. Haemogregarina [BPL75]. Hake [SHD72]. Halawakee [Mat73, MR74b]. Half [MD72]. Half-Life [MD72]. Handling [MLMS74, Mor70a, Sav73, SSG77]. Harbor [PF72, AC72]. Hard [Mac77]. Harengula [SSH72]. harengus [EEW77]. Harvest [Car72, Cla75, GS72a, SEC77, You76]. Harvesting [BS75]. Hatcheries [Spe73]. Hatchery [Eck71, FMS79, Fra74b, HA73, Hav73c, MLL71, MLMS74, Pet77, PB73, Sta79a, Szl74, TB79, War76, WWN76]. Hatchery-Reared [Fra74b, HA73, MLL71, Hav73c]. hatching [KS77]. Hatteras [RH74]. Hawaii [MP75, Wat73]. Hawaiian [Hig70, SN75]. Head [MM76]. Headboat [HCD78]. Headrope [EP73]. Headwater [Hoo75]. Health [Bou72]. Heat [BNS70, Ott74, SH74b, Syl73]. Heat-stressed [Syl73]. Heated [AFS79, BG72, KK78, Kay77, MJN72, MP74, NM74, Sar77, SS75, WW70]. Heavy [Che75]. Hectare [CSG^{+}78]. Held [MLT76]. Hellgrammite [MSS73]. Helminth [ECCC76]. Hematocrit [JP76]. Hematological [BCP76, NOSW75]. Hematopoietic [MA78]. Hemocyanin [RGBL79]. Hemoglobin [Wil75]. Hemorrhagic [CC78]. Henle [Tho73b]. Hepatopancreas [NB72]. Herbicide [WW73]. Heritability [MA78]. Hermaphroditic [Eds70b, RBB75]. Hermaphroditism [SML75]. Herring [AS71b, Bur74, Chi72b, EEW77, LL77a, Mes75, Mes76, MP78, PS77, RSL70, SP71, SC79, SM76b]. heteroclitus [Voy75]. Heterogamety [EC70a]. Heteropneustes [ADSB77]. Heuristic [GH73]. Hexavalent [RS75]. High [AS71a, CH75, RWC73]. Highfin [WW76b]. Hiodon [Gle78, SD78]. Hippoglossoides [MW71]. hippurus [BH74, RH74]. Histochemistry [Bou79]. History [Alt79, Bow70, BLA78, Bur72a, DL71, ER72, HB75, Kir76, KS77, LL77a, OC79, Par72, PS79, RM76, Sim76, SOKEK76, SW70, Tsa72]. Hog [BM78a]. Hogchoker [Kos78]. Holding [BM70, LB76]. Holston [NE74]. homari [RGBL79]. Homarus [Coo70, CCN75, Mor71, RGBL79]. Home [RC76a, Win77]. Homing [KP75, Lab71, MLK77, ML72, SGC^{+}78]. Honolulu [MP75]. Hooked [WJ78]. Hooking [MH70, Rei75, War76, War78, WWN76]. Hormone [Ade78]. Horse [Kai73a, Kai73b]. Hosts [LM76]. Hot [Ano73a]. Hotel [Ano70c, Ano72c, Ano73a, Ano74f, MP75]. Hour [RS75]. hoyi [EB78]. Hudson [Kos78, Wal75a]. Hue [RMG73]. Humboldt [AB70]. Humpback [Alt71, Alt79]. Hundred [Ano72c, Ano73a, Ano74f, MP75]. Hundred-First [Ano72c]. Hundred-Fourth [MP75]. Hundred-Second [Ano73a]. Hundred-Third [Ano74f]. Hundredth [Ano71a]. Huron [EP73]. Hurricane [Hoo75]. Hyacinths [Kil78]. Hybrid [Hei75, LH71, Lew78, Mer71a, Mer71b, Set70, Sta76, ZD76]. Hybridization [Hes70]. Hybrids [Alt71, Etn71, KBR71, Nym70, SS71, Wil75]. Hydrocarbons [Ada75, HWB71]. Hydrocynus [Gri75]. Hydroelectric [TG74]. Hydrogen [AS70, DOB74, KSB78, OS72, OS74, SOKEK76]. Hydrogen-ion [DOB74]. Hydroperla [MSS73]. Hydrophone [GS74]. Hydroponics [LYSB78]. Hydroxide [BCS76, FE71, SS76]. hydroxy [Mar70]. Hypentelium [BM78a]. Hypolimnion [Edw78, Fas73, OFTM77]. Hypothesis [Cai71]. Hypoxia [Bou72, Kir74].

Ice [PF72]. Ichthyomass [Cha76]. Ictalurus [AS71a, BMTD72, BS73, BKS76, CSH74, DB72b, DH70, EIAJ^{+}76, GAR76b, JAR76, JAR78, Kir74, KB75, LL71, Lov78, LL78a, OC79, RI78, Tar71, TK72, Wil70, Wor75]. Ictiobus [MJRW70, SA70b]. Idaho [Bjo71, GC72, GB71, Gri74, PW75b, Pet77]. idella [CSG^{+}78, NM78, Pfl78, Sta74]. Identical [Jen72]. Identifiable [BMS71]. Identification [Coo78, JKS78, PH77, SD78]. II [Sti76]. III [Sar77]. Illinois [BBR76, BM77b, SDC78]. Imidan [JS77]. Immature [GB79]. Immediate [FT76]. Immersion [RD73, Wat70]. Immigration [DLM72]. Immune [GCD^{+}77, RP75, SP76, TSW72]. Immunofluorescent [LA71]. Immunological [Fuj70, HB78]. Impact [Hor77, HGJ75]. Impacts [KB77]. Impairment [Far77]. Impingement [MHM77]. Implanting [HS75a]. Implemented [TL74]. Implications [HM74a, TB74]. Importance [EH76b]. Important [FN71, Qua72]. Impounded [WL75, YHP77]. Impoundment [Erm73, TL76, Sim72]. Impoundments [Ray79, RHW75]. Imprinting [SGC^{+}78]. Improved [AP72, MS70b, Ste75]. Improvement [Bor74]. Improving [Hun76]. Inbreeding [Kin76]. Incidence [GHH77, Vin77]. Inclined [Sla71]. Inconnu [Alt71]. inconstans [Deg70]. Increase [Jen72]. Increased [Jen71, SDW78]. Increments [FB77]. Incubation [Bro75, HB74, Tho75]. Index [BCC78, Sim72, Str79]. Indian [DP73]. Indiana [AMB^{+}77, LRIF78, Zim70]. Indicated [WV73]. Indication [Lar74]. Indicators [OP73]. Indices [HM74b, WK79]. Indigestible [KW72a]. Induced [Abb72, AS79, Car73, GE75b, JWA77, Par73b]. Industry [MWK71]. Inexpensive [Cha76, DD70, DD74, GCD^{+}77, SVL73]. Infection [MLT76, RA72a]. Infections [Ant74]. Infectious [MA78]. Inferred [Nic71]. Infertile [Cla77, Hoo75]. Infestation [BH76b]. Inflow [Hea70]. Influence [Ade78, AMG73, AM75, Cam71, DK75, GMS72, HH70, HBM79, HC76, KS75, MER77, Mus77, PCP73, ROLH78, Sav73, Sch74a, SA71, SM76b, WHW78]. Influences [DM71, Has70]. Influencing [Bow75, Gar73, HD70, HS71b, Mar72b, RC76b, WE76]. Information [Pat73]. Inhabiting [AFS79]. Inhibition [AZM75, Cop72, CA77, GHC^{+}79]. Inhibits [BJ79]. Initial [Bec72, KDS79, PH72, SDKT79]. Injected [SP76]. Inland [MS74a]. Inlet [Bor74]. Inlets [LM71b]. Innovative [Lac73b]. Inorganic [LHL77]. Insect [BM74]. Insecticide [GC72, JS77, KS71, KME70, Kyn74, MM70, MWHH72, NF70, Tsa78]. Insecticides [Kyn74, WW75]. Insects [Neb72, Ste75]. Inshore [DJD79, VD79, Wag72]. Insidious [Odu70]. Instant [Jen72]. Intake [LHK^{+}74, Reh78, Van70, WH71]. Intakes [Lis71, ML71, PF72]. Intensity [EB76, RC76c]. Intensive [Elr71]. Interaction [Bul70, Goo72, LL75b, Tag76]. Interactions [AS72, For74, HBM79, MMC78, Sch70, Smi70a, WHW78]. Intergeneric [Mer71a, Mer71b]. Intermittent [BS77, EL75, EH76b, MMM72, SB78]. Intermittent-flow [MMM72]. Internal [BM72]. International [MV70]. interruptus [IWF75, MG70, MMB74]. Interspecific [Smi76]. Intertidal [Mor70a, Tur77]. Intragravel [Tho73a]. Introduced [Ber78, TB74, UF78]. Introduction [Eve71, Hav73a, Hub70, LMG76, MH78b, Spe74, SL78, Cha73, Roy73a]. Introductions [MG70]. Introductory [Sch76]. Invertebrate [EH76a, GG75, Gri74]. Invertebrates [Hil71a, How72, JS77, MS70a, MWHH72, RK76, ST75, WW70]. Involved [SDW78]. Involving [Kem70]. Iodine [TKM70]. Iodophors [Ame74]. ion [DOB74]. Iowa [BSI76, DB72b, Shi74]. Iraq [ADSB77]. Iron [SS76]. Irradiated [BD72, Gun73]. Irradiation [DB70, HBD78, Hof74]. Island [MW71, Ric76b, Ric75a, SS74]. Islands [Val74, Lev73, SD73]. Isolated [CL75, Sch74b]. Isolating [AP72]. Isopods [LM76]. Isozymes [HB78]. IV [BA72, Hog70]. Ives [RA72a, RB70]. Ivlev [Pow74, Str79].

J. [Kem70]. James [Bur74]. January [Ano74b, Ano75b]. Jaw [MR79]. Jawless [HA75]. Jersey [Ogd70]. Joaquin [KM72, SKK79, Ste77, TC72]. John [BE70]. Johnny [Tsa72]. Joint [MSW79]. Jordanella [SLD78]. Journal [Bri74]. Journals [Ste71]. Juglone [Mar70]. July [Ano74c, Ano75c, Hig70, Ric76b]. Jumping [Ell74]. June [Ano75a]. Juvenile [AS71a, Bar71, BA75, Bea70, Bec72, BA70, BA71a, Bor74, CH76, CA73, Cha78b, CP75, Coo71, CBP76, Dav78, DE78, Ebe70, EK75, Gib78, GZ73, GGB77, HO78, HD70, Hav73b, Hav74a, Hig70, JS73, KH78, KG73, Lev73, Lis71, LCG77, ML71, MS71, Med79, OM75, PF72, PB73, Ray79, Ric75a, Ric75b, RW79, RR73, Sch74a, Sek74, SP71, Sta72, SSE78, Wat70, Wea74, Wil70, YHP77]. Juveniles [Coo78, HSF75].

kadiakensis [NF70]. Kafue [Dud74]. Kansas [KM74, SK71]. Karluk [Gar73]. Katsuwonus [Bat72, ING70]. kawatai [Kaw73]. Kennebec [TG74]. Kentucky [Hoy71]. keta [LL77b]. Kiamichi [ES76]. Kidney [HL75]. Kill [CL75, Kus74a, OC74]. Kinetics [EBM^{+}75]. King [Dev78, MY71, PPSF79, PH72]. kisutch [BS79, Eng76, LHL77, MS74a, Pec70, ZT74]. Kleen [DM79]. km [SA70a].

Known [PKP74]. Kruse [RA72a]. Kuron(R) [WW73].

L [KP75]. L. [Nym70, SN75, Voy75, WL70, Wil75]. Labidesthes [Zim70]. Laboratory [Ben74, CH72, DR75, DD70, HH70, LHL77, LCL78, ML71, PPSF79, PB78, PS79, SSH72, SHF74, Smi72b]. Laboratory-Reared [PPSF79, SSH72, SHF74]. Lac [DME77]. Lacépède [Bea72, CH72, Car73, ED73, Wal73a]. Lactate [Cai71, HB78, Mor70b]. Lacustrine [Ant74]. Ladyfish [Sek74]. Lag [OMS78]. Lake [AS71b, Ano72c, Ano74f, Arm71, Bab71, Bai72, BJ70, BLA78, Bro75, BSI76, CH75, CSG^{+}78, DB72b, ET73, EB76, Eng76, For74, Fra78, GG75, HW77, Han71, Hav73a, HA73, Hav73b, Hav73c, Hav74b, Jen78, Jes76, Kel73, KF74, KC72, Kis74, Koc73, Lac73a, LMG76, Lin76, LH72, MS74a, Moy73, MH78b, Nev75a, Nev75b, Nob72, Par72, Pat77, PS70, Rei72, Sch74b, Spe74, Sta79a, SW70, SM76b, ST77, TB74, WP73, Wal72, WH75, WV73, War78, Wil70, Win77, WM78, WBG75, ALTA77, BK74, BH75, Ber78, BTL75, BH76b, Cha76, CvW78b, Che73, Clo75, Coo79, Del74, DPRC78, DE78, DWB75, ER72, Eds70b, EP73, Ell78, Elr74, EB78, FN77, Gri75, Gri77, HGAF70, HGO^{+}77, Hur73, Jen78, Man72a, ML72, Mou74, NM74, OK78, OC79, Ort79, OKR76]. Lake [Par71, Pec70, Pec74, Pri70, Pri71, Pri75a, Reg71a, Rue77, SEC77, SGC^{+}78, SA77, SA70b, Sta72, Sta79a, Sta79b, Tar74, TG79, Tyu74, VW74, Wag72, Wil78, Wis74, Wol77]. Lake-dwelling [War78]. Lakes [AMB^{+}77, Bur72a, Cla77, Etn71, HP74, Hut71, LC70, Mat78, Mel76, Mom78, NS75, PS70, RWC73, War73, HS71a, Has70, Law70, Man72b, Smi70a, Smi73b]. Lakesucker [Koc73]. Lakeward [Bow75, RC71]. Lamprey [Bea74, Law70, Man72a]. Lampreys [Man72b]. Lampricides [RK76]. LAN [Bou79]. Lance [WAMR79]. Landlocked [AS79, Hav73a, HA73, Hav73c, Hav74b, Lin76, Spe74, War76, War78, WJ78]. Large [GM78, HW76, HKB76, MD71, Mat78, Yeh77]. Large-Scale [HW76]. Largemouth [BH76a, BBR79, BG72, BG75, CH72, Car73, CCM79, Cla75, Cla77, Clo75, Cob72, DE75, ES72, Ell76, Eng77, ED73, FB77, FB78, FL78a, GB72, HL78, KDS79, LHK^{+}74, LCG77, MH78b, Ort79, Sav78, SDKT79, SCR78, SA77, Smi76, WL75, Win77, ZD76]. Larvae [Ant74, ASH75, BH73, CSH74, Coo78, Coo79, Dea78, Eds70a, EEW77, Fui79, FW79, HMJ73, JWA77, KF77, MD71, MUR^{+}76, MP77, PCP73, SSH72, SDC78, SN75]. Larval [BM78a, BM78b, BM79c, CH78a, Coo78, FB72, FL78b, GE75a, GM78, HB77, IWR78, LHBS79, LM71b, MP78, Sie72, SM76b, Tau77]. latipinna [LL70]. Latitudinal [PS77]. Laurentian [Smi73b]. Lawrence [Mes76, Mil74]. LC [RS75]. LDH [WA70]. Lead [Ada75]. Leesville [Sim76]. Leiostomus [HH79c]. Length [BA75, BA70, BA71b, Eng74, FN71, Gri78b, KH78, LL75b, MMS75, MS78, Mci70, MP76, Nic71, RM71, RH77, Sek74, Sto72b]. Length-Frequency [MS78]. Length-Total [MMS75]. Lengths [DDA72]. leniusculus [BGM75]. Lens [Eck71, SC73]. Lentic [HWB71, HW76]. Lepidomeda [RD73]. Lepisosteus [ER72, MW72]. Lepomis [AMB^{+}77, BK74, BM79a, Ben75, BG70, BMC72, BM78b, CMS74, CWW70, DB78, DPRC78, Eat74, Ell74, ECCC76, Etn71, GM72, HD73, HP76, KH72, NS76, Par73a, PW74, PGP77, Ree71a, Sar77, Sav71, Sav73, SS75, Shi74, SOKEK76, Tau77, Wil75]. Lernaea [BWH79, WL70]. Lernaeenicus [VS79]. Lesions [HL75]. Leslie [Hor77, VS76, vWDB78]. LeSueur [MG72b, Ogd70, AS71b, Bow70, Coo71, Hur73]. Lethal [Alc76, Bea72, CS78, HD73, OKR76, RS75, TT76]. leucichthys [Alt71]. Leucine [Bou79]. Level [Cop72, Sav73]. Levels [CS78, MJ72, NF70, OS72, PR77, TGR74, Wor75]. Lewis [Rue77, SW70, Wal72]. Liberated [Mor71]. Life [Alt79, Bow70, BLA78, Bur72a, CH76, DL71, ER72, HB75, Kir76, KS77, LHL77, LL77a, MD72, OC79, Par72, PS79, SA77, SOKEK76, SW70, Tsa72, Tsa73]. Light [Cam71, DR75, EB76, HCS76, RC76c, Smi70b, Syl73, WCH71]. Lighting [DD70]. Lightly [Cla75]. Likelihood [MS78, SMS78]. Lime [SS76]. Lime-Neutralized [SS76]. Limestone [AB75]. Limit [Bab71]. Limiting [WP73]. Limits [Alc76, SCR78]. Limnetic [SDC78]. limosus [Rei72]. Lindane [HW76]. Linear [GS74, HS71b, Str79]. Linear-array [GS74]. Lines [Kin79]. Linnaeus [Alt71, Bai72, Bat72, Cas74, Cas75, HMJ73, MMS75, Ree71a, Sek74]. Lipid [EIAJ^{+}76, ROLH78, dVKP78]. Lipids [YS79]. liquefaciens [LA71]. Lironeca [LM76]. Literature [KB71]. Little [BR76]. Littoral [HW77]. Live [BF76, ECCC76]. Live-Box [ECCC76]. Liver [BCC78, CB72]. Liver-Somatic [BCC78]. Lobster [Ach75, CCN75, OMS78, RGBL79, Coo70]. Lobsters [Dav71, Dav75, Dav78, KS73, Mor71]. Local [Mar72b]. Location [Ebe70, RMG73, SRT73, Vin77]. Locations [VD79]. Loci [Eck73, MSW79, Mor70b, UH72, WA70]. Locomotor [EBM^{+}75, TK72]. Logging [Bur72b, RH75]. Long [CH78b, HBL79, Hun76, OS72, Ric76b, KF74, ST77]. Long-Term [HBL79, Hun76, CH78b, OS72]. Longear [BG70]. Longevity [SEC77]. Longfin [Mou74]. Longnose [BLA78, KM74, Vin77]. Loon [Fra74b]. Lophius [BPL75]. Loss [Leg72, Tar71]. Lota [Bai72]. Louisiana [Ano70c, DPRC78, BA75, BA70, BA71a, BA71b, KH78, RA72a, RHW75, Sek74, YHP77]. Love [Hav73b, Hav74b]. Low [Chi72a, DE75, EEW77, Gri78a, Hei75, Neb72, RCL^{+}78, Sav73]. Lower [KS75, BR76, CBP76, MR74a]. Lowered [CSH74]. Luciolates [CvW78b, Ell78]. lucius [Cas74, Cas75, DME77, HMJ73, MT72]. Ludington [ALTA77, BTL75, BLA78]. Lugworm [SN76]. Lunar [LA71]. Lure [MH70]. Lure-Caught [MH70]. Lutjanus [MBT70, SD73, WAWR79]. lutrensis [FEL79, MGW76]. Lymphocystis [MLT76].

M [RC76c, Wil70]. M. [MUR^{+}76, MP78]. Machiavelli [Reg71b]. machrochirus [Sav73]. Mackerel [Dev78, Kai73a, Kai73b]. Macrobenthic [Bro77, How72]. Macrobenthos [Sim72]. Macrobrachium [RSD74]. macrochirus [AMB^{+}77, BK74, BM79a, Ben75, BMC72, CMS74, CWW70, DB78, Eat74, ECCC76, GM72, HD73, NS76, Par73a, PW74, PGP77, Sar77, Sav71, SS75, Shi74, SOKEK76, Wil75]. Macroinvertebrate [DCA71]. Macroinvertebrates [Her73, Jac71]. macrolepidotum [BM77b, BM79c]. maculatus [Kos78]. Madison [Erm72]. Madtom [HR72]. Magnified [HWB71]. Magnitude [PKP74]. Maine [Ach75, BPL75, CCN75, HD70, Hav73a, HA73, Hav73b, Hav73c, Hav74a, Hav74b, Jen70, Kir76, LH72, MF75, MLK77, Nev75a, OMS78, RM73, Rei72, SP71, Spe74, TM74, TG74, War73]. Mainstream [HS75c]. Maintain [LYSB78]. Making [SRM76, PLZ75]. Malate [MSW79]. Malathion [CP75, Han72, OS72]. Male [CS78, Gue75, TK72]. maliger [MLT76]. Management [Ach75, All73, And73, BLCD75, CLB^{+}73, Cob73, Cru73, Dic73, Dut76, Gul78, HM74a, HL76, HA76, Hun76, Jos70, Mag73, MP71, OE75, Pat77, PLZ75, SH75, Smi73b, Spi73, Str73, TB79, TB74, UF78]. Management-Related [Hun76]. Managing [Mag76]. Manipulation [TG74, Wal71]. Manitoba [Gle78, Kel73]. Manure [BBR78]. Maomao [Med79]. Mapping [SRT73]. March [Ano70f, Ano74b, Ano75b]. Marine [CA73, Cob73, LM76, Mag73, MMD73, PSG^{+}74, Str73]. marinus [Bea74, Man72a]. Marion [ET73]. Mark [And72, Kli74, Reg71a]. Mark-Recapture [Kli74]. Mark-Recovery [Reg71a]. Marked [And72, Kim76, KH78]. Marking [HT70, LM74, MR79, RD73, Wat70]. Marks [Coo70, DM73, Eng77]. Marsh [KH78, TL76, YHP77, RHW75]. Mary [Reg71b, Win77]. Maryland [MB75b]. masquinongy [Gal73, GE75a, GHH77, HH79a, MC79, MLMS74]. Massachusetts [HC75, LPH70, Mor71, MC78, PH77, Ree71a]. Masses [FM70]. Matamek [GG75, PCP73]. mate [Wat73]. Materials [Boy73]. Matrix [Hor77, VS76, vWDB78]. Mattamuskeet [Tyu74]. Maturation [Ber78, Elr74, Kai73b, Mus77]. Mature [Dav75, MLK77, MY71]. Maturity [ALTA77, Bat72, Joh71, Kos78, MH70, Tau76]. maxillingua [FL78b]. Maximum [HB74, Lar77, MS78]. Maximum-Likelihood [MS78]. McConaughy [Van74]. Mean [GCB73]. Means [BCC78, Set70]. Measured [BBAN75]. Measurement [DE71]. Measurements [BJ70, Pit73]. Measuring [HCD78, PH74, RH77]. Mechanical [HKB76]. Media [FE71]. Medication [BJ79]. Meeting [Ano70c, Ano70f, Ano71a, Ano72c, Ano73a, Ano74f, MP75]. Meetings [Cha71]. megalotis [BG70]. Melanophores [Man72b]. melanops [HB77, MG72a]. melas [DB72b, Wil70]. Member [Ano70f]. Memoriam [Ano72b]. Men [Sch70]. Mendota [VW74, Wis74]. Menhaden [FB72, HO78, HC73b, KG72, KG73, KGJ74, Lev73, LM71b, MMS75, Nic71, Reg71b, SH72, WWLC77, Wea74]. Menidia [ASH75, Bar78, CM70a, EB76, HSS71, LMG76, MH78b]. Mephistopheles [Reg71b]. mercedis [SKK79]. Mercenaria [Mac77]. Mercuric [DOB74]. Mercury [DOB74, Har78, HG74, HJB^{+}78, KO74, KF74, LWO76]. Meristic [Wal73a]. Mesh [Jes77, MM76]. Metabolic [Par73b]. Metabolism [Fra73a, GPT70, LC70, MS70b, Spe73]. Metal [AMB^{+}77, Che75, GW77]. Metamorphosis [PS79]. Meter [EP73, KJ78, MS70b]. Metering [MMM72]. Method [AP72, BF76, Che75, DD70, JKS78, KH71, LM75a, MJ76, PH74, Reg71a, SRM76, VS76, WH75, vWDB78]. Methods [Bar71, GM78, Kem70, RW79]. Methoxychlor [EH76a, MBN72]. Methyl [Far78]. Methylated [HJB^{+}78]. Methylating [HJB^{+}78]. Methylmercury [PB78]. Methyltestosterone [FMS79]. Mexico [Bor71, FN71, GG79, Jes76, LM75b, PSG^{+}74, SM76a, TW75]. Michigan [ALTA77, BTL75, BLA78, BH76b, Cla77, DWB75, Eds70b, EB78, FN77, HW77, Hil71a, Men75, OKR76, Pat77, Pec74, Pri77, Ric76a, SEC77, SGC^{+}78, Sta72, Sta79a, Tau76, Wag72, WH75, Wil70]. Microhematocrit [BJ76]. microlepis [MH78a, MBT70]. Micropogon [KH78, MMS75, YHP77]. Micropterus [BG72, BG75, CH72, Car73, CCM79, ED73, FB77, FB78, GH79, GB72, LCG77, MR79, MH78b, Nev75a, OK78, RC76c, Sav78, VR75, Wal73a, WL75]. Microspectrophotometric [BR75]. Microsporidan [RI71]. Microstomus [SM76a]. Mid [HGM78, MM77]. Mid-Atlantic [MM77]. Mid-Columbia [HGM78]. Middle [GS78, HS75c, Mun70]. Midge [KDR74]. Midwater [Hig70, WWO73]. Migrating [Bea70, Gar73, Rei75]. Migration [Arm71, Leg72, Pec74, Sta72, Tho73a, Zau70]. Migrations [BR76, Bow75, Coo70, FM70, KG73, LM70, RC71, Ray79]. Migratory [Ric75a]. Miles [SA70a]. militaris [Ose77]. Milkfish [Tsa78]. Mine [Sim72, SDW78, LC70]. Mineral [DTD76]. Mineralization [HGJ75]. Minimum [Dav75]. Minnesota [Etn71, HGO^{+}77, Moy73, NS75, PS70, PW75a, Plu72, Sta78, Win77]. Minnow [And70, BNS70, Bru71, FL78b, GHC^{+}79, HP74, Hoy71, KP75, LM75b, MVM77, NPD74, Pit73, Wal73b]. Minnows [Bro70, Cop72, HSF75, Jen72, Moy73, RS75, SHF74, SM74]. minutus [Hea71]. Minytrema [HB77, MG72a]. Mirex [HGS^{+}74, SN76, Tag76]. Mississippi [Jud73, CM70a, EB76, LMG76, Mci70, MW72, MH78b]. Missouri [GS78, SW70, Bow70, FF74, MTM72, MS77, Pfl78]. Mitchell [Wil70]. Mitchill [Ben74, Bur74, Eck71, Eck73, LL71, Mil74, Nob73]. mitchilli [MMS75]. Mixtures [EK75, PLCM75, SLD78]. Moapa [Cro76, DB72a]. Model [BM79b, DC79, Fox70, GQ79, GG79, HM74a, HW76, KKM^{+}74, OE75, OMS78, Ort79, Pat75, Pow74, RH73, RR73]. Models [Wal75a]. Moderate [AZM75]. Modification [BR75, DM79]. Modified [ECCC76, Van70]. Modifying [TL78]. mollispinis [RD73]. Molluscs [CJS71]. Molly [LL70]. Molt [Abb74]. Monitoring [CMS74, GCD^{+}77, LHBS79]. Mono [MP71]. Monoaromatic [EEW77]. Monobothrium [Wil74]. Monochloramine [SBCG79]. Monofilament [Han74]. Monogenea [Bec72]. Monona [NM74]. Montana [Reh78]. Montsweag [MF75]. Monumental [BR76]. Mooneye [Gle78, SD78]. mordax [Coo78, FN77, MC78]. Moribund [PR77]. Morning [BRJ78]. Morone [DMT77, HA75, MR74a, MB75b, MKK73, MUR^{+}76, MP78, NK73, OM75, Rei72, Ste77, TC72, VW74, WH71, Wis74]. Morphological [FL78b, Tau77]. Morphology [JAR76, Set70, SM74]. Mortality [BS79, BMC72, BM79b, CH78a, Cai71, CvW78a, GE75a, GE75b, HC75, HCP76, Jen71, JWA77, KDS79, MH77, MH70, Nob72, OHC79, PR77, SA77, TC79, VS76, WW71a, War76, War78, WJ78, You72, vWDB78]. Mosaics [AS79]. Mosquitofish [CRC^{+}76, Han72, Kyn74, Ott74]. mossambica [KS75, Que75]. Motility [Wal75b]. Mottled [FW77, LL75b]. Mountain [PW75b, PP73, Sim76, TD76]. Mouthless [Men74]. Movement [DL78, EL75, HS71a, HGO^{+}77, KW72a, Lab71, Men75, Mun70, PW75b, RR73, SD72b, Tho73a]. Movements [Bjo71, Del74, DME77, DD74, DMT77, GCD^{+}77, HGM78, HC75, LM70, MLK77, MB75b, Mor71, Nic71, NM78, SND71, SHH76, Tyu74, VD79, WL75, Win77]. Moxostoma [Bow70, BM77b, BM79c, FW79]. MS [CB72]. MS-222 [CB72]. Mud [OC74]. Muddy [DB72a, RM76]. Mugil [SN75, TB71]. Müller [Tho73b]. Mullet [SN75, TB71]. Multile [MH76]. Multispecies [Hög73, Tan70]. Mummichog [Voy75]. Munsel [Han71]. murphyi [Kai73a, Kai73b]. Muscle [BKS76, LWO76]. Muscles [RB70]. Muskellunge [Gal73, GE75a, HH79a, MC79, MLMS74]. musty [LS73b]. Mutant [MG72b]. Mycteroperca [MH78a, MBT70]. mydas [BNMC76]. Mysis [DR75, Smi70b]. Myxobolus [GHH77]. Myxosoma [Hof74, O’G79]. Myxosporida [GHH77].

Na-K-ATPase [AZM75]. naphthoquinone [Mar70]. Nassau [Smi72a]. nasus [Alt76]. National [Ano70f, Bea70, Cla74, Erm72, KK78, Kay77, Lab71]. Native [Cro76, Hoo75, KC72, Ril78]. Natural [Eck71, KDS79, Kus74a, NC73, Set70, You72]. Naturalization [SMS78]. Nature [JBR71, Ohn70]. Nauplii [Wea74]. Near [BLA78, ALTA77, BTL75, CA73, MF75]. Neascus [Vin77]. Nebraska [HB71, Van74]. nebulosus [JWA77, RI78]. Necroses [RB70]. Necrosis [MA78]. Needs [Eip73]. Negative [GM72]. Nekton [KTJ75]. Nektonic [HKB76]. nelma [Alt71]. neogaeus [Sta78]. Neomysis [SKK79]. Nereis [CE76]. nerka [DE78, MA78, SHH76]. Net [GW72, Hoa77, KTJ75]. Net-board [GW72]. Nets [Han74, Hög73, Jes77, MM76]. Neutral [FE71, Wat70]. Neutralized [SS76]. Nevada [Cro76, DB72a, Koc73]. nevadensis [RD73]. Newfoundland [VD79]. Niagara [GHH77, HH79a]. Niche [DPRC78]. Nichols [Kai73b]. niger [Mci70, MG72b, War73]. Nightlights [Wic73]. Nighttime [MF75, RBC78]. nigricans [BM78a]. nigrofasciata [Mat73]. nigromaculatus [CH78b, LMG76, SH77, Coo71]. Nine [DJD79]. Ninespine [GS72b]. Ninety [Ano70c]. Ninety-ninth [Ano70c]. ninth [Ano70c]. Nitrification [Spe73]. Nitrite [CA77, SW74, TRS78]. Nitrogen [BE70, BS78, Boy79b, NBS76, Sav73, Sta74, Wor75]. No [Hil71b]. Noise [BH73, GCD^{+}77]. Noise-Immune [GCD^{+}77]. Non [Mar72a].

Non-Persistent [Mar72a]. Nonheated [AFS79]. Nonlicensed [DH78]. Nonrandom [Mor70b]. Nonuniform [MDS78]. Normal [Gal73]. Norman [OC79]. North [Dev78, OC79, RH74, Tyu74, VS79, All72, Bat72, Bri74, BBR78, DL71, Dra70, Gri78b, HP74, LM71b, MS71, Mat78, MG70, MGW76, PW75b, SH78, SA70a]. Northeastern [FM70]. Northern [AS70, Bey71, BM78a, Cas74, Cas75, DME77, Has70, HMJ73, LL75b, MT72, MTM72, WM78, ZD76, Bor71, Bur72b, Fui79, KB75, Mom78, Pat77]. Northwest [PSG^{+}74]. Northwestern [BGM75, MW79]. notatus [Ste73]. Note [Wal71]. Notemigonus [LL71, MSS73]. Notes [All72, HD73, HSS71, LM71b, RSD74]. Notropis [ABW72, DL71, FEL79, GG77, GB78, HB75, HD73, MR74b, MGW76]. Notropus [BCS76]. Noturus [HR72]. Nova [Mes76]. November [Ric76b]. nubila [BNS70]. nubulosus [KB75]. Nuclear [RI78]. Nucleus [SC73]. Nuisance [Mit78]. Number [Kim76, LPH70, Tho73b]. Numbers [HKB76]. Nuptial [Mad71]. Nursery [Bar71, KH78, OP73, WJ78]. Nutrient [Boy70, Hai73, WFS78]. Nutrients [WP73]. Nylon [Hög73]. Nymph [WW71b]. Nymphs [Ell70].

O [Mag73, Str73, RWC73]. O. [LL77b, SH74b]. Oahe [Has70]. Objectives [Cru73]. Observation [Sal73]. Observations [BCSS76, CvW78b, Lew78, MJN72, PH72, Ree71a, TD76]. Observed [SHH76]. Obtain [CSG^{+}78]. Occupying [CA73, PL72]. Occurrence [Alt71, BA70, BPL75, BWH79, CBP76, CJS71, FW77, KBR71, KG72, LM71b, MS71, MGW76, SM76a, SHD72, WL70]. Ocean [All72, CH76, Dra70, FM70, Jos70, MW79, Qua72, RH74]. ocellata [BA75, BA71b]. October [Ano74d, Ano75d]. oculatus [MW72]. Off [LPH70, PH77, DDA72, HC75, IWR78, Kai73a, Kai73b, MH78a, Mor71, Qua72, RH74, SM76a, TW75]. Offal [Boy70]. Offspring [Bul70]. Ohio [WW76b]. ohione [RSD74]. Oil [EK75, EEW77, MRK79, SC79]. Oil-dispersant [EK75]. Oils [EK75]. Oklahoma [ES76, FEL79, Ort79, SA77]. Olencira [LM76]. Olfactory [Jah72]. olivaris [HS75a]. olmstedi [Tsa72]. Omaka [Wat73]. Oncorhynchus [Arm74, Ber78, BS79, DE78, Ebe70, Eng76, GGB77, GH77, GT72, HL75, HT70, KM76, LHL77, LL77b, MA78, MS74a, Pec70, SHH76, SH74b, SSE78, ZT74]. One [PPSF79, RK76, Ano71a, Ano72c, Ano73a, Ano74f, MP75]. Oneida [Che73, CH75, For74, Nob72, Tar74]. onitis [OS78]. Onslow [Dev78]. Ontario [Fra78, Hur73, KME70, Mom78, Reg71a, WM78]. Open [BJ70, NM78]. Opeongo [Reg71a]. Operation [Wor75]. Opposed [EBM^{+}75]. Opsanus [NOSW75]. Optimal [SH75]. Optimization [Spi73]. Optimizing [Fox70]. Oral [HN77]. Orcas [MW71]. Orconectes [Mom78, Rei72]. Order [EBM^{+}75]. Oregon [DWB75, Han71, KM76, Smi73a]. Organisms [BMS71, BGS77, CR73, DCA71, GE75a, GC72, HKB76, Par73a, Rod73].

Organochlorine [KF74]. Organoleptic [WC79]. Organophosphate [Cop72]. Orientation [GB79, LCG77, RI78]. Orifice [Lis71]. Orifices [PF72]. Orleans [Ano70c]. Osmerus [Coo78, FN77, MC78]. Osmolality [LL71]. Osmoregulation [SC71]. Ostracod [Kaw73]. Ostracods [KS71]. Other [Bou72, Cob71, Goo72, Man79, SND71, DB78, GCD^{+}77, Hav74a, Hil71b, MLT76, Tau77]. Otolith [Bar78]. Otoliths [Jen70, Mes75]. Otter [KJ78]. Outfalls [Tsa73]. Outflow [Ell74]. Output [FMD77]. ovalis [LM76]. Over-Fishing [Sil77]. Overwinter [MH76, OHC79]. Owens [MP71]. Oxide [FB79]. Oxygen [AMG73, AM75, Boy73, BRJ78, Boy79b, CSH74, CH78b, CCM79, DE75, HP76, KDR74, Neb72, NBS76, RH75, RGBL79, RBC78, RB79, SH77, TGR74, Voy75, WAWR79]. Oxygen-Nitrogen [NBS76]. Oxygenation [OFTM77, MJ76]. Oyster [MWK71, NOSW75]. Ozark [BNS70, FF74]. Ozarks [MTM72].

p [DZ72, HN77, MNL79]. p-chlorophenyl [DZ72]. P. [CP75, DLM72, GZ73, LMG76]. Pacifastacus [BGM75]. Pacific [All72, FM70, Jos70, BNMC76, CH76, EEW77, Mor70a, NB76, PS77, Qua72, SC79, TW75]. pacificus [SM76a]. Pad [SM74]. Paddlefish [Reh78, RH77]. Pagrus [DDA72, MH77]. Paiute [DL78]. Palaemonetes [Far77, HBM79, NF70]. Pallas [Alt71, Hea71, RM71]. pallasi [EEW77]. Pallette [Eng76]. Panel [Hil71b]. Pantothenic [MA75]. Panulirus [Dav71, Dav75, Dav78]. Papers [Mag73, Str73]. Paralichthys [SW75]. Paralithodes [MY71]. Parameters [Bou72, CC78, vWDB78]. Parasite [BWH79, Clo75]. Parasites [BH74, ECCC76, OP73, Ril78, Val74, VIS70, WL70]. Parasitic [KG72, LM76, VS79]. Parasiticide [RF74]. Parasitized [RBB75]. Parasitizing [RI71]. Parathion [Far77, Far78]. Parental [Wil75]. Park [Bea70, Cla74, Erm72, KK78, Kay77, Lab71]. Parophrys [OP73]. Parr [FMD77]. Participants [KB77]. Participation [DH78]. Particle [Lev74]. Particular [Hut71]. Partitioning [GH79]. Parvin [Lac73a]. Pass [DLM72]. Passage [Dom73, Sla71]. Passing [PF72]. Pastures [Boy76]. Pathogenic [Sch74b]. Pathogenicity [CL75]. patronus [FB72, MMS75]. Patterns [ADSB77, BK74, DD70, Sar77, SH74a]. PCB [NP74]. PCB’s [Sta79a]. Peach [MHM77]. Pelagic [BSI76, Cha76, KW71, RM73, Wic73, WWO73]. pelamis [Bat72, ING70]. Pelleted [WFS78]. Pelvic [OC75]. penaei [RA72a]. Penaeid [FN71, NB72, RHW75, Tur77]. Penaeus [Bea70, CNHT76, CP75, DLM72, GZ73, MECC76, RA72a, RB70, RBB75, RI71, VIS70, Wat76]. Pennsylvania [Hoo75, MHM77, MMN79]. pensacolae [SSH72]. Pentachlorophenate [RS75]. Pentachlorophenol [PGP77]. Peprilus [MW79]. Perca [BTL75, CH75, LWO76, NS75, Nob73, Nob75, TKM70]. Perch [BTL75, BS77, CH76, CH75, Cla77, Eng74, For71, For74, Fra78, HB71, HCS76, IWF75, KF74, LWO76, MR74a, MG70, MUR^{+}76, MP78, MMB74, NS75, Nob73, Nob75, Qua72, Rei72, SH78, Sie72, Sto72a, Sto72b, ST77, TKM70, Tar74]. Perciformes [DDA72]. Percina [Mat73]. Performance [Bea74, DJD79, GMS72, HO78, KS75, Sch74a]. Periodic [TB78]. Periodically [Orc79]. Periods [Cam71]. Permanent [EL75]. Permanganate [Lay71, MB75a, TB77]. Permethrin [JAKG78]. perotteti [Tho73b]. Persistence [Kor73]. Persistent [Mar72a]. Peru [Sch70]. Pesticide [OC74, PR77]. Pesticides [Cop72, HGAF70, MJ72]. petenense [Gri78a, HB74, Joh70, Joh71]. petersoni [DL71]. Petromyzon [Bea74, Man72a]. pH [Bea72, RCL^{+}78, RA72b, SA71]. Phanerodon [Ban73]. Phase [KH72]. Phenol [RS75]. Pheromone [TK72]. Phosphorus [DB76, Ket75, LB77, Lov78, Sta74, Wor75]. Photoperiod [KH72, SKFM72]. Photosynthesis [MLO77]. Phoxinus [KP75]. Phyllodistomum [FW77]. Phylogenetic [Beh70, Hes70]. Physical [FB78, Gre70, Gun73, LDM71, PKP74]. Physicochemical [Her70]. Physiological [MLMS74, WWN76]. Phytoplankton [AB75, LRIF78, TB78]. Pickerel [Mci70, MG72b, War73]. pidschian [Alt71]. Pigment [Eng77, HT70, RD73]. Pigments [And72]. Pike [AS70, Bey71, Cas74, Cas75, DME77, Has70, HMJ73, MT72, WM78]. Pike-Bluegill [Bey71]. Pilot [BM72]. pilsbryi [HD73]. Pimephales [And70, Bru71, HP74, MVM77, RS75, SM74]. Piney [FF74]. Pink [Bea70, CNHT76, Hea72, HT70, PPSF79, RI71, SH74b, VIS70]. Pinnixid [Pri75b]. Pipe [Sla71]. Pirate [SH78]. Pisces [Mar72b, Ree71b, Vin77]. Piscicide [Mar72a]. Piscivorous [Wag72]. Pituitary [Laa71]. Placement [Lis71]. plagiusa [Sti76]. Plain [SH78]. Plankton [Boy76, Kus74a]. Plant [NM74, RI78]. Planted [Fra78, Hav73c, Pec70]. Plants [Nug70]. Plasma [BR75, Gre70]. Plastic [BJ70, Kem70]. Plate [LB76]. Platichthys [PS79]. platorynchus [MS77]. Platte [EC77, Hil71a, Tau76]. platycephalus [OC79]. Platyrhinoides [TSW72]. Pleuronectiformes [Sti76]. Plume [SRT73]. Poecilia [LL70]. Pogonias [Ric73]. Poinsett [HGAF70, SA70b]. Point [KF74, KDS79, SDKT79, ST77, SM76a]. Poisoning [Fre73]. Policy [Dut76]. Political [Cru73]. Polluted [KB75, SDW78, Tsa72]. Pollution [CMS74, OC74]. Polychaete [CE76]. Polychlorinated [NPD74, NP74]. Polyculture [BBR78]. Polymorphic [Koe70]. Polymorphism [RSL70]. Polymorphisms [Eck71, WA70]. Polynomial [OMS78]. Polyodon [RH77]. Polyploid [AS78, AS79]. Pomatomus [LM70, NRMJ74, Ric76b, Sil75]. Pomoxis [AR71, CH78b, Coo71, LMG76, SH77]. Pompano [BA70, BA71a]. Ponar [How71]. Pond [Abb72, BBR76, CLB^{+}73, CSR78, ES72, EC77, Ell74, Elr71, Gre71, HA76, Kir76, Kus74a, LJH76, LL78b, MMN79, MLK77, Orc79, Ree71a, Sav78, VIS70]. Pond-Reared [Abb72, VIS70]. Ponds [AB75, BH76a, BBR79, Boy73, Boy76, BS78, BRJ78, Boy79a, BT79, BTR70, DB76, FA78, FL78a, Kil78, LRIF78, Lew78, LB77, MWHH72, RC76a, RC76b, RBC78, RB79, Shi74, Tan70, Tsa78, TB78, WC79]. Pontchartrain [DPRC78]. Pontoporeia [Smi72b]. Pool [Dom73, Jud73]. Population [Bjo71, Bro70, BWH79, CSG^{+}78, Etn72, For71, Hai73, HW74, HA73, Hav74b, HS75b, Hoo75, Hor77, Jen71, Joh70, Kus74a, MC78, MW79, Ort79, PW75a, Pri77, Ree71a, Sav78, SC73, Van74]. Populations [And73, Bai78, BH76a, Bél74, Bey71, BCC78, Cla75, DC79, Eck71, EH76a, Erm73, FL78a, Fox70, GG75, Kin76, KJ78, KM79, Kus74b, Kyn74, LH71, Mac77, MD71, MKK73, NS76, Nob75, OE75, TB74, UF78, Wal71, WA70]. Porgy [DDA72, MH77]. Porichthys [Ste73]. Portable [Sal73]. Portion [CK71]. Possible [Spa70]. Post [KS77]. Post-hatching [KS77]. Postal [Car72]. Postlarval [DLM72]. Postorbital [HL75]. Potamologica1 [Szl74]. Potassium [DB76, Lay71, MB75a, TB77]. Potential [DB72b, MLO77, TG79, WV73]. Potentials [DD74]. Potomac [CK71]. Pound [MM76]. Pound-Head [MM76]. Power [BA72, MHM77, NM74, Nug70, RI78, TG74]. Powered [HKB76]. Poygan [Pri75a]. praegustator [LM76]. Pre [Sim72]. Pre-impoundment [Sim72]. Precalibrating [MJ76]. Precipitating [Boy79a]. Precision [MDS79]. Predation [BM77a, Cob72, CM70b, ES72, For74, Fra74b, Hut71, KF77, Mac77, Pec74, SCR78, Syl73]. Predator [Dea78, Far78, Goo72, KO74, SA70a, Tag76, WHW78]. Predators [Ell70, MP76]. Predicted [Dud74]. Predicting [Bél74, BRJ78, MDS79, RBC78]. Predictions [Mat78]. Preference [DME77, ED73]. Preferences [Par71, RC76c]. Preferred [MER77, OKR76, RI78]. Preliminary [AC72, Sim72]. Presence [GE75a, Mad71]. Present [Eip73, KB77, Kim76]. Preservation [GAR76a]. Preservatives [How72]. Preserved [Her73]. Prevalence [Sch74b]. Previous [CLB^{+}73]. Prey [DE78, Eng76, Far78, For74, Fra74a, Goo72, KO74, PPSF79, Tag76, WA77, WHW78]. Primary [MMD77, Mel76, Ste71, WP73]. Prior [Que75]. Pristis [Tho73b]. Probability [MDS78]. Probable [MGW76]. Problems [Beh70, Ano70g]. Procambarus [GA77]. Procedures [HS75a, SGC^{+}78]. Process [HJB^{+}78]. Prochristianella [RA72a]. Production [AB75, BBR79, BTR70, Cla75, Cla77, GB71, Gue75, HW74, Hav73b, HA76, HGS^{+}74, Jes76, LL75a, LB76, MLO77, Mom78, Nev75a, Par73a, Par72, PW75a, SA77, Sta76, TBM79, WP73, Wal71]. Production-Type [LL75a]. Production/Biomass [Mom78]. Productivity [AB75, BJ70, Mel76, WV73]. Products [GA77]. Progeny [Gun73]. Program [Hog70]. Programs [Mag73, MC79, Str73]. Project [BM72, RM76, Sim76]. Prolarval [GE75a]. Prolonged [BK77]. promelas [And70, Bru71, HP74, MVM77, RS75, SM74]. Promoter [FMS79]. Promotion [Ade78]. Propagation [Wal76]. Properties [Gre70, HB78, Her70]. Proposal [HWB71]. Proposed [Str79]. Prosopium [ALTA77, PW75b, PP73, TD76]. Prospective [HA76]. Protein [Bou72, DS78, Eck71, Eck73, Laa71, NS76, ROLH78, Sav71, Shi74, SL73, UF78]. Proteins [BR75, Par70a, SC73]. Proteocephalus [Ant74]. Proteus [RP75]. Protolarvae [SD78]. Protozoa [GHH77]. Prudhoe [MRK79, SC79]. pseudoharengus [Coo78, Dom73, DM73, Gan76, Hav73b, Hut71, Kis74, Lin76, OKR76, Ric75a, RW79, SC71, Tyu74]. pseudolimnaeus [OS74, Ose77]. Pseudomonad [Sch74b]. Pseudopleuronectes [BCP76, CE76, Fra73a, Fra73b, Fra74a, Lev74, LPH70, Lux73, UM72, VD79]. Public [Kea71]. Publication [Gun71]. Publications [Ano70d, Ano70e, Ano71b, Ano71c, Ano72d, Ano72e, Ano72f, Ano72g, Ano73b, Ano73c, Ano73d, Ano73e]. Puget [Pat73]. Pugheadedness [Bor71]. pugio [Far77, HBM79]. Pulsating [MLL71]. Pulsator [SVL73]. Pulsed [FMD77]. Pumped [HE76, RM76, Sch76]. Pumpkinseed [Ree71a]. punctatus [AS71a, BMTD72, BS73, BKS76, CSH74, DH70, EIAJ^{+}76, GAR76b, JAR76, JAR78, Kir74, LL71, Lov78, LL78a, Tar71, TK72, Wor75]. punctulatus [OK78]. Pungitius [GS72b]. Pupfish [MP71, Nai75]. purpureus [LM75a]. Purse [JS73]. Pylodictis [HS75a]. Pyramid [Koc73].

Quality [DS78, GM79, Kus74a, Lar74, LYSB78, Tsa73, TB77, TBM79, WK79]. Quantification [BE79]. Quantifying [JKS78]. Quantitative [EH76b, Hud70, Jac71, Kus74b]. Quantitatively [GT75, Hoa77]. Quantity [DS78]. Quebec [GG75, PCP73]. Questionnaire [Dut76]. Quiescent [KH72]. Quillback [Mad71, MLT76, WW76b]. Quinaldine [Mor70a, SA71].

R [Mag73, Str73]. Rabbitfish [EK75]. Radiation [RB79]. radiatus [VS79]. Radio [HGO^{+}77]. Radio-biotelemetry [HGO^{+}77]. Radioactivity [RR73]. Radioisotopic [PH74]. radiosus [MP71]. Rafinesque [AR71, BG70, Bru71, BMC72, CMS74, Eat74, GM72, KH72, Kir74, LL71, MG72a, NK73, Par73a, RS75, Sav71, Sav73, Shi74, SM74, Tar71, VW74, Wil75, Wis74]. Rainbow [Abb72, Ame74, Bor74, Bou79, BBAN75, BS77, BK77, CC78, Coo78, Coo79, CB72, Del74, DM71, DM79, EL75, EH76b, Fas73, Fra78, FE71, Fre73, GMS72, HS71a, HN77, Kay77, Kin76, KBvL77, Kin79, KW72a, MRSK70, MLL71, MER77, MC78, NC73, OFTM77, PB78, Plu72, PKP74, ROLH78, ROH79, SGC^{+}78, SB78, Sie72, SW74, Sta72, Sta79b, Tat73, UH72, Van74, WKN^{+}76, WFS78, WBG75, WWN76]. Rainbow-Steelhead [Del74]. Ramifications [HE76]. Rana [KF77, WL70]. Range [Win77]. Rapid [Mos70]. Rappahannock [KBR71]. Rate [AS71a, Ban73, DE78, Eds70a, HW74, KBvL77, KH78, Pri71, TBM79, Wat73, WKN^{+}76, You76, vWDB78]. Rates [AS72, BM79a, BM79b, DC79, Gan76, HP76, HSS71, Jes77, MJ76, Med79, Nob72, Nob73, Par73b, TL78, VS76, You72]. Rathbun [Abb74]. Ratio [BCC78, Str79]. Ration [DE78]. Rational [Spe73]. Ratios [Bat72, BD72, Cas75, NBS76, RH74, Mom78]. Ray [LPH70]. Rays [WW71a]. Re [MJ76]. Re-oxygenation [MJ76]. Reaction [EBM^{+}75]. Reactions [FT76, LCL78, MJN72, RCL^{+}78]. Read [Hil71b]. Reality [MWK71]. Reared [Abb72, Fra74b, HA73, MLL71, MLMS74, PPSF79, PS79, SSH72, SHF74, VIS70, Hav73c]. Rebuttal [Kem70]. Recapture [DM73, Kli74]. Recaptures [You72]. Received [Ano74c, Ano74d, Ano74b, Ano74a, Ano75c, Ano75d, Ano75b, Ano75a]. Receiving [BG72, MP74, PKP74]. Rechanneling [Etn72]. Recirculated [LB76, LYSB78]. Reclaimed [Orc79]. Recognition [RMG73]. Recolonization [Nev75b]. Recommendations [CR73]. Recorded [RW79]. Recording [SND71]. Records [Bec72, MLT76]. Recoveries [Hav73c]. Recovery [ABW72, Fre73, HS71a, HW74, Reg71a]. Recreational [DH78, MH78a, PLZ75]. Recreationists [Man79]. Recruit [Wis72]. Recruitment [ECCC76, HCP76]. Red [BA75, BA71b, DDA72, FEL79, FA78, MH77, MBT70, MGW76, TL76, WAWR79, CJS71, NS75, PS70, Wat70]. Redbreast [BM78b]. Redd [Hea72]. Redds [Han75, HC76]. Redhorse [Bow70, BM77b, BM79c]. Reduced [BH75, Hea70, Par73b, SH77]. Reduction [CH78b, Qua72, SA70a]. Reference [Hut71, Kin79, RI78, TM74]. Reflection [Ohn70]. Regard [Gib78]. Regeneration [WW75]. Regimes [Ott74]. Region [CCN75]. Regions [HGJ75]. regius [SHD72]. Regulated [Hav74a]. Regulation [Bab71]. Regulations [KB77]. Rehoboth [DP73]. Reintroduction [WV73]. Related [ABW72, Ano70g, Bjo71, BKS76, Coo71, Cop72, Hun76, TC79, Vin77]. Relation [Ant74, BH76a, Bea74, Bro77, Eng76, FM70, GGB77, GH77, Hil70, Hun76, MH70, Men75, NM74, Nob75, OHC79, PT70, SKK79, Ste77, Tar74, TC72]. Relations [FN71]. Relationship [BA75, BA70, BA71b, Gri78b, LL75b, RS75, Sek74, Zau70]. Relationships [Bro77, DDA72, GG75, Hes70, LM76, MMS75, Mci70, RM71, TB77]. Relative [BA72, Hav73c, Hög73, Ogd70, OHC79, PB78, Yeh77]. Release [Ebe70, MGW76]. Released [Del74, MH70, Pet77, TK72]. Releases [Edw78]. Releasing [Rei75]. Reliability [Str79]. relicta [DR75, Smi70b]. Remarks [Sch76]. Removal [CSG^{+}78, DMB76, LRIF78, NC73, WV73]. Removing [PS76]. Reparameterizing [GQ79]. Replicate [BTR70]. Reply [Cha73]. Repopulation [BG70, OC74, Phi75]. Report [Ano70g, Ano72a, Sim72, Ano70f]. Reported [Rad73]. Representatives [MM70]. Reproduction [Bai72, BH76a, Ben75, IWF75, NPD74, NP74, Pec70, Smi72b, Sta78, TD76, Wal73b, dVKP78]. Reproductive [AS78, BG75, DB72b, FEL79, GMS72, Hoy71, Kay77, Kir76, Koc73, MH70, MR74b, Pet77, Rue77, SMS78]. Requirement [Ket75, Lov78]. Requirements [HMJ73, Sav71, SMS78]. rerio [SP76]. Research [All73, And73, Ano70f, CLB^{+}73, Cob73, Hun76, Kem70, Mag73, Smi73b, Str73]. Reserves [dVKP78]. Reservoir [BG72, BG75, Edw78, KDS79, KM74, MP74, NK73, NM78, OE75, Pat75, SDKT79, SDC78, Van74, Yeh77, AR71, VR75]. Reservoirs [AFS79, Joh70, Joh71, MJRW70, Sar77, SS75]. Residency [Cha78a]. Residue [KME70]. Residues [GPT70, GTP71, KF74, SA71]. Resistance [BNS70, BGM75, BGS77, HH79c, NF70, Orc79, OS72, SKFM72, SDW78]. Resistant [Kyn74]. Resort [Ano74f]. Resource [TL74]. Resources [KB71]. Respect [Wal71]. Respiration [WH71]. Respiratory [CCM79, MSS73, NOSW75, SND71]. Respirometer [TGR74]. Response [CMS74, DOB74, MNL79, Par73a, PPSF79, Ric75b, RP75, SP76, TSW72, WWN76]. Responses [CCM79, CRC^{+}76, Chi72a, Cop70, DS78, DW78, FMD77, MLMS74, Mos70, NOSW75, OR77, SSG77, TK72, Van70]. Restocking [BLCD75]. Result [Hoo75, Par73b]. Results [BM72, Jen72, SA70a]. Retardation [BD72, WW75]. Retention [And72, Coo70, DTD76, Eng77, Har78, LWO76, MM76, WWLC77]. Return [HBD78, MBT70]. Returning [You74]. Returns [Del74, TBM79]. Reuse [Spe73]. Revealed [DM79, JH77]. Review [CLB^{+}73, CR73, MMD77]. Reviews [Ano70h, Ano70i, Ano71d, Ano71e, Ano72h, Ano72i, Ano72j, Ano72k, Ano73f, Ano73g, Ano73h, Ano73i, Ano74g, Ano74h, Ano74i, Ano74j, Ano75g, Ano75h, Ano75i, Ano75j, Ano76b, Ano76c, Ano76d, Ano76e, Ano76f, Ano77a, Ano77b, Ano77c, Ano77d, Ano77e, Ano77f, Ano78a, Ano78b, Ano78c, Ano78d, Ano79a, Ano79b, Ano79c, Ano79d, Ano79e, Ano79f, WW76a]. Revolving [LB76]. Rheotropism [DM79]. Rhinichthys [BLA78, FT76, Vin77]. rhinichthysi [Vin77]. Rhode [Ric75a, SS74]. Rhomboplites [Bor71, Gri78b]. rhotheus [Tho73a]. Rhythms [RC76c]. Richardson [Fas73]. Riffle [EH76a]. Rings [Bar78]. Ripe [TK72]. River [Alt71, Bec72, Bur74, CA73, Cro76, DP73, DMT77, ES76, GM78, Gar73, GG75, Gle78, Hil71a, HS75c, JS73, KK78, Kay77, KME70, KBR71, Kos78, LL77b, MS77, MC78, NE74, PW75b, Pet77, PCP73, RM73, RSD74, Ric76a, SML75, SD72a, Ste77, SK71, Tau76, TD76, TG74, TC72, VS79, WJ78, BGS77, BE70, BR76, BCSS76, BM78a, BM78b, BM79c, CK71, CBP76, DB72a, FF74, GHH77, GD79, GS78, HH79a, HGM78, Jud73, KM76, KG72, MR74a, Mun70, Ray79, RR73, SA70a, SHH76, SW70, Wal75a]. River-Back [RM73]. Riverine [SBCG79]. Rivers [Arm74, GT75, KM76, Man79, EC77, Erm72]. rivulatus [EK75]. RNA [BCC78, Mus77]. RNA-DNA [BCC78]. Road [Bur72b, HGJ75]. Roccus [Pri70]. Rock [Bou72, BM77b, GH79, HGJ75, LWO76]. Rockfish [MLT76, Pat73, Pri75b]. Rocky [GT75]. Role [Bur72a]. Roller [GD79]. Roosevelt [Ano70c]. roseipinnis [HB75]. Rostral [Tho73b]. rostrata [Ogd70, Hur73]. Rot [MMD73]. Rotenone [Nev75b, Orc79, Par70b]. Rough [Hai73, Jes76, MR74b]. Round [ALTA77]. Roving [MDS78]. Run [RM76, Arm71]. Runoff [LCL78]. rupestris [Bou72, GH79, LWO76].

S. [Mag73, Str73]. Sable [Ric76a, SA70a]. Sac [AS70, BK77]. Sacramento [SKK79, IWF75, MG70, MMB74, Ste77, TC72]. Sagehen [Erm73, EL75, GS72a, SG72]. Saginaw [EP73]. Sailfin [LL70]. Salamanders [ET73]. salar [AS79, Gib78, HA73, Hav74b, Nym70, Rob70, War76, War78, WJ78]. Salcha [LL77b]. Salicylanilide [Mar72a]. Saline [HP74]. Salinities [Voy75]. Salinity [BN75, Chi72b, Han72, OM75, SS71]. Salmincola [GMS72]. Salmo [Abb72, AZM75, Alc76, AS79, Arm71, Bou79, BK77, CC78, Del74, DW78, DL78, DM71, DK75, Fas73, GMS72, Gib78, Gri74, Han71, HA73, Hav74b, KW72a, Lab71, MH70, MSW79, MER77, ML72, NC73, Nym70, OFTM77, Pet77, PB78, Plu72, ROLH78, Rob70, Tat73, Van74, War76, War78, WJ78, WKN^{+}76, WFS78, WWN76]. salmoides [BG72, BG75, CH72, Car73, CCM79, FB77, FB78, GB72, LCG77, MH78b, RC76c, Sav78, WL75, ED73]. Salmon [All73, AS79, Arm74, BE70, BR76, Ber78, Bjo71, BD72, BS79, BJ79, BH76b, CS78, Cha78a, Cha78b, CBP76, CA77, DE78, DB70, DWB75, Ebe70, Ell70, Eng76, FMD77, FMS79, FM70, Gar73, Gib78, GGB77, GH77, GT72, Gun73, HD70, Hav73a, HA73, Hav73c, Hav74a, Hav74b, Hea72, HL75, HT70, HBD78, JS73, Ket75, KM76, LHL77, LH72, LL77b, ML71, MA78, MS74a, MNL79, NBS76, NB76, Nym70, PF72, Pec70, Pec74, PLCM75, Ray79, Rob70, RR73, Sch74a, SB78, Sla71, SW74, SS76, Spe74, SHH76, SH74b, SSE78, Syl73, Tho75, Van70, WW71a, War76, War78, WJ78, Zau70, ZT74]. Salmonid [Bon76, BJ76, GG75, RMG73]. Salmonidae [Beh70]. Salmonids [Ant74, O’G79, Smi73a, SSG77]. Salt [AZM75, Ano72c, BS79, HSS71, RHW75, Zau70]. Salt-marsh [RHW75]. saltatrix [LM70, NRMJ74, Ric76b, Sil75]. Saltwater [TL76]. Salvelinus [Ada75, AS78, Ben74, BLCF76, DOB74, Eck71, Eck73, Gib78, Gri74, HC76, HBL79, HJB^{+}78, MLK77, Mil74, OC75, WE76]. Same [PL72]. Sample [CH73]. Sampler [GT75, Jac71]. Samples [JP76, LM71a, Sel71]. Sampling [All72, CH78a, DCA71, GT75, GM78, GGG75, Hoa77, Hud70, Kus74b, MDS78]. San [SKK79, Ste77, TC72, KM72]. Sanctuary [HS75b]. Sand [GD79, HC76, Hea71, WAMR79]. sapidissima [Chi72a, Leg72]. sapidus [Abb74, CV72, LCL78]. Sardine [SSH72, Tal73]. Sardine-Anchovy [Tal73]. Satinfin [GB78]. Saturometer [FSM75]. Sauger [Wal72]. saurus [Sek74]. Savannah [DMT77]. Sawfish [Tho73b]. saxatilis [DMT77, HA75, MB75b, MKK73, MUR^{+}76, MP78, OM75, Ste77, TC72]. Scale [Coo71, HW76, Hil70, JKS78, Sil75]. Scaled [SSH72]. Scales [Cob70, HH79b]. Scaphirhynchus [MS77]. Scavengers [Ell70]. Schedules [Har78]. Schoodic [Hav73a, HA73, Hav73c, LH72]. Schooled [Ell76]. Schooling [BIJ70, CH76, Par73b]. Schools [Pit73]. Sciaenops [BA75, BA71b]. Science [Lac73b, Lac75, SE73]. Sciences [Roy73a, Cha73]. Scientific [Cha71]. Scientist [All71, Eve71, Mea71]. Scientists [Hil71b, Ste71]. Scomberomorus [Dev78]. Scope [DK75]. Scotia [Mes76]. Screens [ML71]. Sculpin [FW77, LL75b, NE74, OR77, PW75a, Tho73a]. Sea [Arm71, Bea74, BNMC76, Dav71, Law70, MM77, SC71, Man72a]. Sea-run [Arm71]. Seagrass [Bro77, CA73]. Seaperch [Ban73]. Searobin [MD70]. Season [GH77, JAR78, MLK77]. Seasonal [ALTA77, BA70, BCP76, CCN75, FN77, Gle78, HW77, HGM78, IWR78, Jes77, KME70, Kel73, Laa71, Lev74, Lin76, MM77, SH74a, SP71, Tar71, UM72, VD79, You76, dVKP78]. Seasonality [VS79]. Seatrout [JWA77]. Seawater [BJ79, CA77]. Sebastes [CH76, MLT76, Pri75b]. Second [Abb74, Ano73a]. Secondary [MMD77]. Section [Phi75, SL78]. Sections [BG70]. Securing [BF76]. sedecim [DDA72]. Seeking [BIJ70]. Segregation [DPRC78, MSW79, Mor70b, Moy73, TB74]. Seines [GW72]. Seining [JS73]. Selected [DB78, ED73, SBCG79]. Selection [BM79a, CSR78, Eng76, GZ73, Han72, KBvL77, Lev74, Que75, Str79]. Selective [Bur74, CSG^{+}78, CM70b, Gan76, Han74, Mar74, Par71]. Selectivity [DE78, SA70b, Yeh77]. seleniris [DL78]. Selenium [HG74]. Self [Van74]. Self-sustaining [Van74]. Semi [YHP77]. Semi-impounded [YHP77]. Seminal [Gre70]. Semotilus [BM71, Ree71b]. Sensitivity [MRK79]. Sensors [Byr71]. Separate [Arm74]. Separating [Ste75]. Separation [BMS71, Par70a, Tau77]. September [Ano70c, Ano71a, Ano72c, Ano73a, Ano74c, Ano74f, Ano75c, Hig70, MP75]. Sequential [HA76, Sim72]. Sera [PR77]. Series [OE75]. Seriously [Hea70]. Serum [Arm74, LL71, NS76, Shi74, UM72]. Session [Eve71]. setiferus [CP75, DLM72, GZ73, RBB75]. Setting [GW72]. Several [CC78, JAKG78, WW75]. Sewage [FT76, Tsa72, Tsa73]. Sewage-polluted [Tsa72]. Sex [Bat72, BD72, Cas74, Cas75, RH74]. Sexing [NDN76]. Sexual [BNS70, Bat72, KH72, Tau76, Tho73b]. Shad [BMS71, Chi72a, Coo78, CM70b, Gri78a, HB74, Joh70, Joh71, Jud73, Leg72, MJN72, MP76, Mos70, OFTM77, Wat70]. Shallow [Bar71, KTJ75, Kus74b, Lin76]. Shape [JKS78]. Sharks [Cai71]. Sharpe [Elr74, Has70]. Shaw [WL70]. Shear [MUR^{+}76]. Sheep [TD76]. Sheepscot [RM73]. Sheepshead [Cop72, GHC^{+}79, HSF75, LM75b, SHF74]. Shellfish [KB77]. Shelter [BIJ70, Dav71]. Shelters [VR75]. Sheraton [MP75]. Sheraton-Waikiki [MP75]. Shiner [ABW72, BCS76, FEL79, GG77, GB78, HB75, HD73, LL71, MSS73, MR74b, MGW76, SB78, Sie72]. Shiners [BM77a, Orc79]. Shoals [KP75, VR75]. Shock [Dea78, HH79c, Smi70b]. Shocked [ABW72]. Shocks [PKP74]. Shore [DP73]. Short [DM73]. Short-Term [DM73]. Shorthead [BM77b, BM79c]. Shovelnose [MS77]. Shrimp [Bar71, BM72, Bea70, CNHT76, CP75, DLM72, FN71, GZ73, GG79, HBM79, Kli74, MECC76, NF70, NB72, PPSF79, RA72a, RSD74, RB70, RBB75, RHW75, RI71, Tur77, VIS70, Wat76, Wea74]. sicculus [Zim70]. Siganus [EK75]. Significance [HE76, MT72]. Silk [SD73]. Silverjaw [Hoy71, Wal73b]. Silverside [ASH75, EB76, LMG76, MH78b, Zim70]. Silversides [Bar78, CM70a]. Silvex [WW73]. Simazine [TB78]. Simple [Eis75, GQ79, Goo72, KO74]. Simplified [MMM72]. Simulated [Cla75, Fra73a, GGB77, Van70]. Simulating [DD70]. Simulation [BM79b, HL76, Ort79]. Simultaneous [SRT73]. Simultaneously [BMC72]. Since [BR76]. Single [MH76, Sav78]. Site [VS79]. Sites [WE76]. Situ [LHBS79]. Six [MBT70, UH72]. Size [Bab71, BKS76, CH73, Dav75, DC79, EIAJ^{+}76, HA76, Jes77, MH78a, Mer71b, MM76, OHC79, SCR78, TC79, Vin77, WFS78]. Size-Age [MH78a]. Size-Related [BKS76, TC79]. Sizes [Ohn70]. Skipjack [Bat72, ING70]. Slimy [PW75a]. Small [And72, BF76, Boy70, BS75, BM70, Fra78, GW72, Gri74, HM74b, Kil78, LM75a, MP74, PW75a, TB74, WP73, WP76]. Smallmouth [BBR79, Cla75, Cla77, Cob71, DC79, FF74, GH79, Hai73, KF74, MR79, Mun70, Nev75a, OHC79, Wal73a]. Smelt [Coo78, FN77, Hav73a, Mou74, MC78, SB78, Sie72]. Smith [Clo75, Mag73, OK78, Str73, RSD74, Sim76]. Smokies [HGJ75]. Smolt [Cha78a]. Smolts [BD72, BS79, BJ79, LL77b, NB76]. Snail [ES72]. Snake [BR76, Mun70, Ray79]. Snapper [Bor71, Gri78b, MBT70, SD73, WAWR79]. Snook [MS71]. Snow [PPSF79]. Snowmobile [Ada75]. Social [Ach75]. Society [Ano70b, Ano70c, Ano70f, Ano71a, Ano72a, Ano72c, Ano73a, Ano74f, Gun71, MP75, Ano74e]. Sockeye [All73, Cha78a, DE78, FM70, Gar73, MA78, SHH76, Syl73]. Sodium [RS75]. Soft [AB75]. Soil [Gre71]. Solar [RB79]. Sole [MS74b, MW71, OP73, SM76a]. Solid [DD74]. Solid-State [DD74]. Solids [NOSW75]. Soluble [Eck73, SC79]. Solutions [SDW78]. Solving [MV70]. Somatic [BCC78]. Some [Boy70, Bur72b, CWW70, DC79, EB76, Han75, HE76, Her70, HS75b, HS71b, How72, LL70, MM70, MMD77, PLCM75, Pit73, RC76b, SH75, Wal72, WAMR79, WW70, dVKP78]. Sonic [KW72b]. Sort [LM71a]. Sound [Bro77, MW71, Pat73, SS74, Ric76b]. Sounder [Cha76]. Source [HE76]. Sources [EIAJ^{+}76]. South [Elr74, HGAF70, Has70, MS77, NE74, SA70b, Wal72, BG72, Bur72a, DDA72, EC77, FEL79, Gri78b, Mci70, Nev75a, Tat73, TL76, VIS70]. Southeastern [Arm71, BA71b, CH76, MH78a, RA72a, Wil74]. Southern [ADSB77, LM70, MS74b, Mes76, TW75]. Southwest [HBH74, KH78]. Southwestern [Bai72, YHP77]. sp [BPL75, CL75, PSG^{+}74, RBB75]. Space [Smi76]. Sparidae [DDA72]. spathula [RH77]. Spatial [Cam71]. Spawner [GG77, GB78]. Spawning [Ben74, BM77b, CH72, Car73, CH78b, DM71, EH76b, GG77, GB78, HH70, Hea72, Hil71a, JAR78, KBvL77, Kis74, Koc73, MT72, MG72a, Mou74, RH75, SH77, Smi72a, Smi73a, Smi76, SW70, Tho73a, Tyu74, VR75, WE76]. Special [SL78, TM74]. Speciation [Rob70]. Species [AC72, BS75, BSI76, Dea78, DJD79, ES76, Edw78, ED73, Etn71, Far78, Fui79, LM71b, Moy73, MM77, PL72, Ric76a, SHD72, Smi70a, Smi73c, TK72, Wil74, Wil75, WW70, Tau77]. Specific [Bon76, Cop72]. Speckled [SKFM72]. Speed [FB77, FB78, GT72]. Sperm [BJ76, Gre70, GAR76a]. Spermatology [GAR76b]. Spermatozoa [Wal75b]. Spermatozoal [JAR76, JAR78]. spilopterus [GG77]. Spin [Her73]. Spin-Drying [Her73]. Spiny [Dav71, Dav75, Dav78]. Splake [Fra78]. Spoil [SS74]. Spontaneous [RB70]. Sport [CH73, Car72, GCB73, GS78, WWO73]. Sportfishing [LS73a]. Spot [Bar71, HH79c]. Spotfin [GG77]. Spottail [SB78]. Spotted [HB77, JWA77, MG72a, OK78]. Spotting [Lux73]. spp [MECC76, PCP73]. Sprayed [HT70]. Spring [Arm74, HL75, War73]. Springs [Ano73a]. Squamation [And70]. Squoxin [BM74]. St [EC77, Mes76]. Stage [KS77, PPSF79]. Stage-One [PPSF79]. Stages [Cha78b, LHL77, SOKEK76]. Stain [RD73]. Stained [BM77a]. Staining [KS73, MD71]. Standard [DDA72, EP73, Kin79, MS70b, MMS75]. Standing [HW74, HD70, Hav74a]. Stanozolol [BS73]. stappersi [CvW78b, Ell78]. Starry [PS79]. Starvation [Sav71]. State [DD74, Ant74]. States [GG78, MH78a, Mil72, SMS78, SL78, Wil74]. Static [LSN70].

Station [MHM77]. Stationary [KTJ75]. Statistical [PHF^{+}77]. Status [Cro76, MG70, Pfl78]. Stay [KH78]. Ste [DME77]. Steelhead [AZM75, CS78, Cha78b, Del74, NB76, PF72, Pet77, PLCM75, PB73, Ray79, Rei75, SD72a]. Steindachner [Gue75]. stellatus [PS79]. Stenodus [Alt71]. Sterility [AS78]. Steroid [BS73]. Stickleback [Deg70, GS72b, Val74]. Stimuli [DW78]. Stizostedion [HGO^{+}77, JKS78, Wal75b]. Stock [Hog70, JKS78, SH72]. Stocked [BBR79, HS71a, KC72, Kil78]. Stocking [Bai78, FF74, Han71, KM79, MC79, SGC^{+}78]. Stocks [Qua72, UH72]. Stolephorus [LM75a]. Stolothrissa [CvW78a]. Stomach [KH71, Mci70]. Stonefly [Ell70, MSS73]. Stony [Ste75]. Storage [HE76, RM76, Sch76]. Storer [Tsa72]. Stores [OHC79]. Storm [LCL78]. Strain [GMS72]. Strains [Eng77, PGSP75, ROH79]. Straits [BH74]. Strategies [Gul78]. Strategy [BLCD75]. Straying [Pec70]. Stream [Bjo71, Bor74, BS75, BWH79, DW78, Eis75, EH76b, Etn72, FF74, HW74, HM74b, HD70, Hav74a, Hoo75, KK78, Kay77, Lar74, Men75, Nai75, PW75a, Phi75, SH78, Ste75, WW70, WW76b]. Streamer [MR79, MECC76]. Streams [BG70, Bjo71, Bur72a, Bur72b, GB71, Gri74, Ogd70, Pfl78, Sek74]. Strength [Has70, Ste77, Wal72, WH75]. Streptococcus [CL75, PSG^{+}74]. Stress [BMC72, CC78, CRC^{+}76, MMD77, Sav73, SSG77, WWN76]. stressed [Syl73]. Stresses [Goo72]. striata [MM77]. striatus [Smi72a]. Striking [HL78]. Strip [LC70]. Strip-mine [LC70]. Striped [DMT77, HA75, HC73a, KBR71, Kil78, KW72b, MD70, MB75b, MKK73, MUR^{+}76, MP78, OM75, Ste77, SN75, TB71, TC72, Wal75a]. Structure [Clo75]. Structures [KW71, Wal73a, WWO73]. Studfish [MTM72]. Studies [Bow78, Byr71, CL75, FB79, HH79b, LA71, ML71, ML72, MBN72, Mes76]. Study [AC72, Bey71, BM79b, Cla74, GAR76b, Kli74, PH77, Sal73, SC73]. Studying [OE75]. Stunted [HA75]. Sturgeon [HGM78, MS77]. Sub [KS77]. Sub-gravel [KS77]. Subalpine [WP73]. Subcutaneous [LM74]. Subjected [Elr71, Jen72]. Sublethal [Far77, KO74, KS71, LHL77, OS72]. Subsampling [HKB76]. Subsections [Mil74]. Subspecies [ZD76]. Substrate [Ben74, DCA71, Jac71]. Substrates [GT75, GGG75, Ste75]. Subsystems [Dic73]. Subtropical [dVKP78]. Subunit [Mor70b]. Success [Mer71b, Pet77, SH77]. Successful [CH72, LM75a]. Successive [JP76]. Sucker [Bea72, BM78a, Cam71, HB77, MG72a, Sie72, SML75]. Suckers [EC77]. Suction [GT75]. Sugar [CB72, LM71a]. Suggested [RH77]. Suitability [MR79]. Sulfate [Boy79a, TB78]. Sulfide [AS70, HGJ75, OS72, OS74, SOKEK76]. Sulfonate [HS71b]. Sulfuric [SDW78]. Summary [LHA78]. Summer [BM71, NE74, Sek74, TM74, Wil70, Win77]. Sun [GB79, LCG77]. Sun-Compass [LCG77]. Sunapee [Kir76, MLK77]. Sunfish [BH76a, BG70, BM78b, DPRC78, DB76, Ell74, HP76, KH72, LH71, Lew78, LB77, Par73a, Sav71, Sav73, Tau77, Wil75]. Sunfishes [Etn71, Hei75, Hes70]. Superior [AS71b, Bai72, Ber78, Jen78, Man72a, Pec70, Sta79b]. Supersaturated [GH77, NBS76, NB76]. Supersaturating [PS76].

Supersaturation [FM78]. Supplemental [GA77, LH71]. Supplementation [MA75]. Supply [WP76]. Suppression [Smi76]. Surface [Par73a]. Surfacing [Par70b]. Surfperches [AB70]. Surge [BLA78]. Surgical [HS75a]. Surplus [Fox70]. Surplus-Yield [Fox70]. Survey [Car72, Erm72, MDS78, MDS79, Wal75a]. Surveys [CH73, Dut76]. Survival [AZM75, And72, ASH75, Ben75, Bro75, Che73, Cob71, DM73, DE75, Ebe70, Eds70a, Elr74, For74, Fra73b, Fra78, Gri78a, GS72b, HC76, HD70, IWF75, LHBS79, MLL71, MH76, Neb72, NPD74, NP74, NBS76, NB76, NC73, OM75, PKP74, SSH72, SEC77, Sta79a, SM76b, Tar74, Tho75, TT76, WA77]. Susceptibility [BM77a, MS70a, MM70, O’G79]. Susceptible [Kyn74]. Suspended [NOSW75]. Suspension [BCS76]. Suspensions [SS76]. Susquehanna [BM78a, BM78b, BM79c]. Sustained [Lar77]. sustaining [Van74]. Swamp [FA78, Kus74a]. Swarm [Set70]. Swimmerets [KS73]. Swimming [Bea74, DJD79, FB77, FB78, GS74, GT72, HO78, KS75, OS72, Sch74a, SMDN76, WH71, SND71]. Swimways [RC76a]. Swine [BBR78]. Symbiosis [Spa70]. Symphurus [Sti76]. Symposium [Hub70]. Synthesized [LS73b]. System [BM70, CMS74, Del74, Eis75, KO74, LYSB78, MC78, NM78, PCP73, RSD74, TG79]. Systematics [Beh70]. Systems [Boy79b, LJH76, PS76, Spi73, UF78, WP76, WA77].

Tadpole [HR72]. Tag [Dav78, Del74, MECC76, You72]. Tagged [KG73, KGJ74, LM70, MBT70, MB75b, Mor71, SEC77]. Tagging [BM72, LM75a, ML72]. Tags [BM72, MR79, Sto72a, You74]. Tailwater [Tsa72]. Taiwan [Tan70]. Taken [BMS71]. Talbot [Mag73, Str73]. tanganicae [CvW78a]. Tanganyika [CvW78b, Ell78, Cha76]. Tank [FM78, SW75]. Tanks [BM70, CNHT76, Cob72, SD72b]. Tannin [Wal75b, Wal76]. Tanytarsus [NP74]. Tanzania [CvW78b]. Tapeworm [Ant74]. tau [NOSW75]. Tautog [OS78]. Tautoga [OS78]. Taxonomically [BMS71]. TCDD [HN77, MNL79, WNH79]. Teaching [GH73, Lac73b, SE73, TL74]. Technique [BJ76, ECCC76, Goo72, HA76, KS73, LA71, LM74, NDN76, RH77, SRT73]. Techniques [RMG73, RD73]. Teeth [Tho73b]. Telemetry [NM78, WL75]. Teleostean [EC70a]. Temephos [Tsa78]. Temperature [Ade78, Alc76, AS72, ASH75, Bea74, BM79a, Bjo71, Bru71, BMC72, CMS74, CCM79, Dea76, DR75, DK75, Eds70a, EC70b, Gib78, Gri78a, HBM79, HH79c, HD73, HMJ73, HB74, HCS76, Jen72, KH72, KS75, MH70, MER77, Mos70, OHC79, OS78, OR77, OM75, RH75, SSH72, Smi72b, SC71, SMDN76, SSG77, TM74, TL78, WW71a, Wal73a, WKN^{+}76, WFS78]. Temperatures [AZM75, AS71a, BGM75, Bro75, Chi72a, Hei75, Med79, Nug70, OKR76, RI78]. Temporal [Wea74]. Ten [Hut71]. Tendencies [BIJ70]. Tennessee [WC79]. Tensions [RWC73]. Tentans [KDR74, DZ72]. tergisus [Gle78, SD78]. Term [DM73, HBL79, Hun76, CH78b, OS72]. Terminal [Abb74]. Test [BBAN75, FMS79]. Testicular [JAR78]. Testing [GQ79, WK79]. testudinum [Bro77]. Tetrachlorobiphenyl [BBAN75].

Tetrachlorodibenzo [HN77, MNL79]. Tetragoneuria [WW71b]. Texas [CLB^{+}73, FEL79, FB72, LS73a, Men74, PT70, RSD74, Sar77, SS75, Yeh77]. Texoma [ER72]. TFM [Kaw73, KDR74]. Thalassia [Bro77]. Their [Bul70, Bur72a, CH76, Coo70, EIAJ^{+}76, Etn71, HBD78, Kel73, Pis74, UF78, ZD76]. Thelohania [RBB75, RI71]. Theoretical [Jen71]. Theory [And73, Cob73, Mag73, SH75, Smi73b, Str73]. Thermal [CRC^{+}76, Dea78, GGB77, HH79c, Nai75, Ott74, Smi70b, Smi73c, SRT73, SKFM72, SN75]. Thermocycles [SNDM77]. Thermoregulation [NM74]. Third [Ano74f]. Thornback [TSW72]. Threadfin [Gri78a, HB74, Joh70, Joh71, MP76]. Threatened [Mil72]. Three [AB70, BBR76, BGS77, BLCF76, Deg70, DM73, DTD76, Etn71, FN71, HBH74, Hud70, KBvL77, KM76, Moy73, SGC^{+}78]. Threespine [Val74]. Threshold [RS75]. Thunnus [Dra70]. Thymallus [KS77, RM71]. thynnus [Dra70]. Thyroid [DWB75, LM75b]. Tickler [CV72]. Tidal [DLM72, MF75, TM74]. Tide [CJS71]. Tigerfish [Gri75]. Tilapia [DS78, Dud74, Gue75, KS75, Que75, RS79, RBC78]. Tilesius [MY71]. Time [FB77, Gar73, OE75, PH74, TT76, Sar77]. Times [Rad73]. Tissue [BKS76, Che75, Kir74, RWC73]. Tissues [Dea76, GTP72, KME70, Tar71]. Toadfish [NOSW75]. Tohopekaliga [Wil78]. Tolerance [BJ79, Chi72b, DR75, EC70b, HSS71, MA78, Ott74, Smi70b, Smi72b, Smi73c, SS71, SN75]. Tolerances [Alc76, FM78]. Tomales [Ban73]. Tool [GM79]. Topical [RMG73]. Torrent [Tho73a]. Tortugas [Dav75]. Total [Boy79b, DDA72, Kim76, MMS75, MP76, PT70, RM71, RA72b]. Tow [CV72]. Toxaphene [MMC78]. Toxic [BLCF76, CJS71, DMB76, FT76, MJ72, RK76]. Toxicant [Mar70, Mar74]. Toxicants [Goo72]. Toxicities [Cha78b, EK75, Kaw73, LHL77]. Toxicity [Ame74, CA77, Eat74, FE71, GHC^{+}79, HN77, HS71b, HG74, JAKG78, JS77, KSB78, LSN70, MWHH72, Mar74, MB75a, MBN72, MP77, OS74, SP76, SW74, SOKEK76, SH74b, TRS78, Voy75, WW73]. Toxicology [DM79]. Trace [AMB^{+}77, GW77]. Trachinotus [BA70, BA71a]. Trachurus [Kai73a, Kai73b]. Tracking [JH77, KW72b, ML72, SHH76]. Train [Fra74b]. Traits [EIAJ^{+}76]. Transactions [Ano74e]. Transferrin [WA70]. Transferrins [Her70]. Transformations [Bur74, Wis74]. Transient [AC72]. Transition [Hög73]. Transmit [RI71]. Transmitters [FMD77, GS74, HS75a]. transmontanus [HGM78]. Transplanted [SGC^{+}78]. Trap [SS74]. Trappers [Bab71]. Travel [Gar73]. Trawl [CV72, Cla74, EP73, KJ78, Wat76]. Trawling [Hig70]. Treated [Nev75b, OS72]. Treatment [TB77]. Trematoda [Vin77]. Trends [SS74, Wea74]. triacanthus [MW79]. Tributaries [Coo79, Sta79b]. Tributary [EL75, Pec70, Pec74, Sta72]. Trichogaster [RP75]. trichopterus [RP75]. Trinectes [Kos78]. triseriata [TSW72]. Trophic [Bro77]. Tropical [Jos70, Mel76, Wis72]. Trout [Abb72, Ada75, AZM75, AS78, Ame74, Arm71, AJ75, Bab71, Bél74, Ben74, BLCF76, BE79, Bjo71, Bor74, Bou79, Bow75, BBAN75, BS77, BK77, Bur72a, CC78, CB72, Del74, DW78, DL78, DM71, DM79, DOB74, DK75, Eck71, Eck73, EL75, EH76b, Fas73, Fra74b, Fra78, FE71, Fre73, GMS72, GS72a, Gib78, Gri74, HH70, HL76, HS71a, Han71, Han75, HW74, Har78, HC76, HN77, HB78, Her70, Hog71, HBL79, Hoo75, HJB^{+}78, Hun76, Jah72, Jen71, Jen78, JS73, KK78, Kay77, Kin76, KBvL77, Kin79, KW72a, Kir76, KM79, Lab71, MRSK70, MH70, ML71, MLL71, MSW79, MER77, MLK77, ML72, MH76, Men75, Mil74, Mor70b, NC73, Nym70, OC75, OFTM77, PF72, PW75a, Pet77, PLCM75, PB78, Phi75, Plu72, PB73, PKP74, RWC73, RC71, Rei75]. Trout [ROLH78, ROH79, SGC^{+}78, SG72, SA70a, Sla71, SW74, SS76, Spe73, Spe74, Sta72, Sta79a, Sta79b, SDW78, Tat73, Tau76, TRS78, TB74, UH72, Van74, WE76, WKN^{+}76, WFS78, WA70, WBG75, WWN76]. trutta [DW78, MSW79, Nym70]. Trypanorhyncha [RA72a]. tshawytscha [Arm74, Ber78, Ebe70, GGB77, GH77, GT72, HL75, KM76, LL77b, SH74b, SSE78]. Tubercles [Mad71]. Tuberculosis [BNMC76]. Tumor [LM75b, SM76a]. Tumor-Bearing [SM76a]. Tumors [MS74b, MW71]. Tuna [Bat72, Dra70, ING70, SC73, SMDN76, Wis72]. Tunas [Fuj70, Hig70, Jos70]. Turbidity [Boy79a, DE71, HP76, SM76b]. Turbine [Lis71, ML71, PF72, Van70]. Turbot [LKT79]. Turnover [GTP72]. Turtles [BNMC76]. Tuttle [KM74]. Twelfth [Ano70f]. Twenty [RK76]. Twenty-one [RK76]. Twinning [Ste73]. Two [Ame74, Arm74, AMB^{+}77, BH73, Bey71, Bjo71, BCC78, Cla77, Dea78, Eng77, Far78, GB71, LM71b, LJH76, Mom78, Mor70b, PL72, PS76, ROLH78, RW79, RK76, Sar77, SS75, SHD72, Smi73c, Sta79b, TB71, TB74, Val74, VD79, VS79, Wil78, Wil74, WW70]. Two-Year [Sta79b]. Tylochromis [Gri77]. Type [DCA71, LL75a, LB76]. Types [Deg70, Wil78]. tyrannus [HO78, KG72, LM71b].

U. [SHD72]. ui [Koc73]. Ultrasonic [Byr71, FMD77, GS74, HS75a, JH77, ML72, SHH76]. Ultrastructure [JAR76, MVM77, MLT76]. Ultraviolet [Hof74]. Unanticipated [HGJ75]. Unbalanced [Pri77]. Unchannelized [GS78, MS77]. undecimalis [MS71]. Undergraduate [Roy73b]. Underwater [NM78, Ree71a, WL75]. Underyearling [WBG75]. undulans [FW77]. undulatus [KH78, MMS75, YHP77]. Unfertilized [Boy76]. Unirradiated [Gun73]. United [MH78a, Wil74, GG78, Mil72, SMS78, SL78]. Units [LB76]. Univac [Hog70]. Unmated [MY71]. Unplayed [Pet77]. Untagged [KGJ74]. Unusual [TB71]. upon [LH72]. Upper [Alc76, HD73, HS75c, KS75, CK71, GHH77, HH79a, Man72b, MKK73, NE74, TG74]. Upstream [Erm73]. Uptake [DZ72, DOB74, GPT70, GTP71, Kor73, MRSK70]. Upwelling [WE76]. Urchins [Dav71]. Urinary [FW77]. Urophycis [SHD72]. Use [Ano70g, Bar78, DCA71, Dut76, GGG75, Gue75, HM74a, HL78, HH79b, Hor77, KM72, LM71a, LB76, LYSB78, MP76, Mor70a, Sim72, VR75, Win77, MDS79]. Used [SGC^{+}78, Yeh77, RBC78]. Using [BMS71, BM72, Byr71, ECCC76, HCS76, VS76]. Utah [Ano72c].

Utilization [BBR78, CSG^{+}78, CBP76, Gri74, PW74, Rei72, Sav71, Wag72]. Utilized [LA71].

Valenciennes [BPL75, MMS75]. Validation [Jen70]. Valley [GC72, KM72]. Value [LL75a]. Values [JP76]. var [RGBL79]. Variability [NS76, Shi74, You76]. Variable [SVL73]. Variables [MDS79, SPSF72]. Variants [UF78]. Variation [BTR70, Cla74, Jen72, KM76, Laa71, LPH70, MS70a, PS77, SDKT79, Tar71, UH72, dVKP78]. Variations [EIAJ^{+}76, Jes73, JP76, Man72b, MMN79, ROH79, SP71, ST77]. variegatus [GHC^{+}79, HSF75, LM75b, Mar72b, SHF74]. Various [AM75, LM74, O’G79, SOKEK76, Voy75]. Varying [ROLH78]. Vegetated [CSR78]. Vegetation [LRIF78, MT72, Mit78, Tur77]. velifer [WW76b]. Velocity [FB77, Gib78, Smi73a]. Ventilation [HP76]. Ventilatory [GCD^{+}77]. Verduin [Kem70]. Vermilion [Bor71, Gri78b]. Vermont [Man79]. Versatile [SVL73]. versus [BJ70, LSN70]. Vertebral [MJ72]. Vertical [DLM72]. vetulus [OP73]. via [Cha71]. Vinyl [MECC76]. Viral [SP76]. virens [CE76]. Virgin [SD73]. Virginia [BS75, Bur74, HBH74, KBR71, KB75, Ric73]. Virginian [NRMJ74]. viridans [RGBL79]. virilis [Mom78]. Virus [MLT76, Plu72, PGSP75]. Virus-Caused [Plu72]. Viruses [AP72]. Vista [Ano74f]. Vitamin [JBR71, LL78a, LL78b, MMC78]. vitreum [JKS78, HGO^{+}77, JKS78, Wal75b]. vittatus [Gri75]. vivanus [SD73]. Viviparus [ES72]. Vivo [KDR74, HJB^{+}78]. Voltage [SVL73]. Volumes [Ano74e]. Vrain [EC77]. vulgaris [RP75]. Vulnerability [Cob72, Syl73].

W [Roy73a]. Waikiki [MP75]. Walbaum [Fra73a, LPH70, Lux73, ZT74]. Waldorf [Ano71a]. Waldorf-Astoria [Ano71a]. Walleye [Che73, For74, Gre70, HGO^{+}77, HCS76, JKS78, Kel73, KC72, Nob72, Wal75b, Wal76, Wol77]. Walleyes [BSI76, Par71, Par72, SEC77]. Warmouth [Clo75]. Warmwater [And73, LS73a, Sch74b]. Washing [Sel71]. Washington [Ano70f, Hav73b, OC74, Ant74, Bec72, Del74, DE78, MW71, Mou74, Pat73, SD72a]. Water [AZM75, AB75, BS79, Boy76, BM70, DMB76, FM70, Gib78, GH77, Gre71, GM79, HKB76, Hoa77, Hof74, Kil78, KB75, Kus74a, Lar74, LCL78, LRIF78, LYSB78, MH70, MNL79, NBS76, NB76, PB78, Ric75b, RH75, SC79, Spe73, SC71, Tat73, TL74, TG74, Tsa72, Tsa73, TB77, TBM79, WP76, WE76, WW70, Wor75, Zau70, NF70]. Water-Powered [HKB76]. Waters [Bat72, Boy73, DL71, Gri78b, Hig70, IWR78, KTJ75, LM70, MS71, MP74, Mci70, MB75b, NRMJ74, Pat77, SP71, SDW78, WL75]. Weight [BA75, Bea74, BA70, BA71b, DDA72, EIAJ^{+}76, Eng74, FN71, Gri78b, Leg72, Mci70, MP76, PT70, RM71, Sek74, SNDM77, Sto72b]. Weights [How72, MP76]. Weir [Dom73]. Weiss [FSM75]. Weld [EC77]. West [TB71, BS75, Kel73, KDS79, KB75, SDKT79]. Western [Ott74, AS71b, Dra70, Par72, RH74, SKK79, SD72a, SK71]. Wet [Che75]. Wexford [Wil70]. Whirling [Hof74, O’G79]. White [AR71, Ban73, Bea72, Cam71, CP75, DLM72, EC77, GZ73, HGM78, HB71, KF74, KBR71, Kli74, LMG76, Lux73, MR74a, MMN79, MUR^{+}76, MP78, NK73, Pri70, Pri71, Rei72, RBB75, Rue77, Sie72, SML75, VW74, WH71, Wis74]. Whitefish [Alt71, Alt76, Alt79, ALTA77, Bro75, Pat77, PS70, PW75b, PP73, TD76]. Width [PT70]. Width-Total [PT70]. Wild [BCSS76, KM79, VIS70, WWN76]. williamsoni [PW75b, PP73, TD76]. Wilson [Leg72, Lin76, Tyu74]. Winds [CH75]. Wing [EP73]. Wingra [BK74]. Winnebago [Pri70, Pri71]. Winter [BCP76, CE76, Fra73a, Fra73b, Fra74a, FM70, HC75, HCP76, JH77, KH72, Lev74, LPH70, Lux73, PH77, PR77, SD72a, TC79, UM72, VD79]. Wiscasset [MF75, RM73]. Wisconsin [BK74, Eng76, KC72, MS74a, Mil74, NM74, Pri70, Pri71, Pri75a, VW74, Wis74]. Within [MLK77, Smi76]. Within-Season [MLK77]. Woods [Boy76]. World [Ano74f]. Worms [WJ78]. Write [Hil71b]. Wyoming [BE79, WP73].

X [WW71a]. X-Rays [WW71a]. Xanthochroic [MW72]. xanthurus [HH79c].

Year [Bur74, Che73, EC70b, For71, GH79, Gle78, Has70, HCP76, KDS79, Lux73, NS75, Par71, Par72, Pri75b, RH77, SDKT79, SA77, Sta79b, Ste77, Tar74, TC72, Wal72, WH75, WH71, Wis74, DC79, MH78b]. Year-Class [Has70, HCP76, Lux73, Par72, Wal72]. Yearling [BR76, Pec74]. Yearllng [VW74]. Years [BBR76, MBT70]. Yellow [BTL75, BS77, CH75, Cla77, Eng74, For71, For74, Fra78, HCS76, KF74, LWO76, NS75, Nob73, Nob75, Pri75a, Sie72, Sto72a, Sto72b, ST77, Tar74]. Yellowfin [Lev73, Wis72]. Yellowstone [KK78, Kay77, Lab71, Erm72, ML72]. Yield [Bél74, CNHT76, Fox70, Lar77, TG79, Wis72]. Yield-Per-Recruit [Wis72]. Yields [Mat78, Mel76, Tur77]. York [Ano71a, Bor74, CH75, DH78, For74, Kos78, LM70, MMD73, Nob72]. Young [AR71, Bar78, BSI76, Bur74, Chi72a, Chi72b, Cla77, Dav71, EC70b, Fra73a, Fra73b, Fra74a, GH79, Gle78, Jen71, MJN72, Mes75, Mos70, MH78b, Nob73, Nob75, NRMJ74, Pri75b, RH77, TC72, VW74, WH71, Wis74]. Young-of-the-Year [Bur74, EC70b, Gle78, Pri75b, TC72, WH71, Wis74]. Young-of-year [MH78b].

Zambia [TG79, Dud74, Gri75, Gri77]. Zebrafish [SP76]. Zectran [GC72]. Zinc [BMC72, CS78, Cha78a, Cha78b, HBL79, LL71, SP76, SLD78]. Zoeae [PPSF79]. Zone [BLA78, CA73, DP73, HW77, SKK79]. Zoogeography [GG78]. Zooplankton [CBP76, CM70b, Eng76, Fas71, Gan76, Hut71, Lac73a, Nev75b, Nob75].

References

Anderson:1970:AGT

[AB70] Robert D. Anderson and Charles F. Bryan. Age and growth of three surfperches (Embiotocidae) from Humboldt Bay, California. Transactions of the American Fisheries Society, 99(3):475–482, July 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Arce:1975:EAL

[AB75] Rodolfo G. Arce and Claude E. Boyd. Effects of agricultural limestone on water chemistry, phytoplankton productivity, and fish production in soft water ponds. Transactions of the American Fisheries Society, 104(2):308–312, April 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Abbott:1972:IAP

[Abb72] Robert Riley Abbott. Induced aggregation of pond-reared rainbow trout (Salmo gairdneri) through acoustic conditioning. Transactions of the American Fisheries Society, 101(1):35–43, January 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Abbe:1974:STM

[Abb74] George R. Abbe. Second terminal molt in an adult female blue crab, Callinectes sapidus Rathbun. Transactions of the American Fisheries Society, 103(3):643–644, July 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Adams:1972:RSC

[ABW72] William J. Adams, David J. Behmer, and Warren O. Weingarten. Recovery of shocked common shiner, Notropis cornutus, related to electric energy. Transactions of the American Fisheries Society, 101(3):552–555, July 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anderson:1972:PST

[AC72] Robert D. Anderson and Kenneth K. Chew. Preliminary study on the transient species of fish in Big Beef Harbor. Transactions of the American Fisheries Society, 101(4):726–729, October 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Acheson:1975:FMS

[Ach75] James M. Acheson. Fisheries management and social context: The case of the Maine lobster fishery. Transactions of the American Fisheries Society, 104(4):653–668, October 1975. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Adams:1975:ELH

[Ada75] Evelyn S. Adams. Effects of lead and hydrocarbons from snowmobile exhaust on brook trout (Salvelinus fontinalis). Transactions of the American Fisheries Society, 104(2):363–373, April 1975. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Adams:1976:FEE

[Ada76] S. Marshall Adams. Feeding ecology of eelgrass fish communities. Transactions of the American Fisheries Society, 105(4):514–519, July 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Adelman:1978:ITG

[Ade78] Ira R. Adelman. Influence of temperature on growth promotion and body composition of carp (Cyprinus carpio) due to bovine growth hormone. Transactions of the American Fisheries Society, 107(5):747–750, September 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Al-Daham:1977:DPF

[ADSB77] Najim K. Al-Daham, Abdul Latif Sarker, and Muhammad Nazir Bhatti. Diel patterns of feeding in Heteropneustes fossilis from Southern Iraq. Transactions of the American Fisheries Society, 106(6):614–616, November 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ames:1979:AAC

[AFS79] Lee J. Ames, James D. Felley, and Michael H. Smith. Amounts of asymmetry in centrarchid fish inhabiting heated and nonheated reservoirs. Transactions of the American Fisheries Society, 108 (5):489–495, September 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Atchison:1975:DDB

[AJ75] Gary J. Atchison and Howard E. Johnson. The degradation of DDT in brook trout eggs and fry. Transactions of the American Fisheries Society, 104(4):782–784, October 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Alcorn:1976:TTU

[Alc76] Steven R. Alcorn. Temperature tolerances and upper lethal limits of Salmo apache. Transactions of the American Fisheries Society, 105(2):294–295, March 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Allen:1971:DFS

[All71] J. Frances Allen. Definition of a fisheries scientist. Transactions of the American Fisheries Society, 100(2):379–384, April 1971. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Allen:1972:NSA

[All72] George H. Allen. Notes on sampling amphipods and euphausiids in the North Pacific Ocean. Transactions of the American Fisheries Society, 101(4):577–582, October 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Allen:1973:ARM

[All73] K. Radway Allen. The application of research to the management of British Columbia sockeye salmon. Transactions of the American Fisheries Society, 102(1):172–177, January 1973. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See discussion [Mag73, Str73].

Alt:1971:OHB

[Alt71] Kenneth T. Alt. Occurrence of hybrids between inconnu, Stenodus leucichthys nelma (Pallas), and humpback whitefish, Coregonus pidschian (Linnaeus) in Chatanika River, Alaska. Transactions of the American Fisheries Society, 100(2):362–365, April 1971. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Alt:1976:AGA

[Alt76] Kenneth T. Alt. Age and growth of Alaskan broad whitefish, Coregonus nasus. Transactions of the American Fisheries Society, 105 (4):526–528, July 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Alt:1979:CLH

[Alt79] Kenneth T. Alt. Contributions to the life history of the humpback whitefish in Alaska. Transactions of the American Fisheries Society, 108(2):156–160, March 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Armstrong:1977:AGM

[ALTA77] John W. Armstrong, Charles R. Liston, Peter I. Tack, and Robert C. Anderson. Age, growth, maturity, and seasonal food habits of round whitefish, Prosopium cylindraceum, in Lake Michigan near Ludington, Michigan. Transactions of the American Fisheries Society, 106(2):151–155, March 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Andrews:1975:IVC

[AM75] James W. Andrews and Yoshiaki Matsuda. The influence of various culture conditions on the oxygen consumption of channel catfish. Transactions of the American Fisheries Society, 104(2):322–327, April 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Atchison:1977:TMC

[AMB^{+}77] Gary J. Atchison, Brian R. Murphy, William E. Bishop, Alan W. McIntosh, and Roger A. Mayes. Trace metal contamination of bluegill (Lepomis macrochirus) from two Indiana lakes. Transactions of the American Fisheries Society, 106(6):637–640, November 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Amend:1974:CTT

[Ame74] Donald F. Amend. Comparative toxicity of two iodophors to rainbow trout eggs. Transactions of the American Fisheries Society, 103(1):73–78, January 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Andrews:1973:IDO

[AMG73] James W. Andrews, Takeshi Murai, and George Gibbons. The influence of dissolved oxygen on the growth of channel catfish. Transactions of the American Fisheries Society, 102(4):835–838, October 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Andrews:1970:SCF

[And70] Austin K. Andrews. Squamation chronology of the fathead minnow, Pimephales promelas. Transactions of the American Fisheries Society, 99(2):429–432, April 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Andrews:1972:SMR

[And72] Austin K. Andrews. Survival and mark retention of a small cyprinid marked with fluorescent pigments. Transactions of the American Fisheries Society, 101(1):128–133, January 1972. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anderson:1973:ATR

[And73] Richard O. Anderson. Application of theory and research to management of warmwater fish populations. Transactions of the American Fisheries Society, 102(1):164–171, January 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See discussion [Mag73, Str73].

Anonymous:1970:A

[Ano70a] Anonymous. Addendum. Transactions of the American Fisheries Society, 99(3):613, July 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1970:CAF

[Ano70b] Anonymous. Constitution of the American fisheries society. Transactions of the American Fisheries Society, 99(2):448–458, April 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1970:NNA

[Ano70c] Anonymous. Ninety-ninth Annual Meeting of the American Fisheries Society: Roosevelt Hotel New Orleans, Louisiana, 9–13 September 1969. Transactions of the American Fisheries Society, 99(1):255–290, January 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1970:RPa

[Ano70d] Anonymous. Recent publications. Transactions of the American Fisheries Society, 99(1):253, January 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1970:RPb

[Ano70e] Anonymous. Recent publications. Transactions of the American Fisheries Society, 99(2):447, April 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1970:RAF

[Ano70f] Anonymous. Report of American Fisheries Society Member on the National Research Council on the Twelfth Annual Meeting of the Council, Washington, D. C., March 9–11, 1969. Transactions of the American Fisheries Society, 99(2):459–460, April 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1970:RCU

[Ano70g] Anonymous. Report of the Conference on Use of Exotic Fishes and Related Problems. Transactions of the American Fisheries Society, 99(1):291–295, January 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1970:Ra

[Ano70h] Anonymous. Reviews. Transactions of the American Fisheries Society, 99(1):249–253, January 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1970:Rb

[Ano70i] Anonymous. Reviews. Transactions of the American Fisheries Society, 99(2):445–447, April 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1971:OHA

[Ano71a] Anonymous. One Hundredth Annual Meeting of the American Fisheries Society: The Waldorf-Astoria, New York City, 13–16 September 1970. Transactions of the American Fisheries Society, 100(1):149–205, January 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1971:RPa

[Ano71b] Anonymous. Recent publications. Transactions of the American Fisheries Society, 100(1):147–148, January 1971. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1971:RPb

[Ano71c] Anonymous. Recent publications. Transactions of the American Fisheries Society, 100(3):599–600, July 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1971:Ra

[Ano71d] Anonymous. Reviews. Transactions of the American Fisheries Society, 100(1):141–147, January 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1971:Rb

[Ano71e] Anonymous. Reviews. Transactions of the American Fisheries Society, 100(3):593–599, July 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1972:ARA

[Ano72a] Anonymous. Auditor’s report for the American fisheries society, 1971. Transactions of the American Fisheries Society, 101(2):398– 399, April 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1972:M

[Ano72b] Anonymous. In memoriam. Transactions of the American Fisheries Society, 101(2):397, April 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1972:OHF

[Ano72c] Anonymous. One Hundred-First Annual Meeting of the American Fisheries Society: Hotel Utah, Salt Lake City: 15–18 September 1971. Transactions of the American Fisheries Society, 101(1):145– 217, January 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1972:RPa

[Ano72d] Anonymous. Recent publications. Transactions of the American Fisheries Society, 101(1):143–144, January 1972. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1972:RPb

[Ano72e] Anonymous. Recent publications. Transactions of the American Fisheries Society, 101(2):395–396, April 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1972:RPc

[Ano72f] Anonymous. Recent publications. Transactions of the American Fisheries Society, 101(3):575–576, July 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1972:RPd

[Ano72g] Anonymous. Recent publications. Transactions of the American Fisheries Society, 101(4):763–764, October 1972. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1972:Ra

[Ano72h] Anonymous. Reviews. Transactions of the American Fisheries Society, 101(1):136–143, January 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1972:Rb

[Ano72i] Anonymous. Reviews. Transactions of the American Fisheries Society, 101(2):385–394, April 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1972:Rc

[Ano72j] Anonymous. Reviews. Transactions of the American Fisheries Society, 101(3):568–575, July 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1972:Rd

[Ano72k] Anonymous. Reviews. Transactions of the American Fisheries Society, 101(4):757–763, October 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1973:OHS

[Ano73a] Anonymous. One Hundred-Second Annual Meeting of the American Fisheries Society: Arlington Hotel, Hot Springs, 10–13 September 1972. Transactions of the American Fisheries Society, 102(1): 209–280, January 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1973:RPa

[Ano73b] Anonymous. Recent publications. Transactions of the American Fisheries Society, 102(1):206–207, January 1973. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1973:RPb

[Ano73c] Anonymous. Recent publications. Transactions of the American Fisheries Society, 102(2):508–509, April 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1973:RPc

[Ano73d] Anonymous. Recent publications. Transactions of the American Fisheries Society, 102(3):670–671, July 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1973:RPd

[Ano73e] Anonymous. Recent publications. Transactions of the American Fisheries Society, 102(4):856–858, October 1973. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1973:Ra

[Ano73f] Anonymous. Reviews. Transactions of the American Fisheries Society, 102(1):196–206, January 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1973:Rb

[Ano73g] Anonymous. Reviews. Transactions of the American Fisheries Society, 102(2):500–508, April 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1973:Rc

[Ano73h] Anonymous. Reviews. Transactions of the American Fisheries Society, 102(3):666–670, July 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1973:Rd

[Ano73i] Anonymous. Reviews. Transactions of the American Fisheries Society, 102(4):850–856, October 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1974:BBRd

[Ano74a] Anonymous. Books and bulletins received. Transactions of the American Fisheries Society, 103(4):841–844, October 1974. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1974:BBRc

[Ano74b] Anonymous. Books and bulletins received: (January–March, 1974). Transactions of the American Fisheries Society, 103(3): 658–661, July 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1974:BBRa

[Ano74c] Anonymous. Books and bulletins received: (July–September, 1973). Transactions of the American Fisheries Society, 103(1): 153–154, January 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1974:BBRb

[Ano74d] Anonymous. Books and bulletins received: (October–December, 1973). Transactions of the American Fisheries Society, 103(2): 419–421, April 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1974:CVT

[Ano74e] Anonymous. Corrections to volumes 101 and 102 of Transactions of the American Fisheries Society. Transactions of the American Fisheries Society, 103(2):421, April 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1974:OHT

[Ano74f] Anonymous. One Hundred-Third Annual Meeting of the American Fisheries Society: Contemporary Resort Hotel, Disney World, Lake Buena Vista, Florida, 12–15 September 1973. Transactions of the American Fisheries Society, 103(1):155–208, January 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1974:Ra

[Ano74g] Anonymous. Reviews. Transactions of the American Fisheries Society, 103(1):146–153, January 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1974:Rb

[Ano74h] Anonymous. Reviews. Transactions of the American Fisheries Society, 103(2):411–419, April 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1974:Rc

[Ano74i] Anonymous. Reviews. Transactions of the American Fisheries Society, 103(3):645–658, July 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1974:Rd

[Ano74j] Anonymous. Reviews. Transactions of the American Fisheries Society, 103(4):833–841, October 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1975:BBRd

[Ano75a] Anonymous. Books and bulletins received: (April–June, 1975). Transactions of the American Fisheries Society, 104(4):843–845, October 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1975:BBRc

[Ano75b] Anonymous. Books and bulletins received: (January–March, 1975). Transactions of the American Fisheries Society, 104(3): 649–652, July 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1975:BBRa

[Ano75c] Anonymous. Books and bulletins received: (July–September, 1974). Transactions of the American Fisheries Society, 104(1): 174–178, January 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1975:BBRb

[Ano75d] Anonymous. Books and bulletins received: (October–December, 1974). Transactions of the American Fisheries Society, 104(2): 434–438, April 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1975:Ea

[Ano75e] Anonymous. Erratum. Transactions of the American Fisheries Society, 104(1):178, January 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1975:Eb

[Ano75f] Anonymous. Erratum. Transactions of the American Fisheries Society, 104(2):438, April 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1975:Ra

[Ano75g] Anonymous. Reviews. Transactions of the American Fisheries Society, 104(1):161–174, January 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1975:Rb

[Ano75h] Anonymous. Reviews. Transactions of the American Fisheries Society, 104(2):415–434, April 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1975:Rc

[Ano75i] Anonymous. Reviews. Transactions of the American Fisheries Society, 104(3):635–649, July 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1975:Rd

[Ano75j] Anonymous. Reviews. Transactions of the American Fisheries Society, 104(4):824–843, October 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1976:E

[Ano76a] Anonymous. Errata. Transactions of the American Fisheries Society, 105(2):352, March 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1976:Ra

[Ano76b] Anonymous. Reviews. Transactions of the American Fisheries Society, 105(1):181–189, January 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1976:Rb

[Ano76c] Anonymous. Reviews. Transactions of the American Fisheries Society, 105(3):494–502, May 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1976:Rc

[Ano76d] Anonymous. Reviews. Transactions of the American Fisheries Society, 105(4):567–573, July 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1976:Rd

[Ano76e] Anonymous. Reviews. Transactions of the American Fisheries Society, 105(5):643–646, September 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1976:Re

[Ano76f] Anonymous. Reviews. Transactions of the American Fisheries Society, 105(6):737–738, November 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1977:Ra

[Ano77a] Anonymous. Reviews. Transactions of the American Fisheries Society, 106(1):115–119, January 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1977:Rb

[Ano77b] Anonymous. Reviews. Transactions of the American Fisheries Society, 106(2):196–199, March 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1977:Rc

[Ano77c] Anonymous. Reviews. Transactions of the American Fisheries Society, 106(3):291–294, May 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1977:Rd

[Ano77d] Anonymous. Reviews. Transactions of the American Fisheries Society, 106(4):404–410, July 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1977:Re

[Ano77e] Anonymous. Reviews. Transactions of the American Fisheries Society, 106(5):500–504, September 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1977:Rf

[Ano77f] Anonymous. Reviews. Transactions of the American Fisheries Society, 106(6):649–656, November 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1978:Ra

[Ano78a] Anonymous. Reviews. Transactions of the American Fisheries Society, 107(2):382–384, March 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1978:Rb

[Ano78b] Anonymous. Reviews. Transactions of the American Fisheries Society, 107(4):642–650, July 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1978:Rc

[Ano78c] Anonymous. Reviews. Transactions of the American Fisheries Society, 107(5):757–762, September 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1978:Rd

[Ano78d] Anonymous. Reviews. Transactions of the American Fisheries Society, 107(6):872–876, November 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1979:BRa

[Ano79a] Anonymous. Book reviews. Transactions of the American Fisheries Society, 108(1):110–113, January 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1979:BRb

[Ano79b] Anonymous. Book reviews. Transactions of the American Fisheries Society, 108(2):208–213, March 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1979:BRc

[Ano79c] Anonymous. Book reviews. Transactions of the American Fisheries Society, 108(3):332–334, May 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1979:BRd

[Ano79d] Anonymous. Book reviews. Transactions of the American Fisheries Society, 108(4):422–426, July 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1979:BRe

[Ano79e] Anonymous. Book reviews. Transactions of the American Fisheries Society, 108(5):502–504, September 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1979:BRf

[Ano79f] Anonymous. Book reviews. Transactions of the American Fisheries Society, 108(6):653–659, November 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Antipa:1974:FHL

[Ant74] Ross Antipa. Food habits of Lacustrine salmonids in Washington State in relation to infections with larvae of the bass tapeworm (Proteocephalus ambloplitis). Transactions of the American Fisheries Society, 103(4):811–814, October 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Amend:1972:IMI

[AP72] Donald F. Amend and John P. Pietsch. An improved method for isolating viruses from asymptomatic carrier fish. Transactions of the American Fisheries Society, 101(2):267–269, April 1972. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Andrews:1975:EFF

[AP75] James W. Andrews and Jimmy W. Page. The effects of frequency of feeding on culture of catfish. Transactions of the American Fisheries Society, 104(2):317–321, April 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Associates:1971:FHF

[AR71] Dilip Mathur Ichthyological Associates and Timothy W. Robbins. Food habits and feeding chronology of young white crappie, Pomoxis annularis Rafinesque, in Conowingo Reservoir. Transactions of the American Fisheries Society, 100(2):307–311, April 1971. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Armstrong:1971:AFM

[Arm71] Robert H. Armstrong. Age, food, and migration of sea-run cutthroat trout, Salmo clarki, at Eva Lake, Southeastern Alaska. Transactions of the American Fisheries Society, 100(2):302–306, April 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Armour:1974:SAD

[Arm74] Carl L. Armour. Serum antigenic differences in spring Chinook salmon (Oncorhynchus tshawytscha) from two separate rivers. Transactions of the American Fisheries Society, 103(2):379–382, April 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Adelman:1970:EHS

[AS70] Ira R. Adelman and Lloyd L. Smith, Jr. Effect of hydrogen sulfide on northern pike eggs and sac fry. Transactions of the American Fisheries Society, 99(3):501–509, July 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Allen:1971:RAJ

[AS71a] Kenneth O. Allen and Kirk Strawn. Rate of acclimation of juvenile channel catfish, Ictalurus punctatus, to high temperatures. Transactions of the American Fisheries Society, 100(4):665–671, October 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anderson:1971:FAA

[AS71b] Emory D. Anderson and Lloyd L. Smith, Jr. Factors affecting abundance of lake herring (Coregonus artedii Lesueur) in Western Lake Superior. Transactions of the American Fisheries Society, 100(4):691–707, October 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Andrews:1972:IFR

[AS72] James W. Andrews and Robert R. Stickney. Interactions of feeding rates and environmental temperature on growth, food conversion, and body composition of channel catfish. Transactions of the American Fisheries Society, 101(1):94–99, January 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Allen:1978:RSP

[AS78] Standish K. Allen, Jr. and Jon G. Stanley. Reproductive sterility in polyploid brook trout, Salvelinus fontinalis. Transactions of the American Fisheries Society, 107(3):473–478, May 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Allen:1979:PMI

[AS79] Standish K. Allen, Jr. and Jon G. Stanley. Polyploid mosaics induced by cytochalasin B in landlocked Atlantic salmon Salmo salar. Transactions of the American Fisheries Society, 108(5):462– 466, September 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Austin:1975:ETD

[ASH75] Herbert M. Austin, Allan D. Sosnow, and Clarence R. Hickey, Jr. The effects of temperature on the development and survival of the eggs and larvae of the Atlantic silverside, Menidia menidia. Transactions of the American Fisheries Society, 104(4):762–765, October 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Adams:1975:ISW

[AZM75] B. L. Adams, W. S. Zaugg, and L. R. McLain. Inhibition of salt water survival and Na-k-ATPase elevation in steelhead trout (Salmo gairdneri) by moderate water temperatures. Transactions of the American Fisheries Society, 104(4):766–769, October 1975. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bellinger:1970:SOG

[BA70] John W. Bellinger and James W. Avault, Jr. Seasonal occurrence, growth, and length–weight relationship of juvenile pompano, Trachinotus carolinus, in Louisiana. Transactions of the American Fisheries Society, 99(2):353–358, April 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bellinger:1971:FHJ

[BA71a] John W. Bellinger and James W. Avault, Jr. Food habits of juvenile pompano, Trachinotus carolinus, in Louisiana. Transactions of the American Fisheries Society, 100(3):486–494, July 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boothby:1971:FHL

[BA71b] Rea N. Boothby and James W. Avault, Jr. Food habits, length– weight relationship, and condition factor of the red drum (Sciaenops ocellata) in Southeastern Louisiana. Transactions of the American Fisheries Society, 100(2):290–295, April 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Berude:1972:RFP

[BA72] Catherine L. Berude and Norman J. Abramson. Relative fishing power, CDC 6600, FORTRAN IV. Transactions of the American Fisheries Society, 101(1):133, January 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bass:1975:FHL

[BA75] Robert J. Bass and James W. Avault, Jr. Food habits, length– weight relationship, condition factor, and growth of juvenile red drum, Sciaenops ocellata, in Louisiana. Transactions of the American Fisheries Society, 104(1):35–45, January 1975. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Babcock:1971:ESL

[Bab71] W. H. Babcock. Effect of a size limit regulation on the trout fishery in Trappers Lake, Colorado. Transactions of the American Fisheries Society, 100(1):50–54, January 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bailey:1972:AGR

[Bai72] Merryll M. Bailey. Age, growth, reproduction, and food of the burbot, Lota lota (Linnaeus), in Southwestern Lake Superior. Transactions of the American Fisheries Society, 101(4):667–674, October 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bailey:1978:CFP

[Bai78] William M. Bailey. A comparison of fish populations before and after extensive grass carp stocking. Transactions of the American Fisheries Society, 107(1):181–206, January 1978. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Banerjee:1973:ARG

[Ban73] Tapan Banerjee. Age and rate of growth of the white seaperch (Phanerodon furcatus) from Tomales Bay, California. Transactions of the American Fisheries Society, 102(2):448–450, April 1973. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Barr:1971:MEA

[Bar71] Louis Barr. Methods of estimating the abundance of juvenile spot shrimp in a shallow nursery area. Transactions of the American Fisheries Society, 100(4):781–787, October 1971. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Barkman:1978:UOG

[Bar78] Robert C. Barkman. The use of otolith growth rings to age young Atlantic silversides, Menidia menidia. Transactions of the American Fisheries Society, 107(6):790–792, November 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Batts:1972:SMF

[Bat72] Billy S. Batts. Sexual maturity, fecundity, and sex ratios of the skipjack tuna, Katsuwonus pelamis (Linnaeus), in North Carolina waters. Transactions of the American Fisheries Society, 101(4): 626–637, October 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Branson:1975:BTR

[BBAN75] D. R. Branson, G. E. Blau, H. C. Alexander, and W. B. Neely. Bioconcentration of 2,2’,4,4’-tetrachlorobiphenyl in rainbow trout as measured by an accelerated test. Transactions of the American Fisheries Society, 104(4):785–792, October 1975. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baur:1976:TYE

[BBR76] Richard J. Baur, D. Homer Buck, and C. Russell Rose. Three years of experience with a channel catfish catchout pond in Illinois. Transactions of the American Fisheries Society, 105(2):247–253, March 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Buck:1978:USM

[BBR78] D. Homer Buck, Richard J. Baur, and C. Russell Rose. Utilization of swine manure in a polyculture of Asian and North American fishes. Transactions of the American Fisheries Society, 107(1): 216–222, January 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baur:1979:PAL

[BBR79] Richard J. Baur, D. Homer Buck, and C. R. Rose. Production of age-0 largemouth bass, smallmouth bass, and bluegills in ponds stocked with grass carp. Transactions of the American Fisheries Society, 108(5):496–498, September 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bulow:1978:CTB

[BCC78] Frank J. Bulow, C. B. Coburn, Jr., and Charles S. Cobb. Comparisons of two bluegill populations by means of the RNA-DNA ratio and liver-somatic index. Transactions of the American Fisheries Society, 107(6):799–803, November 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bridges:1976:SHC

[BCP76] D. W. Bridges, J. J. Cech, Jr., and D. N. Pedro. Seasonal hematological changes in winter flounder, Pseudopleuronectes americanus. Transactions of the American Fisheries Society, 105(5):596–600, September 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brenner:1976:EFH

[BCS76] Fred J. Brenner, Scott Corbett, and Richard Shertzer. Effect of ferric hydroxide suspension on blood chemistry in the common shiner, notropus cornutus. Transactions of the American Fisheries Society, 105(3):450–455, May 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bouck:1976:OGB

[BCSS76] Gerald R. Bouck, Gary A. Chapman, Phillip W. Schneider, Jr., and Donald G. Stevens. Observations on gas bubble disease among wild adult Columbia River fishes. Transactions of the American Fisheries Society, 105(1):114–115, January 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bonham:1972:SRR

[BD72] Kelshaw Bonham and Lauren R. Donaldson. Sex ratios and retardation of gonadal development in chronically gamma-irradiated Chinook salmon smolts. Transactions of the American Fisheries Society, 101(3):428–434, July 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beiningen:1970:EJD

[BE70] Kirk T. Beiningen and Wesley J. Ebel. Effect of John Day Dam on dissolved nitrogen concentrations and salmon in the Columbia River, 1968. Transactions of the American Fisheries Society, 99 (4):664–671, October 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Binns:1979:QFT

[BE79] N. Allen Binns and Fred M. Eiserman. Quantification of fluvial trout habitat in Wyoming. Transactions of the American Fisheries Society, 108(3):215–228, May 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beardsley:1970:DMJ

[Bea70] Grant L. Beardsley, Jr. Distribution of migrating juvenile pink shrimp, Penaeus duorarum duorarum Burkenroad, in Buttonwood Canal, Everglades National Park, Florida. Transactions of the American Fisheries Society, 99(2):401–408, April 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beamish:1972:LPW

[Bea72] Richard J. Beamish. Lethal pH for the white sucker Catostomus commersoni (Lacépède). Transactions of the American Fisheries Society, 101(2):355–358, April 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beamish:1974:SPA

[Bea74] F. W. H. Beamish. Swimming performance of adult sea lamprey, Petromyzon marinus, in relation to weight and temperature. Transactions of the American Fisheries Society, 103(2):355–358, April 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Becker:1972:IRM

[Bec72] C. Dale Becker. Initial records of Monogenea from juvenile bass in the Central Columbia River, Washington. Transactions of the American Fisheries Society, 101(4):723–726, October 1972. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Behnke:1970:ACB

[Beh70] Robert J. Behnke. The application of cytogenetic and biochemical systematics to phylogenetic problems in the family Salmonidae. Transactions of the American Fisheries Society, 99(1):237–248, January 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beland:1974:PYB

[Bél74] Pierre Béland. On predicting the yield from brook trout populations. Transactions of the American Fisheries Society, 103(2):

353–355, April 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Benoit:1974:ALS

[Ben74] Duane A. Benoit. Artificial laboratory spawning substrate for brook trout (Salvelinus fontinalis Mitchill). Transactions of the American Fisheries Society, 103(1):144–145, January 1974. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Benoit:1975:CEC

[Ben75] D. A. Benoit. Chronic effects of copper on survival, growth, and reproduction of the bluegill (Lepomis macrochirus). Transactions of the American Fisheries Society, 104(2):353–358, April 1975. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Berg:1978:GMC

[Ber78] Richard E. Berg. Growth and maturation of Chinook salmon, Oncorhynchus tshawytscha, introduced into Lake Superior. Transactions of the American Fisheries Society, 107(2):281–283, March 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beyerle:1971:STN

[Bey71] George B. Beyerle. A study of two northern pike-bluegill populations. Transactions of the American Fisheries Society, 100(1): 69–73, January 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baker:1976:MSG

[BF76] Anne M. Baker and Douglas F. Fraser. A method for securing the gut contents of small, live fish. Transactions of the American Fisheries Society, 105(4):520–522, July 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Berra:1970:RED

[BG70] Tim M. Berra and Gerald E. Gunning. Repopulation of experimentally decimated sections of streams by longear sunfish, Lepomis megalotis megalotis (Rafinesque). Transactions of the American Fisheries Society, 99(4):776–781, October 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bennett:1972:FLB

[BG72] David H. Bennett and J. Whitfield Gibbons. Food of largemouth bass (Micropterus salmoides) from a South Carolina reservoir receiving heated effluent. Transactions of the American Fisheries Society, 101(4):650–654, October 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bennett:1975:RCL

[BG75] David H. Bennett and J. Whitfield Gibbons. Reproductive cycles of largemouth bass (Micropterus salmoides) in a cooling reservoir. Transactions of the American Fisheries Society, 104(1):77–82, January 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Becker:1975:RNC

[BGM75] C. Dale Becker, Robert G. Genoway, and J. A. Merrill. Resistance of a northwestern crayfish, Pacifastacus leniusculus (Dana), to elevated temperatures. Transactions of the American Fisheries Society, 104(2):374–387, April 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Becker:1977:CCR

[BGS77] C. D. Becker, R. G. Genoway, and M. J. Schneider. Comparative cold resistance of three Columbia River organisms. Transactions of the American Fisheries Society, 106(2):178–184, March 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Banner:1973:ENE

[BH73] Arnold Banner and Martin Hyatt. Effects of noise on eggs and larvae of two estuarine fishes. Transactions of the American Fisheries Society, 102(1):134–136, January 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Burnett-Herkes:1974:PGB

[BH74] James Burnett-Herkes. Parasites of the gills and buccal cavity of the dolphin, Coryphaena hippurus, from the Straits of Florida. Transactions of the American Fisheries Society, 103(1):101–106, January 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Benson:1975:ERF

[BH75] Norman G. Benson and Patrick L. Hudson. Effects of a reduced fall drawdown on benthos abundance in Lake Francis case. Transactions of the American Fisheries Society, 104(3):526–528, July 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Barwick:1976:RLB

[BH76a] D. Hugh Barwick and Dennis E. Holcomb. Relation of largemouth bass reproduction to crowded sunfish populations in Florida ponds. Transactions of the American Fisheries Society, 105(2):244–246, March 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Buttner:1976:ECC

[BH76b] Joseph K. Buttner and Robert W. Hamilton. Ergasilus (Copepoda: Cyclopoida) infestation of Coho and Chinook salmon in Lake Michigan. Transactions of the American Fisheries Society, 105(3):491–493, May 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:1970:SSS

[BIJ70] B. E. Brown, I. Inman, and A. Jearld, Jr. Schooling and shelter seeking tendencies in fingerling channel catfish. Transactions of the American Fisheries Society, 99(3):540–545, July 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bender:1970:PEV

[BJ70] Michael E. Bender and Robert A. Jordan. Plastic enclosure versus open lake productivity measurements. Transactions of the American Fisheries Society, 99(3):607–610, July 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bouck:1976:ESS

[BJ76] Gerald R. Bouck and Janet Jacobson. Estimation of salmonid sperm concentration by microhematocrit technique. Transactions of the American Fisheries Society, 105(4):534–535, July 1976. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bouck:1979:MIT

[BJ79] Gerald R. Bouck and David A. Johnson. Medication inhibits tolerance to seawater in Coho salmon smolts. Transactions of the American Fisheries Society, 108(1):63–66, January 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bjornn:1971:TSM

[Bjo71] T. C. Bjornn. Trout and salmon movements in two Idaho streams as related to temperature, food, stream flow, cover, and population density. Transactions of the American Fisheries Society, 100(3): 423–438, July 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baumann:1974:DPD

[BK74] Paul C. Baumann and James F. Kitchell. Diel patterns of distribution and feeding of bluegill (Lepomis macrochirus) in Lake Wingra, Wisconsin. Transactions of the American Fisheries Society, 103(2): 255–260, April 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Burkhalter:1977:EPE

[BK77] Dalton E. Burkhalter and Calvin M. Kaya. Effects of prolonged exposure to ammonia on fertilized eggs and sac fry of rainbow trout (Salmo gairdneri). Transactions of the American Fisheries Society, 106(5):470–475, September 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bulkley:1976:SRF

[BKS76] R. V. Bulkley, R. L. Kellogg, and L. R. Shannon. Size-related factors associated with Dieldrin concentrations in muscle tissue of channel catfish Ictalurus punctatus. Transactions of the American Fisheries Society, 105(2):301–307, March 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brazo:1978:LHL

[BLA78] Dan C. Brazo, Charles R. Liston, and Robert C. Anderson. Life history of the longnose dace, Rhinichthys cataractae, in the surge zone of Eastern Lake Michigan near Ludington, Michigan. Transactions of the American Fisheries Society, 107(4):550–556, July 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bryson:1975:RAF

[BLCD75] William T. Bryson, Robert T. Lackey, John Cairns, Jr., and Kenneth L. Dickson. Restocking after fishkills as a fisheries management strategy. Transactions of the American Fisheries Society, 104(2):256–263, April 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Benoit:1976:TEC

[BLCF76] Duane A. Benoit, Edward N. Leonard, Glenn M. Christensen, and James T. Fiandt. Toxic effects of cadmium on three generations of brook trout (Salvelinus fontinalis). Transactions of the American Fisheries Society, 105(4):550–560, July 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brungs:1970:WDS

[BM70] William A. Brungs and Donald I. Mount. A water delivery system for small fish-holding tanks. Transactions of the American Fisheries Society, 99(4):799–802, October 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Barber:1971:SFC

[BM71] Willard E. Barber and W. L. Minckley. Summer foods of the cyprinid fish Semotilus atromaculatus. Transactions of the American Fisheries Society, 100(2):283–289, April 1971. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bearden:1972:RPS

[BM72] Charles M. Bearden and Michael D. Mckenzie. Results of a pilot shrimp tagging project using internal anchor tags. Transactions of the American Fisheries Society, 101(2):358–362, April 1972. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brusven:1974:ESI

[BM74] M. A. Brusven and C. Macphee. An evaluation of squoxin on insect drift. Transactions of the American Fisheries Society, 103(2):362– 365, April 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baker:1977:SPB

[BM77a] John A. Baker and Timothy Modde. Susceptibility to predation of blacktail shiners stained with Bismarck Brown Y. Transactions of the American Fisheries Society, 106(4):334–338, July 1977. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Burr:1977:SBS

[BM77b] Brooks M. Burr and Michael A. Morris. Spawning behavior of the shorthead redhorse, Moxostoma macrolepidotum, in Big Rock Creek, Illinois. Transactions of the American Fisheries Society, 106(1):80–82, January 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Buynak:1978:LDN

[BM78a] Gerard L. Buynak and Harold W. Mohr, Jr. Larval development of the northern hog sucker (Hypentelium nigricans), from the Susquehanna River. Transactions of the American Fisheries Society, 107 (4):595–599, July 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Buynak:1978:LDR

[BM78b] Gerard L. Buynak and Harold W. Mohr, Jr. Larval development of the redbreast sunfish (Lepomis auritus) from the Susquehanna River. Transactions of the American Fisheries Society, 107(4): 600–604, July 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beitinger:1979:GRT

[BM79a] Thomas L. Beitinger and John J. Magnuson. Growth rates and temperature selection of bluegill, Lepomis macrochirus. Transactions of the American Fisheries Society, 108(4):378–382, July 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Butler:1979:SSC

[BM79b] Susan Ann Butler and Lyman L. McDonald. A simulation study of a catch–effort model for estimating mortality rates. Transactions of the American Fisheries Society, 108(4):353–357, July 1979. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Buynak:1979:LDS

[BM79c] Gerard L. Buynak and Harold W. Mohr, Jr. Larval development of the shorthead redhorse (Moxostoma macrolepidotum) from the Susquehanna River. Transactions of the American Fisheries Society, 108(2):161–165, March 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Burton:1972:MCB

[BMC72] Dennis T. Burton, Eric L. Morgan, and John Cairns, Jr. Mortality curves of bluegills (Lepomis macrochirus Rafinesque) simultaneously exposed to temperature and zinc stress. Transactions of the American Fisheries Society, 101(3):435–441, July 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baker:1971:STI

[BMS71] Claude D. Baker, Duncan W. Martin, and Eugene H. Schmitz. Separation of taxonomically identifiable organisms and detritus taken from shad foregut contents using density-gradient centrifugation. Transactions of the American Fisheries Society, 100(1):138–139, January 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bonneau:1972:FHG

[BMTD72] D. L. Bonneau, J. W. Mcguire, O. W. Tiemeier, and C. W. Deyoe. Food habits and growth of channel catfish fry, Ictalurus punctatus. Transactions of the American Fisheries Society, 101(4):613–619, October 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bahner:1975:SCF

[BN75] Lowell H. Bahner and Delwayne R. Nimmo. A salinity controller for flow-through bioassays. Transactions of the American Fisheries Society, 104(2):388–389, April 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brock:1976:TPG

[BNMC76] J. A. Brock, R. M. Nakamura, A. Y. Miyahara, and E. M. L. Chang. Tuberculosis in Pacific green sea turtles, Chelonia mydas. Transactions of the American Fisheries Society, 105(4):564–566, July 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baker:1970:SDH

[BNS70] Claude D. Baker, William H. Neill, Jr., and Kirk Strawn. Sexual difference in heat resistance of the Ozark minnow, Dionda nubila (Forbes). Transactions of the American Fisheries Society, 99(3): 588–591, July 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bonham:1976:SGS

[Bon76] Kelshaw Bonham. Specific gravity of salmonid eggs. Transactions of the American Fisheries Society, 105(2):351–352, March 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bortone:1971:PVS

[Bor71] Stephen A. Bortone. Pugheadedness in the vermilion snapper, Rhomboplites aurorubens, in the Northern Gulf of Mexico. Transactions of the American Fisheries Society, 100(2):366–368, April 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boreman:1974:ESI

[Bor74] John Boreman. Effects of stream improvement on juvenile rainbow trout in Cayuga Inlet, New York. Transactions of the American Fisheries Society, 103(3):637–641, July 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bouck:1972:EDH

[Bou72] Gerald R. Bouck. Effects of diurnal hypoxia on electrophoretic protein fractions and other health parameters of rock bass (Ambloplites rupestris). Transactions of the American Fisheries Society, 101(3):488–493, July 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bouck:1979:HLA

[Bou79] Gerald R. Bouck. Histochemistry of leucine aminonaphthylamidase (LAN) in rainbow trout (Salmo gairdneri). Transactions of the American Fisheries Society, 108(1):57–62, January 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bowman:1970:LHB

[Bow70] Milton L. Bowman. Life history of the black redhorse, Moxostoma duquesnei (Lesueur), in Missouri. Transactions of the American Fisheries Society, 99(3):546–559, July 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bowler:1975:FIG

[Bow75] Bert Bowler. Factors influencing genetic control in lakeward migrations of cutthroat trout fry. Transactions of the American Fisheries Society, 104(3):474–482, July 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bowen:1978:CAA

[Bow78] Stephen H. Bowen. Chromic acid in assimilation studies — a caution. Transactions of the American Fisheries Society, 107(5):755– 756, September 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boyd:1970:NCO

[Boy70] Claude E. Boyd. Nutrient content of offal and small fish from some freshwater fish cultures. Transactions of the American Fisheries Society, 99(4):809–811, October 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boyd:1973:COD

[Boy73] Claude E. Boyd. The chemical oxygen demand of waters and biological materials from ponds. Transactions of the American Fisheries Society, 102(3):606–611, July 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boyd:1976:WCP

[Boy76] Claude E. Boyd. Water chemistry and plankton in unfertilized ponds in pastures and in Woods. Transactions of the American Fisheries Society, 105(5):634–636, September 1976. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boyd:1979:ASA

[Boy79a] Claude E. Boyd. Aluminum sulfate (alum) for precipitating clay turbidity from fish ponds. Transactions of the American Fisheries Society, 108(3):307–313, May 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boyd:1979:DTA

[Boy79b] Claude E. Boyd. Determination of total ammonia nitrogen and chemical oxygen demand in fish culture systems. Transactions of the American Fisheries Society, 108(3):314–319, May 1979. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bridges:1975:OHS

[BPL75] David W. Bridges, Drusilla R. Pedro, and Marshall Laird. The occurrence of Haemogregarina sp. in the goosefish (Lophius americanus Valenciennes) from the Gulf of Maine. Transactions of the American Fisheries Society, 104(2):284–285, April 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Balsano:1975:MAE

[BR75] J. S. Balsano and E. M. Rasch. Modification of acrylamide electrophoresis for microspectrophotometric analysis of fish plasma proteins. Transactions of the American Fisheries Society, 104(1): 150–157, January 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bentley:1976:DMY

[BR76] Wallace W. Bentley and Howard L. Raymond. Delayed migrations of yearling Chinook salmon since completion of Lower Monumental and Little Goose Dams on the Snake River. Transactions of the American Fisheries Society, 105(3):422–424, May 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brigham:1974:JCN

[Bri74] Warren U. Brigham. Journal coverage by North American fishery biologists. Transactions of the American Fisheries Society, 103(2): 387–389, April 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boyd:1978:PEM

[BRJ78] Claude E. Boyd, Robert P. Romaire, and Ellen Johnston. Predicting early morning dissolved oxygen concentrations in channel catfish ponds. Transactions of the American Fisheries Society, 107 (3):484–492, May 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:1970:EDP

[Bro70] Bradford E. Brown. Exponential decrease in a population of fathead minnows. Transactions of the American Fisheries Society, 99(4):807–809, October 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brooke:1975:EDC

[Bro75] L. T. Brooke. Effect of different constant incubation temperatures on egg survival and embryonic development in lake whitefish (Coregonus clupeaformis). Transactions of the American Fisheries Society, 104(3):555–559, July 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brook:1977:TRS

[Bro77] Iver M. Brook. Trophic relationships in a seagrass community (Thalassia testudinum), in Card Sound, Florida. fish diets in relation to macrobenthic and cryptic faunal abundance. Transactions of the American Fisheries Society, 106(3):219–229, May 1977. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brungs:1971:CEC

[Bru71] William A. Brungs. Chronic effects of constant elevated temperature on the fathead minnow (Pimephales promelas Rafinesque). Transactions of the American Fisheries Society, 100(4):659–664, October 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bulkley:1973:EAS

[BS73] Ross V. Bulkley and Gary L. Swihart. Effects of the anabolic steroid stanozolol on growth of channel catfish, Ictalurus punctatus, and goldfish, Carassius auratus. Transactions of the American Fisheries Society, 102(2):444–446, April 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brandt:1975:EHA

[BS75] Thomas M. Brandt and Carl B. Schreck. Effects of harvesting aquatic bait species from a small West Virginia stream. Transactions of the American Fisheries Society, 104(3):446–453, July 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brooks:1977:EIC

[BS77] Arthur S. Brooks and Gregory L. Seegert. The effects of intermittent chlorination on rainbow trout and yellow perch. Transactions of the American Fisheries Society, 106(3):278–286, May 1977. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boyd:1978:NFP

[BS78] Claude E. Boyd and John W. Sowles. Nitrogen fertilization of ponds. Transactions of the American Fisheries Society, 107(5): 737–741, September 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bouck:1979:MED

[BS79] Gerald R. Bouck and Stanley D. Smith. Mortality of experimentally descaled smolts of Coho salmon (Oncorhynchus kisutch) in fresh and salt water. Transactions of the American Fisheries Society, 108(1):67–69, January 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bulkley:1976:FPY

[BSI76] Ross V. Bulkley, Vernon L. Spykermann, and Lloyd E. Inmon. Food of the pelagic young of walleyes and five cohabiting fish species in Clear Lake, Iowa. Transactions of the American Fisheries Society, 105(1):77–83, January 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boyd:1979:EAF

[BT79] Claude E. Boyd and Craig S. Tucker. Emergency aeration of fish ponds. Transactions of the American Fisheries Society, 108(3): 299–306, May 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brazo:1975:AGF

[BTL75] Dan C. Brazo, Peter I. Tack, and Charles R. Liston. Age, growth, and fecundity of yellow perch, Perca flavescens, in Lake Michigan near Ludington, Michigan. Transactions of the American Fisheries Society, 104(4):726–730, October 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Buck:1970:VCP

[BTR70] D. Homer Buck, Charles F. Thoits III, and C. Russell Rose. Variation in carp production in replicate ponds. Transactions of the American Fisheries Society, 99(1):74–79, January 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bulkley:1970:FIB

[Bul70] Ross V. Bulkley. Feeding interaction between adult bass and their offspring. Transactions of the American Fisheries Society, 99(4): 732–738, October 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bulkley:1972:DCF

[Bul72] Ross V. Bulkley. Diethylstilbestrol in catfish feed. Transactions of the American Fisheries Society, 101(3):537–539, July 1972. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Burns:1972:DLH

[Bur72a] James W. Burns. The distribution and life history of South American freshwater crabs (Aegla) and their role in trout streams and lakes. Transactions of the American Fisheries Society, 101(4):595– 607, October 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Burns:1972:SEL

[Bur72b] James W. Burns. Some effects of logging and associated road construction on Northern California streams. Transactions of the American Fisheries Society, 101(1):1–17, January 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Burbidge:1974:DGS

[Bur74] Richard G. Burbidge. Distribution, growth, selective feeding, and energy transformations of young-of-the-year blueback herring, Alosa aestivalis (Mitchill), in the James River, Virginia. Transactions of the American Fisheries Society, 103(2):297–311, April 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bulow:1979:OCP

[BWH79] Frank J. Bulow, James R. Winningham, and Ronald C. Hooper. Occurrence of the copepod parasite Lernaea cyprinacaea in a stream fish population. Transactions of the American Fisheries Society, 108(1):100–102, January 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Byrne:1971:FCU

[Byr71] John E. Byrne. A further contribution to using ultrasonic sensors for fish activity studies. Transactions of the American Fisheries Society, 100(4):792–794, October 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carr:1973:FHJ

[CA73] William E. S. Carr and Clayton A. Adams. Food habits of juvenile marine fishes occupying seagrass beds in the estuarine zone near Crystal River, Florida. Transactions of the American Fisheries Society, 102(3):511–540, July 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Crawford:1977:SIN

[CA77] Richard E. Crawford and George H. Allen. Seawater inhibition of nitrite toxicity to Chinook salmon. Transactions of the American Fisheries Society, 106(1):105–109, January 1977. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Caillouet:1971:LAC

[Cai71] Charles W. Caillouet, Jr. Lactate acidosis as a cause of mortality in captured sharks: an hypothesis. Transactions of the American Fisheries Society, 100(1):139–140, January 1971. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Campbell:1971:ILD

[Cam71] K. Perry Campbell. Influence of light and dark periods on spatial distribution and activity of the white sucker, Catostomus commersoni. Transactions of the American Fisheries Society, 100(2): 353–355, April 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carline:1972:BHE

[Car72] Robert F. Carline. Biased harvest estimates from a postal survey of a sport fishery. Transactions of the American Fisheries Society, 101(2):262–266, April 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carlson:1973:ISL

[Car73] Anthony R. Carlson. Induced spawning of largemouth bass [Micropterus salmoides Lacépède]. Transactions of the American Fisheries Society, 102(2):442–444, April 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Casselman:1974:ESD

[Cas74] John M. Casselman. External sex determination of northern pike, Esox lucius Linnaeus. Transactions of the American Fisheries Society, 103(2):343–347, April 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Casselman:1975:SRN

[Cas75] John M. Casselman. Sex ratios of northern pike, Esox lucius Linnaeus. Transactions of the American Fisheries Society, 104(1): 60–63, January 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Crowley:1972:EMB

[CB72] G. J. Crowley and D. J. Berinati. Effect of MS-222 on blood sugar and liver glycogen in rainbow trout. Transactions of the American Fisheries Society, 101(1):125–128, January 1972. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Craddock:1976:OUZ

[CBP76] Donovan R. Craddock, Theodore H. Blahm, and William D. Parente. Occurrence and utilization of zooplankton by juvenile Chinook salmon in the Lower Columbia River. Transactions of the American Fisheries Society, 105(1):72–76, January 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cairns:1978:EHS

[CC78] Michael A. Cairns and Alan R. Christian. Effects of hemorrhagic stress on several blood parameters in adult rainbow trout (Salmo gairdneri). Transactions of the American Fisheries Society, 107(2): 334–340, March 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cech:1979:RRL

[CCM79] Joseph J. Cech, Jr., Clain G. Campagna, and Stephen J. Mitchell. Respiratory responses of largemouth bass (Micropterus salmoides) to environmental changes in temperature and dissolved oxygen. Transactions of the American Fisheries Society, 108(2):166–171, March 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cooper:1975:SAA

[CCN75] Richard A. Cooper, Roger A. Clifford, and Clifford D. Newell. Seasonal abundance of the American lobster, Homarus americanus, in the Boothbay Region of Maine. Transactions of the American Fisheries Society, 104(4):669–674, October 1975. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chesney:1976:EWF

[CE76] Edward J. Chesney, Jr. and Jose Iglesias Estevez. Energetics of winter flounder (Pseudopleuronectes americanus) fed the polychaete, Nereis virens, under experimental conditions. Transactions of the American Fisheries Society, 105(5):592–595, September 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carlson:1972:SSL

[CH72] Anthony R. Carlson and John G. Hale. Successful spawning of largemouth bass Micropterus salmoides (Lacépède) under laboratory conditions. Transactions of the American Fisheries Society, 101(3):539–542, July 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Caillouet:1973:SSS

[CH73] Charles W. Caillouet, Jr. and James B. Higman. Sample size in sport fishery surveys. Transactions of the American Fisheries Society, 102(2):466–468, April 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clady:1975:EHW

[CH75] Michael Clady and Brendan Hutchinson. Effect of high winds on eggs of yellow perch, Perca flavescens, in Oneida Lake, New York. Transactions of the American Fisheries Society, 104(3):524–525, July 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carlson:1976:JLP

[CH76] H. Richard Carlson and Richard E. Haight. Juvenile life of Pacific Ocean perch, Sebastes alutus, in coastal fiords of Southeastern Alaska: Their environment, growth, food habits, and schooling behavior. Transactions of the American Fisheries Society, 105(2): 191–201, March 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cada:1978:ASM

[CH78a] Glenn F. Cada and Gary L. Hergenrader. An assessment of sampling mortality of larval fishes. Transactions of the American Fisheries Society, 107(2):269–274, March 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carlson:1978:ELT

[CH78b] Anthony R. Carlson and Lawrence J. Herman. Effect of long-term reduction and diel fluctuation in dissolved oxygen on spawning of black crappie, Pomoxis nigromaculatus. Transactions of the American Fisheries Society, 107(5):742–746, September 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chapman:1971:SCM

[Cha71] D. W. Chapman. Scientific communications via meetings. Transactions of the American Fisheries Society, 100(2):400–402, April 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chapman:1973:IFS

[Cha73] D. W. Chapman. Introduction to the Fishery Sciences: a reply. Transactions of the American Fisheries Society, 102(3):664–665, July 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See [Roy73a].

Chapman:1976:AEP

[Cha76] D. W. Chapman. Acoustic estimates of pelagic ichthyomass in Lake Tanganyika with an inexpensive echo sounder. Transactions of the American Fisheries Society, 105(5):581–587, September 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chapman:1978:ECZ

[Cha78a] Gary A. Chapman. Effects of continuous zinc exposure on sockeye salmon during adult-to-smolt freshwater residency. Transactions of the American Fisheries Society, 107(6):828–836, November 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chapman:1978:TCC

[Cha78b] Gary A. Chapman. Toxicities of cadmium, copper, and zinc to four juvenile stages of Chinook salmon and steelhead. Transactions of the American Fisheries Society, 107(6):841–847, November 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chevalier:1973:CFF

[Che73] John R. Chevalier. Cannibalism as a factor in first year survival of walleye in Oneida Lake. Transactions of the American Fisheries Society, 102(4):739–744, October 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chernoff:1975:MWD

[Che75] Barry Chernoff. A method for wet digestion of fish tissue for heavy metal analyses. Transactions of the American Fisheries Society, 104(4):803–804, October 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chittenden:1972:RYA

[Chi72a] Mark E. Chittenden, Jr. Responses of young American shad, Alosa sapidissima, to low temperatures. Transactions of the American Fisheries Society, 101(4):680–685, October 1972. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chittenden:1972:STY

[Chi72b] Mark E. Chittenden, Jr. Salinity tolerance of young blueback herring, Alosa aestivalis. Transactions of the American Fisheries Society, 101(1):123–125, January 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cummins:1971:OTB

[CJS71] Joseph M. Cummins, Abner C. Jones, and Alan A. Stevens. Occurrence of toxic bivalve molluscs during a gymnodinium breve “Red Tide”. Transactions of the American Fisheries Society, 100 (1):112–116, January 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cairns:1971:CAF

[CK71] John Cairns, Jr. and Roger L. Kaesler. Cluster analysis of fish in a portion of the Upper Potomac River. Transactions of the American Fisheries Society, 100(4):750–756, October 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cook:1975:PSS

[CL75] David W. Cook and Sandra R. Lofton. Pathogenicity studies with a Streptococcus sp. isolated from fishes in an Alabama–Florida fish kill. Transactions of the American Fisheries Society, 104(2):286– 288, April 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clark:1974:SVT

[Cla74] Stephen H. Clark. A study of variation in trawl data collected in Everglades National Park, Florida. Transactions of the American Fisheries Society, 103(4):777–785, October 1974. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clady:1975:ESA

[Cla75] Michael D. Clady. The effects of a simulated angler harvest on biomass and production in lightly exploited populations of smallmouth and largemouth bass. Transactions of the American Fisheries Society, 104(2):270–276, April 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clady:1977:APY

[Cla77] Michael Clady. Abundance and production of young largemouth bass, smallmouth bass, and yellow perch in two infertile Michigan lakes. Transactions of the American Fisheries Society, 106(1):57– 63, January 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Champ:1973:EAA

[CLB^{+}73] M. A. Champ, J. T. Lock, C. D. Bjork, W. G. Klussmann, and J. D. Mccullough, Jr. Effects of anhydrous ammonia on a Central Texas pond, and a review of previous research with ammonia in fisheries management. Transactions of the American Fisheries Society, 102(1):73–82, January 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cloutman:1975:PCS

[Clo75] Donald G. Cloutman. Parasite community structure of largemouth bass, warmouth, and bluegill in Lake Fort Smith, Arkansas. Transactions of the American Fisheries Society, 104(2):277–283, April 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cook:1970:MSM

[CM70a] S. F. Cook, Jr. and R. L. Moore. Mississippi silversides, Menidia audens (Atherinidae), established in California. Transactions of the American Fisheries Society, 99(1):70–73, January 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cramer:1970:SPZ

[CM70b] Joe D. Cramer and G. Richard Marzolf. Selective predation on zooplankton by gizzard shad. Transactions of the American Fisheries Society, 99(2):320–332, April 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cairns:1974:RBL

[CMS74] J. Cairns, Jr., E. L. Morgan, and R. E. Sparks. The response of bluegills (Lepomis macrochirus Rafinesque) in a pollution monitoring system to a diurnal temperature change. Transactions of the American Fisheries Society, 103(1):138–140, January 1974. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Caillouet:1976:GYP

[CNHT76] Charles W. Caillouet, Jamps P. Norris, Eric J. Heald, and Durbin C. Tabb. Growth and yield of pink shrimp (Penaeus duorarum duorarum) in feeding experiments in concrete tanks. Transactions of the American Fisheries Society, 105(2):259–266, March 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Coble:1970:FAF

[Cob70] Daniel W. Coble. False annulus formation in bluegill scales. Transactions of the American Fisheries Society, 99(2):363–368, April 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Coble:1971:EFC

[Cob71] Daniel W. Coble. Effects of fin clipping and other factors on survival and growth of smallmouth bass. Transactions of the American Fisheries Society, 100(3):460–473, July 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Coble:1972:VFC

[Cob72] Daniel W. Coble. Vulnerability of fin clipped bluegill to largemouth bass predation in tanks. Transactions of the American Fisheries Society, 101(3):563–565, July 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Coble:1973:ATR

[Cob73] Daniel W. Coble. Application of theory and research to freshwater and marine fisheries management. Transactions of the American Fisheries Society, 102(1):155, January 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cooper:1970:RMT

[Coo70] Richard A. Cooper. Retention of marks and their effects on growth, behavior, and migrations of the American lobster, Homarus americanus. Transactions of the American Fisheries Society, 99(2): 409–417, April 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cooper:1971:SDR

[Coo71] James A. Cooper. Scale development as related to growth of juvenile black crappie, Pomoxis nigromaculatus Lesueur. Transactions of the American Fisheries Society, 100(3):570–572, July 1971. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cooper:1978:IEL

[Coo78] John E. Cooper. Identification of eggs, larvae, and juveniles of the rainbow smelt, Osmerus mordax, with comparisons to larval alewife, Alosa pseudoharengus, and gizzard shad, Dorosoma cepedianum. Transactions of the American Fisheries Society, 107(1): 56–62, January 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cooper:1979:DEL

[Coo79] John E. Cooper. Description of eggs and larvae of fantail (Etheostoma flabellare) and rainbow (e. caeruleum) darters from Lake Erie tributaries. Transactions of the American Fisheries Society, 108(1):46–56, January 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Copeland:1970:ECR

[Cop70] B. J. Copeland. Estuarine classification and responses to disturbances. Transactions of the American Fisheries Society, 99(4): 826–835, October 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Coppage:1972:OPS

[Cop72] David L. Coppage. Organophosphate pesticides: Specific level of brain AChE inhibition related to death in sheepshead minnows. Transactions of the American Fisheries Society, 101(3):534–536, July 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Conte:1975:EAA

[CP75] Fred S. Conte and Jack C. Parker. Effect of aerially-applied malathion on juvenile brown and white shrimp Penaeus aztecus and P. setiferus. Transactions of the American Fisheries Society, 104(4):793–799, October 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Courtenay:1973:EAO

[CR73] Walter R. Courtenay, Jr. and C. Richard Robins. Exotic aquatic organisms in Florida with emphasis on fishes: a review and recommendations. Transactions of the American Fisheries Society, 102 (1):1–12, January 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cherry:1976:RMG

[CRC^{+}76] Donald S. Cherry, John H. Rodgers, Jr., John Cairns, Jr., Kenneth L. Dickson, and Rufus K. Guthrie. Responses of mosquitofish (Gambusia affinis) to ash effluent and thermal stress. Transactions of the American Fisheries Society, 105(6):686–694, November 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cross:1976:SNF

[Cro76] Jeffrey N. Cross. Status of the native fish fauna of the Moapa River (Clark County, Nevada). Transactions of the American Fisheries Society, 105(4):503–508, July 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Crutchfield:1973:EPO

[Cru73] James A. Crutchfield. Economic and political objectives in fishery management. Transactions of the American Fisheries Society, 102(2):481–491, April 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chapman:1978:ALL

[CS78] Gary A. Chapman and Donald G. Stevens. Acutely lethal levels of cadmium, copper, and zinc to adult male Coho salmon and steelhead. Transactions of the American Fisheries Society, 107 (6):837–840, November 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Colle:1978:USR

[CSG^{+}78] Douglas E. Colle, Jerome V. Shireman, Robert D. Gasaway, Robert L. Stetler, and William T. Haller. Utilization of selective removal of grass carp (Ctenopharyngodon idella) from an 80-hectare Florida lake to obtain a population estimate. Transactions of the American Fisheries Society, 107(5):724–729, September 1978. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carlson:1974:ELD

[CSH74] Anthony R. Carlson, Richard E. Siefert, and Lawrence J. Herman. Effects of lowered dissolved oxygen concentrations on channel catfish (Ictalurus punctatus) embryos and larvae. Transactions of the American Fisheries Society, 103(3):623–626, July 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Colle:1978:FSG

[CSR78] Douglas E. Colle, Jerome V. Shireman, and Roger W. Rottmann. Food selection by grass carp fingerlings in a vegetated pond. Transactions of the American Fisheries Society, 107(1):149–152, January 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chittenden:1972:ETC

[CV72] M. E. Chittenden, Jr. and W. A. Van Engel. Effect of a tickler chain and tow duration on trawl catches of the blue crab, Callinectes sapidus. Transactions of the American Fisheries Society, 101(4): 732–734, October 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chapman:1978:GMS

[CvW78a] D. W. Chapman and P. van Well. Growth and mortality of Stolothrissa tanganicae. Transactions of the American Fisheries Society, 107(1):26–35, January 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chapman:1978:OBL

[CvW78b] D. W. Chapman and P. van Well. Observations on the biology of Luciolates stappersi in Lake Tanganyika (Tanzania). Transactions of the American Fisheries Society, 107(4):567–573, July 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cope:1970:SCE

[CWW70] Oliver B. Cope, George H. Wallen, and Edward M. Wood. Some chronic effects of 2,4-D on the bluegill (Lepomis macrochirus). Transactions of the American Fisheries Society, 99(1):1–12, January 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dahlberg:1970:FAG

[Dah70] Michael D. Dahlberg. Frequencies of abnormalities in Georgia estuarine fishes. Transactions of the American Fisheries Society, 99(1): 95–97, January 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1971:ASS

[Dav71] Gary E. Davis. Aggregations of spiny sea urchins, Diadema antillarum, as shelter for young spiny lobsters, Panulirus argus. Transactions of the American Fisheries Society, 100(3):586–587, July 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1975:MSM

[Dav75] Gary E. Davis. Minimum size of mature spiny lobsters, Panulirus argus, at Dry Tortugas, Florida. Transactions of the American Fisheries Society, 104(4):675–676, October 1975. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1978:FET

[Dav78] Gary E. Davis. Field evaluation of a tag for juvenile spiny lobsters, Panulirus argus. Transactions of the American Fisheries Society, 107(1):100–103, January 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Donaldson:1970:ECE

[DB70] Lauren R. Donaldson and Kelshaw Bonham. Effects of chronic exposure of Chinook salmon eggs and alevins to gamma irradiation. Transactions of the American Fisheries Society, 99(1):112– 119, January 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Deacon:1972:EDF

[DB72a] James E. Deacon and W. Glen Bradley. Ecological distribution of fishes of Moapa (Muddy) River in Clark County, Nevada. Transactions of the American Fisheries Society, 101(3):408–419, July 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dennison:1972:RPB

[DB72b] Samuel G. Dennison and Ross V. Bulkley. Reproductive potential of the black bullhead, Ictalurus melas, in Clear Lake, Iowa. Transactions of the American Fisheries Society, 101(3):483–487, July 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dobbins:1976:PPF

[DB76] Daniel A. Dobbins and Claude E. Boyd. Phosphorus and potassium fertilization of sunfish ponds. Transactions of the American Fisheries Society, 105(4):536–540, July 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1978:CSE

[DB78] John A. Davis and Claude E. Boyd. Concentrations of selected elements and ash in bluegill (Lepomis macrochirus) and certain other freshwater fish. Transactions of the American Fisheries Society, 107(6):862–867, November 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Deangelis:1979:GRS

[DC79] D. L. Deangelis and C. C. Coutant. Growth rates and size distributions of first-year smallmouth bass populations: Some conclusions from experiments and a model. Transactions of the American Fisheries Society, 108(2):137–141, March 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dickson:1971:EUB

[DCA71] Kenneth L. Dickson, John Cairns, Jr., and Jesse C. Arnold. An evaluation of the use of a basket-type artificial substrate for sampling macroinvertebrate organisms. Transactions of the American Fisheries Society, 100(3):553–559, July 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Drummond:1970:IMS

[DD70] Robert A. Drummond and Walter F. Dawson. An inexpensive method for simulating diel patterns of lighting in the laboratory. Transactions of the American Fisheries Society, 99(2):434–435, April 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Drummond:1974:ISS

[DD74] Robert A. Drummond and Walter F. Dawson. An inexpensive solid-state amplifier for detecting movements and electrical potentials of fish. Transactions of the American Fisheries Society, 103(2):391–392, April 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dias:1972:RLS

[DDA72] Robert K. Dias, James K. Dias, and William D. Anderson, Jr. Relationships of lengths (standard, fork, and total) and lengths to weight in the red porgy, Pagrus sedecim (Perciformes, Sparidae), caught off South Carolina. Transactions of the American Fisheries Society, 101(3):503–506, July 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Duchrow:1971:TM

[DE71] Richard M. Duchrow and W. Harry Everhart. Turbidity measurement. Transactions of the American Fisheries Society, 100(4):682– 690, October 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dudley:1975:SLB

[DE75] Richard G. Dudley and A. W. Eipper. Survival of largemouth bass embryos at low dissolved oxygen concentrations. Transactions of the American Fisheries Society, 104(1):122–128, January 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Doble:1978:DFC

[DE78] Bruce D. Doble and Douglas M. Eggers. Diel feeding chronology, rate of gastric evacuation, daily ration, and prey selectivity in Lake Washington juvenile sockeye salmon (Oncorhynchus nerka). Transactions of the American Fisheries Society, 107(1):36–45, January 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dean:1976:TTF

[Dea76] John Mark Dean. Temperature of tissues in freshwater fishes. Transactions of the American Fisheries Society, 105(6):709–711, November 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Deacutis:1978:ETS

[Dea78] Christopher F. Deacutis. Effect of thermal shock on predator avoidance by larvae of two fish species. Transactions of the American Fisheries Society, 107(4):632–635, July 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Degraeve:1970:TTB

[Deg70] G. M. Degraeve. Three types of burrowing behavior of the brook stickleback, Culaea inconstans. Transactions of the American Fisheries Society, 99(2):433, April 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dell:1974:TRM

[Del74] Michael B. Dell. Tag returns and movements of rainbow trout (Salmo gairdneri) and rainbow-steelhead trout released in the Lake Washington system. Transactions of the American Fisheries Society, 103(2):250–254, April 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Devane:1978:FKM

[Dev78] John C. Devane, Jr. Food of king mackerel, Scomberomorus cavalla, in Onslow Bay, North Carolina. Transactions of the American Fisheries Society, 107(4):583–586, July 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dupree:1970:AAE

[DH70] Harry K. Dupree and John E. Halver. Amino acids essential for the growth of channel catfish, Ictalurus punctatus. Transactions of the American Fisheries Society, 99(1):90–92, January 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dunning:1978:PNA

[DH78] Dennis J. Dunning and Wayne F. Hadley. Participation of nonlicensed anglers in recreational fisheries, Erie County, New York. Transactions of the American Fisheries Society, 107(5):678–681, September 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dickie:1973:MFE

[Dic73] L. M. Dickie. Management of fisheries; ecological subsystems. Transactions of the American Fisheries Society, 102(2):470–480, April 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dorn:1979:SPN

[DJD79] Philip Dorn, Llew Johnson, and Cheryl Darby. The swimming performance of nine species of common California inshore fishes. Transactions of the American Fisheries Society, 108(4):366–372, July 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dwyer:1975:ITS

[DK75] William P. Dwyer and Robert H. Kramer. The influence of temperature on scope for activity in cutthroat trout, Salmo clarki. Transactions of the American Fisheries Society, 104(3):552–554, July 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1971:LHE

[DL71] James R. Davis and Darrell E. Louder. Life history and ecology of the cyprinid fish Notropis petersoni in North Carolina waters.

Transactions of the American Fisheries Society, 100(4):726–733, October 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Diana:1978:MDP

[DL78] James S. Diana and E. David Lane. The movement and distribution of Paiute cutthroat trout, Salmo clarki seleniris, in Cottonwood Creek, California. Transactions of the American Fisheries Society, 107(3):444–448, May 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Duronslet:1972:VDP

[DLM72] Marcel J. Duronslet, James M. Lyon, and Frank Marullo. Vertical distribution of postlarval brown, Penaeus aztecus, and white, P. setiferus, shrimp during immigration through a tidal pass. Transactions of the American Fisheries Society, 101(4):748–752, October 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dodge:1971:EIE

[DM71] Douglas P. Dodge and Hugh R. Maccrimmon. Environmental influences on extended spawning of rainbow trout (Salmo gairdneri). Transactions of the American Fisheries Society, 100(2):312–318, April 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dominy:1973:STE

[DM73] C. L. Dominy and G. L. Myatt. Short-term effects of three different marks on the recapture and survival of the alewife, Alosa pseudoharengus. Transactions of the American Fisheries Society, 102(3): 633–637, July 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dodson:1979:DBT

[DM79] Julian J. Dodson and Colin I. Mayfield. The dynamics and behavioral toxicology of aqua-kleen(R) (2,4-D butoxyethanol ester) as revealed by the modification of rheotropism in rainbow trout. Transactions of the American Fisheries Society, 108(6):632–640, November 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dawson:1976:RTC

[DMB76] Verdel K. Dawson, Leif L. Marking, and Terry D. Bills. Removal of toxic chemicals from water with activated carbon. Transactions of the American Fisheries Society, 105(1):119–123, January 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Diana:1977:MHP

[DME77] James S. Diana, W. C. Mackay, and Mark Ehrman. Movements and habitat preference of northern pike (Esox lucius) in Lac Ste. Anne, Alberta. Transactions of the American Fisheries Society, 106(6):560–565, November 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dudley:1977:MAS

[DMT77] Richard G. Dudley, Anthony W. Mullis, and James W. Terrell. Movements of adult striped bass (Morone saxatilis) in the Savannah River, Georgia. Transactions of the American Fisheries Society, 106(4):314–322, July 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Drummond:1974:CRU

[DOB74] Robert A. Drummond, Gayle F. Olson, and Allan R. Batterman. Cough response and uptake of mercury by brook trout, Salvelinus fontinalis, exposed to mercuric compounds at different hydrogenion concentrations. Transactions of the American Fisheries Society, 103(2):244–249, April 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dominy:1973:EEP

[Dom73] C. L. Dominy. Effect of entrance-pool weir elevation and fish density on passage of alewives (Alosa pseudoharengus) in a pool and weir fishway. Transactions of the American Fisheries Society, 102(2):398–404, April 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Derickson:1973:FSZ

[DP73] W. Kenneth Derickson and Kent S. Price, Jr. The fishes of the shore zone of Rehoboth and Indian River Bays, Delaware. Transactions of the American Fisheries Society, 102(3):552–562, July 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Desselle:1978:DFA

[DPRC78] Wayne J. Desselle, Michael A. Poirrier, James S. Rogers, and Robert C. Cashner. A discriminant functions analysis of sunfish (Lepomis) food habits and feeding niche segregation in the Lake Pontchartrain, Louisiana Estuary. Transactions of the American Fisheries Society, 107(5):713–719, September 1978. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Degraeve:1975:FBT

[DR75] G. M. Degraeve and James B. Reynolds. Feeding behavior and temperature and light tolerance of Mysis relicta in the laboratory. Transactions of the American Fisheries Society, 104(2):394–397, April 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dragovich:1970:FBT

[Dra70] Alexander Dragovich. The food of bluefin tuna (Thunnus thynnus) in the Western North Atlantic Ocean. Transactions of the American Fisheries Society, 99(4):726–731, October 1970. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1978:GRT

[DS78] Alan T. Davis and Robert R. Stickney. Growth responses of Tilapia aurea to dietary protein quality and quantity. Transactions of the American Fisheries Society, 107(3):479–483, May 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dove:1976:ETD

[DTD76] G. R. Dove, O. W. Tiemier, and C. W. Deyoe. Effects of three diets on growth and mineral retention of channel catfish fingerlings. Transactions of the American Fisheries Society, 105(3):481–485, May 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dudley:1974:GTK

[Dud74] Richard G. Dudley. Growth of Tilapia of the Kafue Floodplain, Zambia: Predicted effects of the Kafue Gorge Dam. Transactions of the American Fisheries Society, 103(2):281–291, April 1974. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Duttweiler:1976:UQS

[Dut76] Michael W. Duttweiler. Use of questionnaire surveys in forming fishery management policy. Transactions of the American Fisheries Society, 105(2):232–239, March 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

deVlaming:1978:SVR

[dVKP78] Victor L. de Vlaming, Armand Kuris, and Faust R. Parker, Jr. Seasonal variation of reproduction and lipid reserves in some subtropical cyprinodontids. Transactions of the American Fisheries Society, 107(3):464–472, May 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Devore:1978:DRB

[DW78] Philip W. Devore and Ray J. White. Daytime responses of brown trout (Salmo trutta) to cover stimuli in stream channels. Transactions of the American Fisheries Society, 107(6):763–771, November 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Drongowski:1975:TAC

[DWB75] R. A. Drongowski, J. S. Wood, and G. R. Bouck. Thyroid activity in Coho salmon from Oregon and Lake Michigan. Transactions of the American Fisheries Society, 104(2):349–352, April 1975. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Derr:1972:BAC

[DZ72] S. Krogh Derr and Matthew J. Zabik. Biologically active compounds in the aquatic environment: The uptake and distribution of [1,1-dichloro-2,2-bis (p-chlorophenyl) ethylene], DDE by Chironomus tentans Fabricius (Diptera: Chironomidae). Transactions of the American Fisheries Society, 101(2):323–329, April 1972. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eaton:1974:CCT

[Eat74] John G. Eaton. Chronic cadmium toxicity to the bluegill (Lepomis macrochirus Rafinesque). Transactions of the American Fisheries Society, 103(4):729–735, October 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Elston:1976:DFC

[EB76] Ralph Elston and Bruce Bachen. Diel feeding cycle and some effects of light on feeding intensity of the Mississippi silverside, Menidia audens, in Clear Lake, California. Transactions of the American Fisheries Society, 105(1):84–88, January 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Emery:1978:FBC

[EB78] Lee Emery and Edward H. Brown, Jr. Fecundity of the bloater (Coregonus hoyi) in Lake Michigan. Transactions of the American Fisheries Society, 107(6):785–789, November 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ebel:1970:ERL

[Ebe70] Wesley J. Ebel. Effect of release location on survival of juvenile fall Chinook salmon, Oncorhynchus tshawytscha. Transactions of the American Fisheries Society, 99(4):672–676, October 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ellgaard:1975:LAF

[EBM^{+}75] Erik G. Ellgaard, Kerry S. Bloom, Anthony A. Malizia, Jr., Gerald E. Gunning, and Richard E. Jensen. The locomotor activity of fish: an analogy to the kinetics of an opposed first-order chemical reaction. Transactions of the American Fisheries Society, 104 (4):752–754, October 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ebeling:1970:HTF

[EC70a] A. W. Ebeling and T. R. Chen. Heterogamety in teleostean fishes. Transactions of the American Fisheries Society, 99(1):131–138, January 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Edsall:1970:TTY

[EC70b] Thomas A. Edsall and Peter J. Colby. Temperature tolerance of young-of-the-year cisco, Coregonus artedii. Transactions of the American Fisheries Society, 99(3):526–531, July 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eder:1977:FHC

[EC77] Stephen Eder and Clarence A. Carlson. Food habits of carp and white suckers in the South Platte and St. Vrain Rivers and Goosequill Pond, Weld County, Colorado. Transactions of the American Fisheries Society, 106(4):339–346, July 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Esch:1976:RHP

[ECCC76] Gerald W. Esch, James R. Coggins, George C. Campbell, and Russell E. Conners. Recruitment of helminth parasites by bluegills (Lepomis macrochirus) using a modified live-Box technique. Transactions of the American Fisheries Society, 105(3):486–490, May 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eckroat:1971:LPP

[Eck71] Larry R. Eckroat. Lens protein polymorphisms in hatchery and natural populations of brook trout, Salvelinus fontinalis (Mitchill). Transactions of the American Fisheries Society, 100(3):527–536, July 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eckroat:1973:AFA

[Eck73] Larry R. Eckroat. Allele frequency analysis of five soluble protein loci in brook trout, Salvelinus fontinalis (Mitchill). Transactions of the American Fisheries Society, 102(2):335–340, April 1973. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Espinosa:1973:PLB

[ED73] F. A. Espinosa, Jr. and James E. Deacon. The preference of largemouth bass (Micropterus salmoides Lacépède) for selected bait species under experimental conditions. Transactions of the American Fisheries Society, 102(2):355–362, April 1973. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Edsall:1970:ETR

[Eds70a] Thomas A. Edsall. The effect of temperature on the rate of development and survival of alewife eggs and larvae. Transactions of the American Fisheries Society, 99(2):376–380, April 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Edsall:1970:HCL

[Eds70b] Thomas A. Edsall. A hermaphroditic coregonine from Lake Michigan. Transactions of the American Fisheries Society, 99(3):611, July 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Edwards:1978:EHR

[Edw78] Robert J. Edwards. The effect of hypolimnion reservoir releases on fish distribution and species diversity. Transactions of the American Fisheries Society, 107(1):71–77, January 1978. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eldridge:1977:EPH

[EEW77] Maxwell B. Eldridge, Tina Echeverria, and Jeannette A. Whipple. Energetics of Pacific herring (Clupea harengus pallasi) embryos and larvae exposed to low conentrations of benzene, a monoaromatic component of crude oil. Transactions of the American Fisheries Society, 106(5):452–461, September 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eisele:1976:EMR

[EH76a] Paul J. Eisele and Rolf Hartung. The effects of methoxychlor on riffle invertebrate populations and communities. Transactions of the American Fisheries Society, 105(5):628–633, September 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Erman:1976:QII

[EH76b] Don C. Erman and Vernon M. Hawthorne. The quantitative importance of an intermittent stream in the spawning of rainbow trout. Transactions of the American Fisheries Society, 105(6):675–681, November 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

El-Ibiary:1976:SVB

[EIAJ^{+}76] Hussein M. El-Ibiary, James W. Andrews, John A. Joyce, Jimmy W. Page, and Helen L. Deloach. Sources of variations in body size traits, dress-out weight, and lipid content and their correlations in channel catfish, Ictalurus punctatus. Transactions of the American Fisheries Society, 105(2):267–272, March 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eipper:1973:PEN

[Eip73] Alfred W. Eipper. Present educational needs in fishery biology. Transactions of the American Fisheries Society, 102(3):652–655, July 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eisele:1975:SDS

[Eis75] Paul J. Eisele. A simple dosing system for continuous stream dosing. Transactions of the American Fisheries Society, 104(2):359– 362, April 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eisler:1975:TCO

[EK75] Ronald Eisler and George Wm. Kissil. Toxicities of crude oils and oil-dispersant mixtures to juvenile rabbitfish, Siganus rivulatus. Transactions of the American Fisheries Society, 104(3):571–578, July 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Erman:1975:DMR

[EL75] Don C. Erman and George R. Leidy. Downstream movement of rainbow trout fry in a tributary of Sagehen Creek, under permanent and intermittent flow. Transactions of the American Fisheries Society, 104(3):467–473, July 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ellis:1970:ASN

[Ell70] Robert J. Ellis. Alloperla stonefly nymphs: Predators or scavengers on salmon eggs and alevins? Transactions of the American Fisheries Society, 99(4):677–683, October 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ellis:1974:JAB

[Ell74] James E. Ellis. The jumping ability and behavior of green sunfish (Lepomis cyanellus) at the outflow of a 1.6-ha pond. Transactions of the American Fisheries Society, 103(3):620–623, July 1974. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Elliot:1976:DAF

[Ell76] George V. Elliot. Diel activity and feeding of schooled largemouth bass fry. Transactions of the American Fisheries Society, 105(5): 624–627, September 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ellis:1978:BLS

[Ell78] C. M. A. Ellis. Biology of Luciolates stappersi in Lake Tanganyika (Burundi). Transactions of the American Fisheries Society, 107 (4):557–566, July 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Elrod:1971:DFA

[Elr71] Joseph H. Elrod. Dynamics of fishes in an Alabama pond subjected to intensive angling. Transactions of the American Fisheries Society, 100(4):757–768, October 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Elrod:1974:AGS

[Elr74] Joseph H. Elrod. Abundance, growth, survival, and maturation of channel catfish in Lake Sharpe, South Dakota. Transactions of the American Fisheries Society, 103(1):53–58, January 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Engel:1974:EFF

[Eng74] Sandy Engel. Effects of formalin and freezing on length, weight and condition factor of cisco and yellow perch. Transactions of the American Fisheries Society, 103(1):136–138, January 1974. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Engel:1976:FHP

[Eng76] Sandy Engel. Food habits and prey selection of Coho salmon (Oncorhynchus kisutch) and cisco (Coregonus artedii) in relation to zooplankton dynamics in Pallette Lake, Wisconsin. Transactions of the American Fisheries Society, 105(5):607–614, September 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Engelhardt:1977:RFP

[Eng77] William H. Engelhardt. Retention of fluorescent pigment marks by two strains of largemouth bass. Transactions of the American Fisheries Society, 106(1):64–66, January 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ellis:1973:CEM

[EP73] James E. Ellis and Edward N. Pickering. The catching efficiencies of a 21.3-meter (headrope) standard wing trawl and a 21.3-meter electrical wing trawl in the Saginaw Bay area of Lake Huron. Transactions of the American Fisheries Society, 102(1):116–120, January 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Echelle:1972:AEL

[ER72] Anthony A. Echelle and Carl D. Riggs. Aspects of the early life history of gars (Lepisosteus) in Lake Texoma. Transactions of the American Fisheries Society, 101(1):106–112, January 1972. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Erman:1972:FSM

[Erm72] Don C. Erman. A fisherman survey of the Madison–Firehole Rivers in Yellowstone National Park, 1969. Transactions of the American Fisheries Society, 101(2):352–355, April 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Erman:1973:UCF

[Erm73] Don C. Erman. Upstream changes in fish populations following impoundment of Sagehen Creek, California. Transactions of the American Fisheries Society, 102(3):626–629, July 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eckblad:1972:PLB

[ES72] James W. Eckblad and M. H. Shealy, Jr. Predation on largemouth bass embryos by the pond snail Viviparus georgianus. Transactions of the American Fisheries Society, 101(4):734–738, October 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Echelle:1976:FAS

[ES76] Anthony A. Echelle and Gary D. Schnell. Factor analysis of species associations among fishes of the Kiamichi River, Oklahoma. Transactions of the American Fisheries Society, 105(1):17–31, January 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Efford:1973:CFS

[ET73] Ian E. Efford and Kanji Tsumura. A comparison of the food of salamanders and fish in Marion Lake, British Columbia. Transactions of the American Fisheries Society, 102(1):33–47, January 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Etnier:1971:FTS

[Etn71] David A. Etnier. Food of three species of sunfishes (Lepomis, Centrarchidae) and their hybrids in three Minnesota lakes. Transactions of the American Fisheries Society, 100(1):124–128, January 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Etnier:1972:EAR

[Etn72] David A. Etnier. The effect of annual rechanneling on a stream fish population. Transactions of the American Fisheries Society, 101(2):372–375, April 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Everhart:1971:EFS

[Eve71] W. Harry Everhart. Evolution of the fisheries scientist: Introduction to the session. Transactions of the American Fisheries Society, 100(2):378, April 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Forester:1978:EGCa

[FA78] John S. Forester and James W. Avault, Jr. Effects of grass carp on freshwater red swamp crawfish in ponds. Transactions of the American Fisheries Society, 107(1):156–160, January 1978. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Farr:1977:IAB

[Far77] James A. Farr. Impairment of antipredator behavior in Palaemonetes pugio by exposure to sublethal doses of parathion. Transactions of the American Fisheries Society, 106(3):287–290, May 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Farr:1978:EMP

[Far78] James A. Farr. The effect of methyl parathion on predator choice of two estuarine prey species. Transactions of the American Fisheries Society, 107(1):87–91, January 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fasts:1971:EAD

[Fas71] Arlo W. Fasts. Effects of artificial destratification on zooplankton depth distribution. Transactions of the American Fisheries Society, 100(2):355–358, April 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fast:1973:EAH

[Fas73] Arlo W. Fast. Effects of artificial hypolimnion aeration on rainbow trout (Salmo gairdneri Richardson) depth distribution. Transactions of the American Fisheries Society, 102(4):715–722, October 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fore:1972:DFC

[FB72] Paul L. Fore and Kenneth N. Baxter. Diel fluctuations in the catch of larval Gulf menhaden, Brevoortia patronus, at Galveston entrance, Texas. Transactions of the American Fisheries Society, 101(4):729–732, October 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Farlinger:1977:ETV

[FB77] S. Farlinger and F. W. H. Beamish. Effects of time and velocity increments on the critical swimming speed of largemouth bass (Micropterus salmoides). Transactions of the American Fisheries Society, 106(5):436–439, September 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Farlinger:1978:CBC

[FB78] S. Farlinger and F. W. H. Beamish. Changes in blood chemistry and critical swimming speed of largemouth bass, Micropterus salmoides, with physical conditioning. Transactions of the American Fisheries Society, 107(4):523–527, July 1978. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Foltz:1979:COF

[FB79] Jeffery W. Foltz and Stephen H. Bowen. Chromic oxide in food assimilation studies. Transactions of the American Fisheries Society, 108(6):650–652, November 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Freeman:1971:TAH

[FE71] Robert A. Freeman and W. Harry Everhart. Toxicity of aluminum hydroxide complexes in neutral and basic media to rainbow trout. Transactions of the American Fisheries Society, 100(4):644–658, October 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Farringer:1979:RCR

[FEL79] Russel T. Farringer III, Anthony A. Echelle, and Steven F. Lehtinen. Reproductive cycle of the red shiner, Notropis lutrensis, in Central Texas and South Central Oklahoma. Transactions of the American Fisheries Society, 108(3):271–276, May 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Funk:1974:FMO

[FF74] John L. Funk and George G. Fleener. The fishery of a Missouri Ozark stream, Big Piney River, and the effects of stocking fingerling smallmouth bass. Transactions of the American Fisheries Society, 103(4):757–771, October 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fischer:1970:EA

[Fis70] David W. Fischer. An economic addendum. Transactions of the American Fisheries Society, 99(4):849–850, October 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Forester:1978:EGCb

[FL78a] Timothy S. Forester and John M. Lawrence. Effects of grass carp and carp on populations of bluegill and largemouth bass in ponds. Transactions of the American Fisheries Society, 107(1):172–175, January 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fuiman:1978:MCD

[FL78b] Lee A. Fuiman and Jules J. Loos. Morphological changes during the larval development of the cutlips minnow, exoglossum maxillingua. Transactions of the American Fisheries Society, 107(4):605–612, July 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

French:1970:WDS

[FM70] Robert R. French and W. Bruce McAlister. Winter distribution of salmon in relation to currents and water masses in the Northeastern Pacific Ocean and migrations of sockeye salmon. Transactions of the American Fisheries Society, 99(4):649–663, October 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fickeisen:1978:TFD

[FM78] D. H. Fickeisen and J. C. Montgomery. Tolerances of fishes to dissolved gas supersaturation in deep tank bioassays. Transactions of the American Fisheries Society, 107(2):376–381, March 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Facey:1977:RAS

[FMD77] Douglas E. Facey, James D. McCleave, and Glen E. Doyon. Responses of Atlantic salmon parr to output of pulsed ultrasonic transmitters. Transactions of the American Fisheries Society, 106 (5):489–496, September 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fagerlund:1979:TMG

[FMS79] U. H. M. Fagerlund, J. R. McBride, and E. T. Stone. A test of 17αmethyltestosterone as a growth promoter in a Coho salmon hatchery. Transactions of the American Fisheries Society, 108(5):467– 472, September 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fontaine:1971:LWR

[FN71] C. T. Fontaine and R. A. Neal. Length–weight relations for three commercially important penaeid shrimp of the Gulf of Mexico. Transactions of the American Fisheries Society, 100(3):584–586, July 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Foltz:1977:SCF

[FN77] Jeffrey W. Foltz and Carroll R. Norden. Seasonal changes in food consumption and energy content of smelt (Osmerus mordax) in Lake Michigan. Transactions of the American Fisheries Society, 106(3):230–234, May 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Forney:1971:DDY

[For71] John L. Forney. Development of dominant year classes in a yellow perch population. Transactions of the American Fisheries Society, 100(4):739–749, October 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Forney:1974:IBY

[For74] John L. Forney. Interactions between yellow perch abundance, walleye predation, and survival of alternate prey in Oneida Lake, New York. Transactions of the American Fisheries Society, 103(1): 15–24, January 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fox:1970:ESY

[Fox70] William W. Fox, Jr. An exponential surplus-yield model for optimizing exploited fish populations. Transactions of the American Fisheries Society, 99(1):80–88, January 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Frame:1973:BYW

[Fra73a] David W. Frame. Biology of young winter flounder Pseudopleuronectes americanus (Walbaum); metabolism under simulated estuarine conditions. Transactions of the American Fisheries Society, 102(2):423–430, April 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Frame:1973:CES

[Fra73b] David W. Frame. Conversion efficiency and survival of young winter flounder (Pseudopleuronectes americanus) under experimental conditions. Transactions of the American Fisheries Society, 102 (3):614–617, July 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Frame:1974:FHY

[Fra74a] David W. Frame. Feeding habits of young winter flounder (Pseudopleuronectes americanus): Prey availability and diversity. Transactions of the American Fisheries Society, 103(2):261–269, April 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fraser:1974:ATH

[Fra74b] J. M. Fraser. An attempt to train hatchery-reared brook trout to avoid predation by the common loon. Transactions of the American Fisheries Society, 103(4):815–818, October 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fraser:1978:ECY

[Fra78] J. M. Fraser. The effect of competition with yellow perch on the survival and growth of planted brook trout, splake, and rainbow trout in a small Ontario lake. Transactions of the American Fisheries Society, 107(4):505–517, July 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Freeman:1971:CFD

[Fre71] Robert A. Freeman. A constant flow delivery device for chronic bioassay. Transactions of the American Fisheries Society, 100(1): 135–136, January 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Freeman:1973:RRT

[Fre73] Robert A. Freeman. Recovery of rainbow trout from aluminum poisoning. Transactions of the American Fisheries Society, 102(1): 152–154, January 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fickeisen:1975:CEW

[FSM75] D. H. Fickeisen, M. J. Schneider, and J. C. Montgomery. A comparative evaluation of the Weiss saturometer. Transactions of the American Fisheries Society, 104(4):816–820, October 1975. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fava:1976:IBR

[FT76] James A. Fava and Chu-Fa Tsai. Immediate behavioral reactions of blacknose dace, Rhinichthys atratulus, to domestic sewage and its toxic constituents. Transactions of the American Fisheries Society, 105(3):430–441, May 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fuiman:1979:DCCa

[Fui79] Lee A. Fuiman. Descriptions and comparisons of catostomid fish larvae: Northern Atlantic drainage species. Transactions of the American Fisheries Society, 108(6):560–603, November 1979. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fujino:1970:IBG

[Fuj70] Kazuo Fujino. Immunological and biochemical genetics of tunas. Transactions of the American Fisheries Society, 99(1):152–178, January 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fallon:1977:OPU

[FW77] Mary Ellen Fallon and Dale C. Wallace. The occurrence of Phyllodistomum undulans in the urinary bladder of the mottled sculpin, Cottus bairdi. Transactions of the American Fisheries Society, 106(2):189–191, March 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fuiman:1979:DCCb

[FW79] Lee A. Fuiman and David C. Witman. Descriptions and comparisons of catostomid fish larvae: Catostomus catostomus and Moxostoma erythrurum. Transactions of the American Fisheries Society, 108(6):604–619, November 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goyert:1977:APS

[GA77] Jonathan C. Goyert and James W. Avault, Jr. Agricultural byproducts as supplemental feed for crayfish, Procambarus clarkii. Transactions of the American Fisheries Society, 106(6):629–633, November 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Galat:1973:NED

[Gal73] David L. Galat. Normal embryonic development of the muskellunge (Esox masquinongy). Transactions of the American Fisheries Society, 102(2):384–391, April 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gannon:1976:EDD

[Gan76] John E. Gannon. The effects of differential digestion rates of zooplankton by alewife, Alosa pseudoharengus, on determinations of selective feeding. Transactions of the American Fisheries Society, 105(1):89–95, January 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gard:1973:TTI

[Gar73] Richard Gard. Travel time and influencing factors for migrating adult sockeye salmon in Karluk River, Alaska. Transactions of the American Fisheries Society, 102(4):723–727, October 1973. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Guest:1976:PCC

[GAR76a] W. C. Guest, J. W. Avault, Jr., and J. D. Roussel. Preservation of channel catfish sperm. Transactions of the American Fisheries Society, 105(3):469–474, May 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Guest:1976:SSC

[GAR76b] W. C. Guest, J. W. Avault, Jr., and J. D. Roussel. A spermatology study of channel catfish, Ictalurus punctatus. Transactions of the American Fisheries Society, 105(3):463–468, May 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goodnight:1971:FPT

[GB71] W. H. Goodnight and T. C. Bjornn. Fish production in two Idaho streams. Transactions of the American Fisheries Society, 100(4): 769–780, October 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goodyear:1972:ECL

[GB72] C. Phillip Goodyear and Claude E. Boyd. Elemental composition of largemouth bass (Micropterus salmoides). Transactions of the American Fisheries Society, 101(3):545–547, July 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gale:1978:SFF

[GB78] William F. Gale and Gerard L. Buynak. Spawning frequency and fecundity of satinfin shiner (Notropis analostanus) — a fractional, crevice spawner. Transactions of the American Fisheries Society, 107(3):460–463, May 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goodyear:1979:SCO

[GB79] C. Phillip Goodyear and David H. Bennett. Sun compass orientation of immature bluegill. Transactions of the American Fisheries Society, 108(6):555–559, November 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gibson:1972:EZI

[GC72] H. R. Gibson and D. W. Chapman. Effects of zectran insecticide on aquatic organisms in Bear Valley Creek, Idaho. Transactions of the American Fisheries Society, 101(2):330–344, April 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gordon:1973:EES

[GCB73] D. Gordon, D. W. Chapman, and T. C. Bjornn. Economic evaluation of sport fisheries — what do they mean? Transactions of the American Fisheries Society, 102(2):293–311, April 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gruber:1977:INI

[GCD^{+}77] D. Gruber, J. Cairns, Jr., K. L. Dickson, R. Hummel III, A. Maciorowski, and W. H. Van Der Schalie. An inexpensive, noiseimmune amplifier designed for computer monitoring of ventilatory movements of fish and other biological events. Transactions of the American Fisheries Society, 106(5):497–499, September 1977. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gray:1979:BSR

[GD79] Robert H. Gray and Dennis D. Dauble. Biology of the sand roller in the Central Columbia River. Transactions of the American Fisheries Society, 108(6):646–649, November 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Galat:1975:PFO

[GE75a] D. L. Galat and A. W. Eipper. Presence of food organisms in the prolarval environment as a factor in the growth and mortality of larval muskellunge (Esox masquinongy). Transactions of the American Fisheries Society, 104(2):338–341, April 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Grothe:1975:CIM

[GE75b] Donald R. Grothe and John W. Eaton. Chlorine-induced mortality in fish. Transactions of the American Fisheries Society, 104(4): 800–802, October 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gibson:1975:RBI

[GG75] R. John Gibson and D. Galbraith. The relationships between invertebrate drift and salmonid populations in the Matamek River, Quebec, below a lake. Transactions of the American Fisheries Society, 104(3):529–535, July 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gale:1977:SHS

[GG77] William F. Gale and Cynthia A. Gale. Spawning habits of spotfin shiner (Notropis spilopterus) — a fractional, crevice spawner. Transactions of the American Fisheries Society, 106(2):170–177, March 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Guillory:1978:ZGC

[GG78] Vincent Guillory and Robert D. Gasaway. Zoogeography of the grass carp in the United States. Transactions of the American Fisheries Society, 107(1):105–112, January 1978. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Grant:1979:BMG

[GG79] W. E. Grant and W. L. Griffin. A bioeconomic model of the Gulf of Mexico shrimp fishery. Transactions of the American Fisheries Society, 108(1):1–13, January 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gray:1977:BJC

[GGB77] R. H. Gray, R. G. Genoway, and S. A. Barraclough. Behavior of juvenile Chinook salmon (Oncorhynchus tshawytscha) in relation to simulated thermal effluent. Transactions of the American Fisheries Society, 106(4):366–370, July 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goddard:1975:UAS

[GGG75] C. I. Goddard, I. A. Greig, and M. H. Goodwin. The use of artificial substrates in sampling estuarine benthos. Transactions of the American Fisheries Society, 104(1):50–52, January 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Giles:1973:HTE

[GH73] Robert H. Giles, Jr. and Stanley A. Huffman, Jr. Heuristic teaching environments: a design document. Transactions of the American Fisheries Society, 102(3):658–662, July 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gray:1977:DDA

[GH77] Robert H. Gray and James M. Haynes. Depth distribution of adult Chinook salmon (Oncorhynchus tshawytscha) in relation to season and gas-supersaturated water. Transactions of the American Fisheries Society, 106(6):617–620, November 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

George:1979:FHP

[GH79] Elinor L. George and Wayne F. Hadley. Food and habitat partitioning between rock bass (Ambloplites rupestris) and smallmouth bass (Micropterus dolomieui) young of the year. Transactions of the American Fisheries Society, 108(3):253–261, May 1979. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goodman:1979:DCT

[GHC^{+}79] Larry R. Goodman, David J. Hansen, David L. Coppage, James C. Moore, and Edward Matthews. Diazinon(R): Chronic toxicity to, and brain acetylcholinesterase inhibition in, the sheepshead minnow, Cyprinodon variegatus. Transactions of the American Fisheries Society, 108(5):479–488, September 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

George:1977:IMD

[GHH77] Elinor George, Edward J. Harrison, and Wayne F. Hadley. The incidence of Myxobolus dentium (Protozoa: Myxosporida) in Esox masquinongy of the Upper Niagara River. Transactions of the American Fisheries Society, 106(5):476–480, September 1977. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gibson:1978:BJA

[Gib78] R. John Gibson. The behavior of juvenile Atlantic salmon (Salmo salar) and brook trout (Salvelinus fontinalis) with regard to temperature and to water velocity. Transactions of the American Fisheries Society, 107(5):703–712, September 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Glenn:1978:SGD

[Gle78] C. L. Glenn. Seasonal growth and diets of young-of-the-year mooneye (Hiodon tergisus) from the Assiniboine River, Manitoba. Transactions of the American Fisheries Society, 107(4):587–589, July 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gallepp:1972:ENB

[GM72] George W. Gallepp and John J. Magnuson. Effects of negative buoyancy on the behavior of the bluegill, Lepomis macrochirus Rafinesque. Transactions of the American Fisheries Society, 101 (3):507–512, July 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gale:1978:LFD

[GM78] William F. Gale and Harold W. Mohr, Jr. Larval fish drift in a large river with a comparison of sampling methods. Transactions of the American Fisheries Society, 107(1):46–55, January 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Grove:1979:EFT

[GM79] Patricia Grove and Peter Moller. Electric fish as a tool for water quality assessment. Transactions of the American Fisheries Society, 108(4):420–421, July 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gall:1972:EIC

[GMS72] G. A. E. Gall, Emmitt L. Mcclendon, and William E. Schafer. Evidence on the influence of the copepod (Salmincola californiensis) on the reproductive performance of a domesticated strain of rainbow trout (Salmo gairdneri). Transactions of the American Fisheries Society, 101(2):345–346, April 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goodyear:1972:STD

[Goo72] C. Phillip Goodyear. A simple technique for detecting effects of toxicants or other stresses on a predator–prey interaction. Transactions of the American Fisheries Society, 101(2):367–370, April 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Grzenda:1970:UME

[GPT70] Alfred R. Grzenda, Doris Fort Parts, and William J. Taylor. The uptake, metabolism, and elimination of chlorinated residues by goldfish (Carassius auratus) fed a ^{14}C-DDT contaminated diet. Transactions of the American Fisheries Society, 99(2):385–396, April 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gallucci:1979:RFT

[GQ79] Vincent F. Gallucci and Terrance J. Quinn II. Reparameterizing, fitting, and testing a simple growth model. Transactions of the American Fisheries Society, 108(1):14–25, January 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gregory:1970:PCP

[Gre70] Richard W. Gregory. Physical and chemical properties of walleye sperm and seminal plasma. Transactions of the American Fisheries Society, 99(3):518–525, July 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Greene:1971:EEP

[Gre71] George N. Greene. An experiment on the effect of pond soil on calcium in pond water. Transactions of the American Fisheries Society, 100(3):580–582, July 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Griffith:1974:UID

[Gri74] J. S. Griffith, Jr. Utilization of invertebrate drift by brook trout (Salvelinus fontinalis) and cutthroat trout (Salmo clarki) in small streams in Idaho. Transactions of the American Fisheries Society, 103(3):440–447, July 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Griffith:1975:AFG

[Gri75] J. S. Griffith. Annulus formation and growth of tigerfish, Hydrocynus vittatus, in Lake Bangweulu, Zambia. Transactions of the American Fisheries Society, 104(3):499–505, July 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Griffith:1977:GCF

[Gri77] J. S. Griffith. Growth of the cichlid fish tylochromis bangwelensis in Lake Bangweulu, Zambia. Transactions of the American Fisheries Society, 106(2):146–150, March 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Griffith:1978:ELT

[Gri78a] J. S. Griffith. Effects of low temperature on the survival and behavior of threadfin shad, Dorosoma petenense. Transactions of the American Fisheries Society, 107(1):63–70, January 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Grimes:1978:AGL

[Gri78b] Churchill B. Grimes. Age, growth, and length–weight relationship of vermilion snapper, Rhomboplites aurorubens from North Carolina and South Carolina waters. Transactions of the American Fisheries Society, 107(3):454–456, May 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gard:1972:AHT

[GS72a] Richard Gard and Donald W. Seegrist. Abundance and harvest of trout in Sagehen Creek, California. Transactions of the American Fisheries Society, 101(3):463–477, July 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Griswold:1972:ESG

[GS72b] Bernard L. Griswold and Lloyd L. Smith, Jr. Early survival and growth of the ninespine stickleback, Pungitius pungitius. Transactions of the American Fisheries Society, 101(2):350–352, April 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gardella:1974:LAH

[GS74] E. S. Gardella and A. B. Stasko. A linear-array hydrophone for determining swimming depth of fish fitted with ultrasonic transmitters. Transactions of the American Fisheries Society, 103(3): 635–637, July 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Groen:1978:SFU

[GS78] Calvin L. Groen and James C. Schmulbach. The sport fishery of the unchannelized and channelized Middle Missouri River. Transactions of the American Fisheries Society, 107(3):412–418, May 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Greenland:1972:SSF

[GT72] Donald C. Greenland and Allan E. Thomas. Swimming speed of fall Chinook salmon (Oncorhynchus tshawytscha) fry. Transactions of the American Fisheries Society, 101(4):696–700, October 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gale:1975:SSQ

[GT75] William F. Gale and J. Douglas Thompson. A suction sampler for quantitatively sampling benthos on rocky substrates in rivers. Transactions of the American Fisheries Society, 104(2):398–405, April 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Grzenda:1971:UDC

[GTP71] Alfred R. Grzenda, William J. Taylor, and Doris F. Paris. The uptake and distribution of chlorinated residues by goldfish (Carassius auratus) fed a ^{14}C-Dieldrin contaminated diet. Transactions of the American Fisheries Society, 100(2):215–221, April 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Grzenda:1972:ETD

[GTP72] Alfred R. Grzenda, William J. Taylor, and Doris Fort Paris. The elimination and turnover of ^{14}C-Dieldrin by different goldfish tissues. Transactions of the American Fisheries Society, 101(4):686– 690, October 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Guerrero:1975:UAP

[Gue75] Rafael D. Guerrero III. Use of androgens for the production of allmale Tilapia aurea (Steindachner). Transactions of the American Fisheries Society, 104(2):342–348, April 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gulland:1978:FMN

[Gul78] John A. Gulland. Fishery management: New strategies for new conditions. Transactions of the American Fisheries Society, 107 (1):1–11, January 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gunning:1971:AFS

[Gun71] Gerald E. Gunning. American fisheries society publication archives. Transactions of the American Fisheries Society, 100(2):396–397, April 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gunstrom:1973:GPA

[Gun73] Gary K. Gunstrom. Gross physical anomalies in F_{1} progeny of irradiated and unirradiated Coho salmon. Transactions of the American Fisheries Society, 102(2):438–442, April 1973. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Giger:1972:NBS

[GW72] Richard D. Giger and Dewayne A. Williams. A net-board for setting beach seines from small boats. Transactions of the American Fisheries Society, 101(1):133–135, January 1972. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Giesy:1977:FDT

[GW77] John P. Giesy, Jr. and James G. Wiener. Frequency distributions of trace metal concentrations in five freshwater fishes. Transactions of the American Fisheries Society, 106(4):393–403, July 1977. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Giles:1973:CFH

[GZ73] Jeffrey H. Giles and Gilbert Zamora. Cover as a factor in habitat selection by juvenile brown (Penaeus aztecus) and white (P. setiferus) shrimp. Transactions of the American Fisheries Society, 102(1):144–145, January 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Havey:1973:PDH

[HA73] Keith A. Havey and Philip S. Andrews. Population dynamics of hatchery-reared landlocked salmon, Salmo salar, at Schoodic Lake, Maine. Transactions of the American Fisheries Society, 102(4): 728–738, October 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hickey:1975:SGJ

[HA75] Clarence R. Hickey, Jr. and Richard A. Amish. Stunted growth of a jawless striped bass, Morone saxatilis. Transactions of the American Fisheries Society, 104(2):410–412, April 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Huner:1976:SPF

[HA76] Jay V. Huner and James W. Avault, Jr. Sequential pond flooding: a prospective management technique for extended production of bait-size crawfish. Transactions of the American Fisheries Society, 105(5):637–642, September 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Haines:1973:ENE

[Hai73] Terry A. Haines. Effects of nutrient enrichment and a rough fish population (carp) on a game fish population (smallmouth bass). Transactions of the American Fisheries Society, 102(2):346–354, April 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hansen:1971:ESC

[Han71] David J. Hansen. Evaluation of stocking cutthroat trout, Salmo clarki, in Munsel Lake, Oregon. Transactions of the American Fisheries Society, 100(1):55–60, January 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hansen:1972:DME

[Han72] David J. Hansen. DDT and malathion: Effect on salinity selection by mosquitofish. Transactions of the American Fisheries Society, 101(2):346–350, April 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hansen:1974:EDF

[Han74] Randall G. Hansen. Effect of different filament diameters on the selective action of monofilament gill nets. Transactions of the American Fisheries Society, 103(2):386–387, April 1974. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hansen:1975:SEG

[Han75] Edward A. Hansen. Some effects of groundwater on brown trout redds. Transactions of the American Fisheries Society, 104(1):100– 110, January 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hartman:1978:MFS

[Har78] Alan M. Hartman. Mercury feeding schedules: Effects on accumulation, retention, and behavior in trout. Transactions of the American Fisheries Society, 107(2):369–375, March 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hassler:1970:EIE

[Has70] Thomas J. Hassler. Environmental influences on early development and year-class strength of northern pike in Lakes Oahe and Sharpe, South Dakota. Transactions of the American Fisheries Society, 99(2):369–375, April 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Havey:1973:ESI

[Hav73a] Keith A. Havey. Effects of a smelt introduction on growth of landlocked salmon at Schoodic Lake, Maine. Transactions of the American Fisheries Society, 102(2):392–397, April 1973. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Havey:1973:PJA

[Hav73b] Keith A. Havey. Production of juvenile alewives, Alosa pseudoharengus, at Love Lake, Washington County, Maine. Transactions of the American Fisheries Society, 102(2):434–437, April 1973. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Havey:1973:RRH

[Hav73c] Keith A. Havey. Relative recoveries of hatchery-reared landlocked salmon planted at different ages at Schoodic Lake, Maine. Transactions of the American Fisheries Society, 102(1):121–124, January 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Havey:1974:ERF

[Hav74a] Keith A. Havey. Effects of regulated flows on standing crops of juvenile salmon and other fishes at Barrows Stream, Maine. Transactions of the American Fisheries Society, 103(1):1–9, January 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Havey:1974:PDL

[Hav74b] Keith A. Havey. Population dynamics of landlocked salmon, Salmo salar, in Love Lake, Maine. Transactions of the American Fisheries Society, 103(3):448–456, July 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hergenrader:1971:WPN

[HB71] G. L. Hergenrader and Q. P. Bliss. The white perch in Nebraska. Transactions of the American Fisheries Society, 100(4):734–738, October 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hubbs:1974:MIT

[HB74] Clark Hubbs and Chris Bryan. Maximum incubation temperature of the threadfin shad, Dorosoma petenense. Transactions of the American Fisheries Society, 103(2):369–371, April 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Heins:1975:ELH

[HB75] David C. Heins and Gerald I. Bresnick. The ecological life history of the cherryfin shiner, Notropis roseipinnis. Transactions of the American Fisheries Society, 104(3):516–523, July 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hogue:1977:LDS

[HB77] J. J. Hogue, Jr. and J. P. Buchanan. Larval development of spotted sucker (Minytrema melanops). Transactions of the American Fisheries Society, 106(4):347–353, July 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Heckman:1978:IPL

[HB78] J. R. Heckman and M. O. Braune. Immunological properties of lactate dehydrogenase isozymes in trout. Transactions of the American Fisheries Society, 107(2):295–304, March 1978. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hershberger:1978:CEC

[HBD78] William K. Hershberger, Kelshaw Bonham, and Lauren R. Donaldson. Chronic exposure of Chinook salmon eggs and alevins to gamma irradiation: Effects on their return to freshwater as adults. Transactions of the American Fisheries Society, 107(4):622–631, July 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hambrick:1974:NAT

[HBH74] Patrick S. Hambrick, William T. Bryson, and Charles H. Hocutt. New anomalies in three cyprinid genera from Southwest Virginia. Transactions of the American Fisheries Society, 103(3):642, July 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Holcombe:1979:LTE

[HBL79] Gary W. Holcombe, Duane A. Benoit, and Edward N. Leonard. Long-term effects of zinc exposures on brook trout (Salvelinus fontinalis). Transactions of the American Fisheries Society, 108(1): 76–87, January 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hall:1979:IAT

[HBM79] Lenwood W. Hall, Jr., Dennis T. Burton, and Stuart L. Margrey. The influence of acclimation temperature on the interactions of chlorine, elevated temperature, and exposure duration for grass shrimp, Palaemonetes pugio. Transactions of the American Fisheries Society, 108(6):626–631, November 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Humphries:1973:ESB

[HC73a] Earl T. Humphries and Kenneth B. Cumming. An evaluation of striped bass fingerling culture. Transactions of the American Fisheries Society, 102(1):13–20, January 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Huntsman:1973:BDA

[HC73b] Gene R. Huntsman and Robert B. Chapoton. Biostatistical data acquisition in the menhaden fisheries. Transactions of the American Fisheries Society, 102(2):452–456, April 1973. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Howe:1975:WFM

[HC75] Arnold B. Howe and Phillip G. Coates. Winter flounder movements, growth, and mortality off Massachusetts. Transactions of the American Fisheries Society, 104(1):13–29, January 1975. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hausle:1976:ISR

[HC76] Donald A. Hausle and Daniel W. Coble. Influence of sand in redds on survival and emergence of brook trout (Salvelinus fontinalis). Transactions of the American Fisheries Society, 105(1):57–63, January 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Huntsman:1978:MCC

[HCD78] Gene R. Huntsman, David R. Colby, and Robert L. Dixon. Measuring catches in the Carolina headboat fishery. Transactions of the American Fisheries Society, 107(2):241–245, March 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Howe:1976:WFE

[HCP76] Arnold B. Howe, Philip G. Coates, and David E. Pierce. Winter flounder estuarine year-class abundance, mortality, and recruitment. Transactions of the American Fisheries Society, 105(6):647– 657, November 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Huh:1976:ETL

[HCS76] H. T. Huh, H. E. Calbert, and D. A. Stuiber. Effects of temperature and light on growth of yellow perch and walleye using formulated feed. Transactions of the American Fisheries Society, 105(2):254– 258, March 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Havey:1970:FIS

[HD70] Keith A. Havey and Robert M. Davis. Factors influencing standing crops and survival of juvenile salmon at Barrows Stream, Maine. Transactions of the American Fisheries Society, 99(2):297–311, April 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hickman:1973:NUL

[HD73] Gary D. Hickman and Michael R. Dewey. Notes on the upper lethal temperature of the duskystripe shiner, Notropis pilsbryi, and the bluegill, Lepomis macrochirus. Transactions of the American Fisheries Society, 102(4):838–840, October 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hauck:1976:PSS

[HE76] Forrest R. Hauck and Quentin A. Edson. Pumped storage: Its significance as an energy source and some biological ramifications. Transactions of the American Fisheries Society, 105(1):158–164, January 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Heald:1970:DEE

[Hea70] Eric J. Heald. Discussion: The Everglades Estuary: an example of seriously reduced inflow of freshwater. Transactions of the American Fisheries Society, 99(4):847–848, October 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Healey:1971:GDF

[Hea71] M. C. Healey. Gonad development and fecundity of the sand goby, Gobius minutus Pallas. Transactions of the American Fisheries Society, 100(3):520–526, July 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Heard:1972:SBP

[Hea72] William R. Heard. Spawning behavior of pink salmon on an artificial redd. Transactions of the American Fisheries Society, 101(2): 276–283, April 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Heidinger:1975:GHS

[Hei75] Roy C. Heidinger. Growth of hybrid sunfishes and channel catfish at low temperatures. Transactions of the American Fisheries Society, 104(2):333–334, April 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hershberger:1970:SPP

[Her70] William K. Hershberger. Some physicochemical properties of transferrins in brook trout. Transactions of the American Fisheries Society, 99(1):207–218, January 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Herke:1973:SDP

[Her73] William H. Herke. Spin-drying of preserved fishes and macroinvertebrates. Transactions of the American Fisheries Society, 102(3): 643–645, July 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hester:1970:PRS

[Hes70] F. Eugene Hester. Phylogenetic relationships of sunfishes as demonstrated by hybridization. Transactions of the American Fisheries Society, 99(1):100–104, January 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Huckabee:1974:TMS

[HG74] J. W. Huckabee and N. A. Griffith. Toxicity of mercury and selenium to the eggs of carp (Cyprinus carpio). Transactions of the American Fisheries Society, 103(4):822–825, October 1974. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hannon:1970:EDP

[HGAF70] Michael R. Hannon, Yvonne A. Greichus, Richard L. Applegate, and Alfred C. Fox. Ecological distribution of pesticides in Lake Poinsett, South Dakota. Transactions of the American Fisheries Society, 99(3):496–500, July 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Huckabee:1975:ARG

[HGJ75] John W. Huckabee, C. Phillip Goodyear, and Ronald D. Jones. Acid rock in the Great Smokies: Unanticipated impact on aquatic biota of road construction in regions of sulfide mineralization. Transactions of the American Fisheries Society, 104(4):677–684, October 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Haynes:1978:SMW

[HGM78] James M. Haynes, Robert H. Gray, and Jerry C. Montgomery. Seasonal movements of white sturgeon (Acipenser transmontanus) in the Mid-Columbia River. Transactions of the American Fisheries Society, 107(2):275–280, March 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Holt:1977:MWS

[HGO^{+}77] Charles S. Holt, Gary D. S. Grant, George P. Oberstar, Charles C. Oakes, and Dirck W. Bradt. Movement of walleye, Stizostedion vitreum, in Lake Bemidji, Minnesota as determined by radiobiotelemetry. Transactions of the American Fisheries Society, 106 (2):163–169, March 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hyde:1974:EMC

[HGS^{+}74] Kenneth M. Hyde, Jerry B. Graves, Sammy Stokes, James F. Fowler, and Frances L. Bonner. The effect of mirex on channel catfish production. Transactions of the American Fisheries Society, 103(2):366–369, April 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hale:1970:IFS

[HH70] John G. Hale and Donald A. Hilden. The influence of flow on the spawning of brook trout in the laboratory. Transactions of the American Fisheries Society, 99(3):595–597, July 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harrison:1979:BME

[HH79a] Edward J. Harrison and Wayne F. Hadley. Biology of muskellunge (Esox masquinongy) in the Upper Niagara River. Transactions of the American Fisheries Society, 108(5):444–451, September 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harrison:1979:CUC

[HH79b] Edward J. Harrison and Wayne F. Hadley. A comparison of the use of cleithra to the use of scales for age and growth studies. Transactions of the American Fisheries Society, 108(5):452–456, September 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hartwell:1979:TSR

[HH79c] S. Ian Hartwell and Donald E. Hoss. Thermal shock resistance of spot (Leiostomus xanthurus) after acclimation to constant or cycling temperature. Transactions of the American Fisheries Society, 108(4):397–400, July 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Higgins:1970:JTC

[Hig70] Bruce E. Higgins. Juvenile tunas collected by midwater trawling in Hawaiian waters, July–September 1967. Transactions of the American Fisheries Society, 99(1):60–69, January 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hile:1970:BSR

[Hil70] Ralph Hile. Body-scale relation and calculation of growth in fishes. Transactions of the American Fisheries Society, 99(3):468–474, July 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hildebrand:1971:ECS

[Hil71a] Stephen G. Hildebrand. The effect of Coho spawning on the benthic invertebrates of the Platte River, Benzie County, Michigan. Transactions of the American Fisheries Society, 100(1):61–68, January 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hile:1971:COS

[Hil71b] Ralph Hile. Communicating with other scientists — a panel discussion: No read — no write. Transactions of the American Fisheries Society, 100(2):394–395, April 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Huckabee:1978:MMB

[HJB^{+}78] John W. Huckabee, S. A. Janzen, B. G. Blaylock, Yair Talmi, and J. J. Beauchamp. Methylated mercury in brook trout (Salvelinus fontinalis): Absence of an in vivo methylating process. Transactions of the American Fisheries Society, 107(6):848–852, November 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hightower:1976:WPM

[HKB76] Glenn M. Hightower, Kay T. Kimball, and C. A. Bedinger, Jr. A water-powered mechanical device for accurately subsampling large numbers of nektonic organisms. Transactions of the American Fisheries Society, 105(4):509–513, July 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hendricks:1975:KDP

[HL75] Jerry D. Hendricks and Steve L. Leek. Kidney disease postorbital lesions in spring Chinook salmon (Oncorhynchus tshawytscha). Transactions of the American Fisheries Society, 104(4):805–807, October 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hammond:1976:ACT

[HL76] Dennis E. Hammond and Robert T. Lackey. Analysis of catchable trout fisheries management by computer simulation. Transactions of the American Fisheries Society, 105(1):48–56, January 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hackney:1978:SBL

[HL78] Peter A. Hackney and Thomas E. Linkous. Striking behavior of the largemouth bass and use of the binomial distribution for its analysis. Transactions of the American Fisheries Society, 107(5):682– 688, September 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hackney:1974:CMB

[HM74a] Peter A. Hackney and Charles K. Minns. A computer model of biomass dynamics and food competition with implications for its use in fishery management. Transactions of the American Fisheries Society, 103(2):215–225, April 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harima:1974:DIA

[HM74b] Hector Harima and Philip R. Mundy. Diversity indices applied to the fish biofacies of a small stream. Transactions of the American Fisheries Society, 103(3):457–461, July 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hokanson:1973:TRE

[HMJ73] Kenneth E. F. Hokanson, J. Howard Mccormick, and Bernard R. Jones. Temperature requirements for embryos and larvae of the northern pike, Esox lucius (Linnaeus). Transactions of the American Fisheries Society, 102(1):89–100, January 1973. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hawkes:1977:COT

[HN77] Clifford L. Hawkes and Logan A. Norris. Chronic oral toxicity of 2,3,7,8-tetrachlorodibenzo-p-dioxin (TCDD) to rainbow trout. Transactions of the American Fisheries Society, 106(6):641–645, November 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hartwell:1978:SPJ

[HO78] S. Ian Hartwell and Robert G. Otto. Swimming performance of juvenile menhaden (Brevoortia tyrannus). Transactions of the American Fisheries Society, 107(6):793–798, November 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hoagman:1977:FFD

[Hoa77] Walter J. Hoagman. A free-falling drop net for quantitatively sampling a water column. Transactions of the American Fisheries Society, 106(2):140–145, March 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hoffman:1974:DCW

[Hof74] G. L. Hoffman. Disinfection of contaminated water by ultraviolet irradiation, with emphasis on whirling disease (Myxosoma cerebralis) and its effect on fish. Transactions of the American Fisheries Society, 103(3):541–550, July 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hogman:1970:CPS

[Hog70] Walter J. Hogman. A computer program for stock analysis in FOR-TRAN IV, Univac 1108. Transactions of the American Fisheries Society, 99(2):426–427, April 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hogan:1971:BAC

[Hog71] James W. Hogan. Brain acetylcholinesterase from cutthroat trout. Transactions of the American Fisheries Society, 100(4):672–675, October 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hogman:1973:REN

[Hög73] Walter J. Högman. The relative efficiency of nylon gill nets after transition from cotton nets in a multispecies fishery. Transactions of the American Fisheries Society, 102(4):778–785, October 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hoopes:1975:FRH

[Hoo75] Rickalon L. Hoopes. Flooding as the result of Hurricane Agnes, and its effect on a native brook trout population, in an infertile headwater stream in Central Pennsylvania. Transactions of the American Fisheries Society, 104(1):96–99, January 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Horst:1977:ULM

[Hor77] Thomas J. Horst. Use of the Leslie matrix for assessing environmental impact with an example for a fish population. Transactions of the American Fisheries Society, 106(3):253–257, May 1977. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Howmiller:1971:CEE

[How71] Richard P. Howmiller. A comparison of the effectiveness of Ekman and ponar grabs. Transactions of the American Fisheries Society, 100(3):560–564, July 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Howmiller:1972:EPW

[How72] Richard P. Howmiller. Effects of preservatives on weights of some common macrobenthic invertebrates. Transactions of the American Fisheries Society, 101(4):743–746, October 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hoyt:1971:RBS

[Hoy71] Robert D. Hoyt. The reproductive biology of the silverjaw minnow, ericymba buccata cope, in Kentucky. Transactions of the American Fisheries Society, 100(3):510–519, July 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Held:1974:AGF

[HP74] John W. Held and John J. Peterka. Age, growth, and food habits of the fathead minnow, Pimephales promelas, in North Dakota saline lakes. Transactions of the American Fisheries Society, 103(4):743– 756, October 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Horkel:1976:ETV

[HP76] John D. Horkel and William D. Pearson. Effects of turbidity on ventilation rates and oxygen consumption of green sunfish, Lepomis cyanellus. Transactions of the American Fisheries Society, 105(1): 107–113, January 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Holder:1972:CAT

[HR72] Daniel R. Holder and John S. Ramsey. A case of albinism in the tadpole madtom, Noturus gyrinus. Transactions of the American Fisheries Society, 101(3):566–567, July 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hansen:1971:CRC

[HS71a] Martin J. Hansen and Thomas M. Stauffer. Comparative recovery to the creel, movement and growth of rainbow trout stocked in the Great Lakes. Transactions of the American Fisheries Society, 100(2):336–349, April 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hokanson:1971:SFI

[HS71b] Kenneth E. F. Hokanson and Lloyd L. Smith, Jr. Some factors influencing toxicity of linear alkylate sulfonate (LAS) to the bluegill. Transactions of the American Fisheries Society, 100(1):1–12, January 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hart:1975:SPI

[HS75a] Larry G. Hart and Robert C. Summerfelt. Surgical procedures for implanting ultrasonic transmitters into flathead catfish (Pylodictis olivaris). Transactions of the American Fisheries Society, 104(1): 56–59, January 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hill:1975:SES

[HS75b] T. K. Hill and E. W. Shell. Some effects of a sanctuary on an exploited fish population. Transactions of the American Fisheries Society, 104(3):441–445, July 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Holden:1975:DAM

[HS75c] Paul B. Holden and Clair B. Stalnaker. Distribution and abundance of mainstream fishes of the Middle and Upper Colorado River Basins, 1967–1973. Transactions of the American Fisheries Society, 104(2):217–231, April 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hansen:1975:EAE

[HSF75] David J. Hansen, Steven C. Schimmel, and Jerrold Forester. Effects of Aroclor(R) 1016 on embryos, fry, juveniles, and adults of sheepshead minnows (Cyprinodon variegatus). Transactions of the American Fisheries Society, 104(3):584–588, July 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hubbs:1971:DRM

[HSS71] Clark Hubbs, H. Bryan Sharp, and Joel F. Schneider. Developmental rates of Menidia audens with notes on salt tolerance. Transactions of the American Fisheries Society, 100(4):603–610, October 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hennick:1970:EME

[HT70] Daniel P. Hennick and Richard W. Tyler. Experimental marking of emergent pink salmon (Oncorhynchus gorbuscha) fry with sprayed fluorescent pigment. Transactions of the American Fisheries Society, 99(2):397–400, April 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hubbs:1970:ISF

[Hub70] Clark Hubbs. Introduction to symposium on fish cytogenetics. Transactions of the American Fisheries Society, 99(1):98–99, January 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hudson:1970:QST

[Hud70] Patrick L. Hudson. Quantitative sampling with three benthic dredges. Transactions of the American Fisheries Society, 99(3): 603–607, July 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hunt:1976:LTE

[Hun76] Robert L. Hunt. A long-term evaluation of trout habitat development and its relation to improving management-related research. Transactions of the American Fisheries Society, 105(3):361–364, May 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hurley:1973:CFA

[Hur73] Donal A. Hurley. The commercial fishery for American eel, Anguilla rostrata (Lesueur), in Lake Ontario. Transactions of the American Fisheries Society, 102(2):369–377, April 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hutchinson:1971:EFP

[Hut71] Brendan P. Hutchinson. The effect of fish predation on the zooplankton of ten Adirondack lakes, with particular reference to the alewife, Alosa pseudoharengus. Transactions of the American Fisheries Society, 100(2):325–335, April 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hanson:1974:RSC

[HW74] David L. Hanson and Thomas F. Waters. Recovery of standing crop and production rate of a brook trout population in a flooddamaged stream. Transactions of the American Fisheries Society, 103(3):431–439, July 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hamelink:1976:DLL

[HW76] Jerry L. Hamelink and Ronald C. Waybrant. DDE and lindane in a large-scale model lentic ecosystem. Transactions of the American Fisheries Society, 105(1):124–134, January 1976. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hall:1977:SDA

[HW77] Donald J. Hall and Earl E. Werner. Seasonal distribution and abundance of fishes in the littoral zone of a Michigan lake. Transactions of the American Fisheries Society, 106(6):545–555, November 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hamelink:1971:PEE

[HWB71] Jerry L. Hamelink, Ronald C. Waybrant, and Robert C. Ball. A proposal: Exchange equilibria control the degree chlorinated hydrocarbons are biologically magnified in lentic environments. Transactions of the American Fisheries Society, 100(2):207–214, April 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Iversen:1970:CBS

[ING70] Robert T. B. Iversen, Eugene L. Nakamura, and Reginald M. Gooding. Courting behavior in skipjack tuna, Katsuwonus pelamis. Transactions of the American Fisheries Society, 99(1):93, January 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Imler:1975:SRA

[IWF75] Richard L. Imler, Don T. Weber, and O. Leroy Fyock. Survival, reproduction, age, growth, and food habits of Sacramento perch, Archoplites interruptus (Girard), in Colorado. Transactions of the American Fisheries Society, 104(2):232–236, April 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Icanberry:1978:SLF

[IWR78] J. W. Icanberry, J. W. Warrick, and D. W. Rice, Jr. Seasonal larval fish abundance in waters off Diablo Canyon, California. Transactions of the American Fisheries Society, 107(2):225–233, March 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jacobi:1971:QAS

[Jac71] G. Zolton Jacobi. A quantitative artificial substrate sampler for benthic macroinvertebrates. Transactions of the American Fisheries Society, 100(1):136–138, January 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jahn:1972:DOA

[Jah72] Lawrence A. Jahn. Development of the olfactory apparatus of the cutthroat trout. Transactions of the American Fisheries Society, 101(2):284–289, April 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jolly:1978:ATP

[JAKG78] Algie L. Jolly, Jr., James W. Avault, Jr., Kenneth L. Koonce, and Jerry B. Graves. Acute toxicity of permethrin to several aquatic animals. Transactions of the American Fisheries Society, 107(6):825– 827, November 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jaspers:1976:SMU

[JAR76] Edmonde J. Jaspers, James W. Avault, Jr., and Joseph D. Roussel. Spermatozoal morphology and ultrastructure of channel catfish, Ictalurus punctatus. Transactions of the American Fisheries Society, 105(3):475–480, May 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jaspers:1978:TSC

[JAR78] Edmonde J. Jaspers, James W. Avault, Jr., and Joseph D. Roussel. Testicular and spermatozoal characteristics of channel catfish, Ictalurus punctatus, outside the spawning season. Transactions of the American Fisheries Society, 107(2):309–315, March 1978. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jones:1971:ENV

[JBR71] James H. Jones, Elizabeth B. Bullard, and Aida Rodriguez. The essential nature of vitamin A in the diet of goldfish (Carassius auratus). Transactions of the American Fisheries Society, 100(4): 676–681, October 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jensen:1970:VAD

[Jen70] Albert C. Jensen. Validation of ages determined from otoliths of Gulf of Maine cod. Transactions of the American Fisheries Society, 99(2):359–362, April 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jensen:1971:EIM

[Jen71] A. L. Jensen. The effect of increased mortality on the young in a population of brook trout, a theoretical analysis. Transactions of the American Fisheries Society, 100(3):456–459, July 1971. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jensen:1972:VRI

[Jen72] A. L. Jensen. Variation in results of identical bioassays of minnows subjected to instant temperature increase. Transactions of the American Fisheries Society, 101(3):403–407, July 1972. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jensen:1978:ALT

[Jen78] A. L. Jensen. Assessment of the lake trout fishery in Lake Superior: 1929–1950. Transactions of the American Fisheries Society, 107(4): 543–549, July 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jester:1973:VCF

[Jes73] Douglas B. Jester. Variations in catchability of fishes with color of gillnets. Transactions of the American Fisheries Society, 102(1): 109–115, January 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jester:1976:PRC

[Jes76] Douclas B. Jester. Production of rough or commercial fishes in Elephant Butte Lake, New Mexico. Transactions of the American Fisheries Society, 105(2):222–231, March 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jester:1977:ECM

[Jes77] Douglas B. Jester. Effects of color, mesh size, fishing in seasonal concentrations, and baiting on catch rates of fishes in gill nets. Transactions of the American Fisheries Society, 106(1):43–56, January 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnsen:1977:WAC

[JH77] Peter B. Johnsen and Arthur D. Hasler. Winter aggregations of carp (Cyprinus carpio) as revealed by ultrasonic tracking. Transactions of the American Fisheries Society, 106(6):556–559, November 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jarvis:1978:NMQ

[JKS78] Richard S. Jarvis, Harry F. Klodowski, and Sallie P. Sheldon. New method of quantifying scale shape and an application to stock identification in walleye (Stizostedion vitreum vitreum). Transactions of the American Fisheries Society, 107(4):528–534, July 1978. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:1970:AGP

[Joh70] James E. Johnson. Age, growth, and population dynamics of threadfin shad, Dorosoma petenense (Günther), in Central Arizona reservoirs. Transactions of the American Fisheries Society, 99(4):739–753, October 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:1971:MFT

[Joh71] James E. Johnson. Maturity and fecundity of threadfin shad, Dorosoma petenense (Günther), in Central Arizona reservoirs. Transactions of the American Fisheries Society, 100(1):74–85, January 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Joseph:1970:MTT

[Jos70] James Joseph. Management of tropical tunas in the Eastern Pacific Ocean. Transactions of the American Fisheries Society, 99(3):629– 648, July 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jones:1976:VHV

[JP76] B. Juanelle Jones and William D. Pearson. Variations in hematocrit values of successive blood samples from bluegill. Transactions of the American Fisheries Society, 105(2):291–293, March 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnsen:1973:PSJ

[JS73] Richard D. Johnsen and Carl W. Sims. Purse seining for juvenile salmon and trout in the Columbia River Estuary. Transactions of the American Fisheries Society, 102(2):341–345, April 1973. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Julin:1977:TAI

[JS77] Arnold M. Julin and Herman O. Sanders. Toxicity and accumulation of the insecticide imidan in freshwater invertebrates and fishes. Transactions of the American Fisheries Society, 106(4):386–392, July 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jude:1973:FFH

[Jud73] David J. Jude. Food and feeding habits of gizzard shad in Pool 19, Mississippi River. Transactions of the American Fisheries Society, 102(2):378–383, April 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:1977:CIM

[JWA77] Allyn G. Johnson, Theodore D. Williams, and Connie R. Arnold. Chlorine-induced mortality of eggs and larvae of spotted seatrout (Cynoscion nebulosus). Transactions of the American Fisheries Society, 106(5):466–469, September 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kaiser:1973:AGH

[Kai73a] Carlos E. Kaiser. Age and growth of horse mackerel Trachurus murphyi off Chile. Transactions of the American Fisheries Society, 102(3):591–595, July 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kaiser:1973:GMF

[Kai73b] Carlos E. Kaiser. Gonadal maturation and fecundity of horse mackerel Trachurus murphyi (Nichols) off the Coast of Chile. Transactions of the American Fisheries Society, 102(1):101–108, January 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kawatski:1973:ATA

[Kaw73] Joseph A. Kawatski. Acute toxicities of antimycin a, Bayer 73, and TFM to the ostracod cypretta kawatai. Transactions of the American Fisheries Society, 102(4):829–831, October 1973. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kaya:1977:RBR

[Kay77] Calvin M. Kaya. Reproductive biology of rainbow and brown trout in a geothermally heated stream: The Firehole River of Yellowstone National Park. Transactions of the American Fisheries Society, 106 (4):354–361, July 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kelts:1971:FFL

[KB71] Lora I. Kelts and Janet I. Bressler. Fish and fisheries literature resources: an annotated bibliography. Transactions of the American Fisheries Society, 100(2):403–422, April 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Klarberg:1975:FHI

[KB75] David P. Klarberg and Arnold Benson. Food habits of Ictalurus nubulosus in acid polluted water of Northern West Virginia. Transactions of the American Fisheries Society, 104(3):541–547, July 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Katz:1977:ASR

[KB77] Philip L. Katz and L. J. Bledsoe. Alaska shellfish regulations: Present impacts on fishery participants. Transactions of the American Fisheries Society, 106(6):505–529, November 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kerby:1971:OGS

[KBR71] Jerome Howard Kerby, V. G. Burrell, Jr., and C. E. Richards. Occurrence and growth of striped bass × white bass hybrids in the Rappahannock River, Virginia. Transactions of the American Fisheries Society, 100(4):787–790, October 1971. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kincaid:1977:TGS

[KBvL77] H. L. Kincaid, W. R. Bridges, and B. von Limbach. Three generations of selection for growth rate in fall-spawning rainbow trout. Transactions of the American Fisheries Society, 106(6):621–628, November 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kempinger:1972:CNS

[KC72] James J. Kempinger and Warren S. Churchill. Contribution of native and stocked walleye fingerlings to the anglers’ catch, Escanaba Lake, Wisconsin. Transactions of the American Fisheries Society, 101(4):644–649, October 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kawatski:1974:ETB

[KDR74] J. A. Kawatski, V. K. Dawson, and M. L. Reuvers. Effect of TFM and Bayer 73 on in vivo oxygen consumption of the aquatic midge Chironomus tentans. Transactions of the American Fisheries Society, 103(3):551–556, July 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

King:1979:FNM

[KDS79] Terry A. King, William D. Davies, and William L. Shelton. Fishing and natural mortality: Effects on the initial year class of largemouth bass in West Point Reservoir, Albama–Georgia. Transactions of the American Fisheries Society, 108(2):150–155, March 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Keating:1971:CP

[Kea71] Thomas P. Keating. Communicating with the public. Transactions of the American Fisheries Society, 100(2):389–393, April 1971. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kelso:1973:SEC

[Kel73] John R. M. Kelso. Seasonal energy changes in walleye and their diet in West Blue Lake, Manitoba. Transactions of the American Fisheries Society, 102(2):363–368, April 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kemmerer:1970:RJV

[Kem70] Andrew J. Kemmerer. Rebuttal to J. Verduin’s critique of research methods involving plastic bags in aquatic environments. Transactions of the American Fisheries Society, 99(3):612–613, July 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ketola:1975:RAS

[Ket75] H. G. Ketola. Requirement of Atlantic salmon for dietary phosphorus. Transactions of the American Fisheries Society, 104(3): 548–551, July 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kelso:1974:ORM

[KF74] John R. M. Kelso and Richard Frank. Organochlorine residues, mercury, copper and cadmium in yellow perch, white bass and smallmouth bass, Long Point Bay, Lake Erie. Transactions of the American Fisheries Society, 103(3):577–581, July 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kruse:1977:PDL

[KF77] Kipp C. Kruse and Michael G. Francis. A predation deterrent in larvae of the bullfrog, Rana catesbeiana. Transactions of the American Fisheries Society, 106(3):248–252, May 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kroger:1972:OPB

[KG72] Richard L. Kroger and James F. Guthrie. Occurrence of the parasitic branchiuran, Argulus alosae, on dying Atlantic menhaden, Brevoortia tyrannus, in the Connecticut River. Transactions of the American Fisheries Society, 101(3):559–560, July 1972. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kroger:1973:MTJ

[KG73] Richard L. Kroger and James F. Guthrie. Migrations of tagged juvenile Atlantic menhaden. Transactions of the American Fisheries Society, 102(2):417–422, April 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kroger:1974:GFA

[KGJ74] Richard L. Kroger, James F. Guthrie, and Mayo H. Judy. Growth and first annulus formation of tagged and untagged Atlantic menhaden. Transactions of the American Fisheries Society, 103(2): 292–296, April 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kimball:1971:MEF

[KH71] Dan C. Kimball and William T. Helm. A method of estimating fish stomach capacity. Transactions of the American Fisheries Society, 100(3):572–575, July 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kaya:1972:PTE

[KH72] Calvin M. Kaya and Arthur D. Hasler. Photoperiod and temperature effects on the gonads of green sunfish, Lepomis cyanellus (Rafinesque), during the quiescent, winter phase of its annual sexual cycle. Transactions of the American Fisheries Society, 101(2): 270–275, April 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Knudsen:1978:GRM

[KH78] E. Eric Knudsen and William H. Herke. Growth rate of marked juvenile Atlantic croakers, Micropogon undulatus, and length of stay in a coastal marsh nursery in Southwest Louisiana. Transactions of the American Fisheries Society, 107(1):12–20, January 1978. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kilgen:1978:GCC

[Kil78] Ronald H. Kilgen. Growth of channel catfish and striped bass in small ponds stocked with grass carp and water hyacinths. Transactions of the American Fisheries Society, 107(1):176–180, January 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kimura:1976:ETN

[Kim76] Daniel K. Kimura. Estimating the total number of marked fish present in a catch. Transactions of the American Fisheries Society, 105(6):664–668, November 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kincaid:1976:EIR

[Kin76] Harold L. Kincaid. Effects of inbreeding on rainbow trout populations. Transactions of the American Fisheries Society, 105(2): 273–280, March 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kincaid:1979:DSR

[Kin79] H. L. Kincaid. Development of standard reference lines of rainbow trout. Transactions of the American Fisheries Society, 108(5):457– 461, September 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kirk:1974:EHC

[Kir74] William L. Kirk. The effects of hypoxia on certain blood and tissue electrolytes of channel catfish, Ictalurus punctatus (Rafinesque). Transactions of the American Fisheries Society, 103(3):593–600, July 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kircheis:1976:RBE

[Kir76] Frederick W. Kircheis. Reproductive biology and early life history of the Sunapee trout of Floods Pond, Maine. Transactions of the American Fisheries Society, 105(5):615–619, September 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kissil:1974:SAA

[Kis74] George William Kissil. Spawning of the anadromous alewife, Alosa pseudoharengus, in Bride Lake, Connecticut. Transactions of the American Fisheries Society, 103(2):312–317, April 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kjelson:1978:CEM

[KJ78] Martin A. Kjelson and George N. Johnson. Catch efficiencies of a 6.1-meter otter trawl for estuarine fish populations. Transactions of the American Fisheries Society, 107(2):246–254, March 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kaeding:1978:GDT

[KK78] Lynn R. Kaeding and Calvin M. Kaya. Growth and diets of trout from contrasting environments in a geothermally heated stream: The Firehole River of Yellowstone National Park. Transactions of the American Fisheries Society, 107(3):432–438, May 1978. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kitchell:1974:MFB

[KKM^{+}74] James F. Kitchell, Joseph F. Koonce, John J. Magnuson, Robert V. O’Neill, Herman H. Shugart, Jr., and Ray S. Booth. Model of fish biomass dynamics. Transactions of the American Fisheries Society, 103(4):786–798, October 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Klima:1974:WSM

[Kli74] Edward F. Klima. A white shrimp mark-recapture study. Transactions of the American Fisheries Society, 103(1):107–113, January 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kottcamp:1972:UDB

[KM72] Glenn M. Kottcamp and Peter B. Moyle. Use of disposable beverage cans by fish in the San Joaquin Valley. Transactions of the American Fisheries Society, 101(3):566, July 1972. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Klaassen:1974:AGL

[KM74] Harold E. Klaassen and Kenneth L. Morgan. Age and growth of longnose gar in Tuttle Creek Reservoir, Kansas. Transactions of the American Fisheries Society, 103(2):402–405, April 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kristiansson:1976:GVC

[KM76] A. Christine Kristiansson and John D. McIntyre. Genetic variation in Chinook salmon (Oncorhynchus tshawytscha) from the Columbia River and three Oregon coastal rivers. Transactions of the American Fisheries Society, 105(5):620–623, September 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Krueger:1979:ESG

[KM79] Charles C. Krueger and Bruce W. Menzel. Effects of stocking on genetics of wild brook trout populations. Transactions of the American Fisheries Society, 108(3):277–287, May 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kelso:1970:SIR

[KME70] John R. M. Kelso, H. R. Maccrimmon, and D. J. Ecobichon. Seasonal insecticide residue changes in tissues of fish from the Grand River, Ontario. Transactions of the American Fisheries Society, 99(2):423–426, April 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kania:1974:BAS

[KO74] Henry J. Kania and James O’Hara. Behavioral alterations in a simple predator–prey system due to sublethal exposure to mercury. Transactions of the American Fisheries Society, 103(1):134–136, January 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Koch:1973:RCC

[Koc73] David L. Koch. Reproductive characteristics of the cui-ui lakesucker (Chasmistes cujus Cope) and its spawning behavior in Pyramid Lake, Nevada. Transactions of the American Fisheries Society, 102(1):145–149, January 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Koehn:1970:FED

[Koe70] Richard K. Koehn. Functional and evolutionary dynamics of polymorphic esterases in catostomid fishes. Transactions of the American Fisheries Society, 99(1):219–228, January 1970. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Korn:1973:UPC

[Kor73] Sidney Korn. The uptake and persistence of carbaryl in channel catfish. Transactions of the American Fisheries Society, 102(1): 137–139, January 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Koski:1978:AGM

[Kos78] Robert T. Koski. Age, growth, and maturity of the hogchoker, Trinectes maculatus, in the Hudson River, New York. Transactions of the American Fisheries Society, 107(3):449–453, May 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kennedy:1975:EHS

[KP75] G. J. A. Kennedy and T. J. Pitcher. Experiments on homing in shoals of the European minnow, Phoxinus phoxinus (L). Transactions of the American Fisheries Society, 104(3):454–457, July 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kawatski:1971:AIF

[KS71] Joseph A. Kawatski and James C. Schmulbach. Accumulation of insecticide in freshwater ostracods exposed continuously to sublethal concentrations of Aldrin or Dieldrin. Transactions of the American Fisheries Society, 100(3):565–567, July 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Karlsson:1973:TDB

[KS73] John Karlsson and Richard Sisson. A technique for detection of brushed lobsters by staining of cement on swimmerets. Transactions of the American Fisheries Society, 102(4):847–848, October 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kutty:1975:IUL

[KS75] M. N. Kutty and N. Sukumaran. Influence of upper and lower temperature extremes on the swimming performance of Tilapia mossambica. Transactions of the American Fisheries Society, 104 (4):755–761, October 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kratt:1977:PHS

[KS77] L. F. Kratt and R. J. F. Smith. A post-hatching sub-gravel stage in the life history of the Arctic grayling, Thymallus arcticus. Transactions of the American Fisheries Society, 106(3):241–243, May 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kimball:1978:CTH

[KSB78] Gary L. Kimball, Lloyd L. Smith, Jr., and Steven J. Broderius. Chronic toxicity of hydrogen cyanide to the bluegill. Transactions of the American Fisheries Society, 107(2):341–345, March 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kjelson:1975:DSD

[KTJ75] Martin A. Kjelson, William R. Turner, and George N. Johnson. Description of a stationary drop-net for estimating nekton abundance in shallow waters. Transactions of the American Fisheries Society, 104(1):46–49, January 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kushlan:1974:ENF

[Kus74a] James A. Kushlan. Effects of a natural fish kill on the water quality, plankton, and fish population of a pond in the Big Cypress Swamp, Florida. Transactions of the American Fisheries Society, 103(2): 235–243, April 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kushlan:1974:QSF

[Kus74b] James A. Kushlan. Quantitative sampling of fish populations in shallow, freshwater environments. Transactions of the American Fisheries Society, 103(2):348–352, April 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Klima:1971:ACP

[KW71] Edward F. Klima and Donald A. Wickham. Attraction of coastal pelagic fishes with artificial structures. Transactions of the American Fisheries Society, 100(1):86–99, January 1971. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kionka:1972:DMD

[KW72a] B. C. Kionka and J. T. Windell. Differential movement of digestible and indigestible food fractions in rainbow trout, Salmo gairdneri. Transactions of the American Fisheries Society, 101(1):112–115, January 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Koo:1972:STS

[KW72b] Ted S. Y. Koo and John S. Wilson. Sonic tracking striped bass in the Chesapeake and Delaware canal. Transactions of the American Fisheries Society, 101(3):453–462, July 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kynard:1974:ABI

[Kyn74] Boyd Kynard. Avoidance behavior of insecticide susceptible and resistant populations of mosquitofish to four insecticides. Transactions of the American Fisheries Society, 103(3):557–561, July 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lewis:1971:ITD

[LA71] D. H. Lewis and Terry C. Allison. An immunofluorescent technique for detecting Aeromonas liquefaciens in fish utilized in lunar exposure studies. Transactions of the American Fisheries Society, 100 (3):575–578, July 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Laarman:1971:SVP

[Laa71] Percy W. Laarman. Seasonal variation in protein content of bluegill pituitary glands. Transactions of the American Fisheries Society, 100(2):360–362, April 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Labar:1971:MHC

[Lab71] George W. Labar. Movement and homing of cutthroat trout (Salmo clarki) in Clear and Bridge Creeks, Yellowstone National Park. Transactions of the American Fisheries Society, 100(1):41–49, January 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lackey:1973:EAD

[Lac73a] Robert T. Lackey. Effects of artificial destratification on zooplankton in Parvin Lake, Colorado. Transactions of the American Fisheries Society, 102(2):450–452, April 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lackey:1973:ITF

[Lac73b] Robert T. Lackey. Innovative teaching in fisheries science. Transactions of the American Fisheries Society, 102(3):647–648, July 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lackey:1975:CAF

[Lac75] Robert T. Lackey. Computer applications in fisheries science. Transactions of the American Fisheries Society, 104(3):589–590, July 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Larimore:1974:SDI

[Lar74] R. Weldon Larimore. Stream drift as an indication of water quality. Transactions of the American Fisheries Society, 103(3):507–517, July 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Larkin:1977:ECM

[Lar77] P. A. Larkin. An epitaph for the concept of maximum sustained yield. Transactions of the American Fisheries Society, 106(1):1–11, January 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lawrie:1970:SLG

[Law70] A. H. Lawrie. The sea lamprey in the Great Lakes. Transactions of the American Fisheries Society, 99(4):766–775, October 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lay:1971:APP

[Lay71] Bill A. Lay. Applications for potassium permanganate in fish culture. Transactions of the American Fisheries Society, 100(4):813– 816, October 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lewis:1976:ERP

[LB76] William M. Lewis and Gerard L. Buynak. Evaluation of a revolving plate type biofilter for use in recirculated fish production and holding units. Transactions of the American Fisheries Society, 105 (6):704–708, November 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lichtkoppler:1977:PFS

[LB77] Frank Lichtkoppler and Claude E. Boyd. Phosphorus fertilization of sunfish ponds. Transactions of the American Fisheries Society, 106(6):634–636, November 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lind:1970:CMA

[LC70] Owen T. Lind and Robert S. Campbell. Community metabolism in acid and alkaline strip-mine lakes. Transactions of the American Fisheries Society, 99(3):577–582, July 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Loyacano:1977:SCO

[LCG77] Harold A. Loyacano, Jr., Jesse A. Chappell, and Sidney A. Gauthreaux. Sun-compass orientation in juvenile largemouth bass, Micropterus salmoides. Transactions of the American Fisheries Society, 106(1):77–79, January 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Laughlin:1978:FLA

[LCL78] Roger A. Laughlin, Claude R. Cripe, and Robert J. Livingston. Field and laboratory avoidance reactions by blue crabs (Callinectes sapidus) to storm water runoff. Transactions of the American Fisheries Society, 107(1):78–86, January 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lee:1971:EPC

[LDM71] Terrence H. Lee, Philip H. Derse, and Stephen D. Morton. Effects of physical and chemical conditions on the detoxification of antimycin. Transactions of the American Fisheries Society, 100(1): 13–17, January 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Leggett:1972:WLA

[Leg72] William C. Leggett. Weight loss in American shad (Alosa sapidissima, Wilson) during the freshwater migration. Transactions of the American Fisheries Society, 101(3):549–552, July 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Levi:1973:JYM

[Lev73] Eldon J. Levi. Juvenile yellowfin menhaden frown the Bahama Islands. Transactions of the American Fisheries Society, 102(4): 848–849, October 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Levings:1974:SCF

[Lev74] C. D. Levings. Seasonal changes in feeding and particle selection by winter flounder (Pseudopleuronectes americanus). Transactions of the American Fisheries Society, 103(4):828–832, October 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lewis:1978:OGC

[Lew78] William M. Lewis. Observations on the grass carp in ponds containing fingerling channel catfish and hybrid sunfish. Transactions of the American Fisheries Society, 107(1):153–155, January 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lewis:1971:SFH

[LH71] William M. Lewis and Roy Heidinger. Supplemental feeding of hybrid sunfish populations. Transactions of the American Fisheries Society, 100(4):619–623, October 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Locke:1972:EDU

[LH72] David O. Locke and Keith Havey. Effects of DDT upon salmon from Schoodic Lake, Maine. Transactions of the American Fisheries Society, 101(4):638–643, October 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lewis:1978:SE

[LHA78] William M. Lewis, Roy C. Heidinger, and Richard O. Anderson. Summary and evaluation. Transactions of the American Fisheries Society, 107(1):223–224, January 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Laurence:1979:ECM

[LHBS79] Geoffrey C. Laurence, Thomas A. Halavik, Bruce R. Burns, and Alphonse S. Smigielski. An environmental chamber for monitoring “In Situ” growth and survival of larval fishes. Transactions of the American Fisheries Society, 108(2):197–203, March 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lewis:1974:FIL

[LHK^{+}74] William M. Lewis, Roy Heidinger, William Kirk, Wayne Chapman, and Danny Johnson. Food intake of the largemouth bass. Transactions of the American Fisheries Society, 103(2):277–280, April 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Larson:1977:LDA

[LHL77] G. L. Larson, F. E. Hutchins, and L. P. Lamperti. Laboratory determination of acute and sublethal toxicities of inorganic chloramines to early life stages of Coho salmon (Oncorhynchus kisutch). Transactions of the American Fisheries Society, 106(3):268–277, May 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lindenberg:1976:SDD

[Lin76] John G. Lindenberg. Seasonal depth distribution of landlocked alewives, Alosa pseudoharengus (Wilson), in a shallow, eutrophic lake. Transactions of the American Fisheries Society, 105(3):395– 399, May 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Liscom:1971:OPG

[Lis71] Kenneth L. Liscom. Orifice placement in gatewells of turbine intakes for bypassing juvenile fish around dams. Transactions of the American Fisheries Society, 100(2):319–324, April 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lorio:1976:DDF

[LJH76] Wendell J. Lorio, James H. Jenkins, and Melvin T. Huish. Deposition of Dieldrin in four components of two artificial aquatic systems and a farm pond. Transactions of the American Fisheries Society, 105(6):695–699, November 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lane:1979:DFF

[LKT79] E. D. Lane, M. C. S. Kingsley, and D. E. Thornton. Daily feeding and food conversion efficiency of the diamond turbot: an analysis based on field data. Transactions of the American Fisheries Society, 108(6):530–535, November 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lane:1970:SAC

[LL70] Charles E. Lane and Robert J. Livingston. Some acute and chronic effects of Dieldrin on the sailfin molly, Poecilia latipinna. Transactions of the American Fisheries Society, 99(3):489–495, July 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lewis:1971:EZC

[LL71] Sue D. Lewis and William M. Lewis. The effect of zinc and copper on the osmolality of blood serum of the channel catfish, Ictalurus punctatus Rafinesque, and golden shiner, Notemigonus crysoleucas Mitchill. Transactions of the American Fisheries Society, 100(4): 639–643, October 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Leary:1975:VFP

[LL75a] Daniel F. Leary and Richard T. Lovell. Value of fiber in productiontype diets for channel catfish. Transactions of the American Fisheries Society, 104(2):328–332, April 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ludwig:1975:RLA

[LL75b] Gerald M. Ludwig and Eugene L. Lange. The relationship of length, age, and age–length interaction to the fecundity of the northern mottled sculpin, Cottus b. bairdi. Transactions of the American Fisheries Society, 104(1):64–67, January 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Loesch:1977:CLH

[LL77a] Joseph G. Loesch and William A. Lund, Jr. A contribution to the life history of the blueback herring, Alosa aestivalis. Transactions of the American Fisheries Society, 106(6):583–589, November 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Loftus:1977:FHS

[LL77b] William F. Loftus and Herbert L. Lenon. Food habits of the salmon smolts, Oncorhynchus tshawytscha and O. keta, from the Salcha River, Alaska. Transactions of the American Fisheries Society, 106(3):235–240, May 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lovell:1978:EVD

[LL78a] R. T. Lovell and Yin-Pen Li. Essentiality of vitamin D in diets of channel catfish (Ictalurus punctatus). Transactions of the American Fisheries Society, 107(6):809–811, November 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lovell:1978:VCP

[LL78b] R. T. Lovell and C. Lim. Vitamin C in pond diets for channel catfish. Transactions of the American Fisheries Society, 107(2): 321–325, March 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lund:1970:MMB

[LM70] William A. Lund, Jr. and George C. Maltezos. Movements and migrations of the bluefish, Pomatomus saltatrix, tagged in waters of New York and Southern New England. Transactions of the American Fisheries Society, 99(4):719–725, October 1970. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lackey:1971:USF

[LM71a] Robert T. Lackey and Bruce E. May. Use of sugar flotation and dye to sort benthic samples. Transactions of the American Fisheries Society, 100(4):794–797, October 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lewis:1971:OAL

[LM71b] Robert M. Lewis and Walter C. Mann. Occurrence and abundance of larval Atlantic menhaden, Brevoortia tyrannus, at two North Carolina inlets with notes on associated species. Transactions of the American Fisheries Society, 100(2):296–301, April 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lotrich:1974:TEV

[LM74] Victor A. Lotrich and William H. Meredith. A technique and the effectiveness of various acrylic colors for subcutaneous marking of fish. Transactions of the American Fisheries Society, 103(1):140– 142, January 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Leary:1975:SMT

[LM75a] Daniel F. Leary and Garth I. Murphy. A successful method for tagging the small, fragile engraulid, Stolephorus purpureus. Transactions of the American Fisheries Society, 104(1):53–55, January 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lightner:1975:TTS

[LM75b] D. V. Lightner and D. A. Meineke. A thyroid tumor in a sheepshead minnow (Cyprinodon variegatus) from the Gulf of Mexico. Transactions of the American Fisheries Society, 104(1):138–139, January 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lindsay:1976:RPI

[LM76] John A. Lindsay and R. Lynn Moran. Relationships of parasitic isopods Lironeca ovalis and Olencira praegustator to marine fish hosts in Delaware Bay. Transactions of the American Fisheries Society, 105(2):327–332, March 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Li:1976:EIM

[LMG76] Hiram W. Li, Peter B. Moyle, and Ronald L. Garrett. Effect of the introduction of the Mississippi silverside (Menidia audens) on the growth of black crappie (Pomoxis nigromaculatus) and white crappie (P. annularis) in Clear Lake, California. Transactions of the American Fisheries Society, 105(3):404–408, May 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lovell:1978:DPR

[Lov78] R. T. Lovell. Dietary phosphorus requirement of channel catfish (Ictalurus punctatus). Transactions of the American Fisheries Society, 107(4):617–621, July 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lux:1970:GVF

[LPH70] F. E. Lux, A. E. Peterson, Jr., and R. F. Hutton. Geographical variation in fin ray number in winter flounder, Pseudopleuronectes americanus (Walbaum), off Massachusetts. Transactions of the American Fisheries Society, 99(3):483–488, July 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lembi:1978:EVR

[LRIF78] Carole A. Lembi, Brian G. Ritenour, Eric M. Iverson, and Eric C. Forss. The effects of vegetation removal by grass carp on water chemistry and phytoplankton in Indiana ponds. Transactions of the American Fisheries Society, 107(1):161–171, January 1978. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Landry:1973:ACS

[LS73a] Andre M. Landry, Jr. and Kirk Strawn. Annual cycle of sportfishing activity at a warmwater discharge into Galveston Bay, Texas. Transactions of the American Fisheries Society, 102(3):573–577, July 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lovell:1973:ACC

[LS73b] Richard T. Lovell and Lewis A. Sackey. Absorption by channel catfish of earthy-musty flavor compounds synthesized by cultures of blue-green algae. Transactions of the American Fisheries Society, 102(4):774–777, October 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lincer:1970:DEF

[LSN70] Jeffrey L. Lincer, Joseph M. Solon, and John H. Nair III. DDT and endrin fish toxicity under static versus dynamic bioassay conditions. Transactions of the American Fisheries Society, 99(1): 13–19, January 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lux:1973:WSY

[Lux73] Fred E. Lux. White spotting in the 1959 year-class of Georges Bank winter flounder, Pseudopleuronectes americanus (Walbaum). Transactions of the American Fisheries Society, 102(1):83–88, January 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Laarman:1976:RMM

[LWO76] Percy W. Laarman, Wayne A. Willford, and James R. Olson. Retention of mercury in the muscle of yellow perch (Perca flavescens) and rock bass (Ambloplites rupestris). Transactions of the American Fisheries Society, 105(2):296–300, March 1976. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lewis:1978:UHM

[LYSB78] William M. Lewis, John H. Yopp, Harold L. Schramm, Jr., and Alan M. Brandenburg. Use of hydroponics to maintain quality of recirculated water in a fish culture system. Transactions of the American Fisheries Society, 107(1):92–99, January 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Murai:1975:PAS

[MA75] Takeshi Murai and James W. Andrews. Pantothenic acid supplementation of diets for catfish fry. Transactions of the American Fisheries Society, 104(2):313–316, April 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McIntyre:1978:HTI

[MA78] John D. McIntyre and Donald F. Amend. Heritability of tolerance for infectious hematopoietic necrosis in sockeye salmon (Oncorhynchus nerka). Transactions of the American Fisheries Society, 107(2):305–308, March 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mackenzie:1977:PHC

[Mac77] Clyde L. Mackenzie, Jr. Predation on hard clam (Mercenaria mercenaria) populations. Transactions of the American Fisheries Society, 106(6):530–537, November 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Madsen:1971:PNT

[Mad71] Monte L. Madsen. The presence of nuptial tubercles on female quillback (Carpiodes cyprinus). Transactions of the American Fisheries Society, 100(1):132–134, January 1971. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Magnuson:1973:ATR

[Mag73] John J. Magnuson. Application of theory and research in freshwater and marine fisheries management programs — a discussion of the papers given by S. H. Smith, R. O. Anderson, K. R. Allen, and G. B. Talbot. Transactions of the American Fisheries Society, 102(1): 194–195, January 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See [All73, And73, Smi73b, Tal73, Str73].

Magnuson:1976:MEG

[Mag76] John J. Magnuson. Managing with exotics — a game of chance. Transactions of the American Fisheries Society, 105(1):1–9, January 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Manion:1972:FSL

[Man72a] Patrick J. Manion. Fecundity of the sea lamprey (Petromyzon marinus) in Lake Superior. Transactions of the American Fisheries Society, 101(4):718–720, October 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Manion:1972:VMA

[Man72b] Patrick J. Manion. Variations in melanophores among lampreys in the Upper Great Lakes. Transactions of the American Fisheries Society, 101(4):662–666, October 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Manning:1979:BCF

[Man79] Robert E. Manning. Behavioral characteristics of fishermen and other recreationists on four Vermont rivers. Transactions of the American Fisheries Society, 108(6):536–541, November 1979. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marking:1970:JHN

[Mar70] Leif L. Marking. Juglone (5-hydroxy-1,4-naphthoquinone) as a fish toxicant. Transactions of the American Fisheries Society, 99(3): 510–514, July 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marking:1972:SEN

[Mar72a] Leif L. Marking. Salicylanilide i, an effective non-persistent candidate piscicide. Transactions of the American Fisheries Society, 101 (3):526–533, July 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Martin:1972:FIL

[Mar72b] F. Douglas Martin. Factors influencing local distribution of Cyprinodon variegatus (Pisces: Cyprinodontidae). Transactions of the American Fisheries Society, 101(1):89–93, January 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marking:1974:TDE

[Mar74] Leif L. Marking. Toxicity of 2-(digeranylamino)-ethanol, a candidate selective fish toxicant. Transactions of the American Fisheries Society, 103(4):736–742, October 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mathur:1973:FHF

[Mat73] Dilip Mathur. Food habits and feeding chronology of the blackbanded darter, Percina nigrofasciata (Agassiz), in Halawakee Creek, Alabama. Transactions of the American Fisheries Society, 102(1):48–55, January 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Matuszek:1978:EPF

[Mat78] John E. Matuszek. Empirical predictions of fish yields of large North American lakes. Transactions of the American Fisheries Society, 107(3):385–394, May 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marking:1975:TPP

[MB75a] Leif L. Marking and Terry D. Bills. Toxicity of potassium permanganate to fish and its effectiveness for detoxifying antimycin. Transactions of the American Fisheries Society, 104(3):579–583, July 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moore:1975:MSB

[MB75b] Charles J. Moore and Dennis T. Burton. Movements of striped bass, Morone saxatilis, tagged in Maryland waters of Chesapeake Bay. Transactions of the American Fisheries Society, 104(4):703– 709, October 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Merna:1972:EMF

[MBN72] James W. Merna, Michael E. Bender, and James R. Novy. The effects of methoxychlor on fishes I. Acute toxicity and breakdown studies. Transactions of the American Fisheries Society, 101(2): 298–301, April 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moe:1970:RTG

[MBT70] Martin A. Moe, Jr., Dale S. Beaumariage, and Robert W. Topp. Return of tagged gag, Mycteroperca microlepis, and Caribbean red snapper, Lutjanus campechanus, after six years of freedom. Transactions of the American Fisheries Society, 99(2):428–429, April 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Murawski:1978:PDA

[MC78] Steven A. Murawski and Charles F. Cole. Population dynamics of anadromous rainbow smelt, Osmerus mordax, in a Massachusetts river system. Transactions of the American Fisheries Society, 107 (4):535–542, July 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McNeil:1979:FCE

[MC79] Fergus I. McNeil and E. J. Crossman. Fin clips in the evaluation of stocking programs for muskellunge, Esox masquinongy. Transactions of the American Fisheries Society, 108(4):335–343, July 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mcilwain:1970:SCL

[Mci70] T. D. Mcilwain. Stomach contents and length–weight relationships of chain pickerel (Esox niger) in South Mississippi waters. Transactions of the American Fisheries Society, 99(2):439–440, April 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mceachran:1970:AGS

[MD70] John D. Mceachran and Jackson Davis. Age and growth of the striped searobin. Transactions of the American Fisheries Society, 99(2):343–352, April 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Manzi:1971:SLP

[MD71] John J. Manzi and Kathleen A. Donnelly. Staining large populations of bivalve larvae. Transactions of the American Fisheries Society, 100(3):588–590, July 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marking:1972:HLB

[MD72] Leif L. Marking and Verdel K. Dawson. The half-life of biological activity of antimycin determined by fish bioassay. Transactions of the American Fisheries Society, 101(1):100–105, January 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Malvestuto:1978:ERC

[MDS78] Stephen P. Malvestuto, William D. Davies, and William L. Shelton. An evaluation of the roving creel survey with nonuniform probability sampling. Transactions of the American Fisheries Society, 107(2):255–262, March 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Malvestuto:1979:PPC

[MDS79] Stephen P. Malvestuto, William D. Davies, and William L. Shelton. Predicting the precision of creel survey estimates of fishing effort by use of climatic variables. Transactions of the American Fisheries Society, 108(1):43–45, January 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meacham:1971:CFS

[Mea71] Charles H. Meacham. The complete fishery scientist. Transactions of the American Fisheries Society, 100(2):385–388, April 1971. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marullo:1976:VST

[MECC76] Frank Marullo, Dennis A. Emliani, Charles W. Caillouet, and Stephen H. Clark. A vinyl streamer tag for shrimp (Penaeus spp.). Transactions of the American Fisheries Society, 105(6):658–663, November 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Medvick:1979:GRJ

[Med79] Patricia A. Medvick. Growth rates of juvenile maomao, abudefduf abdominalis, at constant and cyclic temperatures. Transactions of the American Fisheries Society, 108(3):293–298, May 1979. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Melack:1976:PPF

[Mel76] John M. Melack. Primary productivity and fish yields in tropical lakes. Transactions of the American Fisheries Society, 105(5):575– 580, September 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Menzel:1974:MCT

[Men74] Bruce W. Menzel. A mouthless carp from Texas. Transactions of the American Fisheries Society, 103(1):142–143, January 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mense:1975:RDB

[Men75] James B. Mense. Relation of density to brown trout movement in a Michigan stream. Transactions of the American Fisheries Society, 104(4):688–695, October 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Merriner:1971:DIC

[Mer71a] John V. Merriner. Development of intergeneric centrarchid hybrid embryos. Transactions of the American Fisheries Society, 100(4): 611–618, October 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Merriner:1971:ESF

[Mer71b] John V. Merriner. Egg size as a factor in intergeneric hybrid success of centrarchids. Transactions of the American Fisheries Society, 100(1):29–32, January 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCauley:1977:IAT

[MER77] R. W. McCauley, J. R. Elliott, and L. A. A. Read. Influence of acclimation temperature on preferred temperature in the rainbow trout Salmo gairdneri. Transactions of the American Fisheries Society, 106(4):362–365, July 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Messieh:1975:GOY

[Mes75] S. N. Messieh. Growth of the otoliths of young herring in the Bay of Fundy. Transactions of the American Fisheries Society, 104(4): 770–772, October 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Messieh:1976:FSA

[Mes76] S. N. Messieh. Fecundity studies on Atlantic herring from the Southern Gulf of St. Lawrence and along the Nova Scotia Coast. Transactions of the American Fisheries Society, 105(3):384–394, May 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCleave:1975:NCF

[MF75] James D. McCleave and Stephen M. Fried. Nighttime catches of fishes in a tidal cove in Montsweag Bay near Wiscasset, Maine. Transactions of the American Fisheries Society, 104(1):30–34, January 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCarraher:1970:ACS

[MG70] D. B. McCarraher and Richard W. Gregory. Adaptability and current status of introductions of Sacramento perch, Archoplites interruptus, in North America. Transactions of the American Fisheries Society, 99(4):700–707, October 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mcswain:1972:SBS

[MG72a] Lawrence E. Mcswain and Richard M. Gennings. Spawning behavior of the spotted sucker Minytrema melanops (Rafinesque). Transactions of the American Fisheries Society, 101(4):738–740, October 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Menzel:1972:CMC

[MG72b] Bruce W. Menzel and David M. Green, Jr. A color mutant of the chain pickerel, Esox niger LeSueur. Transactions of the American Fisheries Society, 101(2):370–372, April 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moore:1976:ORS

[MGW76] Richard H. Moore, Robert A. Garrett, and Paul J. Wingate. Occurrence of the red shiner, Notropis lutrensis, in North Carolina: a probable aquarium release. Transactions of the American Fisheries Society, 105(2):220–221, March 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marnell:1970:HML

[MH70] Leo F. Marnell and Don Hunsaker II. Hooking mortality of lurecaught cutthroat trout (Salmo clarki) in relation to water temperature, fatigue, and reproductive maturity of released fish. Transactions of the American Fisheries Society, 99(4):684–688, October 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mears:1976:OSF

[MH76] Harold C. Mears and Richard W. Hatch. Overwinter survival of fingerling brook trout with single and multile fin clips. Transactions of the American Fisheries Society, 105(6):669–674, November 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Manooch:1977:AGM

[MH77] Charles S. Manooch III and Gene R. Huntsman. Age, growth, and mortality of the red porgy, Pagrus pagrus. Transactions of the American Fisheries Society, 106(1):26–33, January 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Manooch:1978:AGG

[MH78a] Charles S. Manooch III and Manuel Haimovici. Age and growth of the gag, Mycteroperca microlepis, and size-age composition of the recreational catch off the Southeastern United States. Transactions of the American Fisheries Society, 107(2):234–240, March 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moyle:1978:EIM

[MH78b] Peter B. Moyle and Nancy J. Holzhauser. Effects of the introduction of Mississippi silverside (Menidia audens) and Florida largemouth bass (Micropterus salmoides floridanus) on the feeding habits of young-of-year largemouth bass in Clear Lake, California. Transactions of the American Fisheries Society, 107(4):574–582, July 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mathur:1977:IFP

[MHM77] Dilip Mathur, Paul G. Heisey, and Nancy C. Magnusson. Impingement of fishes at Peach Bottom Atomic Power Station, Pennsylvania. Transactions of the American Fisheries Society, 106(3):258– 267, May 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miller:1972:TFF

[Mil72] Robert Rush Miller. Threatened freshwater fishes of the United States. Transactions of the American Fisheries Society, 101(2): 239–252, April 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miller:1974:FBT

[Mil74] John M. Miller. The food of brook trout Salvelinus fontinalis (Mitchill) fry from different subsections of Lawrence Creek, Wisconsin. Transactions of the American Fisheries Society, 103(1): 130–134, January 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mitzner:1978:EBC

[Mit78] Larry Mitzner. Evaluation of biological control of nuisance aquatic vegetation by grass carp. Transactions of the American Fisheries Society, 107(1):135–145, January 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mccann:1972:VDB

[MJ72] John A. Mccann and Robert L. Jasper. Vertebral damage to bluegills exposed to acutely toxic levels of pesticides. Transactions of the American Fisheries Society, 101(2):317–322, April 1972. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McConnell:1976:MPR

[MJ76] William J. McConnell and U. Thomas Jackson. A method for precalibrating re-oxygenation rates in experimental containers. Transactions of the American Fisheries Society, 105(2):343–346, March 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marcy:1972:ORY

[MJN72] Barton C. Marcy, Jr., Paul M. Jacobson, and Richard L. Nankee. Observations on the reactions of young American shad to a heated effluent. Transactions of the American Fisheries Society, 101(4): 740–743, October 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Minckley:1970:FBG

[MJRW70] W. L. Minckley, James E. Johnson, John N. Rinne, and Sandra E. Willoughby. Foods of buffalofishes, genus Ictiobus, in Central Arizona reservoirs. Transactions of the American Fisheries Society, 99(2):333–342, April 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Morgan:1973:EDP

[MKK73] Raymond P. Morgan II, Ted S. Y. Koo, and George E. Krantz. Electrophoretic determination of populations of the striped bass, Morone saxatilis, in the Upper Chesapeake Bay. Transactions of the American Fisheries Society, 102(1):21–32, January 1973. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marquette:1971:LSS

[ML71] Willman M. Marquette and Clifford W. Long. Laboratory studies of screens for diverting juvenile salmon and trout from turbine intakes. Transactions of the American Fisheries Society, 100(3): 439–447, July 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mccleave:1972:FUT

[ML72] James D. Mccleave and George W. Labar. Further ultrasonic tracking and tagging studies of homing cutthroat trout (Salmo clarki) in Yellowstone Lake. Transactions of the American Fisheries Society, 101(1):44–54, January 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCleave:1977:WSH

[MLK77] James D. McCleave, George W. Labar, and Frederick W. Kircheis. Within-season homing movements of displaced mature Sunapee trout (Salvelinus alpinus) in Floods Pond, Maine. Transactions of the American Fisheries Society, 106(2):156–162, March 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Maxfield:1971:SGF

[MLL71] Galen H. Maxfield, Robert H. Lander, and Kenneth L. Liscom. Survival, growth, and fecundity of hatchery-reared rainbow trout after exposure to pulsating direct current. Transactions of the American Fisheries Society, 100(3):546–552, July 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miles:1974:PRH

[MLMS74] Harry M. Miles, Susan M. Loehner, David T. Michaud, and Sandra L. Salivar. Physiological responses of hatchery reared muskellunge (Esox masquinongy) to handling. Transactions of the American Fisheries Society, 103(2):336–342, April 1974. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McConnell:1977:GPE

[MLO77] William J. McConnell, Steven Lewis, and James E. Olson. Gross photosynthesis as an estimator of potential fish production. Transactions of the American Fisheries Society, 106(5):417–423, September 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCosker:1976:ULV

[MLT76] John E. McCosker, Michael D. Lagios, and Thomas Tucker. Ultrastructure of lymphocystis virus in the quillback rockfish, Sebastes maliger, with records of infection in other aquarium-held fishes. Transactions of the American Fisheries Society, 105(2):333–337, March 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Macek:1970:ISS

[MM70] Kenneth J. Macek and William A. Mcallister. Insecticide susceptibility of some common fish family representatives. Transactions of the American Fisheries Society, 99(1):20–27, January 1970. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meyer:1976:REC

[MM76] Henry L. Meyer and John V. Merriner. Retention and escapement characteristics of pound nets as a function of pound-head mesh size. Transactions of the American Fisheries Society, 105(3):370–379, May 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Musick:1977:SDB

[MM77] John A. Musick and Linda P. Mercer. Seasonal distribution of black sea bass, Centropristis striata, in the Mid-Atlantic Bight with comments on the ecology and fisheries of the species. Transactions of the American Fisheries Society, 106(1):12–25, January 1977. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moyle:1974:FHS

[MMB74] Peter B. Moyle, Stephen B. Mathews, and Noel Bonderson. Feeding habits of the Sacramento perch, Archoplites interruptus. Transactions of the American Fisheries Society, 103(2):399–402, April 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mayer:1978:ITV

[MMC78] Foster L. Mayer, Paul M. Mehrle, and P. Logan Crutcher. Interactions of toxaphene and vitamin C in channel catfish. Transactions of the American Fisheries Society, 107(2):326–333, March 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mahoney:1973:FRD

[MMD73] John B. Mahoney, Frederick H. Midlige, and David G. Deuel. A fin rot disease of marine and euryhaline fishes in the New York bight. Transactions of the American Fisheries Society, 102(3):596–605, July 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mazeaud:1977:PSE

[MMD77] Madeleine M. Mazeaud, Frederic Mazeaud, and Edward M. Donaldson. Primary and secondary effects of stress in fish: Some new data with a general review. Transactions of the American Fisheries Society, 106(3):201–212, May 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mcallister:1972:SDM

[MMM72] William A. Mcallister, Jr., Wilbur L. Mauck, and Foster L. Mayer, Jr. A simplified device for metering chemicals in intermittent-flow bioassays. Transactions of the American Fisheries Society, 101(3): 555–557, July 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mathur:1979:VFW

[MMN79] Dilip Mathur, Pauline L. McCreight, and George A. Nardacci. Variations in fecundity of white crappie in Conowingo Pond, Pennsylvania. Transactions of the American Fisheries Society, 108(6):548– 554, November 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Matlock:1975:SLT

[MMS75] Gary C. Matlock, Rocco A. Marcello, Jr., and Kirk Strawn. Standard length-total length relationships of Gulf menhaden, Brevoortia patronus Goode, bay anchovy, Anchoa mitchilli (Valenciennes), and Atlantic croaker, Micropogon undulatus (Linnaeus), from Galveston Bay. Transactions of the American Fisheries Society, 104(2):408–409, April 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miller:1979:RCS

[MNL79] Richard A. Miller, Logan A. Norris, and Bobby R. Loper. The response of Coho salmon and guppies to 2,3,7,8-tetrachlorodibenzo-p-dioxin (TCDD) in water. Transactions of the American Fisheries Society, 108(4):401–407, July 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Momot:1978:APP

[Mom78] Walter T. Momot. Annual production and production/biomass ratios of the crayfish, Orconectes virilis, in two Northern Ontario lakes. Transactions of the American Fisheries Society, 107(6):776– 784, November 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moring:1970:UAQ

[Mor70a] John R. Moring. Use of the anesthetic quinaldine for handling Pacific Coast intertidal fishes. Transactions of the American Fisheries Society, 99(4):802–805, October 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Morrison:1970:NST

[Mor70b] William J. Morrison. Nonrandom segregation of two lactate dehydrogenase subunit loci in trout. Transactions of the American Fisheries Society, 99(1):193–206, January 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Morrissey:1971:MTA

[Mor71] Thomas D. Morrissey. Movements of tagged American lobsters, Homarus americanus, liberated off Cape Cod, Massachusetts. Transactions of the American Fisheries Society, 100(1):117–120, January 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moss:1970:RYA

[Mos70] Sanford A. Moss. The responses of young American shad to rapid temperature changes. Transactions of the American Fisheries Society, 99(2):381–384, April 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moulton:1974:AGS

[Mou74] Lawrence L. Moulton. Abundance, growth, and spawning of the longfin smelt in Lake Washington. Transactions of the American Fisheries Society, 103(1):46–52, January 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moyle:1973:ESA

[Moy73] Peter B. Moyle. Ecological segregation among three species of minnows (Cyprinidae) in a Minnesota lake. Transactions of the American Fisheries Society, 102(4):794–805, October 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miller:1971:MOP

[MP71] Robert Rush Miller and Edwin P. Pister. Management of the owens pupfish, Cyprinodon radiosus, in Mono County, California. Transactions of the American Fisheries Society, 100(3):502–509, July 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McNeely:1974:DCF

[MP74] David L. McNeely and William D. Pearson. Distribution and condition of fishes in a small reservoir receiving heated waters. Transactions of the American Fisheries Society, 103(3):518–530, July 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McNeely:1975:OHF

[MP75] David L. McNeely and William D. Pearson. One Hundred-Fourth Annual Meeting of the American Fisheries Society: Sheraton-Waikiki Hotel, Honolulu, Hawaii, September 9–11, 1974. Transactions of the American Fisheries Society, 104(1):179–210, January 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Minckley:1976:UGW

[MP76] C. O. Minckley and Larry J. Paulson. Use of gizzard weights to determine total length and weight of threadfin shad eaten by predators. Transactions of the American Fisheries Society, 105(3): 409–410, May 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Morgan:1977:CTE

[MP77] Raymond P. Morgan II and Robert D. Prince. Chlorine toxicity to eggs and larvae of five Chesapeake Bay fishes. Transactions of the American Fisheries Society, 106(4):380–385, July 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Morgan:1978:CEL

[MP78] Raymond P. Morgan II and Robert D. Prince. Chlorine effects on larval development of striped bass (Morone saxatilis), white perch (M. americana) and blueback herring (Alosa aestivalis). Transactions of the American Fisheries Society, 107(4):636–641, July 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marcy:1974:AGW

[MR74a] Barton C. Marcy, Jr. and F. Paul Richards. Age and growth of the white perch Morone americana in the Lower Connecticut River.

Transactions of the American Fisheries Society, 103(1):117–120, January 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mathur:1974:RBR

[MR74b] Dilip Mathur and John S. Ramsey. Reproductive biology of the rough shiner, Notropis baileyi, in Halawakee Creek, Alabama. Transactions of the American Fisheries Society, 103(1):88–93, January 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Maccrimmon:1979:SJA

[MR79] Hugh R. Maccrimmon and William H. Robbins. Suitability of jaw and Atkins streamer tags for marking smallmouth bass, Micropterus dolomieui. Transactions of the American Fisheries Society, 108(5):499–501, September 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moles:1979:SAF

[MRK79] Adam Moles, Stanley D. Rice, and Sid Korn. Sensitivity of Alaskan freshwater and anadromous fishes to Prudhoe Bay crude oil and benzene. Transactions of the American Fisheries Society, 108(4): 408–414, July 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Macek:1970:UDE

[MRSK70] Kenneth J. Macek, Charles R. Rodgers, David L. Stalling, and Sidney Korn. The uptake, distribution and elimination of dietary 14C-DDT and ^{14}C-Dieldrin in rainbow trout. Transactions of the American Fisheries Society, 99(4):689–695, October 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Macek:1970:BVS

[MS70a] Kenneth J. Macek and Herman O. Sanders. Biological variation in the susceptibility of fish and aquatic invertebrates to DDT. Transactions of the American Fisheries Society, 99(1):89–90, January 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mathur:1970:IAM

[MS70b] G. B. Mathur and B. D. Shrivastava. An improved activity meter for the determination of standard metabolism in fish. Transactions of the American Fisheries Society, 99(3):602–603, July 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Martin:1971:OJS

[MS71] James R. Martin and Robert L. Shipp. Occurrence of juvenile snook, Centropomus undecimalis, in North Carolina waters. Transactions of the American Fisheries Society, 100(1):131–132, January 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McKnight:1974:FHC

[MS74a] Terrence C. McKnight and Steven L. Serns. Food habits of Coho salmon (Oncorhynchus kisutch) in an inland Wisconsin lake. Transactions of the American Fisheries Society, 103(1):126–130, January 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mearns:1974:EAF

[MS74b] Alan J. Mearns and Marjorie Sherwood. Environmental aspects of fin erosion and tumors in Southern California Dover sole. Transactions of the American Fisheries Society, 103(4):799–810, October 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Modde:1977:FFB

[MS77] Timothy Modde and James C. Schmulbach. Food and feeding behavior of the shovelnose sturgeon, Scaphirhynchus platorynchus, in the unchannelized Missouri River, South Dakota. Transactions of the American Fisheries Society, 106(6):602–608, November 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McNew:1978:EML

[MS78] Ronald W. McNew and Robert C. Summerfelt. Evaluation of a maximum-likelihood estimator for analysis of length-frequency distributions. Transactions of the American Fisheries Society, 107 (5):730–736, September 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Maki:1973:EDR

[MSS73] Alan W. Maki, Kenneth W. Stewart, and J. K. G. Silvey. The effects of dibrom on respiratory activity of the stonefly, Hydroperla crosbyi, Hellgrammite, Corydalus cornutus and the golden shiner, Notemigonus crysoleucas. Transactions of the American Fisheries Society, 102(4):806–815, October 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

May:1979:JSM

[MSW79] Bernie May, Mark Stoneking, and James E. Wright, Jr. Joint segregation of malate dehydrogenase and diaphorase loci in brown trout (Salmo trutta). Transactions of the American Fisheries Society, 108(4):373–377, July 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mccarraher:1972:ESV

[MT72] D. B. Mccarraher and Robert E. Thomas. Ecological significance of vegetation to northern pike, Esox lucius, spawning. Transactions of the American Fisheries Society, 101(3):560–563, July 1972. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mccaskill:1972:FNS

[MTM72] Matt L. Mccaskill, Jamie E. Thomerson, and Paul R. Mills. Food of the northern studfish, Fundulus catenatus, in the Missouri Ozarks. Transactions of the American Fisheries Society, 101(2):375–377, April 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Munther:1970:MDS

[Mun70] Gregory Lennart Munther. Movement and distribution of smallmouth bass in the Middle Snake River. Transactions of the American Fisheries Society, 99(1):44–53, January 1970. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Morgan:1976:ESE

[MUR^{+}76] Raymond P. Morgan II, Robert E. Ulanowicz, V. James Rasin, Jr., Linda A. Noe, and G. Brooke Gray. Effects of shear on eggs and larvae of striped bass, Morone saxatilis, and white perch, M. americana. Transactions of the American Fisheries Society, 105 (1):149–154, January 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mustafa:1977:IMC

[Mus77] Saleem Mustafa. Influence of maturation on the concentrations of RNA and DNA in the flesh of the catfish Clarias batrachus. Transactions of the American Fisheries Society, 106(5):449–451, September 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McKernan:1970:ASI

[MV70] Donald L. McKernan and Wilvan Van Campen. An approach to solving international fishery disputes. Transactions of the American Fisheries Society, 99(3):621–628, July 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Manner:1977:UCF

[MVM77] Harold W. Manner, Mark Vancura, and Carol Muehleman. The ultrastructure of the chorion of the fathead minnow, Pimephales promelas. Transactions of the American Fisheries Society, 106(1): 110–114, January 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miller:1971:ETF

[MW71] Bruce S. Miller and S. R. Wellings. Epizootiology of tumors on flathead sole (Hippoglossoides elassodon) in East Sound, Orcas Island, Washington. Transactions of the American Fisheries Society, 100(2):247–266, April 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mcilwain:1972:XGL

[MW72] T. D. Mcilwain and Richard Waller. A xanthochroic gar, Lepisosteus oculatus, from Mississippi. Transactions of the American Fisheries Society, 101(2):362, April 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Murawski:1979:PAB

[MW79] Steven A. Murawski and Gordon T. Waring. A population assessment of butterfish, Peprilus triacanthus, in the Northwestern Atlantic Ocean. Transactions of the American Fisheries Society, 108 (5):427–439, September 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Macek:1972:TID

[MWHH72] Kenneth J. Macek, David F. Walsh, James W. Hogan, and Delmar D. Holz. Toxicity of the insecticide Dursban(R) to fish and aquatic invertebrates in ponds. Transactions of the American Fisheries Society, 101(3):420–427, July 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Maurer:1971:DOI

[MWK71] Don Maurer, Les Watling, and Richard Keck. The Delaware oyster industry: a reality? Transactions of the American Fisheries Society, 100(1):100–111, January 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mcmullen:1971:DME

[MY71] John C. Mcmullen and Harvey T. Yoshihara. Deposition of mature eggs in unmated king crabs, Paralithodes camtschatica (Tilesius). Transactions of the American Fisheries Society, 100(3):583–584, July 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Naiman:1975:FHA

[Nai75] Robert J. Naiman. Food habits of the Amargosa pupfish in a thermal stream. Transactions of the American Fisheries Society, 104(3):536–538, July 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nimmo:1972:EDC

[NB72] D. R. Nimmo and Robbin R. Blackman. Effects of DDT on cations in the hepatopancreas of penaeid shrimp. Transactions of the American Fisheries Society, 101(3):547–549, July 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nebeker:1976:EAS

[NB76] Alan V. Nebeker and John R. Brett. Effects of air-supersaturated water on survival of Pacific salmon and steelhead smolts. Transactions of the American Fisheries Society, 105(2):338–342, March 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nebeker:1976:CDO

[NBS76] Alan V. Nebeker, Gerald R. Bouck, and Donald G. Stevens. Carbon dioxide and oxygen-nitrogen ratios as factors affecting salmon survival in air-supersaturated water. Transactions of the American Fisheries Society, 105(3):425–429, May 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nicola:1973:EFR

[NC73] Stephen J. Nicola and Almo J. Cordone. Effects of fin removal on survival and growth of rainbow trout (Salmo gairdneri) in a natural environment. Transactions of the American Fisheries Society, 102 (4):753–758, October 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Norton:1976:TSC

[NDN76] V. M. Norton, K. B. Davis, and H. Nishimura. A technique for sexing channel catfish. Transactions of the American Fisheries Society, 105(3):460–462, May 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Novak:1974:SFH

[NE74] John K. Novak and Ray D. Estes. Summer food habits of the black sculpin, Cottus baileyi, in the Upper South Fork Holston River Drainage. Transactions of the American Fisheries Society, 103(2):270–276, April 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nebeker:1972:ELO

[Neb72] Alan V. Nebeker. Effect of low oxygen concentration on survival and emergence of aquatic insects. Transactions of the American Fisheries Society, 101(4):675–679, October 1972. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Neves:1975:FAF

[Nev75a] Richard J. Neves. Factors affecting fry production of smallmouth bass (Micropterus dolomieui) in South Branch Lake, Maine. Transactions of the American Fisheries Society, 104(1):83–87, January 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Neves:1975:ZRL

[Nev75b] Richard J. Neves. Zooplankton recolonization of a lake cove treated with rotenone. Transactions of the American Fisheries Society, 104(2):390–393, April 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Naqvi:1970:LIR

[NF70] Syed M. Naqvi and Denzel E. Ferguson. Levels of insecticide resistance in fresh-water shrimp, Palaemonetes kadiakensis. Transactions of the American Fisheries Society, 99(4):696–699, October 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nicholson:1971:CMA

[Nic71] William R. Nicholson. Coastal movements of Atlantic menhaden as inferred from changes in age and length distributions. Transactions of the American Fisheries Society, 100(4):708–716, October 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Newton:1973:FWB

[NK73] Scott H. Newton and Raj V. Kilambi. Fecundity of the white bass, Morone chrysops (Rafinesque), in Beaver Reservoir, Arkansas. Transactions of the American Fisheries Society, 102(2):446–448, April 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Neill:1974:DEB

[NM74] William H. Neill and John J. Magnuson. Distributional ecology and behavioral thermoregulation of fishes in relation to heated effluent from a power plant at Lake Monona, Wisconsin. Transactions of the American Fisheries Society, 103(4):663–710, October 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nixon:1978:MGC

[NM78] David E. Nixon and Robert L. Miller. Movements of grass carp, Ctenopharyngodon idella, in an open reservoir system as determined by underwater telemetry. Transactions of the American Fisheries Society, 107(1):146–148, January 1978. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Noble:1972:MRW

[Nob72] Richard L. Noble. Mortality rates of walleye fry in a bay of Oneida Lake, New York. Transactions of the American Fisheries Society, 101(4):720–723, October 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Noble:1973:ERY

[Nob73] Richard L. Noble. Evacuation rates of young yellow perch, Perca flavescens (Mitchill). Transactions of the American Fisheries Society, 102(4):759–763, October 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Noble:1975:GYY

[Nob75] Richard L. Noble. Growth of young yellow perch (Perca flavescens) in relation to zooplankton populations. Transactions of the American Fisheries Society, 104(4):731–741, October 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Neumann:1975:RHR

[NOSW75] D. A. Neumann, J. M. O’Connor, J. A. Sherk, and K. V. Wood. Respiratory and hematological responses of oyster toadfish (Opsanus tau) to suspended solids. Transactions of the American Fisheries Society, 104(4):775–781, October 1975. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nebeker:1974:EPBb

[NP74] Alan V. Nebeker and Frank A. Puglisi. Effect of polychlorinated biphenyls (PCB’s) on survival and reproduction of Daphnia, Gammarus, and Tanytarsus. Transactions of the American Fisheries Society, 103(4):722–728, October 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nebeker:1974:EPBa

[NPD74] Alan V. Nebeker, Frank A. Puglisi, and David L. Defoe. Effect of polychlorinated biphenyl compounds on survival and reproduction of the fathead minnow and flagfish. Transactions of the American Fisheries Society, 103(3):562–568, July 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Norcross:1974:DYB

[NRMJ74] J. J. Norcross, S. L. Richardson, W. H. Massmann, and E. B. Joseph. Development of young bluefish (Pomatomus saltatrix) and distribution of eggs and young in Virginian coastal waters. Transactions of the American Fisheries Society, 103(3):477–497, July 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ney:1975:FYG

[NS75] John J. Ney and Lloyd L. Smith, Jr. First-year growth of the yellow perch, Perca flavescens, in the Red Lakes, Minnesota. Transactions of the American Fisheries Society, 104(4):718–725, October 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ney:1976:SPV

[NS76] John J. Ney and Lloyd L. Smith, Jr. Serum protein variability in geographically-defined bluegill (Lepomis macrochirus) populations. Transactions of the American Fisheries Society, 105(2):281–290, March 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nugent:1970:ETE

[Nug70] Richard Nugent. Elevated temperatures and electric power plants. Transactions of the American Fisheries Society, 99(4):848–849, October 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nyman:1970:EAH

[Nym70] O. L. Nyman. Electrophoretic analysis of hybrids between salmon (Salmo salar L.) and trout (Salmo trutta L.). Transactions of the American Fisheries Society, 99(1):229–236, January 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Olmsted:1974:RAF

[OC74] Larry L. Olmsted and Donald G. Cloutman. Repopulation after a fish kill in Mud Creek, Washington County, Arkansas following pesticide pollution. Transactions of the American Fisheries Society, 103(1):79–87, January 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

OConnor:1975:AAP

[OC75] J. F. O’Connor and Brian W. Coad. Anomalies of the adipose and pelvic fins of a brook trout, Salvelinus fontinalis. Transactions of the American Fisheries Society, 104(2):414, April 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Olmsted:1979:LHF

[OC79] Larry L. Olmsted and Donald G. Cloutman. Life history of the flat bullhead, Ictalurus platycephalus, in Lake Norman, North Carolina. Transactions of the American Fisheries Society, 108(1):38–42, January 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Odum:1970:IAE

[Odu70] William E. Odum. Insidious alteration of the estuarine environment. Transactions of the American Fisheries Society, 99(4):836– 847, October 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

OHeeron:1975:CTS

[OE75] Michael K. O’Heeron, Jr. and David B. Ellis. A comprehensive time series model for studying the effects reservoir management on fish populations. Transactions of the American Fisheries Society, 104(3):591–595, July 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Overholtz:1977:HOE

[OFTM77] William J. Overholtz, Arlo W. Fast, Richard A. Tubb, and Richard Miller. Hypolimnion oxygenation and its effects on the depth distribution of rainbow trout (Salmo gairdneri) and gizzard shad (Dorosoma cepedianum). Transactions of the American Fisheries Society, 106(4):371–375, July 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

OGrodnick:1979:SVS

[O’G79] Joseph J. O’Grodnick. Susceptibility of various salmonids to whirling disease (Myxosoma cerebralis). Transactions of the American Fisheries Society, 108(2):187–190, March 1979. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ogden:1970:RAF

[Ogd70] John C. Ogden. Relative abundance, food habits, and age of the American eel, Anguilla rostrata (LeSueur), in certain New Jersey streams. Transactions of the American Fisheries Society, 99(1): 54–59, January 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Oliver:1979:OMF

[OHC79] J. Douglas Oliver, George F. Holeton, and Kian E. Chua. Overwinter mortality of fingerling smallmouth bass in relation to size, relative energy stores, and environmental temperature. Transactions of the American Fisheries Society, 108(2):130–136, March 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ohno:1970:EDG

[Ohn70] Susumu Ohno. The enormous diversity in genome sizes of fish as a reflection of nature’s extensive experiments with gene duplication. Transactions of the American Fisheries Society, 99(1):120– 130, January 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Olmsted:1978:AGS

[OK78] Larry L. Olmsted and Raj V. Kilambi. Age and growth of spotted bass (Micropterus punctulatus) in Lake Fort Smith, Arkansas. Transactions of the American Fisheries Society, 107(1):21–25, January 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Otto:1976:LPT

[OKR76] Robert G. Otto, Max A. Kitchel, and John O’Hara Rice. Lethal and preferred temperatures of the alewife (Alosa pseudoharengus) in Lake Michigan. Transactions of the American Fisheries Society, 105(1):96–106, January 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Otwell:1975:SGJ

[OM75] W. Steven Otwell and John V. Merriner. Survival and growth of juvenile striped bass, Morone saxatilis, in a factorial experiment with temperature, salinity and age. Transactions of the American Fisheries Society, 104(3):560–566, July 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Orach-Meza:1978:APD

[OMS78] Faustino L. Orach-Meza and Saul B. Saila. Application of a polynomial distributed lag model to the Maine lobster fishery. Transactions of the American Fisheries Society, 107(3):402–411, May 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Olson:1973:PIE

[OP73] Robert E. Olson and Ivan Pratt. Parasites as indicators of English sole (Parophrys vetulus) nursery grounds. Transactions of the American Fisheries Society, 102(2):405–411, April 1973. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Otto:1977:RFS

[OR77] Robert G. Otto and John O’Hara Rice. Responses of a freshwater sculpin (Cottus cognatus gracilis) to temperature. Transactions of the American Fisheries Society, 106(1):89–94, January 1977. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Orciari:1979:RRG

[Orc79] Robert D. Orciari. Rotenone resistance of golden shiners from a periodically reclaimed pond. Transactions of the American Fisheries Society, 108(6):641–645, November 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Orth:1979:CSM

[Ort79] Donald J. Orth. Computer simulation model of the population dynamics of largemouth bass in Lake Carl Blackwell, Oklahoma. Transactions of the American Fisheries Society, 108(3):229–240, May 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Oseid:1972:SER

[OS72] Donavon Oseid and Lloyd L. Smith, Jr. Swimming endurance and resistance to copper and malathion of bluegills treated by longterm exposure to sublethal levels of hydrogen sulfide. Transactions of the American Fisheries Society, 101(4):620–625, October 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Oseid:1974:CTH

[OS74] Donavon M. Oseid and Lloyd L. Smith, Jr. Chronic toxicity of hydrogen sulfide to Gammarus pseudolimnaeus. Transactions of the American Fisheries Society, 103(4):819–822, October 1974. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Olla:1978:EET

[OS78] Bori L. Olla and Carol Samet. Effects of elevated temperature on early embryonic development of the tautog, Tautoga onitis. Transactions of the American Fisheries Society, 107(6):820–824, November 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Oseid:1977:CFG

[Ose77] Donavon M. Oseid. Control of fungus growth on fish eggs by Asellus militaris and Gammarus pseudolimnaeus. Transactions of the American Fisheries Society, 106(2):192–195, March 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Otto:1974:EAC

[Ott74] Robert G. Otto. The effects of acclimation to cyclic thermal regimes on heat tolerance of the western mosquitofish. Transactions of the American Fisheries Society, 103(2):331–335, April 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Paris:1970:SFB

[Par70a] Doris Fort Paris. Separation of fish brain proteins on acrylamide gel. Transactions of the American Fisheries Society, 99(1):94–95, January 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Parker:1970:SDF

[Par70b] Richard O. Parker, Jr. Surfacing of dead fish following application of rotenone. Transactions of the American Fisheries Society, 99 (4):805–807, October 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Parsons:1971:SFP

[Par71] John W. Parsons. Selective food preferences of walleyes of the 1959 year class in Lake Erie. Transactions of the American Fisheries Society, 100(3):474–485, July 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Parsons:1972:LHP

[Par72] John W. Parsons. Life history and production of walleyes of the 1959 year-class in Western Lake Erie, 1959–62. Transactions of the American Fisheries Society, 101(4):655–661, October 1972. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pardue:1973:PRB

[Par73a] Garland B. Pardue. Production response of the bluegill sunfish, Lepomis macrochirus Rafinesque, to added attachment surface for fish-food organisms. Transactions of the American Fisheries Society, 102(3):622–626, July 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Parker:1973:RMR

[Par73b] Faust R. Parker, Jr. Reduced metabolic rates in fishes as a result of induced schooling. Transactions of the American Fisheries Society, 102(1):125–131, January 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Patten:1973:BIC

[Pat73] Benjamin G. Patten. Biological information on copper rockfish in Puget Sound, Washington. Transactions of the American Fisheries Society, 102(2):412–416, April 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Patten:1975:RCE

[Pat75] Bernard C. Patten. A reservoir cove ecosystem model. Transactions of the American Fisheries Society, 104(3):596–619, July 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Patriarche:1977:BBM

[Pat77] Mercer H. Patriarche. Biological basis for management of lake whitefish in the Michigan waters of Northern Lake Michigan. Transactions of the American Fisheries Society, 106(4):295–308, July 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pollard:1973:EAH

[PB73] Herbert A. Pollard II and Ted C. Bjornn. The effects of angling and hatchery trout on the abundance of juvenile steelhead trout. Transactions of the American Fisheries Society, 102(4):745–752, October 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Phillips:1978:RCM

[PB78] Glenn R. Phillips and Donald R. Buhler. The relative contributions of methylmercury from food or water to rainbow trout (Salmo gairdneri) in a controlled laboratory environment. Transactions of the American Fisheries Society, 107(6):853–861, November 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pope:1973:IFD

[PCP73] G. F. Pope, J. C. H. Carter, and G. Power. The influence of fish on the distribution of Chaoborus spp. (Diptera) and density of larvae in the Matamek River System, Québec. Transactions of the American Fisheries Society, 102(4):707–714, October 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Peck:1970:SRC

[Pec70] James W. Peck. Straying and reproduction of Coho salmon, Oncorhynchus kisutch, planted in a Lake Superior tributary. Transactions of the American Fisheries Society, 99(3):591–595, July 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Peck:1974:MFH

[Pec74] James W. Peck. Migration, food habits, and predation on yearling Coho salmon in a Lake Michigan tributary and bay. Transactions of the American Fisheries Society, 103(1):10–14, January 1974. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pettit:1977:CRS

[Pet77] Stephen W. Pettit. Comparative reproductive success of caughtand-released and unplayed hatchery female steelhead trout (Salmo gairdneri) from the Clearwater River, Idaho. Transactions of the American Fisheries Society, 106(5):431–435, September 1977. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Park:1972:CJS

[PF72] Donn L. Park and Winston E. Farr. Collection of juvenile salmon and steelhead trout passing through orifices in gatewells of turbine intakes at Ice Harbor Dam. Transactions of the American Fisheries Society, 101(2):381–384, April 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pflieger:1978:DSG

[Pfl78] William L. Pflieger. Distribution and status of the grass carp (Ctenopharyngodon idella) in Missouri streams. Transactions of the American Fisheries Society, 107(1):113–118, January 1978. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pruitt:1977:AEP

[PGP77] Gary W. Pruitt, Billy J. Grantham, and Richard H. Pierce, Jr. Accumulation and elimination of pentachlorophenol by the bluegill, Lepomis macrochirus. Transactions of the American Fisheries Society, 106(5):462–465, September 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Plumb:1975:CCV

[PGSP75] J. A. Plumb, O. L. Green, R. O. Smitherman, and G. B. Pardue. Channel catfish virus experiments with different strains of channel catfish. Transactions of the American Fisheries Society, 104(1): 140–143, January 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Powell:1972:OIK

[PH72] Guy C. Powell and Charles L. Hurd. Observations of initial king crab egg attachment. Transactions of the American Fisheries Society, 101(3):557–559, July 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Peters:1974:RMM

[PH74] D. S. Peters and D. E. Hoss. A radioisotopic method of measuring food evacuation time in fish. Transactions of the American Fisheries Society, 103(3):626–629, July 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pierce:1977:FSW

[PH77] David E. Pierce and Arnold B. Howe. A further study on winter flounder group identification off Massachusetts. Transactions of the American Fisheries Society, 106(2):131–139, March 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pirie:1977:ASA

[PHF^{+}77] Walter R. Pirie, Wayne A. Hubert, D. H. Fickeisen, M. J. Schneider, and J. C. Montgomery. Assumptions in statistical analysis. Transactions of the American Fisheries Society, 106(6):646–648, November 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Phinney:1975:RES

[Phi75] Duane E. Phinney. Repopulation of an eradicated stream section by brook trout. Transactions of the American Fisheries Society, 104(4):685–687, October 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pister:1974:DFT

[Pis74] Edwin P. Pister. Desert fishes and their habitats. Transactions of the American Fisheries Society, 103(3):531–540, July 1974. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pitcher:1973:SFM

[Pit73] T. J. Pitcher. Some field measurements on minnow schools. Transactions of the American Fisheries Society, 102(4):840–843, October 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Post:1974:SRT

[PKP74] George Post, Thomas M. Kloppel, and Dean V. Power. Survival of rainbow trout eggs after receiving physical shocks of known magnitude. Transactions of the American Fisheries Society, 103(4):711– 716, October 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pasch:1972:FHT

[PL72] R. W. Pasch and J. H. Lyford, Jr. The food habits of two species of Cottus occupying the same habitat. Transactions of the American Fisheries Society, 101(2):377–381, April 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Phillips:1975:SEG

[PLCM75] Robert W. Phillips, Richard L. Lantz, Errol W. Claire, and John R. Moring. Some effects of gravel mixtures on emergence of Coho salmon and steelhead trout fry. Transactions of the American Fisheries Society, 104(3):461–466, July 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Plumb:1972:VCE

[Plu72] John A. Plumb. A virus-caused epizootic of rainbow trout (Salmo gairdneri) in Minnesota. Transactions of the American Fisheries Society, 101(1):121–123, January 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Powers:1975:DMR

[PLZ75] Joseph E. Powers, Robert T. Lackey, and James R. Zuboy. Decision-making in recreational fisheries management: an analysis. Transactions of the American Fisheries Society, 104(3):630–634, July 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Powers:1974:CFE

[Pow74] Joseph E. Powers. Competition for food: an evaluation of Ivlev’s model. Transactions of the American Fisheries Society, 103(4): 772–776, October 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pontius:1973:FHM

[PP73] Richard W. Pontius and Michael Parker. Food habits of the mountain whitefish, Prosopium williamsoni (Girard). Transactions of the American Fisheries Society, 102(4):764–773, October 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Paul:1979:PCF

[PPSF79] A. J. Paul, J. M. Paul, P. A. Shoemaker, and H. M. Feder. Prey concentrations and feeding response in laboratory-reared stage-one zoeae of king crab, snow crab, and pink shrimp. Transactions of the American Fisheries Society, 108(5):440–443, September 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Plumb:1977:PLS

[PR77] J. A. Plumb and R. W. Richburg. Pesticide levels in sera of moribund channel catfish from a continuous winter mortality. Transactions of the American Fisheries Society, 106(2):185–188, March 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Priegel:1970:FWB

[Pri70] Gordon R. Priegel. Food of the white bass, Roccus chrysops, in Lake Winnebago, Wisconsin. Transactions of the American Fisheries Society, 99(2):440–443, April 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Priegel:1971:ARG

[Pri71] Gordon R. Priegel. Age and rate of growth of the white bass in Lake Winnebago, Wisconsin. Transactions of the American Fisheries Society, 100(3):567–569, July 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Priegel:1975:AGY

[Pri75a] Gordon R. Priegel. Age and growth of the yellow bass in Lake Poygan, Wisconsin. Transactions of the American Fisheries Society, 104(3):513–515, July 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Prince:1975:PCD

[Pri75b] Eric D. Prince. Pinnixid crabs in the diet of young-of-the-year copper rockfish (Sebastes caurinus). Transactions of the American Fisheries Society, 104(3):539–540, July 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Price:1977:AGU

[Pri77] Robert Price. Age and growth of an unbalanced bluegill population in Michigan. Transactions of the American Fisheries Society, 106 (4):331–333, July 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Peterka:1970:LWC

[PS70] John J. Peterka and Lloyd L. Smith, Jr. Lake whitefish in the commercial fishery of Red Lakes, Minnesota. Transactions of the American Fisheries Society, 99(1):28–43, January 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Penrose:1976:TDR

[PS76] W. R. Penrose and W. R. Squires. Two devices for removing supersaturating gases in aquarium systems. Transactions of the American Fisheries Society, 105(1):116–118, January 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Paulson:1977:LVP

[PS77] Alan C. Paulson and Ronald L. Smith. Latitudinal variation of Pacific herring fecundity. Transactions of the American Fisheries Society, 106(3):244–247, May 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Policansky:1979:ELH

[PS79] David Policansky and Paul Sieswerda. Early life history of the starry flounder, Platichthys stellatus, reared through metamorphosis in the laboratory. Transactions of the American Fisheries Society, 108(3):326–327, May 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Plumb:1974:SSM

[PSG^{+}74] J. A. Plumb, J. H. Schachte, J. L. Gaines, W. Peltier, and B. Carroll. Streptococcus sp. from marine fishes along the Alabama and Northwest Florida Coast of the Gulf of Mexico. Transactions of the American Fisheries Society, 103(2):358–361, April 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pullen:1970:CWT

[PT70] Edward J. Pullen and W. Lee Trent. Carapace width-total weight relation of blue crabs from Galveston Bay, Texas. Transactions of the American Fisheries Society, 99(4):795–798, October 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pierce:1974:ECF

[PW74] Robert J. Pierce and Thomas E. Wissing. Energy cost of food utilization in the bluegill (Lepomis macrochirus). Transactions of the American Fisheries Society, 103(1):38–45, January 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Petrosky:1975:APS

[PW75a] Charles E. Petrosky and Thomas F. Waters. Annual production by the slimy sculpin population in a small Minnesota trout stream. Transactions of the American Fisheries Society, 104(2):237–244, April 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pettit:1975:AGM

[PW75b] Stephen W. Pettit and Richard L. Wallace. Age, growth, and movement of mountain whitefish, Prosopium williamsoni (Girard), in the North Fork Clearwater River, Idaho. Transactions of the American Fisheries Society, 104(1):68–76, January 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Quast:1972:RSP

[Qua72] Jay C. Quast. Reduction in stocks of the Pacific Ocean perch, an important demersal fish off Alaska. Transactions of the American Fisheries Society, 101(1):64–74, January 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Quertermus:1975:PEF

[Que75] Carl John Quertermus, Jr. Prior experience as a factor in habitat selection by the cichlid fish Tilapia mossambica. Transactions of the American Fisheries Society, 104(4):742–751, October 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ragan:1972:IBS

[RA72a] James G. Ragan and David V. Aldrich. Infection of brown shrimp, Penaeus aztecus Ives by Prochristianella penaei Kruse (Cestoda: Trypanorhyncha) in Southeastern Louisiana bays. Transactions of the American Fisheries Society, 101(2):226–238, April 1972. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Roberts:1972:CPT

[RA72b] Raymon F. Roberts and Herbert E. Allen. The control of pH and total alkalinity or total carbonate in aquatic bioassays. Transactions of the American Fisheries Society, 101(4):752–756, October 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Radford:1973:CBA

[Rad73] D. S. Radford. A comparison between actual and reported angling times. Transactions of the American Fisheries Society, 102(3): 641–643, July 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Raymond:1979:EDI

[Ray79] Howard L. Raymond. Effects of dams and impoundments on migrations of juvenile Chinook salmon and steelhead from the Snake River, 1966 to 1975. Transactions of the American Fisheries Society, 108(6):505–529, November 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rigdon:1970:SNM

[RB70] R. H. Rigdon and Kenneth N. Baxter. Spontaneous necroses in muscles of brown shrimp, Penaeus aztecus Ives. Transactions of the American Fisheries Society, 99(3):583–587, July 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Romaire:1979:ESR

[RB79] Robert P. Romaire and Claude E. Boyd. Effects of solar radiation on the dynamics of dissolved oxygen in channel catfish ponds. Transactions of the American Fisheries Society, 108(5):473–478, September 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rigdon:1975:HWS

[RBB75] R. H. Rigdon, Kenneth N. Baxter, and Robert C. Benton. Hermaphroditic white shrimp, Penaeus setiferus, parasitized by Thelohania sp. Transactions of the American Fisheries Society, 104(2):292–295, April 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Romaire:1978:PND

[RBC78] Robert P. Romaire, Claude E. Boyd, and William J. Collis. Predicting nighttime dissolved oxygen decline in ponds used for Tilapia culture. Transactions of the American Fisheries Society, 107(6): 804–808, November 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Raleigh:1971:GCL

[RC71] R. F. Raleigh and D. W. Chapman. Genetic control in lakeward migrations of cutthroat trout fry. Transactions of the American Fisheries Society, 100(1):33–40, January 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Randolph:1976:HAS

[RC76a] Kenneth N. Randolph and Howard P. Clemens. Home areas and swimways in channel catfish culture ponds. Transactions of the American Fisheries Society, 105(6):725–730, November 1976. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Randolph:1976:SFI

[RC76b] Kenneth N. Randolph and Howard P. Clemens. Some factors influencing the feeding behavior of channel catfish in culture ponds. Transactions of the American Fisheries Society, 105(6):718–724, November 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reynolds:1976:ARL

[RC76c] William W. Reynolds and Martha E. Casterlin. Activity rhythms and light intensity preferences of Micropterus salmoides and m. dolomieui. Transactions of the American Fisheries Society, 105(3): 400–403, May 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reynolds:1978:RBC

[RCL^{+}78] William W. Reynolds, Martha E. Casterlin, Roger A. Laughlin, Robert J. Livingston, and Claude R. Cripe. Reactions of blue crabs to low pH. Transactions of the American Fisheries Society, 107(6):868–871, November 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rinne:1973:FPI

[RD73] William E. Rinne and James E. Deacon. Fluorescent pigment and immersion stain marking techniques for Lepidomeda mollispinis and Cyprinodon nevadensis. Transactions of the American Fisheries Society, 102(2):459–462, April 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reed:1971:UOP

[Ree71a] Rocer J. Reed. Underwater observations of the population density and behavior of pumpkinseed, Lepomis gibbosus (Linnaeus) in Cranberry Pond, Massachusetts. Transactions of the American Fisheries Society, 100(2):350–353, April 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reed:1971:BFS

[Ree71b] Roger J. Reed. Biology of the fallfish, Semotilus corporalis (Pisces, Cyprinidae). Transactions of the American Fisheries Society, 100 (4):717–725, October 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Regier:1971:MRM

[Reg71a] Henry A. Regier. A mark-recovery method for estimating Anglers’ catch, with an example from Lake Opeongo, Ontario. Transactions of the American Fisheries Society, 100(3):495–501, July 1971. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Regier:1971:MMM

[Reg71b] Henry A. Regier. Mary, Mephistopheles, Machiavelli, and menhaden. Transactions of the American Fisheries Society, 100(4): 804–812, October 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Regier:1974:EB

[Reg74] Henry A. Regier. Ends and beginnings. Transactions of the American Fisheries Society, 103(2):406–410, April 1974. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rehwinkel:1978:FPI

[Reh78] Bruce J. Rehwinkel. The fishery for paddlefish at intake, Montana during 1973 and 1974. Transactions of the American Fisheries Society, 107(2):263–268, March 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reid:1972:UCO

[Rei72] William F. Reid, Jr. Utilization of the crayfish Orconectes limosus as forage by white perch (Morone americana) in a Maine lake. Transactions of the American Fisheries Society, 101(4):608–612, October 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reinhart:1973:EC

[Rei73] Johanna M. Reinhart. Editorial comment. Transactions of the American Fisheries Society, 102(3):663, July 1973. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reingold:1975:EDH

[Rei75] Melvin Reingold. Effects of displacing, hooking, and releasing on migrating adult steelhead trout. Transactions of the American Fisheries Society, 104(3):458–460, July 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rawson:1974:EAP

[RF74] Mac V. Rawson, Jr. and Alfred C. Fox. An evaluation of antimycin a as a parasiticide. Transactions of the American Fisheries Society, 103(3):618–620, July 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rittenburg:1979:EAV

[RGBL79] James H. Rittenburg, Margie Lee Gallagher, Robert C. Bayer, and Dale F. Leavitt. The effect of Aerococcus viridans (var.) homari on the oxygen binding capacity of hemocyanin in the American lobster (Homarus americanus). Transactions of the American Fisheries Society, 108(2):172–177, March 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Regier:1973:TBE

[RH73] Henry A. Regier and H. Francis Henderson. Towards a broad ecological model of fish communities and fisheries. Transactions of the American Fisheries Society, 102(1):56–72, January 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rose:1974:FHS

[RH74] Curt D. Rose and W. W. Hassler. Food habits and sex ratios of dolphin Coryphaena hippurus captured in the Western Atlantic Ocean off Hatteras, North Carolina. Transactions of the American Fisheries Society, 103(1):94–100, January 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ringler:1975:ELW

[RH75] Neil H. Ringler and James D. Hall. Effects of logging on water temperature, and dissolved oxygen in spawning beds. Transactions of the American Fisheries Society, 104(1):111–121, January 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ruelle:1977:PPS

[RH77] Richard Ruelle and Patrick L. Hudson. Paddlefish (Polyodon spathula): Growth and food of young of the year and a suggested technique for measuring length. Transactions of the American Fisheries Society, 106(6):609–613, November 1977. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rose:1975:ECP

[RHW75] Curt D. Rose, Alva H. Harris, and Burt Wilson. Extensive culture of penaeid shrimp in Louisiana salt-marsh impoundments. Transactions of the American Fisheries Society, 104(2):296–307, April 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Roth:1971:ATE

[RI71] J. N. Roth and E. S. Iversen. Attempts to transmit experimentally the microsporidan Thelohania duorara, parasitizing the pink shrimp, Penaeus duorarum. Transactions of the American Fisheries Society, 100(2):369–370, April 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Richards:1978:PTB

[RI78] F. Paul Richards and Richard M. Ibara. The preferred temperatures of the brown bullhead, Ictalurus nebulosus, with reference to its orientation to the discharge canal of a nuclear power plant. Transactions of the American Fisheries Society, 107(2):288–294, March 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Richards:1973:AGD

[Ric73] C. E. Richards. Age, growth and distribution of the black drum (Pogonias cromis) in Virginia. Transactions of the American Fisheries Society, 102(3):584–590, July 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Richkus:1975:MBG

[Ric75a] William A. Richkus. Migratory behavior and growth of juvenile anadromous alewives, Alosa pseudoharengus, in a Rhode Island drainage. Transactions of the American Fisheries Society, 104(3):

483–493, July 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Richkus:1975:RJA

[Ric75b] William A. Richkus. The response of juvenile alewives to water currents in an experimental chamber. Transactions of the American Fisheries Society, 104(3):494–498, July 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Richards:1976:CFS

[Ric76a] J. Scott Richards. Changes in fish species composition in the Au Sable River, Michigan from the 1920’s to 1972. Transactions of the American Fisheries Society, 105(1):32–40, January 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Richards:1976:AGF

[Ric76b] Sarah W. Richards. Age, growth, and food of bluefish (Pomatomus saltatrix) from East-Central Long Island Sound from July through November 1975. Transactions of the American Fisheries Society, 105(4):523–525, July 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Riley:1978:PGC

[Ril78] Dennis M. Riley. Parasites of grass carp and native fishes in Florida. Transactions of the American Fisheries Society, 107(1):207–212, January 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rye:1976:ATE

[RK76] Robert P. Rye, Jr. and Everett Louis King, Jr. Acute toxic effects of two lampricides to twenty-one freshwater invertebrates. Transactions of the American Fisheries Society, 105(2):322–326, March 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reed:1971:TLW

[RM71] Roger J. Reed and James A. McCann. Total length–weight relationships and condition factors for the Arctic grayling, Thymallus arcticus (Pallas), in Alaska. Transactions of the American Fisheries Society, 100(2):358–359, April 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Recksiek:1973:DPF

[RM73] Conrad W. Recksiek and James D. McCleave. Distribution of pelagic fishes in the Sheepscot River-Back River Estuary, Wiscasset, Maine. Transactions of the American Fisheries Society, 102 (3):541–551, July 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Robbins:1976:MRP

[RM76] Timothy W. Robbins and Dilip Mathur. The Muddy Run Pumped Storage Project: a case history. Transactions of the American Fisheries Society, 105(1):165–172, January 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Raleigh:1973:ETL

[RMG73] R. F. Raleigh, J. B. McLaren, and D. R. Graff. Effects of topical location, branding techniques and changes in hue on recognition of cold brands in centrarchid and salmonid fish. Transactions of the American Fisheries Society, 102(3):637–641, July 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Roberts:1970:ASS

[Rob70] Franklin L. Roberts. Atlantic salmon (Salmo salar) chromosomes and speciation. Transactions of the American Fisheries Society, 99(1):105–111, January 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rode:1973:ADF

[Rod73] Randall Rode. An automated device for feeding aquatic organisms. Transactions of the American Fisheries Society, 102(3):645–646, July 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reinitz:1979:VBC

[ROH79] Gary L. Reinitz, Leo E. Orme, and Frances N. Hitzel. Variations of body composition and growth among strains of rainbow trout. Transactions of the American Fisheries Society, 108(2):204–207, March 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reinitz:1978:IVL

[ROLH78] Gary L. Reinitz, Leo E. Orme, Carol A. Lemm, and Frances N. Hitzel. Influence of varying lipid concentrations with two protein concentrations in diets for rainbow trout (Salmo gairdneri). Transactions of the American Fisheries Society, 107(5):751–754, September 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Royce:1973:CIF

[Roy73a] William F. Royce. Comments on Introduction to the Fishery Sciences (D. W. Chapman). Transactions of the American Fisheries Society, 102(3):663–664, July 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See reply [Cha73].

Royce:1973:UFC

[Roy73b] William F. Royce. An undergraduate fisheries curriculum. Transactions of the American Fisheries Society, 102(3):649–651, July 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Roales:1975:IRB

[RP75] R. R. Roales and A. Perlmutter. The immune response of the blue gourami, Trichogaster trichopterus, to Proteus vulgaris. Transactions of the American Fisheries Society, 104(2):289–291, April 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Romberg:1973:RJC

[RR73] Gerald P. Romberg and William C. Renfro. Radioactivity in juvenile Columbia River salmon: a model to distinguish differences in movement and feeding habits. Transactions of the American Fisheries Society, 102(2):317–322, April 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ruesink:1975:RHL

[RS75] Robert G. Ruesink and Lloyd L. Smith, Jr. The relationship of the 96-hour LC_{50} to the lethal threshold concentration of hexavalent chromium, phenol, and sodium pentachlorophenate for fathead minnows (Pimephales promelas Rafinesque). Transactions of the American Fisheries Society, 104(3):567–570, July 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Redner:1979:AAT

[RS79] Barry D. Redner and Robert R. Stickney. Acclimation to ammonia by Tilapia aurea. Transactions of the American Fisheries Society, 108(4):383–388, July 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reimer:1974:NRS

[RSD74] Rollin D. Reimer, Kirk Strawn, and Allen Dixon. Notes on the river shrimp, Macrobrachium ohione (Smith) 1874, in the Galveston Bay System of Texas. Transactions of the American Fisheries Society, 103(1):120–126, January 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ridgway:1970:PEA

[RSL70] George J. Ridgway, Stuart W. Sherburne, and Robert D. Lewis. Polymorphism in the esterases of Atlantic herring. Transactions of the American Fisheries Society, 99(1):147–151, January 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ruelle:1977:RCF

[Rue77] Richard Ruelle. Reproductive cycle and fecundity of white bass in Lewis and Clark Lake. Transactions of the American Fisheries Society, 106(1):67–76, January 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Richkus:1979:ACA

[RW79] William A. Richkus and Howard E. Winn. Activity cycles of adult and juvenile alewives, Alosa pseudoharengus, recorded by two methods. Transactions of the American Fisheries Society, 108 (4):358–365, July 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rahn:1973:TCT

[RWC73] Hermann Rahn, O. Douglas Wangensteen, and Gerald J. Crowley. Tissue O_{2} and CO_{2} tensions of trout in high altitude lakes. Transactions of the American Fisheries Society, 102(1):132–134, January 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shetter:1970:RPR

[SA70a] David S. Shetter and Gaylord R. Alexander. Results of predator reduction on brook trout and brown trout in 4.2 miles (6.76 km) of the North branch of the Au Sable River. Transactions of the American Fisheries Society, 99(2):312–319, April 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Starostka:1970:FSB

[SA70b] Victor J. Starostka and Richard L. Applegate. Food selectivity of bigmouth buffalo, Ictiobus cyprinellus, in Lake Poinsett, South Dakota. Transactions of the American Fisheries Society, 99(3): 571–576, July 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sills:1971:IPE

[SA71] Joe B. Sills and John L. Allen. The influence of pH on the efficacy and residues of quinaldine. Transactions of the American Fisheries Society, 100(3):544–545, July 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shirley:1977:GPM

[SA77] Kenneth E. Shirley and Austin K. Andrews. Growth, production, and mortality of largemouth bass during the first year of life in Lake Carl Blackwell, Oklahoma. Transactions of the American Fisheries Society, 106(6):590–595, November 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Salli:1973:POC

[Sal73] Arne J. Salli. A portable observation chamber for the study of fish behavior. Transactions of the American Fisheries Society, 102(2): 456–459, April 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sarker:1977:FEB

[Sar77] Abdul Latif Sarker. Feeding ecology of the bluegill, Lepomis macrochirus, in two heated reservoirs of Texas. III. Time of day and patterns of feeding. Transactions of the American Fisheries Society, 106(6):596–601, November 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Savitz:1971:ESB

[Sav71] J. Savitz. Effects of starvation on body protein utilization of bluegill sunfish (Lepomis macrochirus Rafinesque) with a calculation of caloric requirements. Transactions of the American Fisheries Society, 100(1):18–21, January 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Savitz:1973:ILL

[Sav73] J. Savitz. Influence of low level handling stress on nitrogen excretion of bluegill sunfish (Lepomis machrochirus Rafinesque). Transactions of the American Fisheries Society, 102(3):629–630, July 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Savitz:1978:PDG

[Sav78] J. Savitz. Population dynamics and growth of a single age-class of largemouth bass (Micropterus salmoides) in a pond. Transactions of the American Fisheries Society, 107(3):425–431, May 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Seegert:1978:EIC

[SB78] Gregory L. Seegert and Arthur S. Brooks. The effects of intermittent chlorination on Coho salmon, alewife, spottail shiner, and rainbow smelt. Transactions of the American Fisheries Society, 107(2):346–353, March 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Seegert:1979:EMS

[SBCG79] Gregory L. Seegert, Arthur S. Brooks, John R. Vande Castle, and Kenneth Gradall. The effects of monochloramine on selected riverine fishes. Transactions of the American Fisheries Society, 108(1): 88–96, January 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stanley:1971:ETE

[SC71] Jon G. Stanley and Peter J. Colby. Effects of temperature on electrolyte balance and osmoregulation in the alewife (Alosa pseudoharengus) in fresh and sea water. Transactions of the American Fisheries Society, 100(4):624–638, October 1971. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1973:PSP

[SC73] Albert C. Smith and Harold B. Clemens. A population study by proteins from the nucleus of bluefin tuna eye lens. Transactions of the American Fisheries Society, 102(3):578–583, July 1973. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1979:EWS

[SC79] Ronald L. Smith and Jane Anne Cameron. Effect of water soluble fraction of Prudhoe Bay crude oil on embryonic development of Pacific herring. Transactions of the American Fisheries Society, 108(1):70–75, January 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schaefer:1970:MBA

[Sch70] Milner B. Schaefer. Men, birds and anchovies in the Peru Current — dynamic interactions. Transactions of the American Fisheries Society, 99(3):461–467, July 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schiewe:1974:IDA

[Sch74a] Michael H. Schiewe. Influence of dissolved atmospheric gas on swimming performance of juvenile Chinook salmon. Transactions of the American Fisheries Society, 103(4):717–721, October 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schwartz:1974:PPP

[Sch74b] Joseph J. Schwartz. Prevalence of pathogenic pseudomonad bacteria isolated from fish in a warmwater lake. Transactions of the American Fisheries Society, 103(1):114–116, January 1974. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schoumacher:1976:BCP

[Sch76] Roger Schoumacher. Biological considerations of pumped storage development: Introductory remarks. Transactions of the American Fisheries Society, 105(1):155–157, January 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shireman:1978:SLP

[SCR78] Jerome V. Shireman, Douglas E. Colle, and Roger W. Rottmann. Size limits to predation on grass carp by largemouth bass. Transactions of the American Fisheries Society, 107(1):213–215, January 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Southward:1972:EWS

[SD72a] G. Morris Southward and John Douglas. Evaluation of a winter steelhead fishery on a Western Washington river. Transactions of the American Fisheries Society, 101(1):29–34, January 1972. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Spoor:1972:EDM

[SD72b] W. A. Spoor and R. A. Drummond. An electrode for detecting movement in gradient tanks. Transactions of the American Fisheries Society, 101(4):714–715, October 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sylvester:1973:CBS

[SD73] J. R. Sylvester and A. E. Dammann. Contribution to the biology of the silk snapper Lutjanus vivanus from the Virgin Islands. Transactions of the American Fisheries Society, 102(4):843–845, October 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Snyder:1978:DIM

[SD78] Darrel E. Snyder and Susan C. Douglas. Description and identification of mooneye, Hiodon tergisus, protolarvae. Transactions of the American Fisheries Society, 107(4):590–594, July 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Storck:1978:DLF

[SDC78] Ted W. Storck, Don W. Dufford, and Kurt T. Clement. The distribution of limnetic fish larvae in a flood control reservoir in Central Illinois. Transactions of the American Fisheries Society, 107(3): 419–424, May 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shelton:1979:VGI

[SDKT79] William L. Shelton, William D. Davies, Terry A. King, and Tom J. Timmons. Variation in the growth of the initial year class of largemouth bass in West Point Reservoir, Alabama and Georgia. Transactions of the American Fisheries Society, 108(2):142–149, March 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Swarts:1978:GEF

[SDW78] Frederick A. Swarts, William A. Dunson, and James E. Wright. Genetic and environmental factors involved in increased resistance of brook trout to sulfuric acid solutions and mine acid polluted waters. Transactions of the American Fisheries Society, 107(5):651– 677, September 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schreck:1973:CTE

[SE73] Carl B. Schreck and W. Harry Everhart. Computer teaching exercises in fisheries science. Transactions of the American Fisheries Society, 102(3):656–657, July 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schneider:1977:LSH

[SEC77] James C. Schneider, Paul H. Eschmeyer, and Walter R. Crowe. Longevity, survival, and harvest of tagged walleyes in Lake Gogebic, Michigan. Transactions of the American Fisheries Society, 106 (6):566–568, November 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sekavec:1974:SFL

[Sek74] Glenn B. Sekavec. Summer foods, length–weight relationship, and condition factor of juvenile ladyfish, Elops saurus Linnaeus, from Louisiana coastal streams. Transactions of the American Fisheries Society, 103(3):472–476, July 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Selgeby:1971:BWB

[Sel71] James H. Selgeby. A basket for washing benthological samples. Transactions of the American Fisheries Society, 100(3):590–591, July 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Setzer:1970:ANH

[Set70] Paulette Y. Setzer. An analysis of a natural hybrid swarm by means chromosome morphology. Transactions of the American Fisheries Society, 99(1):139–146, January 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Seegrist:1972:EFT

[SG72] Donald W. Seegrist and Richard Gard. Effects of floods on trout in Sagehen Creek, California. Transactions of the American Fisheries Society, 101(3):478–482, July 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Scholz:1978:HRT

[SGC^{+}78] A. T. Scholz, C. K. Gosse, J. C. Cooper, R. M. Horrall, A. D. Hasler, R. I. Daly, and R. J. Poff. Homing of rainbow trout transplanted in Lake Michigan: a comparison of three procedures used for imprinting and stocking. Transactions of the American Fisheries Society, 107(3):439–443, May 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schaaf:1972:EFA

[SH72] W. E. Schaaf and G. R. Huntsman. Effects of fishing on the Atlantic menhaden stock: 1955–1969. Transactions of the American Fisheries Society, 101(2):290–297, April 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schreck:1974:SAE

[SH74a] Carl B. Schreck and M. Lloyd Hopwood. Seasonal androgen and estrogen patterns in the goldfish, Carassius auratus. Transactions of the American Fisheries Society, 103(2):375–378, April 1974. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stober:1974:TCH

[SH74b] Quentin J. Stober and Charles H. Hanson. Toxicity of chlorine and heat to pink (Oncorhynchus gorbuscha) and Chinook salmon (O. tshawytscha). Transactions of the American Fisheries Society, 103 (3):569–576, July 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Saila:1975:SAO

[SH75] Saul B. Saila and Kurt W. Hess. Some applications of optimal control theory to fisheries management. Transactions of the American Fisheries Society, 104(3):620–629, July 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Siefert:1977:SSB

[SH77] Richard E. Siefert and Lawrence J. Herman. Spawning success of the black crappie, Pomoxis nigromaculatus, at reduced dissolved oxygen concentrations. Transactions of the American Fisheries Society, 106(4):376–379, July 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shepherd:1978:AGD

[SH78] Marsha E. Shepherd and Melvin T. Huish. Age, growth, and diet of the pirate perch in a coastal plain stream of North Carolina. Transactions of the American Fisheries Society, 107(3):457–459, May 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sikora:1972:OFH

[SHD72] Walter B. Sikora, Richard W. Heard, and Michael D. Dahlberg. The occurrence and food habits of two species of hake, Urophycis regius and U. floridanus in Georgia estuaries. Transactions of the American Fisheries Society, 101(3):513–525, July 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schimmel:1974:EAL

[SHF74] Steven C. Schimmel, David J. Hansen, and Jerrold Forester. Effects of Aroclor(R) 1254 on laboratory-reared embryos and fry of sheepshead minnows (Cyprinodon variegatus). Transactions of the American Fisheries Society, 103(3):582–586, July 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stasko:1976:CMA

[SHH76] A. B. Stasko, R. M. Horrall, and A. D. Hasler. Coastal movements of adult Fraser River sockeye salmon (Oncorhynchus nerka) observed by ultrasonic tracking. Transactions of the American Fisheries Society, 105(1):64–71, January 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shireman:1974:SPV

[Shi74] Jerome V. Shireman. Serum protein variability in bluegills, Lepomis macrochirus Rafinesque, from Iowa farm ponds. Transactions of the American Fisheries Society, 103(3):630–632, July 1974. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Siefert:1972:FFL

[Sie72] Richard E. Siefert. First food of larval yellow perch, white sucker, bluegill, emerald shiner, and rainbow smelt. Transactions of the American Fisheries Society, 101(2):219–225, April 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Silverman:1975:SDB

[Sil75] Myron J. Silverman. Scale development in the bluefish, Pomatomus saltatrix. Transactions of the American Fisheries Society, 104(4): 773–774, October 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Silvert:1977:EF

[Sil77] W. Silvert. The economics of over-fishing. Transactions of the American Fisheries Society, 106(2):121–130, March 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Simmons:1972:PRU

[Sim72] George M. Simmons, Jr. A preliminary report on the use of the sequential comparison index to evaluate acid mine drainage on the macrobenthos in a pre-impoundment basin. Transactions of the American Fisheries Society, 101(4):701–713, October 1972. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Simmons:1976:SML

[Sim76] George M. Simmons, Jr. The Smith Mountain–Leesville Project: a case history. Transactions of the American Fisheries Society, 105 (1):173–180, January 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stucky:1971:GCC

[SK71] Norman P. Stucky and Harold E. Klaassen. Growth and condition of the carp and the river carpsucker in an altered environment in Western Kansas. Transactions of the American Fisheries Society, 100(2):276–282, April 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Starkey:1972:EPT

[SKFM72] Edward E. Starkey, Joel A. Kazmierski, Milford R. Fletcher, and O. Eugene Maughan. The effect of photoperiod on thermal resistance of speckled dace. Transactions of the American Fisheries Society, 101(4):715–718, October 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Siegfried:1979:DAN

[SKK79] Clifford A. Siegfried, Mark E. Kopache, and Allen W. Knight. The distribution and abundance of Neomysis mercedis in relation to the entrapment zone in the Western Sacramento–San Joaquin Delta.

Transactions of the American Fisheries Society, 108(3):262–270, May 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1973:DAP

[SL73] Barry W. Smith and R. T. Lovell. Determination of apparent protein digestibility in feeds for channel catfish. Transactions of the American Fisheries Society, 102(4):831–835, October 1973. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stanley:1978:SSG

[SL78] Jon G. Stanley and William M. Lewis. Special section: Grass carp in the United States: Introduction. Transactions of the American Fisheries Society, 107(1):104, January 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Slatick:1971:PAS

[Sla71] Emil Slatick. Passage of adult salmon and trout through an inclined pipe. Transactions of the American Fisheries Society, 100(3):448– 455, July 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Spehar:1978:CEC

[SLD78] Robert L. Spehar, Edward N. Leonard, and David L. Defoe. Chronic effects of cadmium and zinc mixtures on flagfish (Jordanella floridae). Transactions of the American Fisheries Society, 107(2):354–360, March 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1974:FMD

[SM74] R. J. F. Smith and B. D. Murphy. Functional morphology of the dorsal pad in fathead minnows (Pimephales promelas Rafinesque). Transactions of the American Fisheries Society, 103(1):65–72, January 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sherwood:1976:OTB

[SM76a] Marjorie J. Sherwood and Alan J. Mearns. Occurrence of tumorbearing Dover sole (Microstomus pacificus) off Point Arguello, California and off Baja California, Mexico. Transactions of the American Fisheries Society, 105(4):561–563, July 1976. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Swenson:1976:ITS

[SM76b] William A. Swenson and Melvin L. Matson. Influence of turbidity on survival, growth, and distribution of larval lake herring (Coregonus artedii). Transactions of the American Fisheries Society, 105(4):541–545, July 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Steffel:1976:TDC

[SMDN76] Sherry Steffel, John J. Magnuson, Andrew E. Dizon, and William H. Neill. Temperature discrimination by captive freeswimming tuna, Euthynnus affinis. Transactions of the American Fisheries Society, 105(5):588–591, September 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1970:SIA

[Smi70a] Stanford H. Smith. Species interactions of the alewife in the Great Lakes. Transactions of the American Fisheries Society, 99(4):754– 765, October 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1970:TMR

[Smi70b] Wesley E. Smith. Tolerance of Mysis relicta to thermal shock and light. Transactions of the American Fisheries Society, 99(2):418– 422, April 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1972:SAN

[Smi72a] C. Lavett Smith. A spawning aggregation of Nassau grouper, Epinephelus striatus (Bloch). Transactions of the American Fisheries Society, 101(2):257–261, April 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1972:CRT

[Smi72b] Wesley E. Smith. Culture, reproduction, and temperature tolerance of Pontoporeia affinis in the laboratory. Transactions of the American Fisheries Society, 101(2):253–256, April 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1973:DAS

[Smi73a] Allan K. Smith. Development and application of spawning velocity and depth criteria for Oregon salmonids. Transactions of the American Fisheries Society, 102(2):312–316, April 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1973:ATR

[Smi73b] Stanford H. Smith. Application of theory and research in fishery management of the Laurentian Great Lakes. Transactions of the American Fisheries Society, 102(1):156–163, January 1973. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See discussion [Mag73, Str73].

Smith:1973:TTT

[Smi73c] Wesley E. Smith. Thermal tolerance of two species of Gammarus. Transactions of the American Fisheries Society, 102(2):431–433, April 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1976:BSS

[Smi76] Stephen Lee Smith. Behavioral suppression of spawning in largemouth bass by interspecific competition for space within spawning areas. Transactions of the American Fisheries Society, 105(6):682– 685, November 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sikstrom:1975:HWS

[SML75] C. B. Sikstrom, D. A. Metner, and W. L. Lockhart. Hermaphroditism in a white sucker (Catostomus commersoni) from the Athabasca River, Alberta. Transactions of the American Fisheries Society, 104(2):413, April 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stanley:1978:RRL

[SMS78] Jon G. Stanley, W. Woodard Miley II, and David L. Sutton. Reproductive requirements and likelihood for naturalization of escaped grass carp in the United States. Transactions of the American Fisheries Society, 107(1):119–128, January 1978. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sylvester:1975:TTE

[SN75] J. R. Sylvester and Colin E. Nash. Thermal tolerance of eggs and larvae of Hawaiian striped mullet, Mugil cephalus L. Transactions of the American Fisheries Society, 104(1):144–147, January 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schoor:1976:EMB

[SN76] W. P. Schoor and S. M. Newman. The effect of mirex on the burrowing activity of the lugworm (Arenicola cristata). Transactions of the American Fisheries Society, 105(6):700–703, November 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Spoor:1971:ECR

[SND71] W. A. Spoor, T. W. Neiheisel, and R. A. Drummond. An electrode chamber for recording respiratory and other movements of free-swimming animals. Transactions of the American Fisheries Society, 100(1):22–28, January 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Spieler:1977:ETB

[SNDM77] Richard E. Spieler, Teresa A. Noeske, Victor Devlaming, and Albert H. Meier. Effects of thermocycles on body weight gain and gonadal growth in the goldfish, Carassius auratus. Transactions of the American Fisheries Society, 106(5):440–444, September 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1976:THS

[SOKEK76] Lloyd L. Smith, Jr., Donavon M. Oseid, Gary L. Kimball, and Sayed M. El-Kandelgy. Toxicity of hydrogen sulfide to various life history stages of bluegill (Lepomis macrochirus). Transactions of the American Fisheries Society, 105(3):442–449, May 1976. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sherman:1971:SVF

[SP71] Kenneth Sherman and Herbert C. Perkins. Seasonal variations in the food of juvenile herring in coastal waters of Maine. Transactions of the American Fisheries Society, 100(1):121–124, January 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sarot:1976:TZI

[SP76] Daniel A. Sarot and Alfred Perlmutter. The toxicity of zinc to the immune response of the zebrafish, Brachydanio rerio, injected with viral and bacterial antigens. Transactions of the American Fisheries Society, 105(3):456–459, May 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Spall:1970:PCC

[Spa70] Richard D. Spall. Possible cases of cleaning symbiosis among freshwater fishes. Transactions of the American Fisheries Society, 99 (3):599–600, July 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Speece:1973:TMC

[Spe73] Richard E. Speece. Trout metabolism characteristics and the rational design of nitrification facilities for water reuse in hatcheries. Transactions of the American Fisheries Society, 102(2):323–334, April 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Speirs:1974:FHL

[Spe74] Garrett D. Speirs. Food habits of landlocked salmon and brook trout in a Maine lake after introduction of landlocked alewives. Transactions of the American Fisheries Society, 103(2):396–399, April 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Spivey:1973:OCM

[Spi73] W. Allen Spivey. Optimization in complex management systems. Transactions of the American Fisheries Society, 102(2):492–499, April 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Saila:1972:CBA

[SPSF72] S. B. Saila, T. T. Polgar, D. J. Sheehy, and J. M. Flowers. Correlations between alewife activity and environmental variables at a fishway. Transactions of the American Fisheries Society, 101(4): 583–594, October 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stira:1976:MMA

[SRM76] R. J. Stira, R. J. Reed, and S. Middleman. A method for making artificial fish eggs. Transactions of the American Fisheries Society, 105(2):349–350, March 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Spigarelli:1973:TSE

[SRT73] S. A. Spigarelli, G. P. Romberg, and R. E. Thorne. A technique for simultaneous, echo location of fish and thermal plume mapping.

Transactions of the American Fisheries Society, 102(2):462–466, April 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stickney:1971:STC

[SS71] Robert R. Stickney and Bill A. Simco. Salinity tolerance of catfish hybrids. Transactions of the American Fisheries Society, 100(4): 790–792, October 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sissenwine:1974:RIS

[SS74] Michael P. Sissenwine and Saul B. Saila. Rhode Island sound dredge spoil disposal and trends in the floating trap fishery. Transactions of the American Fisheries Society, 103(3):498–506, July 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Serns:1975:AGB

[SS75] Steven L. Serns and Kirk Strawn. Age and growth of bluegill, Lepomis macrochirus, in two heated Texas reservoirs. Transactions of the American Fisheries Society, 104(3):506–512, July 1975. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1976:EDE

[SS76] Edward J. Smith and Jan L. Sykora. Early developmental effects of lime-neutralized iron hydroxide suspensions on brook trout and Coho salmon. Transactions of the American Fisheries Society, 105(2):308–312, March 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Strange:1978:CCC

[SSE78] Richard J. Strange, Carl B. Schreck, and Richard D. Ewing. Cortisol concentrations in confined juvenile Chinook salmon (Oncorhynchus tshawytscha). Transactions of the American Fisheries Society, 107(6):812–819, November 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Strange:1977:CSR

[SSG77] Richard J. Strange, Carl B. Schreck, and James T. Golden. Corticoid stress responses to handling and temperature in salmonids. Transactions of the American Fisheries Society, 106(3):213–218, May 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Saksena:1972:ETG

[SSH72] Vishnu P. Saksena, Charles Steinmetz, Jr., and Edward D. Houde. Effects of temperature on growth and survival of laboratory-reared larvae of the scaled sardine, Harengula pensacolae Goode and Bean. Transactions of the American Fisheries Society, 101(4):691–695, October 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Swanson:1975:CCC

[ST75] George A. Swanson and William H. Thornsberry. A circular chamber for counting benthic invertebrates. Transactions of the American Fisheries Society, 104(2):406–407, April 1975. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sztramko:1977:AVF

[ST77] L. Sztramko and G. C. Teleki. Annual variations in the fecundity of yellow perch from Long Point Bay, Lake Erie. Transactions of the American Fisheries Society, 106(6):578–582, November 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stauffer:1972:AGD

[Sta72] Thomas M. Stauffer. Age, growth, and downstream migration of juvenile rainbow trout in a Lake Michigan tributary. Transactions of the American Fisheries Society, 101(1):18–28, January 1972. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stanley:1974:NPB

[Sta74] Jon G. Stanley. Nitrogen and phosphorus balance of grass carp, Ctenopharyngodon idella, fed elodea, Egeria densa. Transactions of the American Fisheries Society, 103(3):587–592, July 1974. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stanley:1976:PHA

[Sta76] Jon G. Stanley. Production of hybrid, androgenetic, and gynogenetic grass carp and carp. Transactions of the American Fisheries Society, 105(1):10–16, January 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stasiak:1978:RAG

[Sta78] Richard H. Stasiak. Reproduction, age, and growth of the finescale dace, chrosomus neogaeus, in Minnesota. Transactions of the American Fisheries Society, 107(5):720–723, September 1978. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stauffer:1979:EDP

[Sta79a] Thomas M. Stauffer. Effects of DDT and PCB’s on survival of lake trout eggs and fry in a hatchery and in Lake Michigan, 1973–1976. Transactions of the American Fisheries Society, 108(2):178–186, March 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stauffer:1979:TYC

[Sta79b] Thomas M. Stauffer. Two-year cycles of abundance of age-0 rainbow trout in Lake Superior tributaries. Transactions of the American Fisheries Society, 108(6):542–547, November 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stevenson:1971:CBS

[Ste71] J. C. Stevenson. Communication between scientists through primary journals. Transactions of the American Fisheries Society, 100(2):398–399, April 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stephens:1973:TPN

[Ste73] Robert Rhea Stephens. Twinning in Porichthys notatus Girard. Transactions of the American Fisheries Society, 102(4):846, October 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stewart:1975:IES

[Ste75] Kenneth W. Stewart. An improved elutriator for separating stream insects from stony substrates. Transactions of the American Fisheries Society, 104(4):821–823, October 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stevens:1977:SBM

[Ste77] Donald E. Stevens. Striped bass (Morone saxatilis) year class strength in relation to river flow in the Sacramento–San Joaquin Estuary, California. Transactions of the American Fisheries Society, 106(1):34–42, January 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stickney:1976:FHG

[Sti76] Robert R. Stickney. Food habits of Georgia estuarine fishes II. Symphurus plagiusa (Pleuronectiformes: Cynoglossidae). Transactions of the American Fisheries Society, 105(2):202–207, March 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stobo:1972:EDT

[Sto72a] Wayne T. Stobo. The effects of dart tags on yellow perch. Transactions of the American Fisheries Society, 101(2):365–366, April 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stobo:1972:EFL

[Sto72b] Wayne T. Stobo. Effects of formalin on the length and weight of yellow perch. Transactions of the American Fisheries Society, 101(2):362–364, April 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stroud:1973:ATR

[Str73] Richard H. Stroud. Application of theory and research in freshwater and marine fisheries management programs — a discussion of the papers given by S. H. Smith, R. O. Anderson, K. R. Allen and G. B. Talbot. Transactions of the American Fisheries Society, 102(1):188–193, January 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See [All73, And73, Smi73b, Tal73, Mag73].

Strauss:1979:REI

[Str79] Richard E. Strauss. Reliability estimates for Ivlev’s electivity index, the forage ratio, and a proposed linear index of food selection. Transactions of the American Fisheries Society, 108(4):344–352, July 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1973:VIV

[SVL73] E. N. Smith, B. J. Venables, and O. T. Lind. A versatile and inexpensive variable voltage pulsator for electrofishing. Transactions of the American Fisheries Society, 102(2):468–469, April 1973. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Swedberg:1970:SEL

[SW70] Donald V. Swedberg and Charles H. Walburg. Spawning and early life history of the freshwater drum in Lewis and Clark Lake, Missouri River. Transactions of the American Fisheries Society, 99(3): 560–570, July 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1974:ENT

[SW74] Charlie E. Smith and Warren G. Williams. Experimental nitrite toxicity in rainbow trout and Chinook salmon. Transactions of the American Fisheries Society, 103(2):389–390, April 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stickney:1975:ATC

[SW75] Robert R. Stickney and David B. White. Ambicoloration in tank cultured flounder, Paralichthys dentatus. Transactions of the American Fisheries Society, 104(1):158–160, January 1975. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sylvester:1973:ELV

[Syl73] J. R. Sylvester. Effect of light on vulnerability of heat-stressed sockeye salmon to predation by Coho salmon. Transactions of the American Fisheries Society, 102(1):139–142, January 1973. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Szluha:1974:PEF

[Szl74] Adam T. Szluha. Potamologica1 effects of fish hatchery discharge. Transactions of the American Fisheries Society, 103(2):226–234, April 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tagatz:1976:EMP

[Tag76] Marlin E. Tagatz. Effect of mirex on predator–prey interaction in an experimental estuarine ecosystem. Transactions of the American Fisheries Society, 105(4):546–549, July 1976. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Talbot:1973:CSA

[Tal73] G. B. Talbot. The California sardine-anchovy fisheries. Transactions of the American Fisheries Society, 102(1):178–187, January 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See discussion [Mag73, Str73].

Tang:1970:EBB

[Tan70] Y. A. Tang. Evaluation of balance between fishes and available fish foods in multispecies fish culture ponds in Taiwan. Transactions of the American Fisheries Society, 99(4):708–718, October 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tarrant:1971:SVA

[Tar71] Robert M. Tarrant, Jr. Seasonal variation in the accumulation and loss of ^{131}I by tissues of adult female channel catfish, Ictalurus punctatus (Rafinesque). Transactions of the American Fisheries Society, 100(2):237–246, April 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tarby:1974:CYP

[Tar74] Martin J. Tarby. Characteristics of yellow perch cannibalism in Oneida Lake and the relation to first year survival. Transactions of the American Fisheries Society, 103(3):462–471, July 1974. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tatum:1973:BWC

[Tat73] Walter M. Tatum. Brackish water cage culture of rainbow trout (Salmo gairdneri), in South Alabama. Transactions of the American Fisheries Society, 102(4):826–828, October 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Taube:1976:SMF

[Tau76] Clarence M. Taube. Sexual maturity and fecundity in brown trout of the Platte River, Michigan. Transactions of the American Fisheries Society, 105(4):529–533, July 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Taubert:1977:EMD

[Tau77] Bruce D. Taubert. Early morphological development of the green sunfish, Lepomis cyanellus, and its separation from other larval Lepomis species. Transactions of the American Fisheries Society, 106(5):445–448, September 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Topp:1971:TUS

[TB71] Robert W. Topp and Dale S. Beaumariage. Two unusual striped mullet, Mugil cephalus, from the Florida West Coast. Transactions of the American Fisheries Society, 100(3):579–580, July 1971. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Trojnar:1974:MIE

[TB74] John R. Trojnar and Robert J. Behnke. Management implications of ecological segregation between two introduced populations of cutthroat trout in a small Colorado lake. Transactions of the American Fisheries Society, 103(3):423–430, July 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tucker:1977:RBP

[TB77] Craig S. Tucker and Claude E. Boyd. Relationships between potassium permanganate treatment and water quality. Transactions of the American Fisheries Society, 106(5):481–488, September 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tucker:1978:CPA

[TB78] Craig S. Tucker and Claude E. Boyd. Consequences of periodic applications of copper sulfate and simazine for phytoplankton control in catfish ponds. Transactions of the American Fisheries Society, 107(2):316–320, March 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tomlinson:1979:DAF

[TB79] J. W. C. Tomlinson and P. S. Brown. Decision analysis in fish hatchery management. Transactions of the American Fisheries Society, 108(2):121–129, March 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tucker:1979:EFR

[TBM79] Luther Tucker, Claude E. Boyd, and Edward W. McCoy. Effects of feeding rate on water quality, production of channel catfish, and economic returns. Transactions of the American Fisheries Society, 108(4):389–396, July 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Turner:1972:DAY

[TC72] Jerry L. Turner and Harold K. Chadwick. Distribution and abundance of young-of-the-year striped bass, Morone saxatilis, in relation to river flow in the Sacramento–San Joaquin Estuary. Transactions of the American Fisheries Society, 101(3):442–452, July 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Toneys:1979:SRF

[TC79] Michael L. Toneys and Daniel W. Coble. Size-related, first winter mortality of freshwater fishes. Transactions of the American Fisheries Society, 108(4):415–419, July 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thompson:1976:OAG

[TD76] Gerald E. Thompson and Ronald W. Davies. Observations on the age, growth, reproduction, and feeding of mountain whitefish (Prosopium williamsoni) in the Sheep River, Alberta. Transactions of the American Fisheries Society, 105(2):208–219, March 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Trotzky:1974:EWF

[TG74] Howard M. Trotzky and Richard W. Gregory. The effects of water flow manipulation below a hydroelectric power dam on the bottom fauna of the Upper Kennebec River, Maine. Transactions of the American Fisheries Society, 103(2):318–324, April 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Toews:1979:EEP

[TG79] D. R. Toews and J. S. Griffith. Empirical estimates of potential fish yield for the Lake Bangweulu System, Zambia, Central Africa. Transactions of the American Fisheries Society, 108(3):241–252, May 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tackett:1974:RCO

[TGR74] Dewey L. Tackett, John J. Giudice, and R. W. Raible. A respirometer with controllable oxygen levels. Transactions of the American Fisheries Society, 103(4):825–828, October 1974. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thomas:1973:SMI

[Tho73a] Allan E. Thomas. Spawning migration and intragravel movement of the Torrent sculpin, Cottus rhotheus. Transactions of the American Fisheries Society, 102(3):620–622, July 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thorson:1973:SDN

[Tho73b] Thomas B. Thorson. Sexual dimorphism in number of rostral teeth of the sawfish, Pristis perotteti Müller and Henle, 1841. Transactions of the American Fisheries Society, 102(3):612–614, July 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thomas:1975:EEC

[Tho75] Allan E. Thomas. Effect of egg concentration in an incubation channel on survival of Chinook salmon fry. Transactions of the American Fisheries Society, 104(2):335–337, April 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Timms:1972:LRM

[TK72] A. M. Timms and H. Kleerekoper. The locomotor responses of male Ictalurus punctatus, the channel catfish, to a pheromone released by the ripe female of the species. Transactions of the American Fisheries Society, 101(2):302–310, April 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Takatalo:1970:EIP

[TKM70] S. Takatalo, S. Kolehmainen, and J. K. Miettinen. Excretion of iodine by perch (Perca fluviatilis). Transactions of the American Fisheries Society, 99(3):515–517, July 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Titlow:1974:DCI

[TL74] Franklin B. Titlow and Robert T. Lackey. Dam: a computerimplemented water resource teaching game. Transactions of the American Fisheries Society, 103(3):601–609, July 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Theiling:1976:AGR

[TL76] Dale L. Theiling and Harold A. Loyacano, Jr. Age and growth of red drum from a saltwater marsh impoundment in South Carolina. Transactions of the American Fisheries Society, 105(1):41–44, January 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thornton:1978:TAM

[TL78] Kent W. Thornton and Allan S. Lessem. A temperature algorithm for modifying biological rates. Transactions of the American Fisheries Society, 107(2):284–287, March 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Targett:1974:SAF

[TM74] Timothy E. Targett and James D. McCleave. Summer abundance of fishes in a Maine tidal cove with special reference to temperature. Transactions of the American Fisheries Society, 103(2):325–330, April 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thurston:1978:ATA

[TRS78] Robert V. Thurston, Rosemarie C. Russo, and Charlie E. Smith. Acute toxicity of ammonia and nitrite to cutthroat trout fry. Transactions of the American Fisheries Society, 107(2):361–368, March 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tsai:1972:LHE

[Tsa72] Chu-Fa Tsai. Life history of the Eastern Johnny darter, Etheostoma olmstedi Storer, in cold tailwater and sewage-polluted water. Transactions of the American Fisheries Society, 101(1):80–88, January 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tsai:1973:WQF

[Tsa73] Chu-Fa Tsai. Water quality and fish life below sewage outfalls. Transactions of the American Fisheries Society, 102(2):281–292, April 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tsai:1978:CCM

[Tsa78] Shan-Ching Tsai. Control of chironomids in milkfish (Chanos chanos) ponds with Abater(R) (Temephos) insecticide. Transactions of the American Fisheries Society, 107(3):493–499, May 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thomas:1972:IRB

[TSW72] Paul M. Thomas, Bob G. Sanders, and Kay L. Wiley. Immune response to bacteriophage antigens in the thornback (Platyrhinoides triseriata). Transactions of the American Fisheries Society, 101 (1):115–121, January 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tompkins:1976:STL

[TT76] James A. Tompkins and Chu-Fa Tsai. Survival time and lethal exposure time for the blacknose dace exposed to free chlorine and chloramines. Transactions of the American Fisheries Society, 105 (2):313–321, March 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Turner:1977:IVC

[Tur77] R. Eugene Turner. Intertidal vegetation and commercial yields of penaeid shrimp. Transactions of the American Fisheries Society, 106(5):411–416, September 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Talbot:1975:FPB

[TW75] Gerald B. Talbot and Paul G. Wares. Fishery for Pacific billfish off Southern California and Mexico, 1903–69. Transactions of the American Fisheries Society, 104(1):1–12, January 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tyus:1974:MSA

[Tyu74] Harold M. Tyus. Movements and spawning of anadromous alewives, Alosa pseudoharengus (Wilson) at Lake Mattamuskeet, North Carolina. Transactions of the American Fisheries Society, 103(2):392–396, April 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Utter:1978:PSG

[UF78] Fred Utter and Leroy Folmar. Protein systems of grass carp: Allelic variants and their application to management of introduced populations. Transactions of the American Fisheries Society, 107 (1):129–134, January 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Utter:1972:BGV

[UH72] Fred M. Utter and Harold O. Hodgins. Biochemical genetic variation at six loci in four stocks of rainbow trout. Transactions of the American Fisheries Society, 101(3):494–502, July 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Umminger:1972:SCS

[UM72] Bruce L. Umminger and John B. Mahoney. Seasonal changes in the serum chemistry of the winter flounder, Pseudopleuronectes americanus. Transactions of the American Fisheries Society, 101 (4):746–748, October 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Valdez:1974:TPT

[Val74] Richard A. Valdez. Two parasites of threespine stickleback from Amchitka, Aleutian Islands, Alaska. Transactions of the American Fisheries Society, 103(3):632–635, July 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vanderwalker:1970:RFC

[Van70] John G. Vanderwalker. Responses of fingerling Coho and Chinook salmon to modified flows in a simulated turbine intake. Transactions of the American Fisheries Society, 99(3):532–539, July 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanVelson:1974:SSR

[Van74] Rodney C. Van Velson. Self-sustaining rainbow trout (Salmo gairdneri) population in McConaughy Reservoir, Nebraska. Transactions of the American Fisheries Society, 103(1):59–64, January 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanGuelpen:1979:SMW

[VD79] Louis Van Guelpen and Charles C. Davis. Seasonal movements of the winter flounder, Pseudopleuronectes americanus, in two contrasting inshore locations in Newfoundland. Transactions of the American Fisheries Society, 108(1):26–37, January 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vinikour:1977:INR

[Vin77] William S. Vinikour. Incidence of Neascus rhinichthysi (Trematoda: Diplostomatidae) on longnose dace, Rhinichthys cataractae (Pisces: Cyprinidae), related to fish size and capture location. Transactions of the American Fisheries Society, 106(1):83–88, January 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Villella:1970:CPP

[VIS70] John B. Villella, Edwin S. Iversen, and Carl J. Sindermann. Comparison of the parasites of pond-reared and wild pink shrimp (Penaeus duorarum Burkenroad) in South Florida. Transactions of the American Fisheries Society, 99(4):789–794, October 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Voyer:1975:EDO

[Voy75] Richard A. Voyer. Effect of dissolved oxygen concentration on the acute toxicity of cadmium to the mummichog, Fundulus heteroclitus (L.), at various salinities. Transactions of the American Fisheries Society, 104(1):129–134, January 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vogele:1975:UBS

[VR75] Louis E. Vogele and William C. Rainwater. Use of brush shelters as cover by spawning black basses (Micropterus) in Bull Shoals Reservoir. Transactions of the American Fisheries Society, 104(2): 264–269, April 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vaughan:1976:MDM

[VS76] Douglas S. Vaughan and Saul B. Saila. A method for determining mortality rates using the Leslie matrix. Transactions of the American Fisheries Society, 105(3):380–383, May 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Voorhees:1979:ASS

[VS79] James T. Voorhees and Frank J. Schwartz. Attachment site, seasonality, and effects of the parasitic copepod Lernaeenicus radiatus on two estuarine fishes in the Cape Fear River, North Carolina. Transactions of the American Fisheries Society, 108(2):191–196, March 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Voigtlander:1974:FHY

[VW74] Clyde W. Voigtlander and Thomas E. Wissing. Food habits of young and yearllng white bass, Morone chrysops (Rafinesque), in Lake Mendota, Wisconsin. Transactions of the American Fisheries Society, 103(1):25–31, January 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

vanWinkle:1978:DDF

[vWDB78] Webster van Winkle, Donald L. Deangelis, and Stephen R. Blum. A density-dependent function for fishing mortality rate and a method for determining elements of a Leslie matrix with density-dependent parameters. Transactions of the American Fisheries Society, 107 (3):395–401, May 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wright:1970:PLT

[WA70] James E. Wright, Jr. and Louisa M. Atherton. Polymorphisms for LDH and transferrin loci in brook trout populations. Transactions of the American Fisheries Society, 99(1):179–192, January 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weithman:1977:SGP

[WA77] A. Stephen Weithman and Richard O. Anderson. Survival, growth, and prey of Esocidae in experimental systems. Transactions of the American Fisheries Society, 106(5):424–430, September 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wagner:1972:UAI

[Wag72] Wilbert C. Wagner. Utilization of alewives by inshore piscivorous fishes in Lake Michigan. Transactions of the American Fisheries Society, 101(1):55–63, January 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Walters:1971:MFP

[Wal71] Carl J. Walters. Manipulation of fish populations with respect to food production: a cautionary note. Transactions of the American Fisheries Society, 100(2):365–366, April 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Walburg:1972:SFA

[Wal72] Charles H. Walburg. Some factors associated with fluctuation in year-class strength of sauger, Lewis and Clark Lake, South Dakota. Transactions of the American Fisheries Society, 101(2):311–316, April 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wallace:1973:ETD

[Wal73a] Charles R. Wallace. Effects of temperature on developing meristic structures of smallmouth bass, Micropterus dolomieui Lacépède. Transactions of the American Fisheries Society, 102(1):142–144, January 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wallace:1973:RSM

[Wal73b] Dale C. Wallace. Reproduction of the silverjaw minnow, ericymba buccata cope. Transactions of the American Fisheries Society, 102 (4):786–793, October 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wallace:1975:CCB

[Wal75a] Deborah N. Wallace. A critical comparison of the biological assumptions of Hudson River striped bass models and field survey data. Transactions of the American Fisheries Society, 104(4):710– 717, October 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Waltemyer:1975:ETM

[Wal75b] David L. Waltemyer. The effect of tannin on the motility of walleye (Stizostedion vitreum) spermatozoa. Transactions of the American Fisheries Society, 104(4):808–810, October 1975. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Waltemyer:1976:TAE

[Wal76] David L. Waltemyer. Tannin as an agent to eliminate adhesiveness of walleye eggs during artificial propagation. Transactions of the American Fisheries Society, 105(6):731–736, November 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Westin:1979:SAB

[WAMR79] D. T. Westin, K. J. Abernethy, L. E. Meller, and B. A. Rogers. Some aspects of biology of the American sand lance, Ammodytes americanus. Transactions of the American Fisheries Society, 108 (3):328–331, May 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Warner:1973:SFC

[War73] Kendall Warner. Spring food of chain pickerel (Esox niger) in Maine lakes. Transactions of the American Fisheries Society, 102 (1):149–151, January 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Warner:1976:HML

[War76] Kendall Warner. Hooking mortality of landlocked Atlantic salmon, Salmo salar, in a hatchery environment. Transactions of the American Fisheries Society, 105(3):365–369, May 1976. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Warner:1978:HML

[War78] Kendall Warner. Hooking mortality of lake-dwelling landlocked Atlantic salmon, Salmo salar. Transactions of the American Fisheries Society, 107(4):518–522, July 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Watson:1970:MJA

[Wat70] Jay F. Watson. Marking juvenile American shad by immersion in Bismarck Brown Y and Neutral Red. Transactions of the American Fisheries Society, 99(3):600–602, July 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Watarai:1973:GRC

[Wat73] Lloyd T. Watarai. Growth rate of a carangid fish, the Omaka Caranx mate, in Hawaii. Transactions of the American Fisheries Society, 102(3):617–620, July 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Watson:1976:EST

[Wat76] John W. Watson, Jr. Electrical shrimp trawl catch efficiency for Penaeus duorarum and Penaeus aztecus. Transactions of the American Fisheries Society, 105(1):135–148, January 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wakeman:1979:OCE

[WAWR79] J. M. Wakeman, C. R. Arnold, D. E. Wohlschlag, and S. C. Rabalais. Oxygen consumption, energy expenditure, and growth of the red snapper (Lutjanus campechanus). Transactions of the American Fisheries Society, 108(3):288–292, May 1979. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wurtsbaugh:1975:FDU

[WBG75] Wayne A. Wurtsbaugh, Robert W. Brocksen, and Charles R. Goldman. Food and distribution of underyearling brook and rainbow trout in Castle Lake, California. Transactions of the American Fisheries Society, 104(1):88–95, January 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wilson:1979:COE

[WC79] J. Larry Wilson and Kirby D. Cottrell. Catchability and organoleptic evaluation of grass carp in East Tennessee ponds. Transactions of the American Fisheries Society, 108(1):97–99, January 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wickham:1971:ACA

[WCH71] Donald A. Wickham, Jimmy B. Cagle, and Frank Hightower, Jr. Apparatus for controlling ambient light cycles in experimental environments. Transactions of the American Fisheries Society, 100 (1):128–131, January 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Webster:1976:UWF

[WE76] Dwight A. Webster and Gudny Eiriksdottir. Upwelling water as a factor influencing choice of spawning sites by brook trout (Salvelinus fontinalis). Transactions of the American Fisheries Society, 105(3):416–421, May 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weaver:1974:TTF

[Wea74] James E. Weaver. Temporal trends in fatty acid composition of juvenile Atlantic menhaden fed brine shrimp nauplii. Transactions of the American Fisheries Society, 103(2):382–386, April 1974. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Windell:1978:EFS

[WFS78] John T. Windell, Jeffrey W. Foltz, and John A. Sarokon. Effect of fish size, temperature, and amount fed on nutrient digestibility of a pelleted diet by rainbow trout, Salmo gairdneri. Transactions of the American Fisheries Society, 107(4):613–616, July 1978. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wissing:1971:ESA

[WH71] Thomas E. Wissing and Arthur D. Hasler. Effects of swimming activity and food intake on the respiration of young-of-the-year white bass, Morone chrysops. Transactions of the American Fisheries Society, 100(3):537–543, July 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Walter:1975:MEY

[WH75] Gilbert Walter and Walter J. Hoagman. A method for estimating year class strength from abundance data with application to the fishery of Green Bay, Lake Michigan. Transactions of the American Fisheries Society, 104(2):245–255, April 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Woltering:1978:PPI

[WHW78] D. M. Woltering, J. L. Hedtke, and L. J. Weber. Predator–prey interactions of fishes under the influence of ammonia. Transactions of the American Fisheries Society, 107(3):500–504, May 1978. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wickham:1973:ACC

[Wic73] Donald A. Wickham. Attracting and controlling coastal pelagic fish with nightlights. Transactions of the American Fisheries Society, 102(4):816–825, October 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Williams:1970:SFJ

[Wil70] Wells E. Williams. Summer food of juvenile black bullheads (Ictalurus m. melas) of Mitchell Lake, Wexford County, Michigan. Transactions of the American Fisheries Society, 99(3):597–599, July 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Williams:1974:TNS

[Wil74] Ernest H. Williams, Jr. Two new species of Monobothrium (Cestoda: Carophyllaeidae) from catostomid fishes of the Southeastern United States. Transactions of the American Fisheries Society, 103 (3):610–615, July 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wilson:1975:HCG

[Wil75] J. Larry Wilson. Hemoglobin comparisons of green sunfish (Lepomis cyanellus Rafinesque) × bluegill (L. macrochirus Rafinesque) hybrids to the parental species. Transactions of the American Fisheries Society, 104(1):148–149, January 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wilbur:1978:TTF

[Wil78] Robert L. Wilbur. Two types of fish attractors compared in Lake Tohopekaliga, Florida. Transactions of the American Fisheries Society, 107(5):689–695, September 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Winter:1977:SHR

[Win77] Jimmy D. Winter. Summer home range movements and habitat use by four largemouth bass in Mary Lake, Minnesota. Transactions of the American Fisheries Society, 106(4):323–330, July 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wise:1972:YPR

[Wis72] John P. Wise. Yield-per-recruit estimates for Eastern tropical Atlantic yellowfin tuna. Transactions of the American Fisheries Society, 101(1):75–79, January 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wissing:1974:ETY

[Wis74] Thomas E. Wissing. Energy transformations by young-of-the-year white bass Morone chrysops (Rafinesque) in Lake Mendota, Wisconsin. Transactions of the American Fisheries Society, 103(1): 32–37, January 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Warner:1978:MLA

[WJ78] Kendall Warner and Paul R. Johnson. Mortality of landlocked Atlantic salmon (Salmo salar) hooked on flies and worms in a river nursery area. Transactions of the American Fisheries Society, 107 (6):772–775, November 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weithman:1979:TFQ

[WK79] A. Stephen Weithman and Shriniwas K. Katti. Testing of fishing quality indices. Transactions of the American Fisheries Society, 108(3):320–325, May 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Windell:1976:TRG

[WKN^{+}76] John T. Windell, James F. Kitchell, David O. Norris, James S. Norris, and Jeffrey W. Foltz. Temperature and rate of gastric evacuation by rainbow trout, Salmo gairdneri. Transactions of the American Fisheries Society, 105(6):712–717, November 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wellborn:1970:OAP

[WL70] Thomas L. Wellborn, Jr. and Robert D. Lindsey. Occurrence of anchor parasites (Lernaea cyprinacea L.) on adult bullfrogs (Rana catesbeiana Shaw). Transactions of the American Fisheries Society, 99(2):443–444, April 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Warden:1975:MLB

[WL75] Robert L. Warden, Jr. and Wendell J. Lorio. Movements of largemouth bass (Micropterus salmoides) in impounded waters as determined by underwater telemetry. Transactions of the American Fisheries Society, 104(4):696–702, October 1975. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wolfert:1978:AGF

[WM78] David R. Wolfert and Terence J. Miller. Age, growth, and food of northern pike in Eastern Lake Ontario. Transactions of the American Fisheries Society, 107(5):696–702, September 1978. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wallace:1979:TE

[WNH79] Deborah Wallace, Logan A. Norris, and Glifford L. Hawkes. TCDD in the environment. Transactions of the American Fisheries Society, 108(1):103–109, January 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wolfert:1977:AGW

[Wol77] David R. Wolfert. Age and growth of the walleye in Lake Erie, 1963–1968. Transactions of the American Fisheries Society, 106 (6):569–577, November 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Worsham:1975:NPL

[Wor75] R. L. Worsham. Nitrogen and phosphorus levels in water associated with a channel catfish (Ictalurus punctatus) feeding operation. Transactions of the American Fisheries Society, 104(4):811–815, October 1975. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wagner:1973:PPL

[WP73] John F. Wagner and Michael Parker. Primary production and limiting nutrients in a small, subalpine Wyoming lake. Transactions of the American Fisheries Society, 102(4):698–706, October 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Watters:1976:EWS

[WP76] Kenneth W. Watters and Thomas E. Prinslow. An emergency water supply for small aquarium systems. Transactions of the American Fisheries Society, 105(2):347–348, March 1976. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Walters:1973:PPA

[WV73] Carl J. Walters and Robert E. Vincent. Potential productivity of an alpine lake as indicated by removal and reintroduction of fish. Transactions of the American Fisheries Society, 102(4):675–697, October 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wojtalik:1970:SEH

[WW70] Thomas A. Wojtalik and Thomas F. Waters. Some effects of heated water on the drift of two species of stream invertebrates. Transactions of the American Fisheries Society, 99(4):782–788, October 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wadley:1971:XRT

[WW71a] Gerald W. Wadley and Arthur D. Welander. X-rays and temperature: Combined effects on mortality and growth of salmon embryos. Transactions of the American Fisheries Society, 100(2): 267–275, April 1971. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wilkes:1971:ADD

[WW71b] Frank G. Wilkes and Charles M. Weiss. The accumulation of DDT by the dragonfly nymph, Tetragoneuria. Transactions of the American Fisheries Society, 100(2):222–236, April 1971. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wilbur:1973:THK

[WW73] Robert L. Wilbur and Eugene W. Whitney. Toxicity of the herbicide Kuron(R) (Silvex) to bluegill eggs and fry. Transactions of the American Fisheries Society, 102(3):630–633, July 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weis:1975:RFR

[WW75] Judith S. Weis and Peddrick Weis. Retardation of fin regeneration in Fundulus by several insecticides. Transactions of the American Fisheries Society, 104(1):135–137, January 1975. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Watt:1976:R

[WW76a] Kenneth E. F. Watt and Deborah N. Wallace. Reviews. Transactions of the American Fisheries Society, 105(2):353–360, March 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Woodward:1976:AGF

[WW76b] Ronald L. Woodward and Thomas E. Wissing. Age, growth, and fecundity of the quillback (Carpiodes cyprinus) and highfin (C. velifer) carpsuckers in an Ohio stream. Transactions of the American Fisheries Society, 105(3):411–415, May 1976. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Warlen:1977:ARD

[WWLC77] Stanley M. Warlen, Douglas A. Wolfe, Curtis W. Lewis, and David R. Colby. Accumulation and retention of dietary 14 C-DDT by Atlantic menhaden. Transactions of the American Fisheries Society, 106(1):95–104, January 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wydoski:1976:PRH

[WWN76] R. S. Wydoski, G. A. Wedemeyer, and N. C. Nelson. Physiological response to hooking stress in hatchery and wild rainbow trout (Salmo gairdneri). Transactions of the American Fisheries Society, 105(5):601–606, September 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wickham:1973:EMA

[WWO73] Donald A. Wickham, John W. Watson, Jr., and Larry H. Ogren. The efficacy of midwater artificial structures for attracting pelagic sport fish. Transactions of the American Fisheries Society, 102(3): 563–572, July 1973. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Yeh:1977:RSF

[Yeh77] Chi Fu Yeh. Relative selectivity of fishing gear used in a large reservoir in Texas. Transactions of the American Fisheries Society, 106(4):309–313, July 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Yakupzack:1977:EJA

[YHP77] Paul M. Yakupzack, W. H. Herke, and W. Guthrie Perry. Emigration of juvenile Atlantic croakers, Micropogon undulatus, from a semi-impounded marsh in Southwestern Louisiana. Transactions of the American Fisheries Society, 106(6):538–544, November 1977. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Youngs:1972:ENF

[You72] William D. Youngs. Estimation of natural and fishing mortality rates from tag recaptures. Transactions of the American Fisheries Society, 101(3):542–545, July 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Youngs:1974:EFA

[You74] William D. Youngs. Estimation of the fraction of anglers returning tags. Transactions of the American Fisheries Society, 103(3):616– 618, July 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Youngs:1976:AES

[You76] W. D. Youngs. An analysis of the effect of seasonal variability of harvest on the estimate of exploitation rate. Transactions of the American Fisheries Society, 105(1):45–47, January 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Yingst:1979:EDL

[YS79] William L. Yingst III and Robert R. Stickney. Effects of dietary lipids on fatty acid composition of channel catfish fry. Transactions of the American Fisheries Society, 108(6):620–625, November 1979. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zaugg:1970:CRB

[Zau70] W. S. Zaugg. Comments on the relationship between gill AT-Pase activities, migration, and salt water adaption of Coho salmon.

Transactions of the American Fisheries Society, 99(4):811–813, October 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zolczynski:1976:GCN

[ZD76] Stephen J. Zolczynski, Jr. and William D. Davies. Growth characteristics of the northern and Florida subspecies of largemouth bass and their hybrid, and a comparison of catchability between the subspecies. Transactions of the American Fisheries Society, 105(2):240–243, March 1976. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zimmerman:1970:GFB

[Zim70] Charles J. Zimmerman. Growth and food of the brook silverside, Labidesthes sicculus, in Indiana. Transactions of the American Fisheries Society, 99(2):435–438, April 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zeitoun:1974:ECS

[ZT74] Ibrahim H. Zeitoun and Peter I. Tack. The embryology of the Coho salmon, Oncorhynchus kisutch (Walbaum). Transactions of the American Fisheries Society, 103(2):371–375, April 1974. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).